A Complete Bibliography of Publications in IEEE Transactions on Emerging Topics in Computing

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 April 2024 Version 1.03

Title word cross-reference

AB + CAB + C [XHWI22]. ϵδ [ZZZY24]. K [SCD^{+}21, LRHL21]. N [BM20, ZWZ18]. O(N) [KK21]. P [TSS16, PSZD21].

-Aided [OKK23a]. -Based [HSzXZ17]. -Bit [PSZD21]. -Means [SCD^{+}21]. -Modular [BM20]. -Oriented [KSKA22]. -Space [KK21]. -Stability [TSS16]. -Utility-Optimized [ZZZY24]. -Version [ZWZ18].

/DAC [MBW23].

1 [TKTP21]. 123.0 [TKTP21]. 123.0-B-1 [TKTP21]. 1X [ZZM^{+}19]. 1X-2X [ZZM^{+}19].

2 [AXAD21, CKC^{+}18, GVF^{+}23]. 2021 [IEE21]. 2022 [IEE22, Mon22]. 28th [IEE21]. 29th [IEE22]. 2X [ZZM^{+}19].

3 [SBD^{+}24]. 3-D [SBD^{+}24]. 3-Phase [FM21a]. 3.3 [TKTP21]. 32-Bit [GWD^{+}23]. 3D [PS21, ADN^{+}21, DBOB20, DZK^{+}23, ELVC18, GHSMM21, GKC21, HEYB22, LK19, LYW^{+}21, NYC^{+}21, PASK21, SLR^{+}17, VC17, WASW22, YJ22]. 3D-ICs [NYC^{+}21]. 3D-NoC [DBOB20]. 3D-Stacked [PS21]. 3DL [SBD^{+}24]. 3DL-PIM [SBD^{+}24]. 3G [GAPG16]. 3PIPs [LRYK14].

4.0 [CDF^{+}22b, JSBM22].

512 [ET23]. 5G [SAJ22, AZA^{+}21, QLT17]. 5G-Network-Connected [SAJ22].

60 [SSL^{+}13]. 65-nm [TAC^{+}19]. 6G [FK22, RSK21, RLK22]. 6T [TAC^{+}19].

754 [FM21b, FM21c].

802.11e [CYBD15]. 802.15.4 [AAEKM13].

Abnormal [HDA^{+}20]. Abstract [FSK20]. Abundance [CKR^{+}22]. Academia [WVC21]. Academic [XCW^{+}14, ZLW^{+}21]. ACBAM [RAD22]. ACBAM-Accuracy-Configurable [RAD22]. Accelerated [CRS17]. Accelerating [AEM22, CKR^{+}22, HKC22, ZLY^{+}21]. Acceleration [AmSYlrm24, DQB23, IPRR19, SBD^{+}21]. Accelerator [ATPR23, AMTR24, KKM17, NAMJ23, QGF^{+}23, SBR^{+}22, TKTP21, WH21, WMAB17, YJ22]. Accelerator-Rich [SBR^{+}22]. Accelerators [CMB18, DBLK23, JHB21, RPL^{+}23, SZA^{+}23]. Acceptance [PTT21]. Access [CPLdFM21, CYBD15, DGP22, FNK^{+}13, GSVA23, HTH^{+}22, IPiR18, LCZ21, OZL15, RZD^{+}19, RLK22, RSM^{+}21, ZMK22]. Accuracy [BAS23, RAD22]. accurate [ISI^{+}19]. Achieving [AKHA22, CLWX18]. Acoustic [LPZ^{+}14]. Acquisition [CDF^{+}22b, GLZ19]. Across [ABTH20, HHM20, PG23]. Acting [Mon20]. Action [FVS24, MDB^{+}23]. Activation [ABC18, KK23, MBW23, SAD23]. Active [CWZ^{+}22, LT21, XWZ^{+}23]. Activities [BC16, CFM^{+}22, CCH^{+}23, GSMGP17, RSH^{+}21, RS20, ZYY^{+}24]. Activity [DTK^{+}23, WLO^{+}21, WC22, XCX^{+}20]. Ad [GZFS18, KNK13, LHB^{+}15, LYS13]. Ad-Hoc [LYS13, KNK13]. Adaptive [AFMM19, HZY21, JDP^{+}21, LYA18, LSP^{+}20, MCB21, NBS16, SAM^{+}20, SSV^{+}20, TFM^{+}19, TMS^{+}19, TAV15, TW22, WLG^{+}21, WHL23a, ZLG^{+}22]. ADC [AKSR24, LVR^{+}24, MBW23]. ADC- [MBW23]. ADC-/DAC-Free [MBW23]. ADC-Less [AKSR24]. Adder [GWD^{+}23, SK23a, TZZB21]. Adders [DPP21a, DPP21b, LZM^{+}21, PP23b, SMMTBM^{+}22, SK23a, VC17]. Addition [HAKL22]. Additional [CGPB21]. Additive [CVP^{+}22, XTXY16]. Address [CLW16]. Addressable [GT22, RFKK23]. Adiabatic [FM21a, KTM19]. ADIP [RMVN22]. Adjusting [SKKN20]. Adopting [FMA24]. Adults [RdPF^{+}23]. Advance [LCLK20]. Advanced [FO21, KVP19]. Advancement [AT22, SKKN20]. Advances [DLRL24, HSS24, RWZ^{+}16, SG15]. Advection [BBW22]. Adversarial [OPIV^{+}23, WJL23, XWZ^{+}23, XYH^{+}23, ZLW^{+}22]. AdvParams [XWZ^{+}23]. AES [EMDE^{+}22]. AFDA [ZC15]. Affecting [OBM22]. Affective [NWSG17]. Affects [CG17]. Aframe [HZS^{+}15]. Against [CWZ^{+}22, JBSS^{+}22, KKY24, LWN^{+}22, XAQ22, XYH^{+}23, LRXW21]. Age [TRGVR^{+}22, YZG^{+}24]. Agent [BBC16, HZS^{+}15]. Agent-Based [BBC16, HZS^{+}15]. Aggregate [BSJ22]. Agile [ZL22]. Aging [AC22, GAWT23, JYZ^{+}23, KK14, OEM18, QZW^{+}23, RRFT16, SKKN20]. Aging-Based [KK14]. Aging-Related [QZW^{+}23]. Aging-Resistant [RRFT16]. Agnostic [YMT22]. Agreement [JCM^{+}21, THTK16]. Agriculture [VKBB22]. Ahead [IP21]. AI [RPL^{+}23]. Aid [TD19]. Aided [OKK23a, TFM^{+}19, ZMK22, KMK22]. AIG [BBMM23]. AIG-Rewriting [BBMM23]. AIO [LZD^{+}22]. Air [LGT^{+}19]. Airline [BBC16]. AlexNet [AJL^{+}21]. Algebraic [BWMM22]. Algorithm [DBOB20, DDD20, GJZ^{+}16, GCBK17, HKC22, LCM^{+}23, LHS^{+}22, LJD^{+}15, MBW23, QCN^{+}15, TA19, WMN13, WJ19, WCC^{+}20, XTXY16, YKP22, YCLW14, ZCZ^{+}15, ZLG^{+}22].

Algorithm-Hardware [HKC22, LCM^{+}23]. Algorithmic [NMP19]. Algorithmic-Level [NMP19]. Algorithms [BPBG18, DDB22, FAT^{+}14, FM21b, FM21c, GZB22, JM16, ZMA15]. Alignment [GDP22]. Alleviating [GVO^{+}23]. Allocation [AFMM19, BPB21, CMMF20, GZFS18, HWH^{+}20, HYZ^{+}16, JSZ18, OKK22, OKK23b, QCN^{+}15, WLC^{+}14, WJ15, WZLK21, WLL^{+}21]. ALOHA [ZC15]. ALOJA [BPC^{+}17]. Alone [CFM^{+}22]. ALP [GVO^{+}23]. ALPHA [HKC22]. Alternative [Mél22]. Always [NVS^{+}14, VSH23]. Ambient [BDL^{+}13, WLO^{+}21]. Among [BTC^{+}21, CH19, LT21, MCB21]. Analog [ASKG21, HTS^{+}23, JBSS^{+}22, VSS18]. Analog/Digital [HTS^{+}23]. Analogous [WJL23]. Analyses [STL^{+}14]. Analysis [AAO^{+}20, AKU19, AMKF21, BBM^{+}17, CSBME17, CPM^{+}17, CFL^{+}21, DLTSNA21, EGGOR20, FAT^{+}14, GMTX14, GBVS21, GAPG16, HWFR15, HB15, HK17, KZT^{+}20, LCM18, LZM^{+}21, LYW^{+}21, MYM20, MVS21, NWSG17, PPF^{+}21, PV15, QZW^{+}23, RB22, SAJ22, TSH^{+}17, WHC16, WLL^{+}22, WMP^{+}24, WHL23b, XSYW20, ZLXL22, ZLW^{+}21]. Analytical [KS18]. Analytics [BPC^{+}17, CTC^{+}17, HHM20, LK22, NLF^{+}22, PRKK24, ZCT^{+}14]. Analyze [HHS23]. Analyzing [RdPF^{+}23]. Anatomy [PG23]. Anchor [HYL^{+}20]. Annealing [BWMM22, WH21]. Annotated [PPF^{+}21]. Anomaly [ADCS22, EAM21, MS19, PJK^{+}19, SFZ^{+}16, XCY^{+}24]. Anomaly-Based [PJK^{+}19]. Anonymity [RMVN22]. Anonymization [SC22]. Anonymous [ABC18, EKO^{+}16, MR21]. Answer [AT22]. Antenna [WHC16]. Anthropogenic [YKW^{+}20]. Anthropomorphic [LK22]. Anti [XLQ^{+}22, YDH21]. Anti-Counterfeit [YDH21]. Anti-Rumor [XLQ^{+}22]. AntiDOTE [LWN^{+}22]. App [LXL^{+}22a, PCG^{+}21]. AppDNA [LXL^{+}22a]. Application [AK24, CH19, DDD20, HDL14, LSO17, PCG^{+}21, PMLT21, QWC^{+}18, RB22, VKBB22]. Application-Specific [LSO17]. Applications [AM22, ACH22, BBM^{+}17, BYZZ22, BBC^{+}22, BDL^{+}13, BDFT24, CP23, CRRS22, DLTX21, DLRL24, DTK^{+}23, DBNBT14, EM23, FABC21, FWC15, GP15b, JSBM22, JM16, MT21a, OZAL13, SBD^{+}24, TDVS18, TD19, WMAB17, YCH^{+}23, ZYZ20, ZLY^{+}21]. Applied [GAWT23, LJX^{+}22, NMP19]. Applying [BKS21, CLW^{+}18]. Approach [BBMM23, CQH17, CMRV21, CKC^{+}18, GAI22, HHS23, HWL15, HDL14, HYZ^{+}16, HLX^{+}17, HJCK21, HWW^{+}17, KAF^{+}16, KHY^{+}14, LGT^{+}19, LWC22, LTL^{+}22, MTFK21, Mél22, NB21, PLA20, PASK21, RGP^{+}21, SAI^{+}19, WGYL20, WZLK21, WLWQ22, YKW^{+}20, ZLX18]. Approaches [NL21, SAAJ22, ZYZ20]. Approximate [AM22, AMTR24, ACH22, AMRCP21, BM20, BBMM23, CLLL23, FLB^{+}19, HAKL22, IPiR18, IPRR19, KKM17, KDKB22, LZM^{+}21, LXL^{+}22b, MVC^{+}22, MDB^{+}22, NHT^{+}19, NMP19, RLE^{+}22, SMMTBM^{+}22, SK23a, WWXL22, WWL^{+}22, WASW22, XAQ22, ZNS^{+}22]. Approximation [BBC^{+}22, DGW^{+}23, ZMA15]. Approximation-Aware [DGW^{+}23]. Approximation-Based [BBC^{+}22]. APUF [GLC^{+}21]. Arbiter [GLC^{+}21]. Architectural [LHKH21]. Architecture [AMG22, BHC^{+}23, BYB20, BK22, CDC^{+}23, DDB22, DQB23, ESO^{+}22, EM23, HSzXZ17, KCS23, LLW22, LVR^{+}24, MMPP15, SBD^{+}21, SBGC22, SNK^{+}14, SBD^{+}24, TDG17, YJ22]. Architectures [AFATAH13, BDFT24, CSZ^{+}20, DPO17, EDLT21, KMD^{+}18, KP14, KVP19, MM14, THTK16, WMP^{+}24, YNA^{+}20]. Area [DTK^{+}23, JCM^{+}21, LNK^{+}15, PS21, SCS^{+}20, ZWWF13]. Areas [OLL^{+}21].

Ariadne [VSS18]. ARITH [IEE21, IEE22]. Arithmetic [DRV22, FM21b, FM21c, IEE21, IEE22, JLN21, LLG^{+}23, OSPN22, PL19, Pla21a, Pla21b, XHWI22, BMP23]. AritPIM [LLG^{+}23]. ARMOR [YDH21]. Array [DBLK23, LXL^{+}22b, RAD22]. Array-Based [DBLK23]. Arrays [CMPT17, GZO^{+}18, SBGC22, TA19]. Artificial [FBL^{+}22, LRLG22, MLLU20]. ASIC [DMRR17]. Assertions [GDP22]. Assessed [CSK^{+}23]. Assessment [ADQ21, CMJ21, LRS^{+}22, OPIV^{+}23]. Assignment [CGPB21, CCW^{+}14, DBNBT14, YCS22, ZYLT19]. assist [LMW^{+}17]. Assistance [IMZ^{+}21]. Assisted [CYL^{+}15, HTH^{+}22, HZY21, SAJ22, WLO^{+}21, WSLL21, LWM^{+}21]. Assistive [OGL^{+}21]. Assists [SKM^{+}23]. Association [CGPB21, ZMA15]. Associative [AMJ22, IPiR18]. Assured [WZRR13]. Assuring [JMLH22]. Aster [NW20]. Asymmetric [AMG22, CDLS13]. Asymmetry [LRG^{+}21]. Asynchronous [CLWX18, LZD^{+}22, ZC15]. At-Risk [HWW^{+}17]. Atrial [DVPQ^{+}21]. Attack [BSJ22, CDM20, CWZ^{+}22, IMZ^{+}21, KKY24, MSAS17, SSL^{+}13, WMP^{+}24, YK18, BKS21]. Attacks [APCM20, BHB^{+}14, BBM^{+}17, BCM^{+}21, BWMM22, HHS23, HLM^{+}13, HYW19, LSP^{+}20, LMK^{+}13, LZLC22, MGR19, OPIV^{+}23, RS21, SBR^{+}22, TM14, XAQ22, XYH^{+}23, YMSR20]. Attention [XCY^{+}24, XYZ^{+}17]. Attribute [GSVA23, LCZ21, RSM^{+}21, YZZ^{+}21, YLY^{+}20]. Attribute-Based [GSVA23, LCZ21, RSM^{+}21, YLY^{+}20]. Attributes [BS15, EAM21]. Auction [JSZ18]. Auction-Based [JSZ18]. Audio [IMZ^{+}21, YKW^{+}20]. Auditing [XZW^{+}21a, YLY^{+}20]. Augmentation [FBL^{+}22]. Augmented [LMS^{+}14]. Authenticated [MLH^{+}20]. Authentication [AK24, AGMP21, ABG^{+}22, NT16, PTM23, RMK^{+}14, RB22, SWL15, SLC^{+}13, SWW^{+}20, THTK16, WSLL21, WWG^{+}22, ZSCY20]. Authority [XSY^{+}21]. Authorized [XSY^{+}21]. Authors [Ano22f, Ano23b, Ano23c, Ano23d, Ano23e, Ano24a]. Autoencoder [LVJ22]. Autoencoder-Based [LVJ22]. Automata [GAI22, KJC22, SMMTBM^{+}22]. Automata-Based [GAI22]. Automated [BCX23, GZO^{+}18, LHS^{+}22, NHT^{+}19, PKK22b]. Automatic [BC18, GDP22, IP21, KPL^{+}21, YKW^{+}20]. Autonomous [CMM^{+}21]. AutoRelax [PKK22b]. AVC [MZY^{+}16]. Avionic [FO21]. Avoidance [JM16]. AVX [ET23]. AVX-512 [ET23]. Aware [AMG22, AR20, AC22, ASYK^{+}22, BBS^{+}15, BMB^{+}18, CCW^{+}14, CTL^{+}23, DFRR23, DF22, DGW^{+}23, GLM^{+}23, GCW20, HLM^{+}13, HMSZ20, HMB^{+}21, HAE22, HZS^{+}15, HLL16, JDL^{+}23, KKM17, LC20, LFL^{+}18, LLS^{+}16, LCM18, MYM20, NWSG17, PPF^{+}21, PCXF19, RDS^{+}22, RLX15, SBM21, SP20, TBG^{+}18, WLC^{+}14, WZH^{+}16, WDO^{+}19, XL19, YCS22, YKAE22, YCLW14, YJ22, ZLW^{+}21, ZDLG13, AFMM19]. Awareness [NAM17, PLA20]. AWS [SFZ^{+}16]. AxRMs [WWXL22].

B [TKTP21, WSLL21]. B-TSCA

[WSLL21]. Back [LLLG23]. Back-End [LLLG23]. Backbone [PJK^{+}19]. Bacterial [TGDC^{+}21]. Bag [CMJ21]. Bal [PCG^{+}21]. Bal-App [PCG^{+}21]. Balanced [LLW^{+}15, LK23]. Balancing [CF19, HWSN13]. Band [TMS^{+}19]. Base [ZYZZ15]. Based [AZA^{+}21, AK24, AEM22, ATA21, AS20, AKU19, ABDL19, AL19, ADQ21, ADN^{+}21, ABG^{+}22, APH^{+}22, AMJ22, AS24, AMCMDM^{+}23, BAS23, BBMM23, BPBG18, BHPE21, BTD24, BYB20, BCD^{+}20, BPB21, BBC^{+}22, BBF18, BNCF14, BBC16, CSK^{+}23, CMB18, CDF^{+}22a, CSM19, CMJ21, CKKO20, CRP^{+}23, CQH17, CRi^{+}19, CLA^{+}22, CLS18, CCZZ21, DLRL24, DMRR17, DDB22, DBLK23, DFRR23, DY13, DZK^{+}23, EM23, FNK^{+}13, FK22, FAP21, FVS24, FKNK21, GZT^{+}22, GTC^{+}21a, GAI22, GBVS21, GAPG16, GLC^{+}21, GWYJ21, GSVA23, GSMGP17, HHS23, HYL^{+}20, HWH^{+}20, HTS^{+}23, HLW14, HDL14, HXL^{+}14, HZS^{+}15, HLL16, HK17, HSzXZ17, HQP^{+}21, HAKL22, HSG19, JYZ^{+}23, JHL21, JCM^{+}21, JSZ18, KKY24, KZT^{+}20, KNK13, KAF^{+}16, KRG^{+}17, KKM17, KK23, KK14, KHY^{+}14, KIM^{+}18, eSKZW^{+}22, LD21, LMS^{+}14, LLW22, LNK^{+}15, LCZ21, LZW21, LZX^{+}22, LQT^{+}23, LZL^{+}21, LWM^{+}21]. Based [LZM^{+}21, LXL^{+}22b, LTL^{+}22, LHS^{+}22, LZLC22, LXX^{+}21, LVJ22, LK23, LVR^{+}24, MIMMY^{+}19, MVC^{+}22, MKH24, MTFK21, MRS21, MSAS17, MLLU20, MHL17, NAMJ23, NYC^{+}21, NWSG17, OMSO20, OKK23a, PJK^{+}19, PS21, PGM^{+}21, PP23b, PPKN23, PPF^{+}21, PBTP21, PLSM20, QGF^{+}23, RM20, RPR^{+}22, RS21, RGP^{+}21, RMSD23, RSM^{+}21, SAAJ22, SBD^{+}21, SBGC22, SBAR21, SK23a, SAD23, SNHN15, SBR^{+}22, SZL^{+}22, SQX^{+}20, SZL^{+}20, SWK^{+}17, SNK^{+}14, SBD^{+}24, SAI^{+}19, TDG17, TCNC16, THTK16, TKK^{+}22, VKBB22, VGP^{+}21, WFB^{+}18, WJ15, WGYL20, Wan22, WHL23a, WXT^{+}24, WWL^{+}22, WLY^{+}21, WLL^{+}21, WTW^{+}15, WJL23, XCW^{+}14, XCX^{+}20, XSYW20, XLQ^{+}22, XHWI22, XYZ^{+}17, XL19, XAQ22, XWZ^{+}23, XYH^{+}23, YSS^{+}22, YLW23, YYX15, YMT22, YZZ^{+}21, YMDJ21, YCLW14, YLY^{+}20, YNA^{+}20, YGLO22, ZXXH13, ZCZY16, ZLZ^{+}17, ZLZ^{+}21, ZYZ20, ZLW^{+}21, ZLX^{+}14, ZZM^{+}19, AGMP21, BTK^{+}20, CSLG22, GTC^{+}21b, ISI^{+}19, WDO^{+}19, AmSYlrm24, CCH^{+}23]. Bases [BFPS22]. Basis [PKK22a]. Battery [TKK^{+}22]. Batteryless [DF22]. Bayes [SNHN15, XZW21b]. Bayesian [FLB^{+}19, KIM^{+}18]. BD2K [AFG^{+}21]. be [LRLG22, MYS17]. Beam [TMS^{+}19]. Beams [IMZ^{+}21]. Behavior [AR20, HWFR15, LXL^{+}22a, LRP^{+}22, SY19, YQZ^{+}15, ZYZZ15]. Behavioral [VS17]. Behaviors [SQX^{+}20, ZYY^{+}24]. Behaviours [CFM^{+}22]. Being [KZT^{+}20, LS21, OGL^{+}21, RSH^{+}21]. Benchmarked [KSKA22]. Benchmarking [BPC^{+}17]. Benefits [BMOS16, BDd^{+}21a, BDD^{+}21b]. Between [CYLJ21, SWK^{+}17, KZ21, RSSE20]. Beyond [BCSF17]. BF16 [OAP^{+}22]. BFV [APA^{+}21]. Bi [ZHC^{+}14]. Bi-Clustering [ZHC^{+}14]. Biasing [CLW^{+}18]. Bibliography [WVC21]. BIC [PPKN23]. Bidirectional [LQT^{+}23]. Big [FAT^{+}14, GLZ19, MMM^{+}21, RGP^{+}21, ZCT^{+}14, ZK21, AKC14, CQH17, DHPL19, GZLG14, HNGZ19, HDL14, HXL^{+}14, KHY^{+}14, LDOG16, RWZ^{+}16, SWW^{+}21, TNKM14, WZR^{+}14, XGLW21, ZLW^{+}21]. Big-Data [ZCT^{+}14]. Bijective [AAA18]. BILSTM [LQT^{+}23]. Binarized [NAMJ23, SBD^{+}21]. Binary [AMTR24, BJ21, GZB22, MBW23, PP23a, SK23b, XHWI22, BMP23]. Binary-Activation [MBW23]. Binary-to-Stochastic [PP23a]. Bio [BMB^{+}18, CMRV21, ZSCY20]. Bio-Features [ZSCY20]. Bio-Inspired [CMRV21]. Bio-Signal [BMB^{+}18]. Biological [TRD^{+}24, ZWH^{+}24]. Biometric [AK24]. Biometrics [NT16]. BIST [KWS^{+}20]. Bit [GWD^{+}23, LRL21, NW20, PP23a, PKK22a, PPKN23, PSZD21, RPL^{+}23, TZZB21, ZMRM19]. Bit-Basis [PKK22a]. Bit-Line [RPL^{+}23, ZMRM19]. Bit-Streams [PP23a]. Bitcoin [WZL21]. Bitslicing [KCS23]. Bitwise [PRKK24, RPM23]. Black [ACM16, CQH17]. Black-Box [CQH17]. Black-Hat [ACM16]. Blind [LLD^{+}15, LZW21]. Block [BAS23, BTK^{+}20, BWMM22, KKA22, MSAS17, YGLO22]. Block-Based [YGLO22, BTK^{+}20]. Block-Level [BAS23]. Blockchain [DZZ^{+}21, HHS23, KKY24, KDM^{+}21, LZL^{+}21, LWM^{+}21, LTL^{+}22, MWW^{+}21, PCAP21, VSH23, WSLL21, Wan22, XZW^{+}21a, ZLZ^{+}21]. Blockchain-Based [KKY24, LZL^{+}21, LWM^{+}21, Wan22, ZLZ^{+}21]. Blockchain-Enabled [XZW^{+}21a]. Blockchain-Powered [MWW^{+}21]. Blockchains [RMVN22]. Blocks [BMB^{+}21, HTS^{+}23, WWXL22, ZNS^{+}22]. BLSD [SBL^{+}21]. BNN [DQB23]. Board [CBG^{+}21, FO21]. Board-Level [CBG^{+}21]. Body [CBZ^{+}21, DTK^{+}23, JCM^{+}21, ZWWF13]. Boosting [AT22, VHFH^{+}22]. Booth [RAD22]. Bound [PSSZ22]. Bounded [WLO^{+}21, WLX^{+}16]. Bounding [BBW22, ZYH^{+}20]. Box [CQH17]. Brain [Ode21]. Brainwave [ZSCY20]. Brainwave-Oriented [ZSCY20]. break [TFK20]. break-resistant [TFK20]. Breeding [CCMA21]. Bridging [RSSE20]. Brightness [JDP^{+}21]. Broadcast [SCFH13]. Broadcast/Multicast [SCFH13]. Broken [RAD22]. Budget [YCS22]. Budgets [KZ21]. Buffering [LZH19]. Bugs [QZW^{+}23, XSYW20]. Building [BMB^{+}21, WWXL22, ZNS^{+}22]. Buildings [SWLG17]. Bulk [PRKK24]. Bulk-Bitwise [PRKK24]. Bundled [LZH19]. Burst [GCW20]. Bus [PKK22a]. Bypassing [MGR19].

Cache [CYLJ21, CLD^{+}24, ESO^{+}22, HAE22, MGR19, SKD20]. Caches [ATA21, AFMM19, GKC21]. Caching [FK22, FSCX17, ZLZ^{+}17, ZLG^{+}22]. Calculation [VGP^{+}21]. Calculus [SAI^{+}19]. Call [LXL^{+}22a, ZYZZ15]. CAM [IPRR19]. Camouflaging [RPK^{+}22, YMSR20]. Can [BMOS16, CSM19, GJZ^{+}16, MYS17, SDZ^{+}21]. Cancer [AYG^{+}21]. CANNON [GLM^{+}23]. Capability [MMD^{+}22a, MMD^{+}22b]. Capacitance [Isl21]. Capacitor [MM14]. Capacity [DMRR17]. Car [SSR^{+}22, VGP^{+}21]. Car-as-a-Service [SSR^{+}22]. Card [KTM19]. Carlo [VGP^{+}21]. Carlo-Based [VGP^{+}21]. Carpooling [ZHL^{+}14]. Carry [GWD^{+}23]. Cascadable [RZAD18]. Cascading [HWSN13]. Case [LRXW21, LHS^{+}22, LSR^{+}21, MHL17, PTD21, RdPF^{+}23, SSR^{+}22, SFZ^{+}16, VHFH^{+}22, WVC21]. Catalog [BBMM23]. Catalog-Based [BBMM23]. Categorization [GBVS21, HWL15]. Causal [BTC^{+}21]. Caused [ZLC16]. CCSDS [TKTP21]. Cell [FKNK21, LRL21, OKK23a, TAC^{+}19, TBG^{+}18, ZMA15]. Cell-Free [OKK23a]. Cells [GCBK17]. Cellular [HYZ^{+}16, KJC22, LLS^{+}16, SMMTBM^{+}22]. Center [AFG^{+}21, RWJZ23]. Centered [QLT17]. Centers [CYP^{+}16, GZLG14, LLL19, WZH^{+}16]. Centrality [NVS^{+}14, SAI^{+}19]. Centric [CP23, CGC^{+}24, GVO^{+}23, HWH^{+}20, LDCY21, MZGT17, SWW^{+}20, WDO^{+}19, LBX^{+}23]. Chain [TBG^{+}18]. Chains [HLX^{+}17]. Challenges [EGGOR20, LMS^{+}14, LPZ^{+}14, PCAP21, SKM^{+}23, VPTH19]. Change [RS20]. Channel [CBZ^{+}21, HCGS24, LYA18, RZD^{+}19, WMP^{+}24]. Chaos [LZLC22]. Chaos-Based [LZLC22]. Chaotic [AS24, EMDE^{+}22]. Characteristic [SWK^{+}17]. Characteristics [LK19, WZR^{+}14]. Characterization [GZO^{+}18, NdCFB^{+}23]. Characterizing [YQZ^{+}15]. Charge [IPiR18, ZYH^{+}20]. Chatbot [DVA21]. Check [HMB^{+}21]. Check-Pointing [HMB^{+}21]. Checking [CPM^{+}17, CH21, RZAD18, VS17]. Checkpointing [LWC22, MGR19, OMTH17]. Chemistry [NB21]. Chemistry-Oriented [NB21]. Chief [Lom16, Lom17, Met18, Met19, Met20, Mon20, Mon21, Mon22]. Child [TRGVR^{+}22]. Chip [BYB20, BK22, BCM^{+}21, DFRR23, EMDE^{+}22, GCW20, KMD^{+}18, KKM17, KS18, LFL^{+}18, MM14, PS21, PG23, RSSE20, ZL22]. Chips [CCZN20, CMMF20, CLLL23, CCZZ21, KWS^{+}20]. CiM [WZR^{+}23]. Cipher [GJZ^{+}16, KKA22]. Ciphers [BWMM22, MSAS17]. Circuit [ACH22, ASKG21, BSY^{+}17, CLS18, HF17, LZL^{+}21, MK15, TMCVH21, WX23]. Circuits [BBM^{+}17, CKKO20, CH19, EGR21, Geu20, KJC22, NHT^{+}19, PIK20, RPR^{+}22, STL^{+}14]. Circulating [GCBK17]. Circumvent [MGR19]. Citation [LSR^{+}21]. Cities [DGP22, PMLT21]. Citizens [RSB13]. City [LWM^{+}21]. Class [DDD20, YMT22]. Class-Agnostic [YMT22]. Classification [CH16, CLLL23, GZB22, LZX^{+}22, LXX^{+}21, PJK^{+}19, PBTP21, Sag19, SAM^{+}20, SZL^{+}20, WLZG22, XZW21b]. Classifiers [AKC14, LRHL21]. Clickstream [EGGOR20]. Client [DBNBT14]. Client-Server [DBNBT14]. Clients [AMvO17]. Clinical [GKSJ21, NLF^{+}22]. Cliques [ÁAPDD24]. Clock [EMDE^{+}22, Isl21, NBRF18]. Clock-Less [NBRF18]. Clone [HLX^{+}17]. Cloud [AZA^{+}21, AAKJJ22, AANN15, BPBG18, CWZ^{+}15, CCMA21, DHGR18, GMTX14, GZFS18, HTH^{+}22, HZQ^{+}18, IP21, JYZ^{+}23, JSZ18, KIM^{+}18, LLD^{+}15, LCZ21, MLG^{+}15, RSK21, SLB^{+}20, TRR^{+}19, TCNC16, WZRR13, WZH^{+}16, YDM^{+}18, YLY^{+}20, ZCT^{+}14, ZWZ18, ZK23, ZSL^{+}15]. Cloud-Assisted [HTH^{+}22]. Cloud-Based [BPBG18, TCNC16]. Cloudlet [TFM^{+}19]. Cloudlet-Aided [TFM^{+}19]. Cloudlets [JSZ18]. Clouds [FSCX17, KRG^{+}17, LDOG16, LLD^{+}18, RWZ^{+}16, SFZ^{+}16]. ClubCF [HDL14]. Cluster [DBOB20, DY13, FAP21]. Cluster-Based [DY13, FAP21]. Clustered [CMB18, YKP22]. Clustering [DDD20, FAT^{+}14, HDL14, HWW^{+}17, MKH24, SCD^{+}21, ZHC^{+}14, ZYZ20]. Clustering-Based [HDL14]. Clusters [Sag19]. CMOS [BBM^{+}17, BCSF17, ELVC18, LYW^{+}21, RHRV16, TAC^{+}19, YLZ^{+}21]. CMOS-Based [HTS^{+}23]. CMOS-Memristor [LYW^{+}21]. CMPs [WCG^{+}21]. CNN [AJL^{+}21, DBLK23, LXX^{+}21, QLL^{+}22, RPL^{+}23, YJ22]. CNNs [JDL^{+}23]. CNTFET [AMJ22]. CNTFET-Based [NAMJ23]. Co [AKSR24, CRS17, HKC22, LCM^{+}23, LPY^{+}13, PKK22b, PIK20, RPL^{+}23, ZYS^{+}22]. Co-Design [AKSR24, HKC22, LCM^{+}23, LPY^{+}13, RPL^{+}23]. Co-Evolutionary [ZYS^{+}22]. Co-Optimization [PKK22b, PIK20]. Co-Processor [CRS17]. COACH [HMB^{+}21]. Coalition [KIM^{+}18]. Code [CLWG15, VHFH^{+}22, YLW23]. Coded [LLL19]. Codes [BDd^{+}21a, BDD^{+}21b, CSM19, LRXW21, LK23, PPKN23]. Codesign [DZD^{+}18]. Coding [NTAL13, OMSO20, PKK22a]. Coflows [RWJZ23]. Cognitive [AT22, CRi^{+}19, KR21, LPZ^{+}14, MMY^{+}22, RZD^{+}19]. Cold [ZHC^{+}14]. Cold-Start [ZHC^{+}14]. Collaboration [ABTH20, ZLW^{+}21, ZK23]. Collaborative [FK22, GSMGP17, HINS21, HDL14, KR21, PLSM20, SSV^{+}20, TCNC16]. Collaborators [XCW^{+}14]. Collection [DAF^{+}22, LGC13, WLX^{+}16]. College [ZYY^{+}24]. Collisions [JM16]. Collusion [LCY20]. Color [CF19]. Combat [ABTH20]. Combating [PLA20]. Combinational [PIK20]. Combinations [SWK^{+}17]. Combinatorial [OKK22]. Combine [AS20]. Combined [BHB^{+}14, GHSMM21]. Combining [AKLC21]. Command [FO21].

Commitment [JCM^{+}21]. Communicating [Gel15, OLL^{+}21]. Communication [AK21, CLLL23, CBZ^{+}21, DY13, EKO^{+}16, GLM^{+}23, GCW20, LGC13, MTFK21, NTAL13, TMS^{+}20, TKK^{+}22, WQG^{+}22, WTW^{+}15, ZCZY16]. Communication-Aware [GLM^{+}23]. Communication-Efficient [WQG^{+}22]. Communications [HYZ^{+}16, WHC16]. Community [ALDK22, LDJ20, LYBZ20, MMM^{+}21, WVC21, ZLX^{+}14]. Compact [ISI^{+}19]. Comparative [MRS21, MVS21]. Comparing [Ode21]. Comparison [RPM23, RS20, SUQKA20]. Compensation [AC22, WQG^{+}22]. Compensations [KPL^{+}21]. Competencies [CTC^{+}17]. Complex [RWD22, WJ15, ZMX^{+}22]. Complexity [CBZ^{+}21, LT21, RGS20]. Components [BBMM23, SSR^{+}22]. Composition [WZY^{+}20, YCLW14]. Comprehensive [LCM18, MJ17]. Compressed [BMB^{+}18, MZY^{+}16, YYX15]. Compression [TKTP21]. Computation [AK21, AL19, BYZZ22, DZK^{+}23, DQB23, MTFK21, NT21, RTVG22, WSLL21, YNA^{+}20, ZYH^{+}20]. Computation-in-Memory [YNA^{+}20]. Computational [CSLG22, CKR^{+}22, EGGOR20, HÁMLS23, MA18, MLRRG20, WK14]. Computations [BBE^{+}24]. Compute [LVR^{+}24]. Compute-In-Memory [LVR^{+}24]. Computer [HÁMLS23, IEE21, IEE22, JLN21, LSR^{+}21, OSPN22, SK19, SO20, SO21, TRGVR^{+}22, WVC21]. Computers [AAO^{+}20, MSLL14, DPO17]. Computing [AmSYlrm24, AKSR24, AMRCP21, BM20, BPB^{+}20, BTC^{+}21, BDFT24, CP23, CWZ^{+}15, CLWG15, Che14, CSZ^{+}20, CGC^{+}24, DLTX21, DHGR18, DZZ^{+}21, DLS21, DPO17, DHPL19, DZK^{+}23, DSS21, DA20, EDLT21, EM23, FK22, GAI22, GZFS18, GZG^{+}17, HSS24, HINS21, HZQ^{+}18, HNGZ19, HSG19, IPiR18, IPRR19, JSBM22, JSZ18, JND14, JKL21, KSU16, KKM17, LMC18, LHMW20, LC20, LCZ21, LQT^{+}23, LBX^{+}23, LWM^{+}21, LSO17, LHJ18, MVC^{+}22, MWW^{+}21, MGMC21, NL21, NHT^{+}19, NMP19, OGL^{+}21, OKK22, OKK23a, OKK23b, PL19, PBKL19, RLE^{+}22, RPL^{+}23, RLK22, SAM^{+}20, SLB^{+}20, SG15, TDZ21, VSS18, WCC^{+}20, WZLK21, Wan22, WMP^{+}24, WX23, WDO^{+}19, WZR^{+}23, XYZ^{+}17, YXF^{+}22, YSS^{+}22, YDM^{+}18, ZYZZ15, ZLZ^{+}17, ZWZ18, ZEL20, ZLG^{+}22, ZMK22, ISI^{+}19, Ano13a, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15c, Ano15d, Ano15e, Ano15f, Ano16a, Ano16b, Ano16c]. Computing [Ano16d, Ano16e, Ano17a, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17k, Ano17j, Ano18a, Ano18b, Ano18c, Ano18h, Ano18d, Ano18e, Ano18f, Ano18g, Ano19a, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano20a, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano21a, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano22a, Ano22b, Ano22c, Ano22e, Ano22d, Ano22g, Ano22f, Ano23b, Ano23c, Ano23d, Ano23e, Ano24a, CP23, DPO17, DO18, FO21, HSS24, KGM15a, KGM15b]. Computing-in-Memory [WZR^{+}23]. Concealment [TS21]. Concept [DVA21, GBVS21, HB15, MMPP15, ZS15]. Conceptual [RS20]. Concurrency [AMKF21]. Condition [CCZN20]. Conditioning [LGT^{+}19]. Cone [RS21]. Cone-Based [RS21]. conference [IEE21, IEE22]. Conferences [AXAD21]. Configurable [BAS23, CMPT17, DMRR17, RAD22, WWG^{+}22]. Configuration [CDM20, KM20]. Configuration-Information [CDM20]. Connected [SAJ22, SCD^{+}21]. Connectivity [KNK13]. Consensus [LYS13]. Considering [CWZ^{+}15, LZH19]. Consistency [BBS^{+}15, HMB^{+}21, HAE22]. constant [BMP23]. Constants [BR16].

Constellations [TKK^{+}22]. Constrained [LHP^{+}18, LJX^{+}22, PPKN23, RSM^{+}21, WCG^{+}21]. Constraints [JSZ^{+}21, WLC^{+}14, XTXY16]. Constructing [QSLG22]. Construction [AMCMDM^{+}23, LK23]. Constructions [KSKA22, KKA22]. CONSUMER [LA13]. Consumption [GLFK18, KNK13, LLL^{+}20, QLL^{+}22, SSVJ14, SY19]. Contact [ZWWF15]. Content [FK22, GT22, HWH^{+}20, PCXF19, RFKK23, SWW^{+}20, WDO^{+}19, ZCZY16, ZLG^{+}22, ZS15]. Content-Aware [WDO^{+}19]. Content-Centric [HWH^{+}20, SWW^{+}20]. Contention [BBS^{+}15, FABC21]. Contents [Ano17n, Ano17o, Ano18j, Ano18m, Ano19m, Ano19n, Ano19o, Ano20k, Ano20l, Ano20m, Ano20n, Ano21k, Ano21l, Ano21m, Ano21n, Ano22j, Ano22k, Ano22l, Ano22m, Ano23j, Ano23k, Ano23l, Ano23m, Ano24c, Ano13d, Ano13e, Ano14i, Ano14j, Ano14k, Ano14l, Ano15k, Ano15l, Ano15m, Ano15n, Ano16j, Ano16k, Ano16l, Ano16m, Ano17l, Ano17m, Ano18k, Ano18l, Ano19l]. Context [ADQ21, AXAD21, BTK^{+}20, HLM^{+}13, HZS^{+}15, TAV15, ZS15]. Context-Adaptive [TAV15]. Context-Aware [HLM^{+}13, HZS^{+}15]. Context-Based [ADQ21]. Continuous [KRG^{+}17]. Continuously [WC22]. Contracts [CFL^{+}21]. Control [ADCS22, BBC16, CPLdFM21, DGP22, FNK^{+}13, FO21, GSVA23, KWS^{+}20, LCZ21, MLH^{+}20, MSLL14, RZD^{+}19, RSK21, RSM^{+}21, TMS^{+}19, TKK^{+}22, WHC16]. Controlled [ADQ21, SWK^{+}17, TCNC16]. Convention [NBRF18]. Convergecast [LLL^{+}20]. Converged [HLL16]. Converters [PP23a]. Convolution [ABDL19]. Convolutional [FVS24, HCGS24, KDKB22, LD21, LVJ22, SBD^{+}21, SZL^{+}22, YMT22]. Cooperation [YDM^{+}18, ZCZY16]. Cooperative [FNK^{+}13, KIM^{+}18, ODCZ15, PCXF19, SLC^{+}13, WLWQ22]. Coordinating [AFG^{+}21]. Coordination [BBC16, YMX22, ZOS16]. Copyright [LZL^{+}21]. Core [ABC18, BUS^{+}21, CTC^{+}17, DVPQ^{+}21, MMD^{+}22a, MMD^{+}22b, MM18, UBMA18]. Correcting [YLW23]. Correction [GHSMM21, LRXW21, PPKN23]. correctly [BMP23]. correctly-rounded [BMP23]. Correlation [HWL15, OMSO20]. Cost [AANN15, BBT^{+}16, DTK^{+}23, DRV22, GZLG14, GZG^{+}17, HWSN13, JSZ^{+}21, LGC13, LLS^{+}16, OEM18, VC17, YLZ^{+}21, YDH21]. Cost-Effective [BBT^{+}16, VC17]. Council [PGM^{+}21]. count [TMCVH21]. Counterfeit [YDH21]. Countermeasure [MSAS17]. Countermeasures [BCM^{+}21, YK18]. Coupled [XCY^{+}24]. Course [SWK^{+}17]. Coursework [HH16]. CoV [GVF^{+}23]. Cover [Ano22h, Ano22i, Ano22e, Ano23f, Ano23g, Ano23h, Ano23i, Ano24b, Ano13b, Ano13c, Ano14e, Ano14f, Ano14g, Ano14h, Ano15g, Ano15h, Ano15i, Ano15j, Ano16f, Ano16g, Ano16h, Ano16i, Ano17k, Ano18i, Ano18h, Ano21h, Ano21i, Ano21j, Ano21f, Ano21g]. Coverage [JM16, RLX15]. CPS [GFKL13a, GFKL13b, RKdN^{+}23]. CPU [GVO^{+}23]. CPU-Memory [GVO^{+}23]. CRAM [AmSYlrm24, CKR^{+}22]. CRAM-Based [AmSYlrm24]. CRAM-Seq [CKR^{+}22]. Crisis [PMC^{+}15]. Critical [AAO^{+}20, GZB22, RDS^{+}22]. Criticality [BRRE22, NSH22]. Cross [CCMA21, GLC^{+}13, WTW^{+}15]. Cross-Breeding [CCMA21]. Cross-Domain [GLC^{+}13]. Cross-Layer [WTW^{+}15]. Crossbar [JDL^{+}23, PAR18, TA19]. Crossbar-Aware [JDL^{+}23]. Crosstalk [GCW20]. Crowdsensing [CCFI16, HCCL13, WLL^{+}21, YZG^{+}24]. Crowdsensing-Oriented [HCCL13]. Crowdsourced [YW22]. Crowdsourcing [LL13, YCS22, ZYLT19]. Crypto [DLTSNA21]. Crypto-Hardware [DLTSNA21]. Cryptographic [BBM^{+}17, CCZZ21, EMDE^{+}22, KKA22]. Cryptography [CSK^{+}23, eSKZW^{+}22, XHWI22]. Cryptosystem [YYX15]. Cryptosystems [LZLC22]. Cube [LXX^{+}21]. Cube-CNN-Based [LXX^{+}21]. cuFE [HLK^{+}24]. Current [BSY^{+}17]. Curriculum [CTC^{+}17]. Custom [BDd^{+}21a, BDD^{+}21b]. Custom-Precision [BDd^{+}21a, BDD^{+}21b]. Customizable [LRXW21]. Cyber [AR20, AAEKM13, BY13, CYL^{+}15, DY13, DA20, GLC^{+}13, GZG^{+}17, GFKL13a, GFKL13b, HCCL13, JKSC21, KZT^{+}20, KNK13, KZ21, LL13, LPY^{+}13, LDJ20, LMK^{+}13, LCLK20, LCY20, LCY^{+}19, NWSG17, NBS16, SAAJ22, SBAR21, SCFH13, SLC^{+}13, SGH13, SQX^{+}20, TSH^{+}17, VHFH^{+}22, WCC^{+}20, WZY^{+}20, ZWWF13, ZLZ^{+}17, ZYZ20]. Cyber-Based [NWSG17]. Cyber-Defence [NBS16]. Cyber-Enabled [KZT^{+}20]. Cyber-Physical [AR20, BY13, CYL^{+}15, DY13, DA20, GLC^{+}13, GZG^{+}17, GFKL13a, GFKL13b, HCCL13, KNK13, LL13, LPY^{+}13, LMK^{+}13, LCY20, LCY^{+}19, SAAJ22, SCFH13, SLC^{+}13, SGH13, TSH^{+}17, VHFH^{+}22, ZWWF13]. Cyber-Physical-Social [LDJ20, WZY^{+}20, ZYZ20]. Cybercrime [ACM16]. Cybersecurity [ALDK22]. Cybersickness [MRS21]. Cycling [PIK20].

D [SBD^{+}24]. D2D [HLL16]. DAC-Free [MBW23]. DAICS [ADCS22]. Daily [CFM^{+}22, KZT^{+}20]. Damage [CSM19]. Dangerous [NVS^{+}14, SCS^{+}20]. DApp [LTL^{+}22]. Dark [CPH^{+}15]. Data [AS20, AK21, AJL^{+}21, AFG^{+}21, AAA18, BTK^{+}20, CLL21, CWK^{+}21, CYP^{+}16, CDF^{+}22b, DAF^{+}22, DDD20, EM23, FAT^{+}14, FMA24, FBL^{+}22, FO21, GVO^{+}23, GBVS21, GZLG14, GWYJ21, GLZ19, HTH^{+}22, HSG19, KZT^{+}20, KK23, LGC13, LLD^{+}15, LLD^{+}18, LGT^{+}19, LLL19, LWC22, LCR^{+}21, LK22, LK23, MLG^{+}15, MLH^{+}20, MYS17, MMM^{+}21, MVS21, MLRRG20, MGR19, NT21, Ode21, PBTP21, PBKL19, PMLT21, PCXF19, QCN^{+}15, RMVN22, RGP^{+}21, RWJZ23, SWW^{+}21, SBHL21, SCD^{+}21, SNK^{+}14, SBD^{+}24, TRR^{+}19, TFK20, TCNC16, TSS16, WLC^{+}14, WLX^{+}16, WZH^{+}16, WGYL20, Wan22, WLZG22, WLZ^{+}13, WHL23b, XLQ^{+}22, XZW^{+}21a, YKW^{+}20, YLY^{+}20, ZCT^{+}14, ZK21, ZYC^{+}22, ZYY^{+}24, ZSL^{+}15, BKS21, AKC14, CQH17, DHPL19, GZLG14, HNGZ19, HDL14, HXL^{+}14, KHY^{+}14, LDOG16, RWZ^{+}16, SWW^{+}21, TNKM14, WZR^{+}14, XGLW21, ZLW^{+}21]. Data-Driven [MLRRG20]. Data-Intensive [EM23, SBD^{+}24, WZH^{+}16]. Database [LLLG23, PRKK24, SLB^{+}20]. Databases [GSVA23, SC22, XSY^{+}21]. Datacentre [RWZ^{+}16]. Datacentres [PLA20]. Dataflow [AMKF21]. Dataset [MCUDCD22, TRGVR^{+}22]. Datasets [AKU19, AS24]. Day [WJL23]. DB [MMM^{+}21]. DBLP [WVC21]. DBNs [PSZD21]. Deadline [DFRR23, YCS22]. Deadline-Aware [DFRR23]. Deadlock [CCZN20]. Deadlock-Freedom [CCZN20]. Dealing [ZLC16]. Debug [LWN^{+}22, RM20, RMSD23]. Decentralized [JM16, MWW^{+}21, YDM^{+}18]. Decimators [GZT^{+}22]. Decision [RFKK23, RGP^{+}21, RLK22]. Decline [KR21]. Decoding [NMP19, ZZM^{+}19]. Decomposition [JYZ^{+}23, WHL23a]. Decoupling [MM14]. Dedicated [FLB^{+}19]. Deduplicatable [XZW^{+}21a]. Deep [ADCS22, AYG^{+}21, AS24, BEG^{+}24, CRRS22, DO18, HWH^{+}20, HHT23, KK22, LXL^{+}22a, MTFK21, PKK22b, SWY^{+}22, WZLK21, WLL^{+}22, WXT^{+}24, WZL21, XYDJ22, YGLO22, ZMRM19, ZXLS21, ZLW^{+}22, ZYS^{+}22, ZLG^{+}22]. Deep-Learning [LXL^{+}22a]. DeepEcho [KK22]. Deeply [CLW^{+}18, KSU16]. Deeply-Embedded [KSU16]. Deeply-Scaled [CLW^{+}18]. Defect [CVP^{+}22, LYW^{+}21, MPK18, TA19, WRT^{+}21]. Defective [TS21]. Defects [DBOB20, SMMTBM^{+}22]. Defence [NBS16]. Defense [AAKJJ22, LCLK20, XAQ22]. Defenses [HLM^{+}13]. Defined [CMK^{+}16, HCWL16, MTFK21, SBM21, TCNC16]. Defragmentation [HZQ^{+}18]. Degradation [CKKO20]. Delay [AEM22, BMC17, BYZZ22, CH21, LCM18, PP23b, ZC15]. Delay-Based [AEM22]. Delay-Sensitive [BYZZ22]. Delegated [BBE^{+}24]. Deletion [CH23]. Delivery [CPH^{+}15, CYP^{+}16, LYS13]. Delphi [NBS16]. Demand [BY13, CKC^{+}18, LYA18, LWM^{+}21, MZGT17, TMS^{+}20, ZYS^{+}22]. Demand-Side [BY13]. Dementia [CFM^{+}22]. Demographic [QWC^{+}18]. Demonstrations [RMB^{+}19]. Denial [SBR^{+}22]. Denial-of-Service [SBR^{+}22]. Denoising [HAKL22]. Dense [ASKG21]. Densely [TNKM14, UBMA18]. Dependability [MS20]. Dependable [LRLG22]. Dependencies [ÁAPDD24]. Dependency [QZW^{+}23]. Dependent [BGS^{+}21, VGP^{+}21]. Deployment [AZA^{+}21, DZK^{+}23, RSK21, SAI^{+}19, ZXXH13]. Deployments [BPC^{+}17]. Depth [SWY^{+}22, TMCVH21]. Design [AAO^{+}20, AS20, AM19, AMJ22, AC22, ACH22, AKSR24, AMRCP21, ASKG21, BBMM23, BHB^{+}14, BTK^{+}20, BPBG18, BMC17, BBT^{+}16, CDF^{+}22a, CH16, DAF^{+}22, DBOB20, DLTSNA21, DO18, EDLT21, FLB^{+}19, FABC21, FAP21, Geu20, GLC^{+}21, HKC22, JSZ^{+}21, LC20, LCM^{+}23, LPY^{+}13, LZM^{+}21, LRG^{+}21, LDCY21, LXL^{+}22b, MSAS17, MR21, MHL17, MYM20, MK15, RGS20, RPL^{+}23, RAD22, SK23a, SK19, SO20, SO21, TH16, WHC16, YATR18, YPS^{+}22, ZYMG16, ZLXL22, ZZH^{+}22, ZL22]. Design-for-Trust [RGS20]. Design-Time [AC22, BHB^{+}14]. Designated [SFL17]. Designing [APH^{+}22, CRS17, SSVJ14]. Designs [BSY^{+}17, CP23, LSO17, TMCVH21, WT20, YLZ^{+}21, YXF^{+}22, YCH^{+}23]. Detail [ZYZZ15]. Detect [ADN^{+}21]. Detectability [CPM^{+}17]. Detecting [CFM^{+}22, CCH^{+}23, SCS^{+}20]. Detection [ADCS22, APCM20, CVP^{+}22, CSK^{+}23, EAM21, FWC15, GDP22, HLX^{+}17, KPL^{+}21, KSKA22, KKA22, LYK^{+}22, LHP^{+}18, LRL21, LJX^{+}22, LA13, LYBZ20, MCUDCD22, MDB^{+}23, MYT^{+}18, MMPP15, MS19, MMM^{+}21, MDB^{+}21, OPIV^{+}23, PJK^{+}19, PKR22, SBGC22, SFZ^{+}16, TS21, TRGVR^{+}22, TPM16, VS17, WMAB17, WJL23, XCY^{+}24, YGLO22, ZCZ^{+}15, BKS21]. Detector [LD21]. Determinants [LS21]. Determined [MYS17]. Deterministic [NL21, WX23]. Develop [MLRRG20]. Developing [NLF^{+}22]. Development [ASA^{+}22, MT21a, SPCB16]. Device [CBG^{+}21, HYZ^{+}16, NdCFB^{+}23, ZCZY16]. Device-Level [CBG^{+}21]. Device-to-Device [HYZ^{+}16, ZCZY16]. Devices [BHB^{+}14, BMC17, BCSF17, CPLdFM21, DPO17, DLTSNA21, DF22, HKZH16, HF17, HAE22, MDB^{+}21, PTM23, PSZD21, RPK^{+}22, SWL15, TRGVR^{+}22]. DevOps [SFZ^{+}16]. DFRFT [WTW^{+}15]. DFT [NBRF18]. DFT-Less [NBRF18]. DiabDeep [YMDJ21]. Diabetes [YMDJ21]. Diagnosing [CQH17]. Diagnosis [HMSZ20, KR21, TAV15, YMDJ21]. DICO [MMM^{+}21]. Dielectric [NS15]. Differential [FMA24, PKR22, WHL23a, WXT^{+}24, WLY^{+}21, XZW21b, ZZZY24]. Difficulty [CSZ^{+}20]. Diffusion [FKNK21, KMK22]. Digital [CDF^{+}22a, DAF^{+}22, HTS^{+}23, HHM20, KKM17, MYM20, MMY^{+}22, RWD22, SSR^{+}22, STL^{+}14, SPCB16, WX23, ISI^{+}19]. Digitally [TGDC^{+}21]. Dimension [PJK^{+}19, WLZG22, WX23]. Dimensional [WHL23b]. Dimensionality [KHY^{+}14]. Diodes [BMB^{+}21]. Direct [FKNK21]. Directional [IMZ^{+}21, SSL^{+}13]. Disaggregation [CFM^{+}22]. Disaster [HYL^{+}20, KG20, KMK22, LNK^{+}15, TMS^{+}20, XOD20, YKW^{+}20, ZLX18, ZYS^{+}22, LMW^{+}17]. Discharging [ZMRM19]. Discipline [DLS21, LSR^{+}21]. Discover [TAV15]. Discovering [BTC^{+}21, CFL^{+}21]. Discovery [ÁAPDD24, AFG^{+}21, CLWX18, GVF^{+}23, SSL^{+}13, XSYW20]. Discrete [LCM^{+}23, WTW^{+}15]. Disease [GKSJ21]. Disorders [CMJ21]. Dispatching [ZWZ18]. Displays [JDP^{+}21]. Disruptions [SNK^{+}14]. Dissemination [LLL19, XSZ^{+}15, ZCZY16]. Dissipation [BBM^{+}17]. Distance [ZYH^{+}20]. Distance-Bounding [ZYH^{+}20]. Distillation [CTL^{+}23]. Distributed [ÁAPDD24, AS24, BPBG18, DLTX21, DBNBT14, GHSA15, GZLG14, JMBR^{+}17, LJD^{+}15, TNKM14, TSH^{+}17, VKBB22, WFB^{+}18, WHC16, WQG^{+}22, YDM^{+}18, ZK21]. Distributing [MCB21]. Distribution [HCWL16, Isl21, KRG^{+}17, LA13, NVS^{+}14, WMN13, XZW21b]. Diverse [AKHA22]. Diversity [GJZ^{+}16, MKH24, TMS^{+}19, ZC15]. Dividers [LXL^{+}22b]. Division [ET23, Geu20, TMCVH21, WHC16, WLL^{+}21, BMP23]. DKG [QSLG22]. DNA [HKC22, LK23, PPKN23, TGDC^{+}21, YLW23]. DNA-Based [YLW23]. DNN [OAP^{+}22, SZA^{+}23, XWZ^{+}23]. Document [BC16, HWL15]. Documenting [ADQ21]. Documents [BC18]. Domain [CDC^{+}23, DLS21, GLC^{+}13, QSLG22, WZR^{+}14]. Domain-Specific [CDC^{+}23]. Doors [RPK^{+}22]. Dot [SMMTBM^{+}22]. Double [HLX^{+}17, WT20, YLZ^{+}21, YCH^{+}23]. Double-and-Triple-Node-Upset-Tolerant [YLZ^{+}21]. Double-Node-Upset-Hardened [YCH^{+}23]. Double-Track [HLX^{+}17]. Down [NL21]. Down-Sampling [NL21]. Downloading [ODCZ15]. DPA [BSY^{+}17, EMDE^{+}22, KTM19]. DPA-Resilient [BSY^{+}17]. Draggable [SCFH13]. DRAM [CYLJ21, PS21]. Dream [NB21]. Driven [BDL^{+}13, EMDE^{+}22, HHC^{+}23, LGT^{+}19, LRYK14, LK22, MLRRG20, PPF^{+}21, RWD22, WLWQ22, ZCZY16]. Driver [FSK20]. Drives [CH23, ZZM^{+}19]. Drone [YMX22]. Drug [BEG^{+}24, BTC^{+}21, GVF^{+}23, ZWH^{+}24]. DSP [ACH22]. DT2CAM [RFKK23]. DTD [HLX^{+}17]. Dual [BCD^{+}20, HCGS24, SK23a]. Dual-Channel [HCGS24]. Dual-Issue [BCD^{+}20]. Dual Port [ELVC18]. Due [OBM22]. Duplexing [WHC16]. durable [TFK20]. DWES [CSLG22]. Dynamic [AYL^{+}21, AMKF21, CSLG22, CCZN20, CPH^{+}15, CMMF20, DFRR23, GHSA15, GAI22, HBY^{+}23, HLW14, KKC21, LLW22, LDCY21, RPK^{+}22, SLB^{+}20, SDW^{+}21, WWG^{+}22, XYDJ22, YZZ^{+}21, YCLW14, ZCT^{+}14, ZCZ^{+}15, ZMK22, SLR^{+}17]. Dynamic-Key [LLW22]. Dynamically [WWG^{+}22].

e-Governance [PTD21]. e-Government [MVS21]. e-Learning [BC16]. e-LocGov [PTD21]. e-Participation [AKLC21]. E-Voting [PGM^{+}21]. Early [EGGOR20, FABC21, HWW^{+}17, KR21, LQT^{+}23, PMC^{+}15]. Earth [TKK^{+}22]. Earthquake [WGYL20]. ECCs [AFMM19]. Echoacoustic [KK22]. Echoes [KK22]. Economic [AAKJJ22, JND14]. Economical [LLW^{+}15]. Ecosystems [LHKH21]. eDemocracy [MT19]. Edge [BCX23, BYZZ22, DZK^{+}23, HHT23, LQT^{+}23, PCXF19, RPL^{+}23, RSK21, RLK22, TDZ21, WZLK21, Wan22, YSS^{+}22, ZLG^{+}22, ZK23, ZMK22]. Edge-Optimized [HHT23]. Edgeless [STSC24]. Edgeless-GNN [STSC24]. Edit [YLW23]. Editor [GP15b, KG20, Lom16, Lom17, Met18, Met19, MS20, Met20, MYM20, Mon20, Mon21, Mon22, ZEL20]. Editor-in-Chief [Lom16, Lom17, Met18, Met19, Met20, Mon20, Mon22]. Editorial [BDFT24, CP23, DLTX21, DLRL24, DPO17, DHPL19, DSS21, DZ16, DO18, DA20, FO21, GAWT23, HSS24, JSBM22, KGM15a, KSU16, LMC18, LHMW20, LDLN16, LDLN17, LSO17, LCLK20, LRLG22, LHJ18, MA18, MT19, MT21a, MT21b, MPK18, MK15, Mon20, Mon21, MGMC21, OSPN22, OGL^{+}21, PBKL19, SG15, SO21, TH16, XGLW21]. Editors [BCSF17, HINS21, SK19, SO20]. eDRAM [LRG^{+}21]. Educated [ALDK22]. Education [CSBME17, LMC18, LDLN16, LDLN17]. Educational [AFG^{+}21, LHMW20, SWK^{+}17]. EE [KTM19]. EE-SPFAL [KTM19]. EEG [LZX^{+}22, SZL^{+}22]. Effect [BNCF14, RS20, SBD^{+}21]. Effective [BBT^{+}16, LZD^{+}22, VC17]. Effectiveness [SWK^{+}17]. Effects [AAKJJ22, BS15, BDd^{+}21a, BDD^{+}21b, CKC^{+}18, HQP^{+}21, KDKB22, LRS^{+}22, PG23, SC22]. Efficiency [DZK^{+}23, LFL^{+}18, LFBL14, SKM^{+}23]. Efficient [AM22, BAS23, BJ21, BK22, BPB21, CLY^{+}23, CCH^{+}23, CLWX18, CYP^{+}16, CDC^{+}23, CGC^{+}24, DYJ22, DFRR23, DRV22, GTC^{+}21a, GTC^{+}21b, GP15a, GZG^{+}17, HF17, HHC^{+}23, HJCK21, KKC21, KK23, KTM19, LLD^{+}15, LNK^{+}15, LYA18, LLD^{+}18, LLL19, LZW21, MLG^{+}15, MLLU20, MM18, NT21, NMP19, PKK22b, PKK22a, Pla21a, Pla21b, PSZD21, RPM23, SAD23, SLB^{+}20, SZA^{+}23, TNKM14, TFK20, TPM16, VGP^{+}21, WQG^{+}22, WASW22, WZR^{+}23, XHWI22, YMDJ21, YW22, YCL^{+}22, ZWZ18, ZL22, ZYH^{+}20]. eGovernment [MT21a, MT21b, MT19]. EIC [Lom13, Lom15, Lom14]. Elastic [RB22, ZCT^{+}14]. Elasticity [DHGR18]. ELDC [MLLU20]. Elderly [RSB13]. eLearning [CG17, JMBR^{+}17]. Electricity [LGC13, LGT^{+}19]. Electronic [BC18]. Elementary [FM21b, FM21c]. Elite [LK22]. Elite-Data-Driven [LK22]. Embedded [ASYK^{+}22, AMRCP21, BHPE21, BCD^{+}20, BBT^{+}16, CH21, EAM21, KSU16, MYT^{+}18, PPF^{+}21, SZL^{+}22, WMAB17, YKAE22]. Embedding [TRD^{+}24, YZZ^{+}21]. Embeddings [CMJ21, RMB^{+}19]. Emergency [OLL^{+}21, SBL^{+}21]. Emerging [AM19, Ano13a, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15c, Ano15d, Ano15e, Ano15f, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano17a, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17k, Ano17j, Ano18a, Ano18b, Ano18c, Ano18h, Ano18d, Ano18e, Ano18f, Ano18g, Ano19a, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano20a, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano21a, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano22a, Ano22b, Ano22c, Ano22e, Ano22d, Ano22g, Ano22f, Ano23b, Ano23c, Ano23d, Ano23e, Ano24a, BS15, BDFT24, CP23, DLRL24, DPO17, DLTSNA21, DSS21, DO18, FO21, GZO^{+}18, HSS24, HINS21, JSBM22, JHB21, JLN21, KP14, KGM15a, KGM15b]. Emerging [KG20, KSU16, LMS^{+}14, LDLN16, LDLN17, LSO17, LRL21, LRLG22, MA18, MK15, OSPN22, PLJ15, RPR^{+}22, SK19, SO20, SO21, TH16, VPTH19, ZC15]. Emotion [BDL^{+}13, SZL^{+}22]. Emotions [DVA21]. Empirical [FAT^{+}14, KZ21, TRD^{+}24]. Emulating [LRS^{+}22]. Emulator [TSH^{+}17]. Enabled [AZA^{+}21, DZK^{+}23, JKSC21, KZT^{+}20, XZW^{+}21a, ZL22, HLX^{+}17]. Enabling [GTC^{+}21a, GTC^{+}21b, GSVA23, LLD^{+}15, LHKH21, RPM23, RSM^{+}21, TM14, XSY^{+}21]. Encapsulation [CSK^{+}23]. Encoded [TGDC^{+}21]. Encrypted [AJL^{+}21, BJ21, LLD^{+}15, LLD^{+}18, MLG^{+}15, NT21, TRR^{+}19, WLZ^{+}13, XSY^{+}21]. Encryption [APA^{+}21, AJL^{+}21, CWK^{+}21, CRS17, HLK^{+}24, KKC21, LCZ21, LZL^{+}21, RGS20, SZL^{+}20, TRR^{+}19, XWZ^{+}23]. End [GTC^{+}21a, GTC^{+}21b, LLLG23]. EnDAS [MLG^{+}15]. Endorsement [MR21]. Energy [AM22, AR20, AANN15, BM20, BAS23, BK22, BYZZ22, BMB^{+}18, CLWX18, Che14, CDC^{+}23, DAF^{+}22, DFRR23, DF22, GTC^{+}21a, GTC^{+}21b, GMTX14, GLFK18, HF17, HMB^{+}21, HLW14, HLL16, JDP^{+}21, KK23, KTM19, LNK^{+}15, LYA18, LFL^{+}18, LLL^{+}20, LLS^{+}16, LPY^{+}13, LFBL14, LLW^{+}15, LHP^{+}18, MLLU20, MSLL14, NSH22, NMP19, OMTH17, PLA20, PSZD21, RLX15, RZD^{+}19, SBD^{+}21, SAD23, SSVJ14, SY19, TNKM14, TFK20, TPM16, WMN13, WLC^{+}14, WZH^{+}16, WASW22, WZR^{+}23, XL19, ZWZ18, ZCZ^{+}15, ZYH^{+}20]. Energy- [AM22, WZH^{+}16]. Energy-Aware [BMB^{+}18, DF22, HLL16, WLC^{+}14, XL19]. Energy-Awareness [PLA20]. Energy-Efficiency [LFBL14]. Energy-Efficient [BAS23, BK22, CLWX18, CDC^{+}23, DFRR23, KK23, KTM19, LNK^{+}15, LYA18, MLLU20, SAD23, TFK20, TPM16, WZR^{+}23, ZWZ18, ZYH^{+}20]. Energy-Error-Product [PSZD21]. Energy-Harvesting [OMTH17]. Energy-Minimized [BYZZ22]. Energy-Source-Aware [LLS^{+}16]. Enforcement [RMVN22]. Engaging [HÁMLS23]. Engineering [CSBME17, PPF^{+}21, RMK^{+}14, STL^{+}14]. Enhance [CPM^{+}17, YK18]. Enhanced [CH16, GLFK18, GLZ19, LZW21, NTAL13, SWY^{+}22, TMS^{+}20]. Enhancement [BMC17, CF19, HJCK21, XLQ^{+}22].

Enhancing [FWC15, JHB21]. Enriched [TD19]. Enrollment [AEM22]. Ensemble [AKC14, SR14]. Ensembles [BCX23]. Entanglement [AAA18]. Enterprise [ATA21, CMK^{+}16]. Entities [LWN^{+}22]. Entropy [GLC^{+}21]. Enumeration [GCBK17]. Environment [AZA^{+}21, AC22, BTK^{+}20, BPBG18, GMTX14, PMC^{+}15, TCNC16, YCS22]. Environmentally [JND14]. Environments [LFBL14, PLSM20, QLT17, YXF^{+}22]. eNVM [MBW23]. eParticipation [PTT21]. EPIC [SGH13]. Epidemic [XSZ^{+}15]. Epidemics [ZWWF15]. Epidural [SUQKA20]. Equality [LJX^{+}22]. Equality-Constrained [LJX^{+}22]. Equalizer [TDVS21]. Equally [ÁAPDD24]. Equally-Distributed [ÁAPDD24]. Equations [ZYC^{+}22]. Equipment [GZO^{+}18, RWD22]. Equivalence [CH19]. Era [DO18, LRLG22]. Erasure [CSM19, LLL19]. Erasure-Coded [LLL19]. Erlang [ZYZZ15]. Error [BBW22, BMP23, CSK^{+}23, CH21, GHSMM21, GCW20, KSKA22, KKA22, LRXW21, LRG^{+}21, LRHL21, NW20, PKK22a, PPKN23, PSZD21, SBGC22, SK23a, VHFH^{+}22, WLX^{+}16, YLW23]. Error-Bounded [WLX^{+}16]. Error-Correcting [YLW23]. Error-Resilient [PKK22a]. Error-Tolerant [LRG^{+}21, LRHL21]. Errors [AFMM19, Gel15, LRXW21, LRL21, LRG^{+}21, LRS^{+}22]. ERuDIte [AFG^{+}21]. Escalation [TM14]. Establishment [ZXXH13]. Estimate [RLE^{+}22]. Estimating [WZR^{+}14, WCF23]. Estimation [LS21, LCY^{+}19, RPM23, SWY^{+}22, XZW21b, ZZZY24]. Estimator [PV15]. Estonian [PTD21]. Europe [PGM^{+}21]. Evaluate [YPS^{+}22]. Evaluating [ATA21, AAKJJ22]. Evaluation [AMvO17, APA^{+}21, BTK^{+}20, CSLG22, DMRR17, DAS22, FM21a, FKNK21, KWS^{+}20, LHB^{+}15, MM18, OZAL13, PGM^{+}21, PAR18, SG20, SBL^{+}21, SDW^{+}21, TRD^{+}24, XCX^{+}20]. Event [CLA^{+}22, EAM21]. Events [XYZ^{+}17]. Evil [HDA^{+}20]. Evoking [RS20]. Evolutionary [HZY21, WCC^{+}20, XLQ^{+}22, ZCT^{+}14, ZLX18, ZYS^{+}22]. Example [XYH^{+}23]. examples [TFK20]. Exchange [RMK^{+}14]. Execution [YCL^{+}22]. Execution-Efficient [YCL^{+}22]. Exercise [XCX^{+}20]. Exomars [SDZ^{+}21]. Expansion [PL19]. Expenditures [KZ21]. Experience [JHL21]. Experimental [BMOS16, LGT^{+}19, MIMMY^{+}19]. Experimentation [SGH13]. Expert [RGP^{+}21, TAV15]. Expertise [AKLC21]. Explainability [JMLH22]. Exploitability [Dul20]. Exploiting [AXAD21, BBM^{+}17, DHGR18, LRG^{+}21, RMVN22, XCW^{+}14, YK18]. Exploration [AC22, CDF^{+}22a, CLW^{+}18, EDLT21, FAP21]. Exploratory [GSMGP17]. Exploring [EGGOR20, KDM^{+}21]. Explosion [TMS^{+}20]. Expression [ZYC^{+}22]. Expressions [CDC^{+}23]. Expressive [MSS21]. EXSCALATE [GVF^{+}23]. Extended [LCZ21, SNHN15]. Extra [VHFH^{+}22]. Extra-Functional [VHFH^{+}22]. Extracting [RMB^{+}19]. Extraction [AEM22, LZX^{+}22, YGLO22]. Extreme [GVF^{+}23]. Extreme-Scale [GVF^{+}23].

F [SBGC22]. F/F [SBGC22]. Fabric [MR21]. Fabrication [LZH19, TM14]. Fabrics [DPO17]. Face [SDW^{+}21]. Faceted [TD19]. FacetsBase [SBHL21]. Facility [CBG^{+}21]. Factor [ACCL23, NWSG17, RGP^{+}21, WHL23b]. Factor-Aware [NWSG17]. Factories [LDJ20]. Factors [XCW^{+}14]. Factory [RDS^{+}22]. Failure [GMTX14, PSSZ22]. Failure-Related [GMTX14]. Failures [GLZ19, HWSN13]. Fairness [ZDLG13].

Fairness-Aware [ZDLG13]. Fall [MDB^{+}21]. FarSight [GKSJ21]. Fast [AK24, ET23, LYK^{+}22, LHP^{+}18, TA19, XOD20]. Fault [AMRCP21, BHB^{+}14, BBC^{+}22, CCZN20, GZT^{+}22, GAI22, GCW20, HMSZ20, LYA18, LWC22, MPK18, NYC^{+}21, PG23, WFB^{+}18, WRT^{+}21]. Fault-Tolerance [LYA18]. Fault-Tolerant [AMRCP21, CCZN20, LWC22]. Faults [GJZ^{+}16, NSH22, OBM22]. FCSS [WDO^{+}19]. Feasibility [IMZ^{+}21]. Feature [HCGS24, LZX^{+}22, RM20, RdPF^{+}23, WXT^{+}24]. Feature-Based [RM20]. Features [MMPP15, ZSCY20]. Federated [CLY^{+}23, CCH^{+}23, CTL^{+}23, FK22, KKY24, WLWQ22, YKP22, ZZH^{+}22]. Feedback [KTM19, TW22]. FeFET [LVR^{+}24]. FeFET-Based [LVR^{+}24]. Femtocell [ZMA15]. FerroCoin [CRP^{+}23]. Ferroelectric [CRP^{+}23]. FET [BS15, BSY^{+}17]. Fi [FKNK21]. Fibrillation [DVPQ^{+}21]. Field [CKKO20, CCZZ21, LLW22, WZR^{+}23, XHWI22]. Field-Free [WZR^{+}23]. Fight [GVF^{+}23]. File [CH23, LCZ21]. Filter [KK21, LZW21]. Filtering [HKC22, HDL14, WLX^{+}16, WDO^{+}19]. Filters [CDF^{+}22a, GZT^{+}22, ISI^{+}19]. Filters-Based [GZT^{+}22]. Find [HDA^{+}20]. Finding [RGP^{+}21, ZLXL22, ZMX^{+}22]. Fine [KK23, MLH^{+}20, SLR^{+}17, ZYY^{+}24]. Fine-Grained [MLH^{+}20, SLR^{+}17, ZYY^{+}24]. FinFETs [CLW^{+}18]. Fingerprint [AGMP21, ABG^{+}22]. FINISH [CCH^{+}23]. Finite [KJC22, XHWI22, ZLXL22]. Finite-State-Machine [KJC22]. Finite-Time [ZLXL22]. FireNN [DAS22]. Firmware [GDP22]. First [AJL^{+}21, CBG^{+}21]. Fitness [LT21]. FiWi [FNK^{+}13, GLFK18, GLZ19]. Flash [ZZM^{+}19]. Flash-Based [ZZM^{+}19]. Flexibility [CLD^{+}24, TMS^{+}20]. Flexibility-Enhanced [TMS^{+}20]. Flexible [BUS^{+}21, CPM^{+}17, EM23, GTC^{+}21a, GTC^{+}21b, HTH^{+}22, KWS^{+}20]. Flip [GLC^{+}21, YCH^{+}23]. Flip-Flop [GLC^{+}21, YCH^{+}23]. Flipped [ZC15]. Floating [FM21b, FM21c, BMP23]. Floating-Point [FM21b, FM21c, BMP23]. Flooding [SBR^{+}22]. Flooding-Based [SBR^{+}22]. Floorplanning [CWZ^{+}15]. Flop [GLC^{+}21, YCH^{+}23]. Flow [HMSZ20, HCWL16, YNA^{+}20]. FMA [OAP^{+}22]. Fog [GAI22, GZG^{+}17, JKSC21, WDO^{+}19, ZLG^{+}22]. Fog-Computing-based [WDO^{+}19]. Fog-Enabled [JKSC21]. Fog/Edge [ZLG^{+}22]. Fogs [GWYJ21]. Follow [RMB^{+}19]. Forecasting [WZL21, YKP22, ZYS^{+}22]. FORGE [JMBR^{+}17]. Formal [EKO^{+}16, GBVS21, HB15]. Formats [KK23]. Forming [NB21]. Formulation [DMRR17, RPVWRS^{+}20]. Forward [CWK^{+}21]. Four [ASA^{+}22, MHL17]. Four-Terminal [ASA^{+}22]. Fourier [WTW^{+}15]. FPGA [CSK^{+}23, CH16, CRS17, GZT^{+}22, GLC^{+}21, KSKA22, SP20, TDG17, TKTP21, WH21]. FPGA-Accelerated [CRS17]. FPGAs [eSKZW^{+}22, QGF^{+}23]. Fraction [WX23]. Fractional [WTW^{+}15]. Framework [AFATAH13, BPC^{+}17, BBF18, CVP^{+}22, EAM21, FSCX17, GWYJ21, HKZH16, HZQ^{+}18, HNGZ19, HZS^{+}15, HZY21, JND14, LZX^{+}22, LMK^{+}13, LLLG23, LXX^{+}21, MMM^{+}21, PMLT21, PCXF19, QSLG22, RFKK23, SBAR21, SG20, VSH23, WLX^{+}16, WLL^{+}22, WRT^{+}21, ZLY^{+}21, ZK23]. Fransform [WTW^{+}15]. Free [KS18, LRG^{+}21, OKK23a, THTK16, WZR^{+}23, MBW23]. Freedom [CCZN20]. Frequency [TDVS18, YDH21, ZZZY24]. Frequent [HDA^{+}20]. Friend [ZDLG13]. Friendly [SWL15]. Front [Ano18i, Ano18h, Ano21f, Ano21g, Ano22h, Ano22i, Ano22e, Ano23f, Ano23g, Ano23h, Ano23i, Ano24b, Ano17k, Ano13b, Ano13c, Ano14e, Ano14f, Ano14g, Ano14h, Ano15g, Ano15h, Ano15i, Ano15j, Ano16f, Ano16g, Ano16h, Ano16i, Ano21h, Ano21i, Ano21j]. Frontier [NSKRJ16]. Frontiers [MGMC21]. FTxAC [AMRCP21]. Fully [SZL^{+}20, XYDJ22]. Function [GLC^{+}21, LXL^{+}22a, YLW23, BMP23]. Functional [HLK^{+}24, HNGZ19, RdPF^{+}23, STL^{+}14, VHFH^{+}22]. Functions [KK14, PL19, SAD23, YQ14]. FunkR [ZXLS21]. FunkR-pDAE [ZXLS21]. Fusing [ZWH^{+}24]. Fusion [ZHC^{+}14]. Future [DSS21, KDM^{+}21, LMS^{+}14]. Fuzzy [JCM^{+}21, LYBZ20, PBTP21, YKP22, ZYS^{+}22].

Gain [RMSD23, TMS^{+}19]. Galois

[LCM^{+}23]. Game

[BY13, CSBME17, HYZ^{+}16, KIM^{+}18, MHL17, SBAR21, SUQKA20, XLQ^{+}22]. Game-Theoretic [HYZ^{+}16]. Game-Theoretical [BY13]. Games [CSBME17, GP15b, HWFR15, MHL17, SG20]. Gamification [BMOS16]. Gaming [PLSM20]. GAN [WJL23]. Gap [RSSE20]. Gas [CFL^{+}21]. Gas-Inefficient [CFL^{+}21]. GasChecker [CFL^{+}21]. Gases [SCS^{+}20]. Gate [CMPT17, CLW^{+}18, EGR21, JYZ^{+}23, KKC21, LK19, LCM18, MYM20, PIK20, RZAD18]. Gate-Delay [LCM18]. Gate-Length-Biasing [CLW^{+}18]. Gate-Level [LK19]. Gates [AMCMDM^{+}23]. Gateway [LJD^{+}15, OZL15, BMB^{+}18]. Gather [CLD^{+}24, CDM20]. Gathering [TNKM14]. Gbps [TKTP21]. Gender [BDL^{+}13]. Gender-Driven [BDL^{+}13]. Gene [ZYC^{+}22]. General [TW22]. Generalization [XTXY16]. Generalized [AAA18]. Generating [BFPS22]. Generation [BTD24, FSK20, GDP22, LHS^{+}22, LHJ18, MGMC21, NHT^{+}19, PLSM20, RRFT16, TAT^{+}22, YKP22].

Generative [HWL15, WJL23]. Generator [CRP^{+}23]. Generic [BAS23]. Genes [BTC^{+}21]. Genetic [QCN^{+}15, WLY^{+}21, WLZG22]. Geo [FMA24, GZLG14, TDZ21, YDM^{+}18]. Geo-Distributed [GZLG14, YDM^{+}18]. Geo-Indistinguishability [FMA24]. Geo-Partitioning [TDZ21]. Get [LSR^{+}21]. GHz [SSL^{+}13, TDVS18]. GIS [SCD^{+}21]. GNN [STSC24]. GNNUnlock [APH^{+}22]. Goal [LK22]. Gong [KSKA22]. Google [HH16]. Governance [PTD21]. Government [KZ21, MVS21]. Governments [PTD21]. GPS [MYS17]. GPU [AKHA22, BBF18, CLW16, LXX^{+}21]. GPUs [AJL^{+}21]. Grade [TKTP21]. Gradients [CRi^{+}19]. Gradually [CO16]. Grained [MLH^{+}20, SLR^{+}17, ZYY^{+}24]. Granularity [HHC^{+}23, SSR^{+}22]. Granularity-Driven [HHC^{+}23]. Graph [APH^{+}22, BEG^{+}24, BTD24, DLRL24, FVS24, HCGS24, LXX^{+}24, LTL^{+}22, MMY^{+}22, QSLG22, RGP^{+}21, SZL^{+}22, TRD^{+}24, WHL23b, ZCZY16, ZWH^{+}24, ZK23, MMM^{+}21]. Graph-Based [DLRL24, RGP^{+}21]. Graph-DB [MMM^{+}21]. Graph-Embedded [SZL^{+}22]. Graph-Incorporated [WHL23b]. Graphene [BS15]. GRAPHIC [CLD^{+}24]. Graphical [BSJ22]. Graphs [AMKF21, LXL^{+}22a, TSS16]. GRAVITAS [BSJ22]. Gravity [Sag19]. Green [HWH^{+}20, LLL^{+}20, LLS^{+}16, XCX^{+}20, ZEL20]. Grid [AAEKM13, BY13, DHPL19, GLZ19, HNGZ19, HWSN13, LGC13, LMK^{+}13, LA13, MZGT17, MS19, NTAL13, SY19, TSH^{+}17, WMN13, WLZ^{+}13]. Group [HLL16, HSzXZ17, SWW^{+}20, Wan22]. Grouping [GSMGP17]. Groups [RMB^{+}19]. Grow [DYJ22, YCL^{+}22]. Grow-and-Prune [DYJ22, YCL^{+}22]. Growable [CO16]. Guarantee [YW22]. Guaranteed [CDLS13, CMMF20]. Guest [BCSF17, BDFT24, CP23, DLTX21, DLRL24, DPO17, DHPL19, DSS21, DO18, DA20, FO21, GAWT23, HSS24, HINS21, JSBM22, KGM15a, KSU16, LMC18, LHMW20, LDLN16, LDLN17, LSO17, LCLK20, LRLG22, LHJ18, MA18, MT19, MT21a, MT21b, MPK18, MK15, MGMC21, OSPN22, OGL^{+}21, PBKL19, SG15, SO21, TH16, XGLW21, GP15b, KG20, MS20, MYM20, SK19, SO20, ZEL20]. Guidance [OLL^{+}21]. Guided [KJC22, YCL^{+}22].

H [ASA^{+}22, LQT^{+}23]. H-BILSTM

[LQT^{+}23]. H.264 [MZY^{+}16]. H.264/AVC [MZY^{+}16]. Habit [SWL15]. Hacking [SAAJ22]. Hadoop [BPC^{+}17]. Hall [SBD^{+}21]. Handling [VHFH^{+}22]. Handoff [HLL16, HSzXZ17, QLT17]. Hard [JSZ^{+}21]. Hard/Soft [JSZ^{+}21]. Hardened [WT20, YCH^{+}23]. Hardware [AKSR24, BHC^{+}23, CPM^{+}17, DLTSNA21, DZD^{+}18, HKC22, HYW19, KP14, KSKA22, KKA22, LCM^{+}23, LRP^{+}22, LCY20, MM14, NAMJ23, PG23, PKR22, SK23b, TKTP21, VS17, WMAB17, XHWI22, YCL^{+}22, YGLO22]. Hardware-Guided [YCL^{+}22]. Hardware-Software [DZD^{+}18]. Hardware/Software [AKSR24]. Harmoniously [CLD^{+}24]. Harnessing [RPK^{+}22]. Harsh [YXF^{+}22]. Harvesting [Che14, HMB^{+}21, HLW14, OMTH17, RLX15, RZD^{+}19]. HBM [QGF^{+}23]. HBM-Based [QGF^{+}23]. HCCA [CYBD15]. HCNN [AJL^{+}21]. HCP [FK22]. HDGCN [HCGS24]. Health [DVA21, LHKH21, NSKRJ16, RSH^{+}21]. Healthcare [JMLH22, KKY24, KDM^{+}21, MGMC21]. Healthy [DTK^{+}23]. Helios [PGM^{+}21]. Helperless [BR16]. Heterogeneity [BRRE22, DHGR18, PLA20]. Heterogeneous [AFATAH13, ASYK^{+}22, BYZZ22, DZD^{+}18, FK22, HB15, HCWL16, LRYK14, NAM17, SZA^{+}23, WLC^{+}14, YATR18, YMX22, ZC15].

HetNets [GLFK18]. Heuristic [FAP21]. HEVC [NMP19]. Hierarchical [CCW^{+}14, SBAR21, WFB^{+}18]. Hierarchy [LCZ21]. High [BHPE21, CCZN20, CMMF20, CLD^{+}24, CMRV21, FAP21, GT22, GLC^{+}21, HLK^{+}24, eSKZW^{+}22, LLG^{+}23, MS20, NL21, NAMJ23, NHT^{+}19, QGF^{+}23, SDW^{+}21, TZZB21, TDG17, TH16, VC17, WHC16, WCC^{+}20, WWXL22, WWL^{+}22, WLZG22, WHL23b, YCH^{+}23]. High-Dimensional [WHL23b]. High-Level [NHT^{+}19]. High-Performance [BHPE21, CCZN20, CMRV21, NAMJ23, QGF^{+}23, TDG17, VC17, WWXL22, WWL^{+}22, YCH^{+}23]. High-Speed [GT22, eSKZW^{+}22, WHC16]. High-Throughput [LLG^{+}23]. Higher [HCGS24, ZWH^{+}24]. Higher-Order [HCGS24]. Highlighter [BC18]. Highlighting [BC18]. Highly [DBOB20, LZW21, WH21]. Highly-Parallel [WH21]. Highway [ODCZ15]. Hijackers [RMB^{+}19]. Histograms [CRi^{+}19]. Hit [CYLJ21]. HMCKRAutoEncoder [WLL^{+}22]. Hoc [GZFS18, LHB^{+}15, LYS13, KNK13]. Holistic [RPVWRS^{+}20]. Holo [SBL^{+}21]. Holo-BLSD [SBL^{+}21]. Holographic [SBL^{+}21]. Home [BGS^{+}21, LLW^{+}15, PCG^{+}21, WCF23, ZCZ^{+}15]. Homes [KR21, WC22]. Homomorphic [APA^{+}21, AJL^{+}21, BJ21, CRS17, LZL^{+}21, SZL^{+}20]. Honeycomb [NYC^{+}21]. Horse [CPM^{+}17]. Hot [MM18]. Hotspot [IMZ^{+}21]. Households [SY19]. HPC [BDd^{+}21a, BDD^{+}21b, CCMA21]. HREN [BK22]. HTS [TMS^{+}20]. Human [AS20, AR20, ACCL23, BTC^{+}21, DLTX21, DTK^{+}23, DA20, MSLL14, OGL^{+}21, WCF23]. Human-in-the-Loop [MSLL14]. Humans [AS20]. Hunting [HDA^{+}20]. HW [PKK22b]. HW-SW [PKK22b]. Hybrid [BK22, DAS22, ELVC18, HTS^{+}23, KMD^{+}18, LYW^{+}21, LXL^{+}22b, LA13, PS21, QLT17, QCN^{+}15, SSV^{+}20, WWL^{+}22]. HyperFET [DLTSNA21]. Hypergraphs [ÁAPDD24]. Hyperledger [MR21]. HyperspaceFlow [ZYMG16]. Hyperspectral [LXX^{+}21, TKTP21].

I/O [ATA21, LZD^{+}22]. IC

[CWZ^{+}22, LLW22, SLR^{+}17]. ICs

[LK19, NYC^{+}21, PASK21]. ID [THTK16]. ID-Based [THTK16]. Ideal [DQB23]. Idempotent [NW20]. Identification [CBZ^{+}21, RWJZ23, RB22, YKW^{+}20, YDH21]. Identifier [ZLC16]. Identifier/Locator [ZLC16]. Identify [RGP^{+}21, RMB^{+}19]. Identifying [HWW^{+}17]. Identity [KMW^{+}21, LTL^{+}22]. IEEE [IEE21, IEE22, AAEKM13, CYBD15, FM21b, FM21c, SSVJ14, Ano13a, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15c, Ano15d, Ano15e, Ano15f, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano17a, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17k, Ano17j, Ano18a, Ano18b, Ano18c, Ano18h, Ano18d, Ano18e, Ano18f, Ano18g, Ano19a, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano20a, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano21a, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano22a, Ano22b, Ano22c, Ano22e, Ano22d, Ano22g, Ano22f, Ano23b, Ano23c, Ano23d, Ano23e, Ano24a, BDFT24, CP23, DPO17, DO18, FO21, HSS24, KGM15a, KGM15b]. IGA [CMPT17]. II [GFKL13b, KGM15a, LDLN17]. IID [AS24]. IIoT [CWK^{+}21]. Image [AM22, BBC^{+}22, CLLL23, HJCK21, HAKL22, LJX^{+}22, LXX^{+}21, LVJ22, LK22, SZL^{+}22, SWY^{+}22, TS21, TKTP21, WZRR13, WASW22, YMT22]. Image-Based [SZL^{+}22]. Images [LD21, Ode21]. Imaging [ATPR23, CMM^{+}21, SDW^{+}21]. Imitation [YMX22]. Immersive [BPBG18]. Immune [LJX^{+}22]. Impact [GKC21, KMM15, LHB^{+}15, OEM18, RGP^{+}21, SMMTBM^{+}22]. Impacting [JLN21, OSPN22]. Implant [LRP^{+}22]. Implantable [CPLdFM21]. Implementation [APA^{+}21, BPBG18, CDF^{+}22a, EKO^{+}16, GLC^{+}21, GJZ^{+}16, MHL17, SDZ^{+}21, SK23b, XHWI22]. Implementations [GZT^{+}22]. Implemented [EDLT21]. Implication [MIMMY^{+}19]. Implicit [ZYC^{+}22]. Improve [BMOS16, BTC^{+}21, KNK13]. Improved [ZZM^{+}19]. Improvement [NS15, RMSD23, TKK^{+}22]. Improvements [AMvO17]. Improving [KVP19, MCUDCD22]. In-Field [CKKO20, CCZZ21, LLW22]. In-Memory [AKSR24, BDFT24, CP23, EM23, LLG^{+}23, PRKK24, WMP^{+}24, LHZS20]. In-Processor [OMTH17]. In-Situ [CVP^{+}22]. In-Vehicle [OZL15]. Incentive [SQX^{+}20, ZZH^{+}22]. Incident [LLLG23]. Inclusive [RAD22]. Incomplete [AS24]. Incorporated [WHL23b]. Incorporating [GZB22]. Incremental [DYJ22, HHT23, LWC22]. Independent [FSCX17]. Index [AFG^{+}21, Ano13a, Ano15a, Ano16a, Ano17a, Ano18a, Ano19a, Ano20a, Ano21a, Ano22a]. Indistinguishability [FMA24]. Indoor [FMA24, WCF23]. Induced [LRS^{+}22]. Industrial [ADCS22, CYL^{+}15, LCY^{+}19, PLJ15, RPVWRS^{+}20]. Industry [CDF^{+}22b, JSBM22, XSYW20]. Inefficient [CFL^{+}21]. Inequalities [ZLXL22]. Inexact [HAKL22]. Inference [CGC^{+}24, GTC^{+}21a, GTC^{+}21b, KK23, LTL^{+}22, MBW23, RPL^{+}23, SZA^{+}23, XYDJ22, XYH^{+}23, YJ22]. Inferences [FLB^{+}19, LVJ22]. Influence [TW22, ZLW^{+}21]. Information [Ano22g, Ano22f, Ano23b, Ano23c, Ano23d, Ano23e, Ano24a, BPB21, CDM20, FNK^{+}13, FKNK21, HCGS24, HYW19, HB15, KMK22, KMW^{+}21, NSKRJ16, QWC^{+}18, RDS^{+}22, TGDC^{+}21, WDO^{+}19, XSZ^{+}15, YZG^{+}24, YQ14, YGLO22, ZLZ^{+}17, ZWH^{+}24, Ano14a, Ano14b, Ano14c, Ano14d, Ano15c, Ano15d, Ano15e, Ano15f, Ano16b, Ano16c, Ano16d, Ano16e, LMW^{+}17]. Information-Centric [WDO^{+}19]. Infrastructure [BY13, CDF^{+}22b, LLW^{+}15]. Inherently [ACCL23]. Inkjet [CMPT17]. Inkjet-Configurable [CMPT17]. Inner [HLK^{+}24]. Inner-Product [HLK^{+}24]. Innovation [DPO17, LHMW20, MT21b]. Innovative [AKLC21]. Input [PIK20, ZWPL23]. Insensitive [BNCF14]. Inserting [CSZ^{+}20]. Insertion [NS15, PPKN23]. Insertion-Based [PPKN23]. Insights [RSH^{+}21]. Inspired [CMRV21, LHS^{+}22, PBTP21]. Instruction [BHPE21]. Instrument [SDZ^{+}21]. Integer [AAA18, ET23, TMCVH21]. Integral [ZLXL22]. Integrate [HB15]. Integrated [BMC17, CVP^{+}22, DO18, EGR21, EDLT21]. Integrating [DA20, HHM20]. Integration [FSK20, GKC21, MCUDCD22, ZS15]. Integrative [PKK22a]. Integrity [RMVN22, YLY^{+}20]. Intel [ET23]. Intellectual [VS17, XWZ^{+}23]. Intelligence [BCX23, BDL^{+}13, FBL^{+}22, HDA^{+}20, LRLG22]. Intelligent [HINS21, LQT^{+}23, MTFK21, OKK22, OKK23a]. Intelligibility [CMJ21]. Intensive [EM23, SBD^{+}24, WZH^{+}16]. Inter [DLS21]. Inter-Domain [DLS21]. Interaction [MRS21, TRGVR^{+}22]. Interactive [RS20]. Intercell [HYZ^{+}16]. Interconnect [BHC^{+}23, PS21]. Interconnected [UBMA18]. Interconnection [AMG22, DBLK23]. Interconnects [KMD^{+}18, KS18, LFL^{+}18, LZH19, NS15, PASK21]. Interest [AS20, HQP^{+}21, RMB^{+}19]. Interface [LDJ20, TDVS21]. Interference [TMS^{+}19, ZWWF13]. Interlaced [CH23]. Interlayer [NS15]. Intermittent [AmSYlrm24, WQG^{+}22]. Internal [Mél22].

Internet [HSzXZ17, XSYW20, AMvO17, BSJ22, HSG19, KM20, LLL^{+}20, LFBL14, MJ17, OZL15, PLJ15, PBKL19, RSK21, SAAJ22, TH16, XL19, YQZ^{+}15, ZEL20, ZLZ^{+}21]. Internet-of-Things [BSJ22, LFBL14, MJ17, SAAJ22]. Interpolation [PSZD21]. Interpretable [WLL^{+}22]. Interrelated [YMX22]. Interventions [HWW^{+}17]. Intra [TBG^{+}18]. Intra-Cell [TBG^{+}18]. Introducing [PTD21]. Introduction [BCSF17, DHPL19, GP15b, HINS21, KG20, KSU16, LSO17, LCLK20, MS20, MYM20, PBKL19, SG15, SK19, SO20, ZOS16, ZEL20]. Intrusion [LA13, MCUDCD22, PJK^{+}19, TPM16]. Intrusive [SFZ^{+}16]. Invariant [CQH17]. Invariant-Based [CQH17]. InvarNet [CQH17]. InvarNet-X [CQH17]. Inversion [XAQ22]. Investigation [KZ21]. IoMT [HTH^{+}22]. IOT [HSzXZ17, AFATAH13, BKS21, DF22, GTC^{+}21a, GTC^{+}21b, HWH^{+}20, LHKH21, PJK^{+}19, PTM23, PMC^{+}15, RMVN22, SBM21, SSF^{+}22, ZSCY20]. IoTility [LHKH21]. IoTRec [SSF^{+}22]. IoV [LWM^{+}21]. IoV-Assisted [LWM^{+}21]. IP [ABC18, CLW16, FSK20, KJC22, LZL^{+}21, PASK21, VS17]. IPFE [AK24]. Irradiation [CBG^{+}21]. IRS [OKK22, OKK23a, OKK23b, ZMK22]. IRS-Aided [ZMK22]. iSenior [RSB13]. Isn’t [NVS^{+}14]. Isolation [BBE^{+}24, HYW19]. Issue [BCD^{+}20, DO18, FO21, GFKL13a, GFKL13b, KP14, KGM15a, KGM15b, KSU16, LHMW20, LDLN16, LDLN17, LSO17, MT21a, SG15, SO20, SO21, WK14]. Issues [BCSF17]. Iterated [SCD^{+}21]. Iterative [AKC14].

Jamming [WLWQ22]. January [Mon22]. Jitter [RPVWRS^{+}20]. Job [KAF^{+}16].

Jobs [Che14]. Joint

[DPO17, GCW20, HINS21, LHMW20, LGC13, RZD^{+}19, VKBB22, XZW21b]. Journal [Mon20, Mon22]. Journaling [CH23]. Journals [RGP^{+}21]. Junction [CRP^{+}23]. Junction-Based [CRP^{+}23]. June [IEE21]. Just [CLWG15]. Just-in-Time [CLWG15].

Kernel [EAM21]. Kernels [AKHA22]. Key [CSK^{+}23, CWK^{+}21, CCZZ21, JCM^{+}21, KKC21, LLW22, LHZS20, RRFT16, RMK^{+}14, SBHL21, SSR^{+}22, THTK16, ZXXH13, ZMX^{+}22]. Key-Components [SSR^{+}22]. Key-Exchange [RMK^{+}14]. Key-Gate [KKC21]. Key-Speaker [ZMX^{+}22]. Key-Value [LHZS20, SBHL21]. Keys [NT21]. Keyword [LLD^{+}15, LVR^{+}24]. Know [HDA^{+}20]. Knowledge [CTL^{+}23, HZY21, KAF^{+}16, NVS^{+}14, QSLG22, RMVN22, ZXXH13]. Known [XSYW20]. KWS [LVR^{+}24]. KyberKEM [LCM^{+}23].

L1 [GKC21]. Label [HWL15]. Landmark [CBZ^{+}21]. Language [CO16, SPCB16]. LARAC [XTXY16]. Large [AKC14, ABDL19, APCM20, BFPS22, CLY^{+}23, CCFI16, GMTX14, HNGZ19, NdCFB^{+}23, Ode21, SR14, ZOS16, ZHC^{+}14]. Large-Scale [ABDL19, CLY^{+}23, CCFI16, GMTX14, HNGZ19, SR14, ZOS16, ZHC^{+}14]. Laser [CMM^{+}21, LFL^{+}18, SDZ^{+}21]. Latch [WT20, YLZ^{+}21, YXF^{+}22]. Latches [OEM18]. Latency [CYLJ21, SK23a, TPM16, WLWQ22]. Latent [PV15, WHL23b]. Lattice [CSK^{+}23, eSKZW^{+}22]. Lattice-Based [CSK^{+}23, eSKZW^{+}22]. Lattices [ASA^{+}22]. Layer [CCW^{+}14, MMPP15, PJK^{+}19, WTW^{+}15]. Layers [PG23]. LBS [HLL16]. LBSNs [TFM^{+}19]. LDPC [ZZM^{+}19]. Leakage [ELVC18, HYW19, LLLG23, NSKRJ16, PIK20]. Learned [ACM16]. Learning [ADCS22, AEM22, AYG^{+}21, AS24, BC16, BC18, BMOS16, BPBG18, BSG^{+}16, CLY^{+}23, CSBME17, CO16, CCH^{+}23, CLA^{+}22, CTL^{+}23, CTC^{+}17, DYJ22, FK22, GZB22, HWH^{+}20, HCGS24, HHM20, HLW14, HHT23, JHB21, JMLH22, JHL21, KKY24, KPL^{+}21, KMW^{+}21, LMC18, LXL^{+}22a, LXX^{+}24, LDCY21, LRHL21, LRLG22, LTL^{+}22, LZLC22, LVJ22, LK22, MTFK21, MHL17, MLRRG20, MMY^{+}22, NLF^{+}22, OPIV^{+}23, PKK22b, QLL^{+}22, RM20, RWD22, RSK21, RS20, SAJ22, SAAJ22, SKM^{+}23, STSC24, SR14, SZL^{+}20, SWK^{+}17, TAT^{+}22, VKBB22, WZLK21, WLL^{+}22, WQG^{+}22, WLWQ22, WXT^{+}24, WZL21, WLL^{+}21, YSS^{+}22, YMX22, YZZ^{+}21, YKP22, YGLO22, ZZH^{+}22, ZXLS21, ZWH^{+}24, ZL22, ZYS^{+}22, ZLG^{+}22, ZWPL23, BKS21]. Learning-Based [HLW14, LZLC22, YSS^{+}22, YGLO22]. Learning-Enabled [ZL22]. Length [CLW^{+}18, LK19, PP23a]. LEoNIDS [TPM16]. Less [APH^{+}22, AKSR24, NBRF18, NBRF18]. Lesson [ACM16]. Level [BAS23, CBG^{+}21, CTC^{+}17, FAP21, HMSZ20, LK19, MBW23, MYM20, NHT^{+}19, NMP19, PIK20, TDG17, TPM16, ZYMG16]. Leveraging [AMRCP21, CH23, JMBR^{+}17, LRP^{+}22, MS19, OLL^{+}21, ZCZ^{+}15]. LEXACT [BM20]. Libraries [BDd^{+}21a, BDD^{+}21b]. Library [AK24]. LibreKV [LHZS20]. Life [DTK^{+}23]. Lifecycle [RWD22]. Lifetime [EGR21, LC20, MKAR22, PAR18, TKK^{+}22]. Lightweight [CLL21, CWK^{+}21, CRRS22, KSKA22, KKA22, RS21, TRR^{+}19]. Like [AMCMDM^{+}23]. Likelihood [LS21]. Limitation [CPH^{+}15]. Limited [LRL21]. Limiting [ACCL23]. Limits [BR16]. Line [BUS^{+}21, CKC^{+}18, DAF^{+}22, RPL^{+}23, ZMRM19]. Linear [AAA18, ZS15]. Link [HXL^{+}14]. Links [TRD^{+}24, XSZ^{+}15].

Liquid [MDB^{+}23]. List

[Ano17b, Ano17c, Ano19c, Ano19b, Ano20b, Ano21b, THTK16, Ano15b]. List* [Ano23a]. List-Free [THTK16]. Literature [GPT^{+}21]. Lithography [CF19]. Little [NVS^{+}14]. Live [CBZ^{+}21]. Living [CFM^{+}22, KZT^{+}20, WLO^{+}21]. Load [CFM^{+}22, KMM15]. Local [AMTR24, KZ21, MMPP15, PTD21, WQG^{+}22, WHL23a, XZW21b, ZZZY24]. Localization [CRi^{+}19, HYL^{+}20]. Location [AMvO17, FMA24, GZFS18, HKC22, HKZH16, LZW21, MYS17, MHL17, TS21]. Location-Based [LZW21, MHL17]. Locator [ZLC16]. LocGov [PTD21]. Lock [CCZZ21]. Locked [CH21]. Locking [APH^{+}22, BNCF14, JBSS^{+}22, RS21, YMSR20]. Locomotion [MRS21]. Logarithm [MDB^{+}22]. Logarithm-Approximate [MDB^{+}22]. Logarithmic [ACH22, LXL^{+}22b]. Logic [APH^{+}22, AMCMDM^{+}23, BSY^{+}17, CH19, FM21a, HF17, KKC21, KWS^{+}20, KTM19, LZM^{+}21, NBRF18, PL19, RGS20, RS21, SBD^{+}21, SLR^{+}17, TZZB21, TA19, YMSR20]. Logic-Based [LZM^{+}21]. Logic-in-Memory [SBD^{+}21, TZZB21]. Logical [DMRR17]. Logistic [LS21]. Logistics [YCLW14]. Logs [LLLG23, SFZ^{+}16]. Long [GKSJ21, LQT^{+}23, ZC15]. Long-Term [GKSJ21]. Look [SBD^{+}24]. Look-Up [SBD^{+}24]. Lookup [CLW16]. Loop [MSLL14]. Loops [CH21]. Losses [KZ21]. Lossless [MZY^{+}16]. Lottery [JDL^{+}23]. Low [AC22, ACH22, BM20, BHB^{+}14, CBZ^{+}21, DTK^{+}23, ELVC18, JDP^{+}21, NSH22, NHT^{+}19, OEM18, QLL^{+}22, RGS20, SBD^{+}21, SK23a, TPM16, TDVS18, TKK^{+}22, WLZG22, YLZ^{+}21, YDH21, ZNS^{+}22]. Low-Complexity [CBZ^{+}21]. Low-Cost [DTK^{+}23, OEM18]. Low-Earth-Orbit [TKK^{+}22]. Low-Energy [NSH22, SBD^{+}21]. Low-Latency [TPM16]. Low-Leakage [ELVC18]. Low-Overhead [JDP^{+}21]. Low-Power [ACH22, NHT^{+}19, TDVS18]. Lower [ZWH^{+}24]. Lower-Order [ZWH^{+}24]. LSTM [WGYL20]. LSTMs [YCL^{+}22]. LTE [GLFK18]. LTE-A [GLFK18]. LTRT [KNK13]. Luminance [HJCK21]. LWE [SK23b, XHWI22].

M [TDVS21]. M-PHY [TDVS21]. M2M [HLL16]. M2M-Based [HLL16]. M3D [GKC21]. MA [KNK13]. MA-LTRT [KNK13]. MAC [KMM15]. Machine [CLA^{+}22, GZB22, GCBK17, HLK^{+}24, JHB21, JMLH22, KJC22, KPL^{+}21, LRHL21, LRLG22, LZLC22, MDB^{+}23, RM20, RWD22, RSK21, SAJ22, SAAJ22, SZL^{+}20, TAT^{+}22, WZH^{+}16, BKS21]. Machine-Learning [KPL^{+}21]. Machine-Learning-Driven [RWD22]. Machines [AS20, Dul20, FLB^{+}19]. Made [VSS18]. Magnetic [FBL^{+}22, WZR^{+}23]. Magnitude [LRL21]. Magnonic [GWD^{+}23]. Maintain [NTAL13]. Maintenance [LMS^{+}14]. Majority [LZM^{+}21]. Making [RGP^{+}21]. Malicious [GJZ^{+}16, TM14]. Malware [CCH^{+}23, LD21, WJL23]. Management [AR20, BBS^{+}15, BY13, CMK^{+}16, CMRV21, CDF^{+}22b, DGP22, GZG^{+}17, HYL^{+}20, HHC^{+}23, HLW14, HK17, JKSC21, KG20, KIM^{+}18, LC20, LGT^{+}19, MZGT17, MT21b, PMC^{+}15, RWD22, TMS^{+}20, YKAE22, ZLZ^{+}21, LMW^{+}17]. Management-Oriented [HYL^{+}20]. Manager [YATR18]. MANET [SSVJ14]. Manifold [LHS^{+}22]. Manifold-Inspired [LHS^{+}22]. Manipulatives [SPCB16]. Manufacturing [CVP^{+}22, JSBM22, LLW22, MMY^{+}22, PASK21]. Manuscripts [MSS21]. Many [AMG22, UBMA18]. Many-Core [UBMA18]. Manycore [CMB18]. Mapping [BRRE22, CSBME17, DFRR23, SAM^{+}20, TA19]. Mappings [ZLC16]. MapReduce [MS19, ZLY^{+}21].

Maps [AS24]. Maps-Based [AS24]. Margin [LRHL21]. Margining [RPVWRS^{+}20]. Marketeers [ACM16]. Marketplace [PLJ15]. Markets [ACM16]. Mass [FBL^{+}22]. Massively [GP15b, HWFR15, VSS18]. Matching [RMK^{+}14, SDW^{+}21, WWG^{+}22, WLY^{+}21, ZDLG13]. Material [MIMMY^{+}19]. Materials [KK22]. Mathematical [BDd^{+}21a, BDD^{+}21b, GSMGP17]. Matrix [YZZ^{+}21, ZLXL22, ZXXH13]. Matrix-Based [ZXXH13]. Maturity [ACM16]. MAX [WJ19]. MAX-3-SAT [WJ19]. Maximization [DZK^{+}23, TW22]. Maximize [TMS^{+}19]. Maximizing [DHGR18]. Maximum [LFL^{+}18, LS21]. Mean [KS18]. Means [BC16, CGPB21, SCD^{+}21]. Measurement [ZYZZ15]. Measurements [SNHN15]. Measuring [SKKN20]. MEC [LQT^{+}23]. Mechanics [CSBME17, LZX^{+}22]. Mechanics-Based [LZX^{+}22]. Mechanism [BYB20, GCW20, NT16, WHC16, XZW^{+}21a, YDH21, ZZH^{+}22]. Mechanisms [CSK^{+}23, HLL16, KKA22, ZZZY24]. Media [NWSG17, SCFH13]. Medical [CPLdFM21, DTK^{+}23, GZG^{+}17, TAV15, XSY^{+}21, YMDJ21]. Medicine [NLF^{+}22, WLY^{+}21]. Membership [XYH^{+}23]. Memories [GHSMM21, HHC^{+}23, LRXW21, LRL21]. Memory [AM19, AMJ22, AKSR24, ASKG21, AMCMDM^{+}23, BHPE21, BDFT24, CP23, DQB23, DRV22, ELVC18, EM23, GVO^{+}23, GT22, HMSZ20, HTS^{+}23, HEYB22, IPiR18, JHB21, LLG^{+}23, LZD^{+}22, LQT^{+}23, LBX^{+}23, LHZS20, LYW^{+}21, LVR^{+}24, MBW23, PG23, PRKK24, PAR18, QCN^{+}15, RFKK23, RLE^{+}22, RTVG22, SBD^{+}21, SBGC22, SBD^{+}24, TZZB21, TFK20, VPTH19, WMP^{+}24, WZR^{+}23, XAQ22, YNA^{+}20, GVO^{+}23]. Memory- [CP23]. Memory-Centric [GVO^{+}23]. Memristive [GT22, PAR18]. Memristor [MIMMY^{+}19, LYW^{+}21]. Memristor-Based [MIMMY^{+}19]. Memristors [VSS18]. Mental [DVA21, RSH^{+}21]. Mesh [LNK^{+}15, ZXXH13]. Message [Lom14, Lom15, Lom16, Lom17, Met18, Met19, Met20, SLC^{+}13, Lom13]. Metadata [MVS21]. Metal [CVP^{+}22]. Meter [CFM^{+}22, LGC13]. Metering [WLZ^{+}13, CWZ^{+}22]. Method [CLY^{+}23, CYBD15, EGR21, FBL^{+}22, HAKL22, HHT23, JHL21, KWS^{+}20, KNK13, MRBS22, NBS16, OKK22, PKK22a, YMX22]. Methodologies [MK15]. Methodology [APH^{+}22, CDF^{+}22a, DBOB20, ZYMG16]. Methods [DLRL24, MRS21, MYM20, RWZ^{+}16]. Metrics [BS15, NVS^{+}14, SFZ^{+}16]. MFDS [ZYY^{+}24]. MFDS-STGCN [ZYY^{+}24]. Micro [SC22]. Micro-Databases [SC22]. Microarchitecture [KM20]. Microelectronics [CBG^{+}21]. Microfluidic [GCBK17]. Microprocessor [TM14]. Microprocessors [LCM18]. Microworld [GSMGP17]. Migration [DDB22]. MigSpike [DDB22]. Minimal [JSZ^{+}21]. Minimal-Cost [JSZ^{+}21]. Minimization [CYP^{+}16, GZLG14, WLWQ22]. Minimized [BYZZ22]. Minimizing [GLFK18, LLL^{+}20]. Minimum [TAC^{+}19]. Mining [HDA^{+}20, KR21, SY19, SNK^{+}14, WGYL20, YKW^{+}20, ZK21, ZYC^{+}22]. Missing [TRD^{+}24]. Mission [SDZ^{+}21]. Mist [LLL19]. Mitigate [OEM18, TMS^{+}19]. Mitigating [CPH^{+}15, SKD20]. Mitigation [ZWWF13]. Mixed [BRRE22, JBSS^{+}22, NSH22, RDS^{+}22]. Mixed-Critical [RDS^{+}22]. Mixed-Criticality [BRRE22, NSH22]. Mixed-Signal [JBSS^{+}22]. Mixture [HWL15]. MLC [ZZM^{+}19]. MLGNRs [NS15]. mm [AMG22]. mm-Wave [AMG22]. MMCD [ODCZ15]. MMOGs [BBS^{+}15, GHSA15]. Mobile [CLWX18, DZK^{+}23, FSCX17, GZFS18, HYL^{+}20, HKZH16, HZQ^{+}18, HCCL13, HLL16, HSzXZ17, IP21, JSZ18, JND14, KGM15a, KGM15b, KNK13, KIM^{+}18, LHB^{+}15, LLD^{+}15, LLD^{+}18, LQT^{+}23, MLG^{+}15, RLK22, SAM^{+}20, SWL15, TRGVR^{+}22, TDZ21, WZLK21, WLL^{+}21, XSZ^{+}15, YSS^{+}22, YQZ^{+}15, YZG^{+}24, YDM^{+}18, ZWZ18, ZDLG13, ZSL^{+}15]. Mobility [ACCL23, WCF23, ZLC16]. Mobility-Caused [ZLC16]. Modalities [ZS15]. Mode [BSY^{+}17]. Model [AEM22, AS20, AAEKM13, EKO^{+}16, GLC^{+}13, GWD^{+}23, HWL15, HXL^{+}14, HQP^{+}21, JYZ^{+}23, KKY24, KPL^{+}21, KS18, LSP^{+}20, LCM18, PJK^{+}19, PTT21, PTD21, PPF^{+}21, Sag19, SSV^{+}20, SWL15, VKBB22, WTW^{+}15, WVC21, WHL23b, XLQ^{+}22, XAQ22]. Model-Driven [PPF^{+}21]. Modeling [ASA^{+}22, BBC16, BSG^{+}16, FABC21, GAPG16, HWSN13, LMK^{+}13, SSR^{+}22, WRT^{+}21, ZWWF15]. Modelling [KRG^{+}17]. Models [AYG^{+}21, AANN15, CMB18, CMJ21, CTC^{+}17, FSK20, GPT^{+}21, HK17, MVS21, PP23b, PPF^{+}21, TAT^{+}22, TW22]. Modern [PASK21, TM14]. Modifications [TM14]. Modified [Sag19]. Modular [BM20, Mél22, Pla21a, Pla21b, SK23b]. Molecular [TGDC^{+}21]. Moment [AJL^{+}21]. Monitor [BGS^{+}21, CLA^{+}22, CH21, SKKN20, WLG^{+}21]. Monitored [WC22]. Monitoring [AAEKM13, CYL^{+}15, DAF^{+}22, LCY^{+}19, MLLU20, PKR22]. Monolithic [GKC21, LK19, YJ22]. Monte [VGP^{+}21]. Morphus [GWHG17]. Most [XCW^{+}14]. mother [TFK20]. mother-of-pearl [TFK20]. Motor [PCG^{+}21]. Mounted [TFM^{+}19]. Movement [GVO^{+}23, KPL^{+}21]. Moving [AAKJJ22]. MPSoCs [LRYK14, RMSD23, WFB^{+}18]. MRAM [CGC^{+}24, GZO^{+}18, LBX^{+}23, SKM^{+}23, SAD23, WLG^{+}21, WRT^{+}21].

MRAM-Based [SAD23]. MRAM-Centric [LBX^{+}23, CGC^{+}24]. MTFC [LXX^{+}21]. MTJ [WLG^{+}21]. Multi [BEG^{+}24, BUS^{+}21, DVPQ^{+}21, DZK^{+}23, FVS24, HMSZ20, HZY21, JKL21, LLD^{+}15, LDOG16, LK19, LXX^{+}21, LVJ22, MBW23, MRBS22, MM18, NLF^{+}22, NW20, QLT17, RLK22, TZZB21, TDVS18, TDVS21, WCC^{+}20, WC22, XSY^{+}21, ZMX^{+}22, ZMK22]. Multi-Access [RLK22, ZMK22]. Multi-Authority [XSY^{+}21]. Multi-Bit [NW20, TZZB21]. Multi-Clouds [LDOG16]. Multi-Core [BUS^{+}21, DVPQ^{+}21, MM18]. Multi-GHz [TDVS18]. Multi-GPU [LXX^{+}21]. Multi-Keyword [LLD^{+}15]. Multi-Level [HMSZ20, MBW23]. Multi-Objective [JKL21, QLT17, WCC^{+}20]. Multi-Person [WC22, ZMX^{+}22]. Multi-Protein [BEG^{+}24]. Multi-Rate [TDVS21]. Multi-Stage [FVS24]. Multi-Task [HZY21, LVJ22]. Multi-Technique [MRBS22]. Multi-Tier [LK19]. Multi-Tiered [RLK22]. Multi-UAV-Enabled [DZK^{+}23]. Multi-View [NLF^{+}22]. Multiagent [WJ15]. Multiagent-Based [WJ15]. Multicast [CYP^{+}16, KMD^{+}18, SCFH13]. Multichip [AMG22]. Multiclass [WLZG22]. Multicore [BM20, CCW^{+}14, DFRR23, FABC21, NSH22, NAM17]. Multicue [FWC15]. Multidimensional [ÁAPDD24, ZLW^{+}21]. Multihop [OZL15]. Multilabel [HWL15]. Multilayer [HZS^{+}15, KKM17]. Multilevel [LRL21, RGP^{+}21]. Multimedia [HXL^{+}14]. Multimodal [SG20]. Multiobjective [ZLX18]. Multiplayer [GP15b, HWFR15]. Multiple [ABTH20, CMB18, CF19, CLW^{+}18, ET23, GWYJ21, IPiR18, MZGT17, NT21, TA19, XTXY16, ZYZ20]. Multiple-Access [IPiR18]. Multiple-Precision [ET23]. Multiple-Type-Defect [TA19]. Multiplexers [TDG17]. Multiplexing [PKK22a]. Multiplication [KDKB22, SK23b, BMP23]. Multiplications [eSKZW^{+}22]. Multiplier [BAS23, MDB^{+}22, RAD22, TDVS18]. Multipliers [AM22, ABDL19, CSK^{+}23, LZM^{+}21, WWXL22, WWL^{+}22, ZNS^{+}22]. Multiprocessing [RDS^{+}22]. Multiprocessor [PS21, WLC^{+}14]. Multiregional [CRi^{+}19]. Multiserver [THTK16]. Multispectral [TKTP21]. Multitask [HBY^{+}23]. Multitasking [ZCT^{+}14]. Multitier [AKC14]. Multivariate [XCY^{+}24]. Mutability [PCAP21]. Mutual [THTK16, ZLX^{+}14]. Mutual-Relationship-Based [ZLX^{+}14]. MVCWalker [XCW^{+}14].

Nacre**Nacre [TFK20]. Naive

[SNHN15, XZW21b]. Named [PCXF19]. NAND [MIMMY^{+}19, ZZM^{+}19]. Nano [SMMTBM^{+}22, TA19, YLZ^{+}21]. Nano-Crossbar [TA19]. Nano-scale [YLZ^{+}21]. Nanometer [BBM^{+}17]. Nanonetworks [TGDC^{+}21]. Nanophotonic [LFL^{+}18]. Nanoribbon [BS15]. Nanoscale [BMC17, KP14, MM14]. Nanotechnology [MPK18, RPR^{+}22]. Nanotechnology-Based [RPR^{+}22]. National [WZH^{+}16]. Natural [GJZ^{+}16, PMC^{+}15]. nature [TFK20]. Navigation [VGP^{+}21, ZLW^{+}21]. NBTI [PIK20]. Near [AMTR24, BPB^{+}20, CP23, HEYB22, LBX^{+}23]. Near-Memory [CP23]. Near-Sensor [AMTR24]. Near-Threshold [BPB^{+}20]. Nearest [LRHL21]. Need [OAP^{+}22]. Negative [Isl21]. Neighbor [CLWX18, SSL^{+}13]. Neighbors [LRHL21]. NEMFET [ELVC18]. NEMFET-CMOS [ELVC18]. NEMS [MSAS17]. Netlist [YGLO22]. Network [AZA^{+}21, AM22, BCX23, BK22, BCM^{+}21, CMK^{+}16, CMMF20, CLLL23, CRRS22, CGC^{+}24, DAF^{+}22, DFRR23, DTK^{+}23, DBNBT14, FNK^{+}13, GLM^{+}23, HCGS24, HXL^{+}14, HH18, JYZ^{+}23, JCM^{+}21, KMD^{+}18, KNK13, KMM15, LD21, LQT^{+}23, LJX^{+}22, MCUDCD22, MLLU20, NAMJ23, NTAL13, OKK23a, PBTP21, QZW^{+}23, RLK22, RWJZ23, SAJ22, SBAR21, SAD23, SSL^{+}13, SZL^{+}22, SNK^{+}14, SAI^{+}19, TPM16, UBMA18, WGYL20, WZL21, XZW^{+}21a, YMT22, YZZ^{+}21, YCLW14, ZWWF13, ZL22, ZLW^{+}21, ZSL^{+}15, LMW^{+}17]. Network-Based [HXL^{+}14]. Network-Level [TPM16]. Network-on-Chip [BK22, BCM^{+}21, DFRR23, KMD^{+}18, ZL22]. Network-on-Chips [CMMF20, CLLL23]. Networking [GHSA15, XOD20]. Networks [AKU19, APH^{+}22, AMTR24, AKSR24, APCM20, AMCMDM^{+}23, BEG^{+}24, BTD24, BYB20, BPB21, BYZZ22, CCZN20, CDLS13, CYL^{+}15, CLWX18, CMRV21, DYJ22, DAS22, DY13, ESO^{+}22, FK22, FVS24, GTC^{+}21a, GTC^{+}21b, GP15a, GAPG16, HYL^{+}20, HZS^{+}15, HLL16, HYZ^{+}16, HCWL16, HHT23, KK22, KDKB22, KK23, LNK^{+}15, LLS^{+}16, LYS13, LA13, LPZ^{+}14, LJD^{+}15, LRS^{+}22, MDB^{+}21, NB21, NVS^{+}14, NT16, OZAL13, PJK^{+}19, QZW^{+}23, RSSE20, RLX15, RZD^{+}19, SBD^{+}21, SWW^{+}20, SCS^{+}20, TNKM14, TMS^{+}19, TRD^{+}24, TSS16, WJ15, WJL23, WDO^{+}19, XYDJ22, XCY^{+}24, XSZ^{+}15, XAQ22, YMDJ21, ZWWF13, ZXXH13, ZWWF15, ZLC16, ZLZ^{+}17, ZC15, ZMA15, ZSCY20, ZDLG13, ZLX^{+}14, ZMK22]. Networks-Based [APH^{+}22]. Networks-on-Chip [BYB20, RSSE20]. Networks-on-Chips [CCZN20]. NeuE [BCX23]. Neural [AM22, APH^{+}22, AKSR24, AMCMDM^{+}23, BCX23, BTD24, CRRS22, CGC^{+}24, DYJ22, DAS22, ESO^{+}22, FVS24, GTC^{+}21a, GTC^{+}21b, GLM^{+}23, HHT23, JYZ^{+}23, KK22, KDKB22, KK23, LD21, LJX^{+}22, LRS^{+}22, MLLU20, MDB^{+}21, NAMJ23, PBTP21, SBD^{+}21, SAD23, SZL^{+}22, UBMA18, XYDJ22, XAQ22, YMT22, YMDJ21]. Neural-Like [AMCMDM^{+}23]. Neurological [CMJ21]. Neuromorphic [BHC^{+}23, BUS^{+}21, CLS18, DDB22, KK21, UBMA18]. Neurons [AMJ22]. NeuSB [BHC^{+}23]. Neutron [CBG^{+}21]. NewHope [PSSZ22]. Newton [HAKL22]. Next [HQP^{+}21, LHJ18, MGMC21, PLSM20]. Next-Generation [MGMC21, PLSM20]. NFV [AZA^{+}21]. NLP [AYG^{+}21]. nm [TAC^{+}19]. NMF [CCH^{+}23]. NMF-Based [CCH^{+}23]. NoC [DBOB20, LYA18, RMSD23, WFB^{+}18]. NoC-Based [WFB^{+}18]. Node [HYL^{+}20, HLW14, WT20, YLZ^{+}21, YXF^{+}22, YCH^{+}23]. Node-Based [HYL^{+}20]. Nodes [GTC^{+}21a, GTC^{+}21b, KMM15, STSC24]. NodeTrix [LDJ20]. Noise [KK21, LJX^{+}22, MBW23, PP23a]. Noise-Immune [LJX^{+}22]. Noise-Resilient [MBW23]. Noise-Shaping [PP23a]. Noisy [SNHN15]. Non [ATPR23, HEYB22, LSP^{+}20, SKD20, SFZ^{+}16, TFK20, AS24]. Non-IID [AS24]. Non-Intrusive [SFZ^{+}16]. Non-Submodular [LSP^{+}20]. Non-Volatile [ATPR23, HEYB22, SKD20, TFK20]. Nondestructive [ZMRM19]. Nonlinearly [ZWPL23]. Nonvolatile [AMJ22, HMB^{+}21, OMTH17]. NoSQL [GWHG17, GSVA23]. Notes [GKSJ21]. Notice [SSVJ14]. Novel [ABG^{+}22, BYB20, CDM20, CMRV21, FBL^{+}22, FKNK21, HKC22, HLX^{+}17, KNK13, KTM19, LQT^{+}23, LA13, NYC^{+}21, PBTP21, RGS20, TFM^{+}19, WMAB17, YLZ^{+}21, YXF^{+}22, ZMRM19, ZLXL22]. NP [SLR^{+}17]. NP-Dynamic [SLR^{+}17]. Nuclear [FBL^{+}22]. Null [NBRF18]. Number [CDF^{+}22a, CRP^{+}23, Mél22]. Numbers [BJ21, WX23]. Numerical [AAA18, GWD^{+}23]. Nursing [GKSJ21]. NVM [BHPE21]. NVM-Based [BHPE21].

O [ATA21, LZD^{+}22]. Obfuscatable [SFL17]. Obfuscation [AYL^{+}21, KKC21, PASK21]. Object [LHP^{+}18, MYT^{+}18, NT16, OMSO20, OPIV^{+}23, WMAB17]. Object-Based [OMSO20]. Objective [JKL21, QLT17, WCC^{+}20]. OBP [LRL21]. Obtaining [YQ14]. Off [BBW22, MKH24, PS21, CYLJ21]. Offering [ZSL^{+}15]. Offload [CMB18]. Offload-Based [CMB18]. Offloading [AANN15, BYZZ22, CLWG15, HLL16, RLK22]. Offset [LVR^{+}24]. Older [RdPF^{+}23]. OLED [JDP^{+}21]. OLT [CKC^{+}18]. On-Board [FO21]. On-Chip [KKM17, KS18, LFL^{+}18, MM14, PG23, GCW20]. On-Demand [CKC^{+}18, LYA18, LWM^{+}21]. On-Line [BUS^{+}21, CKC^{+}18]. On-Time [LYS13]. One [AMG22, LZW21, LRL21, TDG17, TFK20]. One-Bit [LRL21]. One-Level [TDG17]. One-Round [LZW21]. One-To-Many [AMG22]. Online [CPM^{+}17, CH21, DAF^{+}22, DBOB20, DBNBT14, GWHG17, GP15b, HH16, HWFR15, HH18, MKAR22, MDB^{+}21, RWJZ23, TSS16, WMN13, XYZ^{+}17]. Ontologies [HB15]. Ontology [KAF^{+}16]. Ontology-Based [KAF^{+}16]. Open [BSG^{+}16, CTC^{+}17, GBVS21, MMD^{+}22a, MMD^{+}22b, MVS21, PMLT21]. Open-Source [MMD^{+}22a, MMD^{+}22b]. Opening [RPK^{+}22]. Operand [VC17]. Operating [CLW^{+}18]. Operation [HLW14, MIMMY^{+}19]. Operations [AAA18, BJ21, BBC16, FM21b, FM21c, PKK22b, SLB^{+}20, WX23]. Operator [SAI^{+}19]. Opinion [AKLC21]. Opportunistic [XSZ^{+}15]. Opportunities [BCSF17, LMS^{+}14, LPZ^{+}14]. Optics [EDLT21]. Optimal [Che14, DF22, LS21, WMN13, WCG^{+}21, YW22, YDM^{+}18]. Optimise [PLA20]. Optimization [CCW^{+}14, DDD20, DZK^{+}23, EGR21, FWC15, GLM^{+}23, HBY^{+}23, HZY21, LL13, LGC13, LLS^{+}16, LJX^{+}22, MKH24, OKK22, OKK23a, OKK23b, PKK22b, PIK20, RPVWRS^{+}20, RSM^{+}21, SSVJ14, WCC^{+}20, YZG^{+}24, ZLX18]. Optimized [HHT23, JCM^{+}21, QGF^{+}23, SK23b, SBHL21, YXF^{+}22, ZZZY24]. Optimizing [CYLJ21, TMCVH21, UBMA18]. OR-AND-XOR-PUF [YPS^{+}22]. Oracle [APH^{+}22]. Oracle-Less [APH^{+}22]. Orbit [TKK^{+}22, WZR^{+}23]. Order [HCGS24, ZWH^{+}24]. Orders [YW22]. Organizing [HXL^{+}14]. Oriented [BPB21, CRi^{+}19, CMRV21, CGC^{+}24, HYL^{+}20, HKZH16, HCCL13, KZT^{+}20, LCY^{+}19, NB21, SWL15, SBD^{+}24, ZSCY20, KSKA22]. Orthogonal [PKK22a]. Oscillator [ADN^{+}21, BNCF14]. Outage [OMTH17]. Outcomes [CG17]. Outdated [ZLC16]. Outfit [BTD24]. Outgoing [Mon20]. Output [TZZB21]. Outsourced [CWK^{+}21, LWN^{+}22, NT21, YLY^{+}20]. Outsourcing [WZRR13]. Overbuilding [CWZ^{+}22]. Overhead [JDP^{+}21, NSH22, TM14, YXF^{+}22]. Overheads [AMRCP21, GVO^{+}23]. Overlapping [LDJ20]. Overlay [SNK^{+}14]. Overlay-Based [SNK^{+}14]. Overproduction [JBSS^{+}22]. Overscaling [BAS23].

P4 [NLF^{+}22]. Packet [CH16, OZL15]. PADS [BPBG18]. Pairs [Sag19]. Pairwise [ZXXH13]. Paper [LSR^{+}21]. Papers [KMW^{+}21]. Paradigm [AMRCP21, BPB^{+}20, DYJ22, MWW^{+}21, PV15, RDS^{+}22, YCL^{+}22]. Paradigms [AM19, AL19, DSS21, MA18]. Parallel [BJ21, DLTX21, JKL21, LYW^{+}21, PP23b, RTVG22, VSS18, WH21, BKS21]. Parallel-Prefix [PP23b]. Parallelism [CLD^{+}24]. Parallelizable [AmSYlrm24]. Parallelized [MDB^{+}23]. Parameter [XWZ^{+}23]. Parameters [AAO^{+}20]. Pareto [YW22]. Pareto-Optimal [YW22]. Parity [LRL21, RZAD18]. Parity-Preserving [RZAD18]. Parking [SSF^{+}22]. PaRQ [WLZ^{+}13]. Part [KP14, LDLN16, LDLN17, GFKL13a, GFKL13b, KGM15a, KGM15b]. Partial [BYZZ22, WWL^{+}22, YW22]. ParticipAct [CCFI16]. Participation [AKLC21]. Particle [CLA^{+}22, HBY^{+}23]. Partitioning [TDZ21, ZLX^{+}14]. Passive [RS20]. Past [DDD^{+}14]. Patches [XSYW20]. Path [HYL^{+}20, KS18, PP23b, SAI^{+}19]. Path-Based [PP23b]. Paths [MLRRG20]. Patient [LDCY21]. Patient-Centric [LDCY21]. Patients [CFM^{+}22, KPL^{+}21]. Pattern [AMTR24, HDA^{+}20, WCG^{+}21]. Patterning [CF19]. Patterns [CMM^{+}21, SY19]. PCB [PKR22]. pDAE [ZXLS21]. Peak [CWZ^{+}15, YKAE22]. Peak-Power-Aware [YKAE22]. pearl [TFK20]. People [BGS^{+}21]. Perception [HJCK21]. Perceptions [HH16]. Perceptron [KKM17]. PERCIVAL [MMD^{+}22a, MMD^{+}22b]. Performance [APA^{+}21, BHPE21, BMC17, CKKO20, CCZN20, CDLS13, CMMF20, CMRV21, FKNK21, HLK^{+}24, KMM15, MM18, NAMJ23, NS15, OZAL13, PS21, QGF^{+}23, RdPF^{+}23, RMSD23, SDZ^{+}21, SFZ^{+}16, TZZB21, TDG17, TAT^{+}22, TH16, VC17, WCC^{+}20, WWXL22, WWL^{+}22, YCH^{+}23, ZL22, ZZM^{+}19]. Performance-Efficient [ZL22]. Performing [WX23]. Permanent [CKC^{+}18, NSH22]. Persistent [LZD^{+}22, LHZS20]. Person [WC22, ZMX^{+}22]. Personal [DGP22, KZT^{+}20, SWK^{+}17]. Personalized [BSG^{+}16, DVPQ^{+}21, HQP^{+}21, LLD^{+}18, WLY^{+}21, ZXLS21]. Perspective [PLJ15]. Perturbation [XWZ^{+}23]. Pervasive [YMDJ21]. PESKEA [EAM21]. Phase [CRi^{+}19, CH21, FM21a, QGF^{+}23]. Phase-Smoothing [CRi^{+}19].

Phenomenon [BNCF14]. Photonic [BYB20]. PHY [TDVS21]. Physical [AEM22, AR20, AAEKM13, BY13, CYL^{+}15, DY13, DA20, GLC^{+}13, GZG^{+}17, GLC^{+}21, GFKL13a, GFKL13b, HMSZ20, HHM20, HCCL13, JKSC21, KNK13, KK14, LL13, LPY^{+}13, LDJ20, LMK^{+}13, LCY20, LCY^{+}19, SAAJ22, SCFH13, SLC^{+}13, SGH13, SQX^{+}20, SNK^{+}14, TSH^{+}17, VHFH^{+}22, WCC^{+}20, WZY^{+}20, YQ14, ZWWF13, ZYZ20]. Physically [BR16]. Physiological [NSKRJ16]. PIM [SBD^{+}24]. Pin [CPH^{+}15]. Pipeline [JSZ^{+}21]. PISA [ATPR23]. Pivoting [APCM20]. Pixels [TS21]. Placement [KKC21, OZL15, SP20, WZH^{+}16]. PLAM [MDB^{+}22]. Planning [HYL^{+}20]. Platform [AKLC21, CCFI16, CCW^{+}14, CMK^{+}16, FK22, GVF^{+}23, LTL^{+}22, PPF^{+}21, WX23]. Platform-Aware [PPF^{+}21]. Platforms [BUS^{+}21, DAS22, JMBR^{+}17, LC20, MVS21, NAM17, UBMA18]. Plausible [WJL23]. PlausMal [WJL23]. PlausMal-GAN [WJL23]. Player [HWFR15]. Playing [HWFR15]. PLSM [MDB^{+}23]. PM [LZD^{+}22]. PM-AIO [LZD^{+}22]. PMNS [DRV22]. Point [FM21b, FM21c, HQP^{+}21, BMP23]. Point-of-Interest [HQP^{+}21]. Pointing [HMB^{+}21]. POIs [MCB21]. Poisoning [KKY24]. Policies [TCNC16]. Policy [CYBD15, HAE22, ZWZ18]. Polling [KMM15]. Pollution [MLLU20]. Polymorphic [LRP^{+}22, RPK^{+}22]. Polynomial [AGMP21, Geu20, eSKZW^{+}22, Mél22]. Polypharmacology [GVF^{+}23]. Polyphase [GZT^{+}22]. PolyWorm [LRP^{+}22]. Pooling [LXX^{+}24]. Population [WVC21]. Portrait [QWC^{+}18]. Posit [CRRS22, MMD^{+}22a, MMD^{+}22b, MDB^{+}22]. Positive [KTM19]. Possible [OBM22]. Post [CSK^{+}23, RM20, RPVWRS^{+}20, SBGC22, XHWI22]. Post-Quantum [CSK^{+}23, XHWI22]. Post-Silicon [RM20, RPVWRS^{+}20, SBGC22]. Power [AAEKM13, AC22, ACH22, BBM^{+}17, BY13, CPH^{+}15, Gel15, HLW14, KWS^{+}20, KNK13, KKM17, MSS21, NHT^{+}19, OMTH17, PKR22, RPK^{+}22, RSK21, SSVJ14, TDVS18, YKAE22, YKP22, YK18, ZNS^{+}22, ZL22]. Power-Aware [KKM17]. Power-Outage [OMTH17]. Powered [HAE22, MWW^{+}21]. Powering [CPH^{+}15]. Practical [NLF^{+}22]. Pragmatic [CYBD15]. Pre [ZXXH13]. Precise [GCBK17]. Precision [BDd^{+}21a, BDD^{+}21b, ET23, LCR^{+}21, SZA^{+}23, YJ22]. Predict [AS20]. Predictability [WCF23]. Predictable [BHPE21]. Predicting [QZW^{+}23, TRD^{+}24, ZYY^{+}24]. Prediction [ACCL23, CLA^{+}22, DVPQ^{+}21, GKSJ21, HWFR15, JYZ^{+}23, MKAR22, QWC^{+}18, SR14, WGYL20]. Predictive [BPC^{+}17]. Preference [TD19]. Preference-Enriched [TD19]. Preferences [MYS17]. Prefix [PP23b]. Preparation [SUQKA20]. Preschoolers [SPCB16]. Presence [ADN^{+}21, DLS21, MRS21]. Present [DDD^{+}14]. Presentation [CG17]. Preserving [AK24, AYG^{+}21, AS24, ABTH20, CLL21, HLK^{+}24, HSS24, LZW21, LDCY21, PTM23, RZAD18, WXT^{+}24, WLZ^{+}13, XYH^{+}23, YZG^{+}24, ZDLG13]. PresQ [ÁAPDD24]. Prevention [PASK21]. Price [DHGR18]. Priced [AFATAH13]. Pricing [LDOG16]. Primary [HÁMLS23]. Primitive [BBF18]. Primitive-Based [BBF18]. Principles [SSVJ14]. Printability [CF19]. Prioritization [APCM20, PMLT21]. Privacy [AK24, AYG^{+}21, AS24, ABTH20, CLL21, CWK^{+}21, DGP22, FMA24, GZFS18, HLK^{+}24, HSS24, LSP^{+}20, LZW21, LDCY21, MZY^{+}16, MYS17, MKH24, NTAL13, PTM23, RMVN22, WZRR13, WHL23a, WXT^{+}24, WLY^{+}21, WLZ^{+}13, XZW21b, XYH^{+}23, YZG^{+}24, ZZZY24, ZDLG13]. Privacy-Assured [WZRR13].

Privacy-Based [WLY^{+}21].

Privacy-Preserving [AK24, AYG^{+}21, AS24, CLL21, HSS24, LDCY21, PTM23, WLZ^{+}13, XYH^{+}23, YZG^{+}24, ZDLG13]. Privacy-Utility [MKH24]. Private [BBE^{+}24, SZL^{+}20]. Privilege [HTH^{+}22, TM14]. privy [ABTH20]. Pro [BBF18]. Proactive [WLX^{+}16]. Probabilistic [HHS23, PSZD21, SAD23, VGP^{+}21, ZCZ^{+}15]. Probability [PSSZ22]. Problem [JM16, TDZ21, WJ19]. Problems [RdPF^{+}23]. proceedings [IEE21, IEE22]. Process [ADN^{+}21, AC22, BRRE22, CLD^{+}24, PPF^{+}21, SKD20]. Processing [AM22, ABDL19, ATPR23, AMTR24, BBC^{+}22, CP23, CRRS22, FO21, GZLG14, HTS^{+}23, HEYB22, LDOG16, LWC22, LCR^{+}21, MBW23, NW20, PRKK24, RDS^{+}22, RWZ^{+}16, SBD^{+}24, ZK23, BKS21]. Processing-In-Memory [HTS^{+}23, CP23, MBW23]. Processing-in-Sensor [ATPR23]. Processor [CRS17, EMDE^{+}22, HMB^{+}21, KK23, KK14, OMTH17]. Processor-Based [KK14]. Processors [BM20, BCD^{+}20, BBT^{+}16, CRRS22, KM20, MM18, OMTH17]. Product [FWC15, HLK^{+}24, PSZD21, WWL^{+}22]. Product-Based [WWL^{+}22]. Production [DAF^{+}22, VHFH^{+}22]. Profile [RdPF^{+}23]. Profiling [BBF18, EAM21, LXL^{+}22a]. Profits [DHGR18]. Programmable [KK23, SBD^{+}24, TDVS21]. Programmers [EGGOR20]. Programming [BBF18, CMB18, CO16, HHM20, SSV^{+}20, WLZG22, ZYC^{+}22, ZCZ^{+}15]. Progression [EGGOR20]. Project [SWK^{+}17, ZXLS21]. Project-Based [SWK^{+}17]. Promises [RPR^{+}22]. Promoted [LSR^{+}21]. Promoter [LSR^{+}21]. Prompt [CH23]. Proof [RMVN22]. Propagate [DPP21a, DPP21b]. Propagation [WJ19, XLQ^{+}22]. Properness [GZB22]. Properties [VS17]. Property [VS17, XWZ^{+}23]. Proposal [FKNK21, WFB^{+}18]. Proposed [PCAP21]. Protect [AFMM19]. Protecting [GZFS18, LRXW21, LWN^{+}22]. Protection [FMA24, KJC22, LZL^{+}21, MZY^{+}16, PASK21, SWW^{+}21, TSS16, XWZ^{+}23, ZWZ18]. Protein [BEG^{+}24]. Protocol [ABC18, FSCX17, JCM^{+}21, SSL^{+}13, SSVJ14, THTK16, Wan22, WWG^{+}22, ZDLG13]. Protocols [CDLS13, HHS23, LHB^{+}15, PTM23]. Prototype [AS20, EKO^{+}16]. Provable [Dul20]. Provably [APH^{+}22, JBSS^{+}22]. Provably-Secure [JBSS^{+}22]. Provenance [HKZH16]. Provide [HSG19]. Provident [HZQ^{+}18]. Providers [ABTH20, MZGT17]. Provision [PLA20, WZH^{+}16]. Provisioning [HLW14]. PROWL [HAE22]. Proximity [HSzXZ17]. Proxy [HLL16]. Prune [DYJ22, YCL^{+}22]. Pruning [JDL^{+}23, KK23, QLL^{+}22]. Public [AFATAH13, AKLC21, CWK^{+}21, GPT^{+}21, SFZ^{+}16, XYZ^{+}17]. Public-Key [CWK^{+}21]. Publication [Ano22g, PMLT21, SSVJ14, Ano14a, Ano14b, Ano14c, Ano14d, Ano15c, Ano15d, Ano15e, Ano15f, Ano16b, Ano16c, Ano16d, Ano16e]. Publishing [CLL21]. PUF [BMB^{+}21, BMC17, BNCF14, CCZZ21, PTM23, RRFT16, RMK^{+}14, WWG^{+}22, YPS^{+}22]. PUFs [AEM22]. Pulling [WQG^{+}22]. Purpose [HQP^{+}21]. PVMC [BRRE22].

Q [TMS^{+}19]. QARMA [KKA22]. QCA [RZAD18, SMMTBM^{+}22]. QoE [ZCZY16]. QoE-Driven [ZCZY16]. QoI [LFBL14]. QoS [FNK^{+}13, GP15a, KIM^{+}18, SBM21, WZH^{+}16, XTXY16, YCLW14]. QoS-Aware [WZH^{+}16, YCLW14]. Quadrotor [BGS^{+}21]. Quadruple [YXF^{+}22]. Quadruple-Node-Upset-Tolerant [YXF^{+}22]. Quadtree [WHL23a]. Quadtree-Based [WHL23a]. Quality [AM22, ADQ21, NL21, RGP^{+}21, RLX15, XCX^{+}20, YW22]. Quality-Aware [RLX15]. Quality-Efficient [AM22]. Quantification [CKR^{+}22]. Quantifying [GKC21, LSP^{+}20, SC22]. Quantization [QLL^{+}22]. Quantized [GTC^{+}21a, GTC^{+}21b]. Quantum [BWMM22, CSK^{+}23, LZX^{+}22, OKK22, OKK23a, OKK23b, PBTP21, SMMTBM^{+}22, TMCVH21, WH21, XHWI22]. Quantum-Dot [SMMTBM^{+}22]. Quantum-Inspired [PBTP21]. Quasi [ÁAPDD24]. Quasi-Cliques [ÁAPDD24]. Queries [KRG^{+}17, WHL23a, YW22]. Query [GWYJ21, WLZ^{+}13]. Querying [SBHL21]. QuesTInSitu [MHL17]. Queueing [HK17]. Queueing-Theory-Based [HK17]. Quick [TAT^{+}22]. Quire [MMD^{+}22a, MMD^{+}22b].

Racetrack [HHC^{+}23, RTVG22]. Radiation [CKC^{+}18, LRS^{+}22, WT20, YXF^{+}22]. Radiation-Induced [LRS^{+}22]. Radio [RZD^{+}19, YDH21]. Radix [LCM^{+}23]. Rail [WHC16]. RAM [AFMM19, CKR^{+}22]. Raman [SDZ^{+}21]. RAMs [WZR^{+}23, ZMRM19]. Random [CRP^{+}23, CLS18, XCW^{+}14, YQ14]. Range [CH16, SDW^{+}21, WHL23a, WLZ^{+}13]. Range-Enhanced [CH16]. Ranked [LLD^{+}15]. Ranking [BPB21, HQP^{+}21, RGP^{+}21]. Ransomware [HDA^{+}20]. Rate [CYLJ21, RZD^{+}19, SKKN20, TDVS21]. Rational [WX23]. Raw [CLL21]. RC [ESO^{+}22]. RC-RNN [ESO^{+}22]. RDAM [YZZ^{+}21]. RDMA [SSL^{+}13]. RE [CKC^{+}18, YLW23]. Re-Synchronization [YLW23]. Read [WLG^{+}21]. Reader [JM16]. Reading [HLM^{+}13]. Real [ASYK^{+}22, BM20, BJ21, BHPE21, CP23, Che14, DVPQ^{+}21, JSZ^{+}21, MYT^{+}18, MTFK21, MS19, RSSE20, SBGC22, WMN13, YKAE22]. Real-Time [ASYK^{+}22, BM20, BHPE21, Che14, DVPQ^{+}21, JSZ^{+}21, MYT^{+}18, MTFK21, MS19, RSSE20, SBGC22, WMN13, YKAE22]. Real-World [CP23]. Realistic [LHB^{+}15]. Reality [BGS^{+}21, LMS^{+}14, MRS21, RS20, SUQKA20]. Reality-Based [LMS^{+}14]. Realized [ZWPL23]. ReaLPrune [JDL^{+}23]. Reasoning [HTS^{+}23]. Reasons [SG20]. Receiver [LYS13, XL19]. Recharging [KMM15]. Recoder [PSZD21]. Recognition [BDL^{+}13, DTK^{+}23, KK22, SZL^{+}22, WLO^{+}21, WC22, XCX^{+}20]. Recommendation [AXAD21, GLC^{+}13, HQP^{+}21, KMW^{+}21, MCB21, NWSG17, TFM^{+}19, XCW^{+}14, ZHL^{+}14, ZHC^{+}14, ZXLS21, ZLW^{+}22]. Recommendations [PGM^{+}21]. Recommender [SSF^{+}22, ZHC^{+}14]. Recommending [RSH^{+}21]. Recomposited [YPS^{+}22]. Recomputation [AK21]. Reconfigurable [ASKG21, AFMM19, CKC^{+}18, DBLK23, DPO17, ESO^{+}22, HF17, LXX^{+}24, RPR^{+}22, TA19]. Reconfiguration [BUS^{+}21, PAR18]. Reconfigurations [GWHG17]. Reconstruction [BMB^{+}18, LLLG23, WZRR13, WASW22]. Record [ZYZZ15]. Recording [CH23]. Recovery [CKKO20, DBOB20, NW20, SK23a, XOD20]. Recurrent [ESO^{+}22, JYZ^{+}23, KK23, MDB^{+}21, SAD23]. Recursive [WWXL22, WWL^{+}22, ZNS^{+}22]. Redaction [MLH^{+}20]. ReDeSIGN [RMSD23]. Redirected [MRBS22]. Reduce [AMRCP21]. Reduced [LCR^{+}21, PP23a, YJ22]. Reduced-Length [PP23a]. Reduced-Precision [YJ22]. Reducing [KK21, MSLL14, SMMTBM^{+}22]. Reduction [CWZ^{+}15, JDP^{+}21, KHY^{+}14, Mél22, PJK^{+}19]. Redundancy [AKHA22, BM20, CYP^{+}16, LRG^{+}21, LCR^{+}21, MM18]. Redundancy-Free [LRG^{+}21]. Reflecting [OKK22, OKK23a]. Reflection [CTC^{+}17]. Reformative [LJX^{+}22]. Regimes [CLW^{+}18, LDCY21]. Region [WXT^{+}24]. Registries [AYG^{+}21]. Regression [CLY^{+}23, MKAR22, RdPF^{+}23, SWY^{+}22]. Regular [CDC^{+}23]. Regulating [DVA21]. Regulators [YK18]. Rehabilitation [PCG^{+}21]. Rei [DBLK23]. Reinforcement [HWH^{+}20, HLW14, JHL21, LDCY21, WZLK21, YSS^{+}22, YZZ^{+}21, ZLG^{+}22]. Rejuvenation [GAWT23, RKdN^{+}23]. Related [GMTX14, QZW^{+}23]. Relation [SWK^{+}17]. Relations [LYBZ20]. Relationship [KZ21, ZK21, ZLX^{+}14]. Relationships [BTC^{+}21]. Relaxation [PKK22b]. Relay [CYL^{+}15, HLM^{+}13, MSAS17]. Relay- [CYL^{+}15]. Relevance [WXT^{+}24]. Reliability [ATA21, CPM^{+}17, CSM19, DAS22, DSS21, EGR21, HK17, JHB21, KP14, KVP19, LC20, LPY^{+}13, LCM18, LRS^{+}22, MCUDCD22, MYM20, OBM22, SBGC22, YKAE22]. Reliability-Aware [LCM18, MYM20]. Reliable [AAEKM13, AAA18, AC22, BK22, CDF^{+}22b, DBOB20, EM23, GCW20, HHC^{+}23, LCR^{+}21, LRLG22, RRFT16, SP20, TAC^{+}19, YXF^{+}22, ZYLT19]. Reliable-Aware [AC22]. Reliably [ZSL^{+}15]. Relief [ZYS^{+}22]. Remaining [MKAR22]. ReMap [YKAE22]. remarkable [TFK20]. Remote [ABC18, NT16, WZR^{+}14]. Removal [YMSR20]. Renewable [HAE22, LNK^{+}15]. Repetitive [VSH23]. Replacement [HAE22]. Replay [JHL21]. Replication [YKAE22]. Report [Mon22]. Repositioning [BTC^{+}21, ZWH^{+}24]. Representation [KAF^{+}16, KMW^{+}21, KHY^{+}14, LXX^{+}24, STSC24, WZL21, YZZ^{+}21, ZWH^{+}24]. Reproducibility [HH18]. Repurchasing [LDOG16]. Repurposing [BEG^{+}24]. Requirements [LHKH21, NBS16]. ReRAM [BMC17, JDL^{+}23]. ReRAMs [OBM22]. Rescheduling [JKL21]. Rescue [ZLX18]. Research [DLS21, HH18, KDM^{+}21, LT21, MT21b, ZLW^{+}21]. Research-Active [LT21]. Researchers [RGP^{+}21]. Residential [AR20, LGT^{+}19]. Residual [ZYH^{+}20]. Residue [CDF^{+}22a]. Resilience [SKD20]. Resiliency [RZAD18]. Resilient [BSY^{+}17, MBW23, OMTH17, PKK22a, RSSE20, RMK^{+}14, LMW^{+}17]. Resistant [CWZ^{+}22, EMDE^{+}22, KTM19, RRFT16, SSL^{+}13, TFK20]. Resistive [ASKG21, DQB23, IPRR19]. Resistivity [KS18]. Resolving [SKM^{+}23]. Resonance [FBL^{+}22]. Resonant [BMB^{+}21]. Resonant-Tunnelling [BMB^{+}21]. Resource [AFG^{+}21, CMK^{+}16, CMMF20, CTL^{+}23, GZG^{+}17, HWH^{+}20, HZQ^{+}18, HYZ^{+}16, JSZ18, KK23, LWM^{+}21, OKK23b, PLA20, QLL^{+}22, TKK^{+}22, WZLK21, YDM^{+}18]. Resource-Aware [CTL^{+}23]. Responding [TMS^{+}20]. Response [BMOS16, KK14, MZGT17, SBL^{+}21]. Responses [SG20]. Restoring [LXL^{+}22b]. Results [LLW^{+}15]. Reticulated [BSJ22]. Retrieval [PCXF19, ZS15]. Returning [KK22]. Reusability [VHFH^{+}22]. Reuse [RMSD23]. Revealing [ACCL23]. Reverse [RMK^{+}14, STL^{+}14]. Reverse-Engineering [RMK^{+}14]. Reversible [CSZ^{+}20, RZAD18]. Review [GPT^{+}21]. Reviewer [CGPB21, RGP^{+}21, Ano15b]. Reviewers [Ano17b, Ano17c, Ano19c, Ano19b, Ano20b, Ano21b, Ano23a]. Rewriting [BBMM23]. RF [CKKO20]. RFID [HLM^{+}13, HLX^{+}17, JM16, KIM^{+}18, KTM19]. RFID-Based [KIM^{+}18]. RFID-Enabled [HLX^{+}17]. RGB [SWY^{+}22]. Rich [SBR^{+}22]. Ridge [CLY^{+}23, CRi^{+}19]. Rigorous [SGH13]. Ring [ADN^{+}21, BNCF14, SK23b, XHWI22]. Ring-LWE [XHWI22]. Ripple [GWD^{+}23]. RISC [CRRS22, GTC^{+}21a, GTC^{+}21b, MMD^{+}22a, MMD^{+}22b]. RISC-V [CRRS22, GTC^{+}21a, GTC^{+}21b, MMD^{+}22a, MMD^{+}22b]. Risk [HWW^{+}17, SAAJ22].

Risks [OBM22]. RLS [SDZ^{+}21]. RLWE [AK24]. RLWE-Based [AK24]. RNA [CKR^{+}22]. RNA-Seq [CKR^{+}22]. RNN [ESO^{+}22, MBW23]. RNS [APA^{+}21, BFPS22]. RO [RRFT16]. RO-PUF [RRFT16]. Roadmap [KDM^{+}21]. Robot [PLSM20]. Robotics [PLSM20]. Robotics-Based [PLSM20]. Robust [AZA^{+}21, ABDL19, AGMP21, DDB22, HCGS24, KKC21, LYK^{+}22, LVR^{+}24, MLLU20, NAMJ23, NT16, PLSM20, RS21, RMK^{+}14]. Robustness [OEM18]. ROI [MZY^{+}16]. Role [HWFR15, JMLH22, XL19]. Role-Based [XL19]. Role-Playing [HWFR15]. Room [SUQKA20]. Room-Scale [SUQKA20]. ROSETTA [KK23]. Rough [KS18]. Round [BBW22, LZW21]. Round-Off [BBW22]. Rounded [FM21b, FM21c, BMP23]. Route [MTFK21]. Router [LYA18]. Routers [TH16]. Routine [CFM^{+}22]. Routing [CCZN20, CDLS13, FSCX17, LHB^{+}15, MLLU20, OZL15, SBM21, SSVJ14, VGP^{+}21, XTXY16]. Routing-Protocol-Independent [FSCX17]. RRAM [EM23, GZO^{+}18, TDG17]. RRAM-Based [EM23, TDG17]. RRC [GAPG16]. RRC-Based [GAPG16]. RTD [SBGC22]. Rule [SWLG17]. Rules [CGPB21]. Rumor [XLQ^{+}22]. Runtime [CMB18, RKdN^{+}23].

S [HZS^{+}15, HSzXZ17]. S-Aframe [HZS^{+}15]. S-IOT [HSzXZ17]. SACC [ZLG^{+}22]. Safe [LRLG22]. SafeProtect [TCNC16]. Safety [AAO^{+}20, CH21, GZB22, JMLH22, OZAL13, XYZ^{+}17]. Safety-Critical [AAO^{+}20, GZB22]. Saliency [OMSO20]. Salient [OPIV^{+}23]. Sample [WLZG22]. Samples [CLY^{+}23]. Sampling [CLS18, NL21, RZD^{+}19]. SARS [GVF^{+}23]. SARS-CoV-2 [GVF^{+}23]. SARVE [AXAD21]. SARVE-2 [AXAD21]. SAT [CWZ^{+}22, DMRR17, WJ19]. SAT-Based [DMRR17]. SATAM [CWZ^{+}22]. Satellite [TMS^{+}19, TKK^{+}22]. Satisfying [JSZ^{+}21]. SC-MLGNRs [NS15]. SCA [FM21a]. Scaffolds [SPCB16]. Scalable [BHC^{+}23, BYB20, CCH^{+}23, CFL^{+}21, DZZ^{+}21, DBOB20, DDB22, HTS^{+}23, WSLL21]. Scale [AKU19, ABDL19, CLY^{+}23, CCFI16, GVF^{+}23, GMTX14, HNGZ19, NdCFB^{+}23, Ode21, SMMTBM^{+}22, SUQKA20, SR14, ZOS16, ZHC^{+}14, YLZ^{+}21]. Scaled [CLW^{+}18]. Scaling [BS15, JDP^{+}21]. Scan [AYL^{+}21, CMM^{+}21, KKC21, KWS^{+}20, LLW22, TBG^{+}18]. Scan-Chain [TBG^{+}18]. Scan-in [KWS^{+}20]. Scanning [SAAJ22, ZWZ18]. ScanSAT [AYL^{+}21]. Scenario [IP21]. Scenarios [KMK22]. Scenes [ZMX^{+}22]. Scheduling [BRRE22, CYBD15, Che14, DFRR23, DF22, FNK^{+}13, GAI22, LPY^{+}13, LRYK14, OKK22, RWJZ23, WLC^{+}14, XL19, YSS^{+}22, ZCT^{+}14, ZLX18, ZMA15, ZMK22]. Scheme [AZA^{+}21, APA^{+}21, BBW22, BY13, CLL21, FNK^{+}13, HSG19, LNK^{+}15, LCZ21, LQT^{+}23, LRHL21, MLLU20, QLT17, SWW^{+}21, SSL^{+}13, SFL17, SQX^{+}20, TRR^{+}19, WLZ^{+}13, WZR^{+}23, YCS22, ZMRM19, ZLXL22, ZXXH13]. Schemes [APH^{+}22, CSK^{+}23, HSzXZ17]. Scholar [ZK21]. Scholarly [MMM^{+}21, RGP^{+}21, SWW^{+}21, SBHL21, XGLW21, ZLW^{+}21]. School [HÁMLS23]. Science [AFG^{+}21, BTK^{+}20, HÁMLS23, LSR^{+}21, WVC21]. Scientific [MSS21]. Scientifically [SGH13]. Scratch [CSLG22, MLRRG20, QSLG22, QSLG22]. Scratch-DKG [QSLG22]. Screening [GVF^{+}23]. Scrubbing [SP20]. Scrubbing-Aware [SP20]. SDLSC [WLL^{+}21]. SDLSC-TA [WLL^{+}21]. SDN [AZA^{+}21, CDM20]. Search [LLD^{+}15, LLD^{+}18, LHS^{+}22, MLG^{+}15, TD19, XSY^{+}21, YMT22]. Search-Based [LHS^{+}22]. Searchable [CWK^{+}21, SWW^{+}21, TRR^{+}19]. SEARE [XCX^{+}20]. Seated [SUQKA20]. Secondary [HÁMLS23]. Secrecy [WLWQ22]. Section [AM19, BDFT24, CP23, DLTX21, DZZ^{+}21, DPO17, DHPL19, DSS21, DA20, GAWT23, HSS24, HINS21, JSBM22, JLN21, KG20, LMC18, LCLK20, LRLG22, MA18, MT19, MT21b, MS20, MPK18, MYM20, MK15, MGMC21, OSPN22, PBKL19, SK19, TH16, ZOS16, ZEL20]. Secure [AGMP21, ABG^{+}22, APH^{+}22, CH23, CCZZ21, DLTSNA21, EKO^{+}16, GWYJ21, HTH^{+}22, HKZH16, HINS21, JBSS^{+}22, KIM^{+}18, KTM19, LLW22, LLD^{+}18, MWW^{+}21, NT21, SWL15, SWW^{+}20, SLB^{+}20, TRR^{+}19, TFK20, WTW^{+}15, YLY^{+}20]. Securing [BBT^{+}16, LCY20, SBR^{+}22]. Security [AKC14, AZA^{+}21, AAKJJ22, BPB^{+}20, BCSF17, BSJ22, DDD^{+}14, DSS21, DGW^{+}23, KP14, KSU16, LRYK14, LCLK20, MM14, MJ17, MGR19, NSKRJ16, RPR^{+}22, SAJ22, SAM^{+}20, SBAR21, SGH13, WDO^{+}19, YYX15, YDH21]. Security-Driven [LRYK14]. Seed [HKC22]. Segmentation [FVS24, WXT^{+}24]. Segmented [YLW23]. Segmented-Edit [YLW23]. Selecting [KM20]. Selection [LJD^{+}15, RM20, RdPF^{+}23, TAV15]. Selectors [OBM22]. Self [BCD^{+}20, RDS^{+}22, RZAD18, SKKN20, SBL^{+}21]. Self-Adjusting [SKKN20]. Self-Aware [RDS^{+}22]. Self-Checking [RZAD18]. Self-Evaluation [SBL^{+}21]. Self-Test [BCD^{+}20]. Self-training [SBL^{+}21]. Semantic [DGP22, HXL^{+}14, HZS^{+}15, KAF^{+}16, MMPP15, PV15, PMC^{+}15, SG15, SNHN15, ZS15]. Semantically [BBS^{+}15]. SEMKC [NT21]. Sender [XL19]. Sender-Receiver [XL19]. Sensible [DLTX21]. Sensing [AFATAH13, BMB^{+}18, DGP22, HTS^{+}23, HSG19, WZR^{+}14, XCX^{+}20, YYX15]. Sensing-Based [YYX15]. Sensitive [BYZZ22]. Sensitization [RS21]. Sensor [ATPR23, AMTR24, BPB21, CDLS13, CYL^{+}15, CLWX18, DY13, GCBK17, HYL^{+}20, HLW14, PBKL19, RLX15, RZD^{+}19, SCS^{+}20, SAI^{+}19, TNKM14, WLX^{+}16, ZSL^{+}15]. Sensors [CRi^{+}19, DVPQ^{+}21, KK21, TS21, WCF23, XYZ^{+}17, YMDJ21]. Sensory [LFBL14]. Separation [ZLC16]. September [IEE22]. Seq [CKR^{+}22]. Sequence [GWYJ21]. Sequences [GDP22]. Sequential [KJC22]. Serial [TDVS21]. Series [HWW^{+}17, JYZ^{+}23, PL19, WLL^{+}22, XCY^{+}24]. Serious [CSBME17, SUQKA20, SG20]. SERMO [DVA21]. Server [DBNBT14, RSK21, VGP^{+}21]. Server-Side [VGP^{+}21]. Service [ABTH20, BPB21, CMMF20, CMRV21, GPT^{+}21, HZS^{+}15, HSzXZ17, HSG19, KIM^{+}18, MLG^{+}15, PBKL19, SBR^{+}22, SSR^{+}22, VKBB22, WZRR13, WZY^{+}20, YCLW14, ZWZ18]. Service-Based [VKBB22]. Service-Oriented [CMRV21]. Services [JYZ^{+}23, LZW21, WZH^{+}16, WDO^{+}19, XZW^{+}21a, ZHL^{+}14]. Sesquilinear [AAA18]. Set [EGGOR20]. Shape [RB22]. Shaped [ASA^{+}22]. Shaping [PP23a]. Sharded [GWHG17]. Sharding [HHS23]. Sharding-Based [HHS23]. Shared [OKK23b]. Sharing [HTH^{+}22, HHT23, JSZ18, TCNC16]. SHARKS [SAAJ22]. Shield [XYH^{+}23]. Shielding [LRYK14]. Shift [BPB^{+}20]. Shifting [BYB20]. Shopping [FWC15]. Short [LQT^{+}23, SNHN15]. Should [RMB^{+}19]. Side [BY13, VGP^{+}21, WMP^{+}24]. Side-Channel [WMP^{+}24]. SIEDEG [MT19]. SIEGDA [MT21a]. SIEGRMI [MT21b]. Sign [RAD22]. Signage [SCFH13]. Signal [BMB^{+}18, JBSS^{+}22, LZX^{+}22, RM20, YYX15]. Signalling [GAPG16]. Signals [BDL^{+}13]. Signature [IP21, RPM23, SFL17, Wan22]. Silicon [CPH^{+}15, GHSMM21, RM20, RPVWRS^{+}20, SBGC22, YQ14]. Similarity [GSMGP17, LZLC22, MKH24, RLE^{+}22, RPM23, SNHN15, YMT22]. Similarity-Based [GSMGP17]. Simplified [Wan22]. SIMT [MM18]. Simulated [WH21]. Simulation [BGS^{+}21, BBC16, LHB^{+}15]. Simulations [UBMA18]. Single [CLA^{+}22, HJCK21, IPiR18, SWY^{+}22]. Single-Charge [IPiR18]. SIRDAM4.0 [CDF^{+}22b]. SiRF [PTM23]. Situ [CVP^{+}22]. Size [Pla21a, Pla21b, TAC^{+}19, WLZG22, ZLG^{+}22]. Sizing [EGR21]. Skeleton [FVS24, YNA^{+}20]. Skeleton-Based [FVS24, YNA^{+}20]. Skills [SBL^{+}21]. Skybridge [SLR^{+}17]. Skyrmion [HHC^{+}23]. Slicing [LYA18]. Small [DRV22]. Smart [AXAD21, BY13, CFM^{+}22, CFL^{+}21, DGP22, JSBM22, KR21, KTM19, LDJ20, LWM^{+}21, LLW^{+}15, MDB^{+}21, PMLT21, SAAJ22, SSF^{+}22, SCFH13, SWLG17, VKBB22, WZLK21, WC22, WCF23, ZYMG16, ZCZ^{+}15, ZK23, BY13, DHPL19, GLZ19, HNGZ19, HWSN13, LGC13, LMK^{+}13, LA13, MZGT17, MS19, NTAL13, SY19, TSH^{+}17, WMN13, WLZ^{+}13]. Smart-Homes [KR21]. Smartphone [QWC^{+}18]. Smartphones [AANN15]. Smoothing [CRi^{+}19]. SMPC [KKY24]. Social [AKU19, AXAD21, BSG^{+}16, DA20, DBNBT14, HZS^{+}15, HH18, LDJ20, NB21, NWSG17, PCXF19, SWW^{+}20, SQX^{+}20, TSS16, WJ15, WCC^{+}20, WZY^{+}20, WDO^{+}19, XSZ^{+}15, XYZ^{+}17, YCLW14, ZWWF13, ZHC^{+}14, ZWWF15, ZLZ^{+}17, ZYZ20, ZDLG13, ZLX^{+}14, HSzXZ17]. Social-Aware [PCXF19]. Socialbot [LSP^{+}20]. Society [ZK23]. Socio [ZOS16]. Socio-Technical [ZOS16]. Soft [JSZ^{+}21, LRXW21, LRS^{+}22, NW20, ZZM^{+}19]. Soft-Errors [LRS^{+}22]. Software [AKSR24, BCD^{+}20, CMK^{+}16, DZD^{+}18, GAWT23, HCWL16, HK17, JYZ^{+}23, MTFK21, NAM17, PPF^{+}21, RKdN^{+}23, SBM21, TSH^{+}17]. Software-Based [BCD^{+}20]. Software-Defined [CMK^{+}16, HCWL16, SBM21]. Solar [CLA^{+}22, YKP22]. Solid [ZZM^{+}19]. Solution [ADCS22, CCZZ21, ZLXL22]. Solutions [OBM22, PCAP21, VPTH19]. Sorting [QGF^{+}23]. SOT [WZR^{+}23]. Sound [FM21a, IMZ^{+}21]. Source [LLS^{+}16, MMD^{+}22a, MMD^{+}22b]. Space [CDF^{+}22a, EDLT21, FAP21, FO21, HHM20, KK21, TKTP21, ZYMG16, ZWPL23, ZMX^{+}22]. Space-Grade [TKTP21]. Spams [ABTH20]. Spare [GHSMM21]. Sparing [ASYK^{+}22]. Sparse [GLM^{+}23, LS21, WLL^{+}21, WHL23b, ZS15]. Sparsity [CGC^{+}24, YJ22]. Sparsity-Aware [YJ22]. Sparsity-Oriented [CGC^{+}24]. Spatial [FVS24, HNGZ19, Sag19, WHL23a, ZYLT19, ZYY^{+}24]. Spatial-Temporal [FVS24, ZYY^{+}24]. Spatio [HQP^{+}21, WGYL20]. Spatio-Temporal [HQP^{+}21, WGYL20]. Spatiotemporal [KK21]. Speaker [ZMX^{+}22]. Spear [XYH^{+}23]. Special [AM19, BDFT24, DLTX21, DZZ^{+}21, DPO17, DHPL19, DSS21, DO18, DA20, FO21, GAWT23, GFKL13a, GFKL13b, HSS24, HINS21, JLN21, KP14, KGM15a, KGM15b, KG20, KSU16, LMC18, LHMW20, LDLN16, LDLN17, LSO17, LCLK20, LRLG22, MA18, MT19, MS20, MPK18, MYM20, MK15, MGMC21, OSPN22, PBKL19, SG15, SK19, SO20, SO21, TH16, WK14, ZOS16, ZEL20, MT21a, MT21b]. Specific [CDC^{+}23, LSO17, Sag19]. Spectrometer [SDZ^{+}21]. Spectrometry [FBL^{+}22]. Spectroscopy [FBL^{+}22]. Speech [BDL^{+}13, CMJ21]. Speed [GT22, eSKZW^{+}22, WHC16]. SPFAL [KTM19]. SpGEMM [PKK22b]. Spike [AMCMDM^{+}23, CLS18]. Spike-Based [AMCMDM^{+}23, CLS18]. Spiking [AKSR24, AMCMDM^{+}23, BHC^{+}23, UBMA18]. Spin [HTS^{+}23, MVC^{+}22, SBD^{+}21, WZR^{+}23]. Spin-Orbit [WZR^{+}23]. Spin/CMOS [HTS^{+}23]. Spin/CMOS-Based [HTS^{+}23]. SpiNNaker [AMCMDM^{+}23]. Spins [Gel15]. Spintronic [AMJ22, HF17, NAMJ23]. Spintronic/CNTFET [NAMJ23]. Spintronic/CNTFET-Based [NAMJ23]. Split [LCM^{+}23, PASK21]. Split-Radix [LCM^{+}23]. Spoken [HWL15]. Spotting [LVR^{+}24]. Spreading [ZWWF15]. SPRING [YJ22]. Sprinting [WCG^{+}21]. Square [ASA^{+}22]. Squaring [ACH22]. SRAM [CLA^{+}22, GZT^{+}22, KVP19, TAC^{+}19, TKTP21]. SRAM-Based [CLA^{+}22, GZT^{+}22]. SSD [ATA21, CSM19, TAT^{+}22]. SSD-Based [ATA21, CSM19]. Stability [LL13, LLS^{+}16, TSS16]. Stabilizers [BR16]. Stack [ADN^{+}21]. Stacked [ELVC18, PS21, SBD^{+}24, VC17]. Stacking [MMPP15]. Stage [FABC21, FVS24, SWY^{+}22]. Staging [Ode21]. Standalone [OKK23b]. Standby [ASYK^{+}22]. Standby-Sparing [ASYK^{+}22]. Start [ZHC^{+}14]. State [BBS^{+}15, KJC22, LCY^{+}19, MDB^{+}23, Mon20, Mon22, NBS16, ZZM^{+}19]. Static [AYL^{+}21, BS15, BBM^{+}17]. Station [ZYZZ15]. Statistical [EGR21, LCM18, PV15, WZR^{+}14]. Statistically [YQ14]. Stealthy [CDM20]. Steganographic [NT16]. STGCN [ZYY^{+}24]. Stimulation [ZCZY16]. Stochastic [ABDL19, AL19, CCW^{+}14, CH19, EDLT21, FLB^{+}19, LL13, LZH19, NL21, PP23a, PL19, TDZ21, WX23, ISI^{+}19]. Stochastic-Based [ABDL19, AL19]. Stochastically [FM21b, FM21c]. Storage [ATA21, AM19, CSM19, ESO^{+}22, LLD^{+}15, LK23, PPKN23, TGDC^{+}21, Wan22, XZW^{+}21a, YLW23, YLY^{+}20]. Store [LHZS20, SBHL21]. Storms [GAPG16].

STPR [HQP^{+}21]. Strategies [BBT^{+}16, PAR18]. Strategy [KJC22, MTFK21, NBRF18, OEM18, ZLZ^{+}17, ZMK22]. Stream [CYBD15, GJZ^{+}16, LWC22]. Streamcipher [KSKA22]. StreamFlow [CCMA21]. Streaming [LK22, LVR^{+}24, RWZ^{+}16, SFZ^{+}16]. Streaming-Image [LK22]. Streams [AAA18, PP23a]. Stroke [KPL^{+}21]. Strong [BBE^{+}24, KK14, LL13, LLS^{+}16]. Structural [KMW^{+}21, STL^{+}14]. Structure [JHL21, LWC22, PCG^{+}21, TFK20]. Structured [DMRR17]. Structures [ADN^{+}21, LDJ20, RMSD23]. STT [AFMM19, SKM^{+}23, WRT^{+}21, ZMRM19]. STT-MRAM [SKM^{+}23, WRT^{+}21]. STT-RAM [AFMM19]. STT-RAMs [ZMRM19]. STTRAM [MGR19]. Student [BMOS16, BSG^{+}16, CTC^{+}17, HH16]. Students [HÁMLS23, HWW^{+}17, ZYY^{+}24]. Studies [MHL17, RSH^{+}21]. Study [BHB^{+}14, BMOS16, BDd^{+}21a, BDD^{+}21b, LSR^{+}21, MRS21, MJ17, NBS16, PTD21, RdPF^{+}23, SFZ^{+}16, VHFH^{+}22, WVC21]. Sub [BHB^{+}14, SK23a]. Sub-Adders [SK23a]. Sub-Threshold [BHB^{+}14]. Subarea [WLL^{+}21]. Subjective [SDW^{+}21]. Submicrometer [ZMRM19]. Submicron [DO18]. Submodular [LSP^{+}20]. Substrate [GHSA15]. Substring [RMK^{+}14]. Sudden [OMTH17]. Sufficient [CCZN20]. Sum [DPP21a, DPP21b]. Summaries [BC16]. Summarization [BC16]. Sunway [ZLY^{+}21]. Supervised [CLA^{+}22, HWL15, YMT22]. Supply [HLX^{+}17]. Support [CMB18, CDF^{+}22b, GP15a, GCBK17, GSMGP17, HLK^{+}24, KZT^{+}20, KR21, KCS23, MKAR22, RSB13, RKdN^{+}23]. Supported [CO16, GZG^{+}17]. Supporting [BC16, GWHG17, OZAL13, SLB^{+}20]. Surface [KMD^{+}18, OKK22, OKK23a].

Surface-Wave [KMD^{+}18]. Surfaces [KS18]. Surrogate [FABC21, HZY21]. Surrogate-Assisted [HZY21]. Surveillance [MZY^{+}16, YSS^{+}22]. Survey [AL19, FAT^{+}14, HWFR15, IP21, LBX^{+}23, PLJ15, ZZH^{+}22]. Survivability [HWSN13]. Surviving [HYW19]. Sustainability [WK14]. Sustainable [HLW14, JND14]. Sustaining [RS20]. SW [PKK22b]. Swarm [HBY^{+}23, YMX22]. Sway [SBM21]. SWCNT [LZH19]. Sweeper [WLX^{+}16]. Switches [ASA^{+}22]. Switching [ASA^{+}22, BS15, CPH^{+}15, LMK^{+}13]. swMR [ZLY^{+}21]. Sybil [HHS23]. Symbol [GHSMM21]. Symbolic [GDP22, OPIV^{+}23]. Symposium [IEE21, IEE22]. Synapses [AMJ22]. Synchronization [DY13, YLW23]. Synchrophasors [MS19]. Synthesis [CH19, FAP21, GPT^{+}21, YNA^{+}20]. System [BTC^{+}21, CDF^{+}22a, CBG^{+}21, CSLG22, CH23, CBZ^{+}21, GZG^{+}17, HCCL13, HWSN13, JKSC21, LZD^{+}22, LA13, MR21, Mél22, MYM20, MCB21, MK15, PGM^{+}21, PMC^{+}15, RPVWRS^{+}20, RSB13, RKdN^{+}23, Sag19, SSF^{+}22, SCFH13, SBR^{+}22, SWLG17, TFM^{+}19, TMS^{+}20, TGDC^{+}21, TPM16, VGP^{+}21, XCX^{+}20, ZYMG16, ZWWF13, ZHL^{+}14, ZSCY20, LMW^{+}17]. System-Level [CBG^{+}21, ZYMG16]. Systematic [APH^{+}22, GPT^{+}21]. Systems [ADCS22, ATA21, AMG22, AM19, AR20, AAEKM13, ATPR23, ASYK^{+}22, AMRCP21, BRRE22, BMOS16, BPBG18, BHPE21, CSM19, CYL^{+}15, CQH17, Che14, CKC^{+}18, DZZ^{+}21, DBOB20, DDB22, DFRR23, DY13, DZK^{+}23, DSS21, DO18, DA20, DZD^{+}18, ESO^{+}22, EAM21, FO21, GLC^{+}13, GZB22, GVO^{+}23, GWHG17, GFKL13a, GFKL13b, HINS21, HMB^{+}21, HEYB22, IMZ^{+}21, JSZ^{+}21, JMBR^{+}17, JKSC21, KKY24, KGM15a, KGM15b, KNK13, KSU16, LL13, LGT^{+}19, LPY^{+}13, LDJ20, LWC22, LCY20, LHJ18, LCY^{+}19, MCUDCD22, MYT^{+}18, MTFK21, MS20, MYM20, MGMC21, MMY^{+}22, NSH22, RSSE20, RSM^{+}21, SAJ22, SAAJ22, SP20, SSV^{+}20, SLC^{+}13, SQX^{+}20, VHFH^{+}22, WLC^{+}14, WHC16, WCC^{+}20, WZY^{+}20, YLW23, YKAE22, YDH21, ZOS16, ZHC^{+}14].

T [TMCVH21]. T-count [TMCVH21]. T-depth [TMCVH21]. Table [Ano13d, Ano13e, Ano14i, Ano14j, Ano14k, Ano14l, Ano15k, Ano15l, Ano15m, Ano15n, Ano16j, Ano16k, Ano16l, Ano16m, Ano17l, Ano17m, Ano17n, Ano17o, Ano18j, Ano18k, Ano18l, Ano18m, Ano19l, Ano19m, Ano19n, Ano19o, Ano20k, Ano20l, Ano20m, Ano20n, Ano21k, Ano21l, Ano21m, Ano21n, Ano22j, Ano22k, Ano22l, Ano22m, Ano23j, Ano23k, Ano23l, Ano23m, Ano24c, HCWL16, SBD^{+}24]. Taihulight [ZLY^{+}21]. TAMTAMS [RHRV16]. Target [AAKJJ22, LJX^{+}22, RLX15]. Targeting [GVF^{+}23, SZA^{+}23, VGP^{+}21]. Targets [BEG^{+}24]. Task [AANN15, BRRE22, DHGR18, DFRR23, DF22, GAI22, GZFS18, HZY21, LRYK14, LVJ22, PLA20, WLC^{+}14, WLL^{+}21, YCS22, YKAE22, ZYLT19, ZLX18, ZMK22]. Tasks [AmSYlrm24, WJ15]. Taxicab [ZHL^{+}14]. Taxonomy [AK21, FAT^{+}14, YKW^{+}20]. TBP [LRL21]. TDMA [NYC^{+}21]. TDMA-Based [NYC^{+}21]. Teachers [GSMGP17]. Teams [NB21]. Technical [ZOS16]. Technique [ABG^{+}22, ASYK^{+}22, CPM^{+}17, CLLL23, EKO^{+}16, FKNK21, MRBS22, NYC^{+}21, RGS20, RS21, TMS^{+}19, XWZ^{+}23, XYH^{+}23]. Techniques [AK21, AAKJJ22, AC22, BCD^{+}20, CLW^{+}18, HWFR15, LMS^{+}14, LRLG22, Ode21, RWZ^{+}16, SC22, TRD^{+}24, VS17, YMSR20]. Technocrats [AKLC21]. Technologies [GP15b, JSBM22, KG20, LMC18, LHMW20, LSO17, MK15, OGL^{+}21, SK19, SO20, SO21]. Technology [ASA^{+}22, BS15, DO18, JHB21, RHRV16, SLR^{+}17]. TEG [KWS^{+}20].

Telecom [ABTH20]. Telecommunication [AKU19]. Temperature [CWZ^{+}15]. Temporal [FVS24, HNGZ19, HQP^{+}21, NWSG17, OMSO20, WGYL20, ZYY^{+}24]. Tensor [KHY^{+}14, ZYZ20]. Tensor-Based [KHY^{+}14, ZYZ20]. Tera [AKU19]. Tera-Scale [AKU19]. Term [GKSJ21, LQT^{+}23]. Terminal [ASA^{+}22, ASKG21]. Ternary [GT22]. Terrestrial [TMS^{+}19]. Test [ADN^{+}21, BHB^{+}14, BCD^{+}20, DDD^{+}14, DVA21, GZO^{+}18, LWN^{+}22, LHS^{+}22, NBRF18]. Test-Time [BHB^{+}14]. Testability [KVP19]. Testbed [SGH13]. Testing [CKC^{+}18, CCZZ21, DSS21, DGW^{+}23, LLW22, LYW^{+}21, TBG^{+}18, VPTH19, WRT^{+}21]. Tests [CBG^{+}21]. TETC [Mon22]. Text [KMW^{+}21]. Texts [SNHN15]. Textual [AT22]. Thank [Mon20]. Thank-you [Mon20]. Theft [ZCZ^{+}15]. Thematic [CP23, JSBM22]. Theoretic [HYZ^{+}16]. Theoretical [BR16, BY13]. Theory [Geu20, HK17, ZCZY16]. Theory-Based [ZCZY16]. Therefore [ZSCY20]. Thermal [ASYK^{+}22, LFL^{+}18, WCG^{+}21]. Thermal-Aware [ASYK^{+}22, LFL^{+}18]. Thing [HSzXZ17]. Things [BSJ22, HSG19, KM20, LLL^{+}20, LFBL14, MJ17, PLJ15, PBKL19, RSK21, SAAJ22, XSYW20, XL19, ZEL20]. Think [ZSCY20]. Thinking [CSLG22, EGGOR20, HÁMLS23, MLRRG20]. Thread [VSS18]. Threat [APCM20, HDA^{+}20]. Threats [LCLK20]. Threshold [BHB^{+}14, BPB^{+}20, HF17]. Through-Silicon [GHSMM21]. Throughput [LLG^{+}23, LNK^{+}15]. Throughput-Based [LNK^{+}15]. Thwart [GJZ^{+}16, RS21]. Ticket [JDL^{+}23]. Tier [LK19, PJK^{+}19]. Tiered [RLK22]. Tight [PSSZ22]. Time [AC22, ASYK^{+}22, BM20, BHB^{+}14, BHPE21, BMC17, CLWG15, Che14, DVPQ^{+}21, HWW^{+}17, JYZ^{+}23, JSZ^{+}21, LLL^{+}20, LYS13, MYT^{+}18, MTFK21, MS19, RSSE20, SBGC22, VGP^{+}21, WLO^{+}21, WMN13, WLC^{+}14, WHC16, WLL^{+}22, WX23, XCY^{+}24, YKAE22, ZLXL22, BHB^{+}14]. Time-Bounded [WLO^{+}21]. Time-Delay [BMC17]. Time-Dependent [VGP^{+}21]. Time-Series [HWW^{+}17]. Time-Varying [ZLXL22]. Timing [EGR21, LCM18, MSAS17]. TITAN [BPB^{+}20]. TLC [ZZM^{+}19]. TMR [LYA18]. ToD [HLW14]. Tolerance [BBC^{+}22, GCW20, LYA18, MPK18, NYC^{+}21, RPVWRS^{+}20, TA19]. Tolerant [AMRCP21, CCZN20, GZT^{+}22, GAI22, LWC22, LRG^{+}21, LRHL21, SNK^{+}14, WFB^{+}18, YLZ^{+}21, YXF^{+}22]. Tolerating [NSH22]. Tool [SBL^{+}21]. Toolchain [NAM17]. Toolkit [JMBR^{+}17]. Tools [HH16]. Topics [Ano13a, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15c, Ano15d, Ano15e, Ano15f, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano17a, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17k, Ano17j, Ano18a, Ano18b, Ano18c, Ano18h, Ano18d, Ano18e, Ano18f, Ano18g, Ano19a, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano20a, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano21a, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano22a, Ano22b, Ano22c, Ano22e, Ano22d, Ano22g, Ano22f, Ano23b, Ano23c, Ano23d, Ano23e, Ano24a, CP23, DLS21, DPO17, DO18, FO21, HSS24, KGM15a, KGM15b, TH16, BDFT24]. Topology [NYC^{+}21]. Topper [DDD20]. TopSort [QGF^{+}23]. Torque [WZR^{+}23]. Total [HAKL22]. Tourism [HB15]. Tourists [MCB21]. Toxic [SCS^{+}20]. Track [HLX^{+}17]. Tracking [LYK^{+}22, PLSM20]. Tractable [HHS23]. Trade [CYLJ21, MKH24, PS21]. Trade-Off [MKH24, PS21, CYLJ21]. Trading [AK21, LWM^{+}21]. Traffic [AMG22, SBM21, UBMA18]. Traffic-Aware [AMG22, SBM21]. Train [JKL21]. Training [AFG^{+}21, HÁMLS23, KKM17, LXX^{+}21, MBW23, OAP^{+}22, WJL23, YJ22, ZLW^{+}22, SBL^{+}21]. Trajectory [KR21]. Transaction [WZL21]. Transactions [Ano13a, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15c, Ano15d, Ano15e, Ano15f, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano17a, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17k, Ano17j, Ano18a, Ano18b, Ano18c, Ano18h, Ano18d, Ano18e, Ano18f, Ano18g, Ano19a, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano20a, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano21a, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano22a, Ano22b, Ano22c, Ano22e, Ano22d, Ano22g, Ano22f, Ano23b, Ano23c, Ano23d, Ano23e, Ano24a, BDFT24, CP23, DPO17, DO18, FO21, HSS24, KGM15a, KGM15b]. Transceivers [DZD^{+}18]. Transfer [HZY21, LVJ22, ZYS^{+}22, ZWPL23]. Transferring [TGDC^{+}21]. Transform [WZR^{+}14]. Transformation [DAF^{+}22, HBY^{+}23, HJCK21, LCM^{+}23]. Transformed [ZWPL23]. Transformer [BTD24]. Transformer-Based [BTD24]. Transient [BNCF14]. Transistors [TAC^{+}19]. Transmission [LCY^{+}19, RSK21, YSS^{+}22, YATR18]. Transparent [ZSCY20]. Treatment [LDCY21]. Tree [JHL21, RFKK23]. Trends [AM19, DLTX21, DLRL24, DSS21, JLN21, KSU16, LMS^{+}14, LDLN16, LDLN17, OSPN22]. Triangular [LDJ20]. Triggering [AT22]. Trip [MCB21]. Triple [WT20, YLZ^{+}21]. Trojan [ADN^{+}21, CPM^{+}17, HYW19, LCY20, PASK21, PKR22, VS17, YGLO22]. Trojans [CSZ^{+}20, LRP^{+}22]. True [CRP^{+}23]. Trust [JKSC21, KP14, RGS20, ZLZ^{+}21]. Trusted [LRLG22]. Trustee [JKSC21]. Trustworthiness [WSLL21]. TSCA [WSLL21]. TSV [DBOB20]. TSV-Cluster [DBOB20]. TSVD [PV15]. TSVs [NYC^{+}21]. Tumor [GCBK17]. Tunable [IPRR19]. Tuning [KK14, LFL^{+}18]. Tunnel [BSY^{+}17, CRP^{+}23]. Tunnelling [BMB^{+}21]. Tutorial [LLW^{+}15]. Tutorials [CG17]. TV [FWC15]. Tweets [RMB^{+}19]. Twin [RWD22, SSR^{+}22]. Twins [MMY^{+}22]. Two [ASKG21, LRL21, MMPP15, PJK^{+}19, QGF^{+}23, RSH^{+}21, SWY^{+}22, YCH^{+}23]. Two-Bit [LRL21]. Two-Layer [MMPP15, PJK^{+}19]. Two-Phase [QGF^{+}23]. Two-Stage [SWY^{+}22]. Two-Tier [PJK^{+}19]. Type [TA19]. Types [SWK^{+}17].

UAV [DZK^{+}23, KMK22, LYK^{+}22, LJD^{+}15, TFM^{+}19]. UAV-Aided [KMK22]. UAV-Mounted [TFM^{+}19]. UAVs [OLL^{+}21]. Ubiquitous [CYL^{+}15, GP15a, KGM15a, KGM15b, LCY^{+}19]. ulps [BMP23]. Ultra [eSKZW^{+}22]. Unauthorized [HLM^{+}13]. Unclonable [BR16, GLC^{+}21, KK14, YQ14]. Unconstrained [HAKL22]. Uncovering [BPB^{+}20]. Understanding [DLS21, NVS^{+}14, PRKK24, RHRV16]. Underwater [LPZ^{+}14, XL19]. Unexploitability [Dul20]. Unidimensional [LZH19]. Unified [LCM18]. Unintentional [MDB^{+}23]. Uniqueness [GLC^{+}21]. Unit [CRRS22, DBLK23, JYZ^{+}23]. Univariate [BBM^{+}17]. Universities [LT21]. Unknown [XSYW20]. Unlabeled [LLLG23]. Unlocking [AYL^{+}21, APH^{+}22]. Unsigned [LXL^{+}22b]. Unstructured [GKSJ21]. Unsupervised [STSC24]. Untrustworthy [LRYK14]. Update [HTH^{+}22]. Updates [LLD^{+}18]. Upset [CLA^{+}22, YLZ^{+}21, YXF^{+}22, YCH^{+}23]. Upsets [WT20]. Uptimes [NdCFB^{+}23]. Upwind [BBW22]. Urban [OLL^{+}21]. Usability [DVA21, MRS21, PCG^{+}21].

Usage [LLW^{+}15]. Use

[OKK23b, SBGC22, SSR^{+}22, XYH^{+}23].

Used [VKBB22]. Useful [MKAR22, ZSL^{+}15]. User [ABG^{+}22, HTH^{+}22, MYS17, MZGT17, QLT17, RSH^{+}21, RB22, SWL15, SWW^{+}20, SQX^{+}20, TCNC16, YQZ^{+}15]. User-Centric [MZGT17]. User-Defined [TCNC16]. User-Friendly [SWL15]. User-Habit-Oriented [SWL15]. Users [QWC^{+}18]. Using [ASA^{+}22, ASKG21, AMCMDM^{+}23, BM20, BBE^{+}24, BWMM22, CFM^{+}22, CMJ21, CTC^{+}17, CKR^{+}22, CLW16, DYJ22, DLTSNA21, ESO^{+}22, ET23, EDLT21, FBL^{+}22, GKSJ21, GCBK17, HH16, HTS^{+}23, HYW19, HB15, HLL16, HJCK21, IMZ^{+}21, IPiR18, JYZ^{+}23, KK22, KMD^{+}18, LD21, LNK^{+}15, LYA18, LCM18, MYS17, MKH24, MKAR22, MDB^{+}21, NVS^{+}14, NW20, NYC^{+}21, NS15, OKK22, PL19, PKK22a, PTM23, QZW^{+}23, RM20, RLE^{+}22, SBD^{+}21, SNHN15, STL^{+}14, TDG17, TAT^{+}22, VS17, WCC^{+}20, WWXL22, WLZG22, WX23, WVC21, WZR^{+}23, ZNS^{+}22, ZWWF13, ZXXH13, ZHC^{+}14, ZYZZ15, ZXLS21, ZYC^{+}22, ZLG^{+}22]. Utility [HSG19, MKH24, ZZZY24]. Utilization [SZA^{+}23]. Utilizing [BMC17, TFK20].

V [CRRS22, GTC^{+}21a, GTC^{+}21b,

MMD^{+}22a, MMD^{+}22b]. V2I [WSLL21]. V2V [KMK22]. Validation [RPVWRS^{+}20, SBGC22]. Valuable [XCW^{+}14]. Value [BPB21, LHZS20, SBHL21]. VANETs [ODCZ15, WSLL21]. Variability [NAM17, SKD20]. Variable [Sag19]. Variable-Specific [Sag19]. Variants [APA^{+}21]. Variation [BRRE22, CCW^{+}14, HAKL22, LZH19, PP23b, SCD^{+}21, WLG^{+}21]. Variation-Aware [CCW^{+}14].

Variation-Based [HAKL22]. Variations [ADN^{+}21]. Various [YK18]. Varying [ZLXL22]. Vault [AGMP21]. VBR [CYBD15]. Vector [GCBK17, HLK^{+}24, KMW^{+}21, MKAR22, PIK20]. Vectors [BSJ22]. Vehicle [KMK22, LMW^{+}17, OZL15, OLL^{+}21]. Vehicle-assist [LMW^{+}17]. Vehicle-to-Vehicle [KMK22]. Vehicles [ZLZ^{+}21]. Vehicular [HZS^{+}15, LYS13, OZAL13, SBAR21, SLC^{+}13, ZMK22]. VeMAC [OZAL13]. Venue [AXAD21]. Verifiable [SLB^{+}20, SWW^{+}21]. Verification [AMvO17, FSK20, IP21, KKY24, MIMMY^{+}19, SWLG17]. Verifier [SFL17]. Version [ZWZ18]. Versioning [TFK20]. Versus [DDD^{+}14, RS20]. Vertical [CLY^{+}23, PASK21]. Very [BFPS22]. Via [HWSN13, SZA^{+}23, AEM22, ÁAPDD24, CPH^{+}15, CH21, EGGOR20, LXL^{+}22a, NT16, OPIV^{+}23, Sag19, WLWQ22, XWZ^{+}23, YCL^{+}22, ZWH^{+}24, DMRR17]. VIA-Configurable [DMRR17]. Vias [GHSMM21]. Video [LYK^{+}22, MMPP15, NWSG17, NT16, OMSO20, YSS^{+}22]. Video-Object [NT16]. Videos [MZY^{+}16]. View [NLF^{+}22]. Violation [SSVJ14]. Virtual [BGS^{+}21, FKNK21, GVF^{+}23, MRS21, RS20, SUQKA20, WZH^{+}16, YZZ^{+}21, IEE21, IEE22]. Virtual-Cell-Based [FKNK21]. Virtualization [CMK^{+}16]. Virus [ZWZ18]. Vision [KK21]. Visual [OMSO20]. Visualization [LD21, LDJ20]. Vita [HCCL13]. VLSI [BBM^{+}17, CWZ^{+}15, CLS18, DSS21, MPK18, WX23]. Vocabularies [ADQ21]. Voice [IMZ^{+}21]. Vol [Ano13a, Ano15a, Ano16a, Ano17a, Ano18a, Ano19a, Ano20a, Ano21a, Ano22a]. Volatile [ATPR23, HEYB22, SKD20, TFK20]. Voltage [BAS23, BHB^{+}14, CMM^{+}21, CLW^{+}18, YK18]. VOPRec [KMW^{+}21].

Voting [LRHL21, PGM^{+}21, TD19, VSH23]. VulHunter [XSYW20]. Vulnerabilities [DGW^{+}23, GDP22, RPR^{+}22]. Vulnerability [BHB^{+}14, BPB^{+}20, LSP^{+}20].

WAGE [KSKA22]. Walk [XCW^{+}14]. Walk-Based [XCW^{+}14]. Walking [MRBS22]. Warning [LQT^{+}23, LYS13, PMC^{+}15, WJ19]. Was [VSS18]. Waste [GMTX14, MSLL14]. Water [NVS^{+}14]. Watermarking [KJC22]. Watersheds [SCD^{+}21]. Wave [AMG22, KMD^{+}18, MVC^{+}22]. Wavelength [BYB20]. Wavelength-Shifting [BYB20]. Wavelet [WZR^{+}14]. Way [AFMM19, IP21]. WBSN [BMB^{+}18]. WBSN-Gateway [BMB^{+}18]. Weakly [YMT22]. Wearable [CLWG15, DVPQ^{+}21, MDB^{+}21, YMDJ21]. Wearout [CKKO20]. Web [LLLG23, RHRV16]. Weight [MBW23]. Weighted [CGPB21, CSLG22]. Weights [LK23]. Weird [Dul20]. Welch [KSKA22]. Welch-Gong [KSKA22]. Well [KZT^{+}20, LS21, MYS17, OGL^{+}21, RSH^{+}21]. Well-Being [KZT^{+}20, LS21, OGL^{+}21, RSH^{+}21]. WG [KSKA22]. Who [RMB^{+}19]. Wi [FKNK21]. Wi-Fi [FKNK21]. Wide [VC17]. Wide-Operand [VC17]. Wideband [NBS16]. Wikipedia [SNHN15]. Wikipedia-Based [SNHN15]. Will [LSR^{+}21]. Windows [LD21]. Wire [CBZ^{+}21, KMD^{+}18, LK19]. Wireless [AMvO17, AMG22, BPB21, BCM^{+}21, CYL^{+}15, CYP^{+}16, DFRR23, DTK^{+}23, DZD^{+}18, HYL^{+}20, HLW14, JCM^{+}21, KMM15, LNK^{+}15, LHJ18, LCY^{+}19, NT16, SCS^{+}20, SAI^{+}19, ZWWF13, ZXXH13, ZLZ^{+}17, ZC15, ZSL^{+}15]. WiSE [SKM^{+}23]. Without [MYS17, ABDL19]. Witness [HKZH16]. WORAL [HKZH16]. Word [Pla21a, Pla21b, RMB^{+}19]. Workforce [RSM^{+}21]. Workload [KRG^{+}17, SR14]. Workloads [ZCT^{+}14]. Works [TRGVR^{+}22]. World [CP23, LT21]. Wormhole [SSL^{+}13]. Write [AFMM19, WLG^{+}21]. WSNs [MLLU20, YATR18].

X [CQH17]. XOR [YPS^{+}22]. XpulpNN [GTC^{+}21a, GTC^{+}21b].

Yield [EGR21, SKD20].

Zero [RMVN22, TM14, WJL23]. Zero-Day [WJL23]. Zero-Overhead [TM14]. Zones [Sag19]. Zynq [DZD^{+}18].

References

Anam:2018:GNE

[AAA18] Mohammad Ashraful Anam, Ijeoma Anarado, and Yiannis Andreopoulos. Generalized numerical entanglement for reliable linear, sesquilinear and bijective operations on integer data streams. IEEE Transactions on Emerging Topics in Computing, 6(4):474–487, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Al-Anbagi:2013:RIM

[AAEKM13] Irfan Al-Anbagi, Melike Erol-Kantarci, and Hussein T. Mouftah. A reliable IEEE 802.15.4 model for cyber physical power grid monitoring systems. IEEE Transactions on Emerging Topics in Computing, 1(2): 258–272, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Alavizadeh:2022:ESE

[AAKJJ22] Hooman Alavizadeh, Samin Aref, Dong Seong Kim, and Julian Jang-Jaccard. Evaluating the security and economic effects of moving target defense techniques on the cloud. IEEE Transactions on Emerging Topics in Computing, 10(4):1772– 1788, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Altamimi:2015:ECM

[AANN15] Majid Altamimi, Atef Abdrabou, Kshirasagar Naik, and Amiya Nayak. Energy cost models of smartphones for task offloading to the cloud. IEEE Transactions on Emerging Topics in Computing, 3(3):384– 398, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ahangari:2020:ADP

[AAO^{+}20] Hamzeh Ahangari, Funda Atik, Yusuf Ibrahim Ozkok, Asil Yildirim, Serdar Oguz Ata, and Ozcan Ozturk. Analysis of design parameters in safety-critical computers. IEEE Transactions on Emerging Topics in Computing, 8(3):712–723, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Alvarez-Ayllon:2024:PDM

[ÁAPDD24] Alejandro Álvarez-Ayllón, Manuel Palomo-Duarte, and Juan-Manuel Dodero. PresQ: Discovery of multidimensional equally-distributed dependencies via quasicliques on hypergraphs. IEEE Transactions on Emerging Topics in Computing, 12(1):224–239, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Amelino:2018:ICR

[ABC18] Domenico Amelino, Mario Barbareschi, and Alessandro Cilardo. An IP core remote anonymous activation protocol. IEEE Transactions on Emerging Topics in Computing, 6(2):258–268, April/ June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Alawad:2019:RLS

[ABDL19] Mohammed Alawad, Yu Bai, Ronald F. DeMara, and Mingjie Lin. Robust and large-scale convolution through stochastic-based processing without multipliers. IEEE Transactions on Emerging Topics in Computing, 7(1):80–97, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ali:2022:NTF

[ABG^{+}22] Syed Sadaf Ali, Vivek Singh Baghel, Iyyakutti Iyappan Ganapathi, Surya Prakash, Ngoc-Son Vu, and Naoufel Werghi. A novel technique for fingerprint based secure user authentication. IEEE Transactions on Emerging Topics in Computing, 10(4):1918–1931, October/ December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Azad:2020:PPP

[ABTH20] Muhammad Ajmal Azad, Samiran Bag, Shazia Tabassum, and Feng Hao. privy: Privacy preserving collaboration across multiple service providers to combat telecom spams. IEEE Transactions on Emerging Topics in Computing, 8(2): 313–327, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anghel:2022:DTE

[AC22] Lorena Anghel and Florian Cacho. Design-time exploration for process, environment and aging compensation techniques for low power reliable-aware design. IEEE Transactions on Emerging Topics in Computing, 10(2):581– 590, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Amichi:2023:RIL

[ACCL23] Licia Amichi, Aline Viana Carneiro, Mark Crovella, and Antonio Loureiro. Revealing an inherently limiting factor in human mobility prediction. IEEE Transactions on Emerging Topics in Computing, 11(3):635–649, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ansari:2022:LPA

[ACH22] Mohammad Saeed Ansari, Bruce F. Cockburn, and Jie Han. Low-power approximate logarithmic squaring circuit design for DSP applications. IEEE Transactions on Emerging Topics in Computing, 10(1):500–506, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Allodi:2016:TNM

[ACM16] Luca Allodi, Marco Corradin, and Fabio Massacci. Then and now: On the maturity of the cybercrime markets the lesson that black-hat marketeers learned. IEEE Transactions on Emerging Topics in Computing, 4(1):35–46, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Abdelaty:2022:DDL

[ADCS22] Maged Abdelaty, Roberto Doriguzzi-Corin, and Domenico Siracusa. DAICS: a deep learning solution for anomaly detection in industrial control systems. IEEE Transactions on Emerging Topics in Computing, 10 (2):1117–1129, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Alhelaly:2021:ROB

[ADN^{+}21] Soha Alhelaly, Jennifer Dworak, Kundan Nepal, Theodore Manikas, Ping Gui, and Alfred L. Crouch. 3D ring oscillator based test structures to detect a Trojan die in a 3D die stack in the presence of process variations. IEEE Transactions on Emerging Topics in Computing, 9(2):774–786, April/ June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Albertoni:2021:DCB

[ADQ21] Riccardo Albertoni, Monica De Martino, and Alfonso Quarati. Documenting context-based quality assessment of controlled vocabularies. IEEE Transactions on Emerging Topics in Computing, 9(1):144–160, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Aghaie:2022:PPL

[AEM22] Anita Aghaie, Maik Ender, and Amir Moradi.

PUFs physical learning: Accelerating the enrollment via delay-based model extraction. IEEE Transactions on Emerging Topics in Computing, 10(3):1621– 1632, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Al-Fagih:2013:PPS

[AFATAH13] Ashraf E. Al-Fagih, Fadi M.

Al-Turjman, Waleed M. Alsalih, and Hossam S. Hassanein. A priced public sensing framework for heterogeneous IoT architectures. IEEE Transactions on Emerging Topics in Computing, 1(1):133–147, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ambite:2021:BTC

[AFG^{+}21] José Luis Ambite, Lily Fierro, Jonathan Gordon, Gully A. P. C. Burns, Florian Geigl, Kristina Lerman, and John D. Van Horn. BD2K training coordinating center’s ERuDIte: The educational resource discovery index for data science. IEEE Transactions on Emerging Topics in Computing, 9(1):316–328, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Azad:2019:AAW

[AFMM19] Zahra Azad, Hamed Farbeh, Amir Mahdi Hosseini Monazzah, and Seyed Ghassem Miremadi. AWARE: Adaptive way allocation for reconfigurable ECCs to protect write errors in STT-RAM caches. IEEE Transactions on Emerging Topics in Computing, 7(3):481–492, July/September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ali:2021:PVS

[AGMP21] Syed Sadaf Ali, Iyyakutti Iyappan Ganapathi, Sajid Mahyo, and Surya Prakash. Polynomial vault: a secure and robust fingerprint based authentication. IEEE Transactions on Emerging Topics in Computing, 9(2):612– 625, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

AlBadawi:2021:TAM

[AJL^{+}21] Ahmad Al Badawi, Chao Jin, Jie Lin, Chan Fook Mun, Sim Jun Jie, Benjamin Hong Meng Tan, Xiao Nan, Khin Mi Mi Aung, and Vijay Ramaseshan Chandrasekhar. Towards the AlexNet moment for homomorphic encryption: HCNN, the first homomorphic CNN on encrypted data with GPUs. IEEE Transactions on Emerging Topics in Computing, 9(3): 1330–1343, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Akturk:2021:TCC

[AK21] Ismail Akturk and Ulya R.

Karpuzcu. Trading computation for communication: a taxonomy of data recomputation techniques. IEEE Transactions on Emerging Topics in Computing, 9(1): 496–506, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Adhikary:2024:FRB

[AK24] Supriya Adhikary and Angshuman Karmakar. A fast RLWE-based IPFE library and its application to privacy-preserving biometric authentication. IEEE Transactions on Emerging Topics in Computing, 12(1): 344–356, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Abawajy:2014:LIM

[AKC14] Jemal H. Abawajy, Andrei Kelarev, and Morshed Chowdhury. Large iterative multitier ensemble classifiers for security of Big Data. IEEE Transactions on Emerging Topics in Computing, 2(3):352– 363, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Alcaide:2022:ADR

[AKHA22] Sergi Alcaide, Leonidas Kosmidis, Carles Hernandez, and Jaume Abella. Achieving diverse redundancy for GPU kernels. IEEE Transactions on Emerging Topics in Computing, 10(2):618– 634, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Androutsopoulou:2021:CTE

[AKLC21] Aggeliki S. Androutsopoulou, Nikos I. Karacapilidis, Euripidis N. Loukis, and Yannis K. Charalabidis. Combining technocrats expertise with public opinion through an innovative e-participation platform. IEEE Transactions on Emerging Topics in Computing, 9(1):174–187, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Apolinario:2024:HSC

[AKSR24] Marco Paul E. Apolinario, Adarsh Kumar Kosta, Utkarsh Saxena, and Kaushik Roy. Hardware/software codesign with ADC-less inmemory computing hardware for spiking neural networks. IEEE Transactions on Emerging Topics in Computing, 12(1):35–47, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Aksu:2019:ASN

[AKU19] Hidayet Aksu, Ibrahim Korpeoglu, and Özgür Ulusoy. An analysis of social networks based on tera-scale telecommunication datasets. IEEE Transactions on Emerging Topics in Computing, 7(2):349– 360, April/June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Alawad:2019:SSB

[AL19] Mohammed Alawad and Mingjie Lin. Survey of stochastic-based computation paradigms. IEEE Transactions on Emerging Topics in Computing, 7 (1):98–114, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ahmad:2022:CEC

[ALDK22] Norita Ahmad, Phillip A.

Laplante, Joanna F. De-Franco, and Mohamad Kassab. A cybersecurity educated community. IEEE Transactions on Emerging Topics in Computing, 10(3): 1456–1463, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Aitken:2019:SSE

[AM19] Rob Aitken and Cecilia Metra. Special section on emerging trends and design paradigms for memory systems and storage. IEEE Transactions on Emerging Topics in Computing, 7(3): 433–434, July/September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ahmadinejad:2022:EQE

[AM22] Mohammad Ahmadinejad and Mohammad Hossein Moaiyeri. Energy- and quality-efficient approximate multipliers for neural network and image processing applications. IEEE Transactions on Emerging Topics in Computing, 10 (2):1105–1116, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ayuso-Martinez:2023:CSB

[AMCMDM^{+}23] Alvaro Ayuso-Martinez, Daniel Casanueva-Morato, J. P. Dominguez-Morales, Angel Jimenez-Fernandez, and Gabriel Jimenez-Moreno. Construction of a spikebased memory using neurallike logic gates based on spiking neural networks on SpiNNaker. IEEE Transactions on Emerging Topics in Computing, 11(4):868–881, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ahmed:2022:AOM

[AMG22] M Meraj Ahmed, Naseef Mansoor, and Amlan Ganguly. An asymmetric, oneto-many traffic-aware mmwave wireless interconnection architecture for multichip systems. IEEE Transactions on Emerging Topics in Computing, 10(1):324– 338, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Amirany:2022:NAM

[AMJ22] Abdolah Amirany, Mohammad Hossein Moaiyeri, and Kian Jafari. Nonvolatile associative memory design based on spintronic synapses and CNTFET neurons. IEEE Transactions on Emerging Topics in Computing, 10(1):428–437, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Alves:2021:CAD

[AMKF21] Tiago A. O. Alves, Leandro A. J. Marzulo, Sandip Kundu, and Felipe M. G. França. Concurrency analysis in dynamic dataflow graphs. IEEE Transactions on Emerging Topics in Computing, 9(1):44–54, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Aponte-Moreno:2021:FLA

[AMRCP21] Alexander Aponte-Moreno, Felipe Restrepo-Calle, and Cesar Pedraza. FTxAC:

Leveraging the approximate computing paradigm in the design of fault-tolerant embedded systems to reduce overheads. IEEE Transactions on Emerging Topics in Computing, 9(2):797– 810, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Akhunov:2024:CBA

[AmSYlrm24] Khakim Akhunov and Kas m Sinan Y ld r m. CRAM-Based acceleration for intermittent computing of parallelizable tasks. IEEE Transactions on Emerging Topics in Computing, 12(1):48–59, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Angizi:2024:NSP

[AMTR24] Shaahin Angizi, Mehrdad Morsali, Sepehr Tabrizchi, and Arman Roohi. A near-sensor processing accelerator for approximate local binary pattern networks. IEEE Transactions on Emerging Topics in Computing, 12(1):73–83, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Abdou:2017:LVW

[AMvO17] AbdelRahman Abdou, Ashraf Matrawy, and Paul C. van Oorschot. Location verification of wireless Internet clients: Evaluation and improvements. IEEE Transactions on Emerging Topics in Computing, 5(4):563–575, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2013:IIT

[Ano13a]  Anonymous.  2013 in-

dex IEEE Transactions on Emerging Topics in Computing vol. 1. IEEE Transactions on Emerging Topics in Computing, 1(2):394– 400, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2013:FCa

[Ano13b] Anonymous. [Front cover].

IEEE Transactions on Emerging Topics in Computing, 1(1):1, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2013:FCb

[Ano13c] Anonymous. [Front cover].

IEEE Transactions on Emerging Topics in Computing, 1(2):C1, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2013:TCa

[Ano13d] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 1(1):1–2, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2013:TCb

[Ano13e] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 1(2):201–202, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2014:ITEa

[Ano14a] Anonymous. IEEE Transactions on Emerging Topics in Computing publication information. IEEE Transactions on Emerging Topics in Computing, 2(1): C2, March 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2014:ITEb

[Ano14b] Anonymous. IEEE Transactions on Emerging Topics in Computing publication information. IEEE Transactions on Emerging Topics in Computing, 2(2):C2, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2014:ITEc

[Ano14c] Anonymous. IEEE Transactions on Emerging Topics in Computing publication information. IEEE Transactions on Emerging Topics in Computing, 2(3): C2, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2014:ITEd

[Ano14d] Anonymous. IEEE Transactions on Emerging Topics in Computing publication information. IEEE Transactions on Emerging Topics in Computing, 2(4): C2, December 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2014:FCa

[Ano14e] Anonymous. [Front cover].

IEEE Transactions on Emerging Topics in Computing, 2 (1):C1, March 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2014:FCb

[Ano14f] Anonymous. [Front cover].

IEEE Transactions on Emerging Topics in Computing, 2 (2):C1, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2014:FCc

[Ano14g] Anonymous. [Front cover].

IEEE Transactions on Emerging Topics in Computing, 2(3):C1, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2014:FCd

[Ano14h] Anonymous. [Front cover].

IEEE Transactions on Emerging Topics in Computing, 2(4):C1, December 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2014:TCa

[Ano14i] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 2(1):1, March 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2014:TCb

[Ano14j] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 2(2):94–95, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2014:TCc

[Ano14k] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 2(3):251, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2014:TCd

[Ano14l] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 2(4):398, December 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2015:IIT

[Ano15a]  Anonymous.  2014 in-

dex IEEE Transactions on Emerging Topics in Computing vol. 2. IEEE Transactions on Emerging Topics in Computing, 3(1):

1–9, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2015:RL

[Ano15b] Anonymous. 2014 reviewer list. IEEE Transactions on Emerging Topics in Computing, 3(1):139–145, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2015:ITEa

[Ano15c] Anonymous. IEEE Transactions on Emerging Topics in Computing publication information. IEEE Transactions on Emerging Topics in Computing, 3(1): C2, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2015:ITEb

[Ano15d] Anonymous. IEEE Transactions on Emerging Topics in Computing publication information. IEEE Transactions on Emerging Topics in Computing, 3(2):C2, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2015:ITEc

[Ano15e] Anonymous. IEEE Transactions on Emerging Topics in Computing publication information. IEEE Transactions on Emerging Topics in Computing, 3(3): C2, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2015:ITEd

[Ano15f] Anonymous. IEEE Transactions on Emerging Topics in Computing publication information. IEEE Transactions on Emerging Topics in Computing, 3(4): C2, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2015:FCa

[Ano15g] Anonymous. [Front cover]. IEEE Transactions on Emerging Topics in Computing, 3 (1):C1, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2015:FCb

[Ano15h] Anonymous. [Front cover]. IEEE Transactions on Emerging Topics in Computing, 3 (2):C1, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2015:FCc

[Ano15i] Anonymous. [Front cover]. IEEE Transactions on Emerging Topics in Computing, 3(3):C1, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2015:FCd

[Ano15j] Anonymous. [Front cover]. IEEE Transactions on Emerging Topics in Computing, 3(4):C1, December 2015.

ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2015:TCa

[Ano15k] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 3(1):1–2, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2015:TCb

[Ano15l] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 3(2):146–147, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2015:TCc

[Ano15m]  Anonymous.  Table of

contents. IEEE Transactions on Emerging Topics in Computing, 3(3):303– 304, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2015:TCd

[Ano15n] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 3(4):454–455, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2016:IIT

[Ano16a]  Anonymous.  2015 in-

dex IEEE Transactions on Emerging Topics in Computing vol. 3. IEEE Transactions on Emerging Topics in Computing, 4(1):1–10, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2016:ITEa

[Ano16b] Anonymous. IEEE Transactions on Emerging Topics in Computing publication information. IEEE Transactions on Emerging Topics in Computing, 4(1):C2, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2016:ITEb

[Ano16c] Anonymous. IEEE Transactions on Emerging Topics in Computing publication information. IEEE Transactions on Emerging Topics in Computing, 4(2):C2, April/ June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2016:ITEc

[Ano16d] Anonymous. IEEE Transactions on Emerging Topics in Computing publication information. IEEE Transactions on Emerging Topics in Computing, 4(3):C2, July/September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2016:ITEd

[Ano16e] Anonymous. IEEE Transactions on Emerging Topics in Computing publication information. IEEE Transactions on Emerging Topics in Computing, 4(4):C2, October/December 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2016:FCa

[Ano16f] Anonymous. [Front cover].

IEEE Transactions on Emerging Topics in Computing, 4(1):C1, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2016:FCb

[Ano16g] Anonymous. [Front cover].

IEEE Transactions on Emerging Topics in Computing, 4(2):C1, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2016:FCc

[Ano16h] Anonymous. [Front cover].

IEEE Transactions on Emerging Topics in Computing, 4(3):C1, July/September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2016:FCd

[Ano16i] Anonymous. [Front cover].

IEEE Transactions on Emerging Topics in Computing, 4 (4):C1, October/December 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2016:TCa

[Ano16j] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 4(1):1–2, January/ March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2016:TCb

[Ano16k] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 4(2):175–176, April/ June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2016:TCc

[Ano16l] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 4(3):316–317, July/ September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2016:TCd

[Ano16m] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 4(4):474, October/ December 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:IIT

[Ano17a]  Anonymous.  2016 in-

dex IEEE Transactions on Emerging Topics in Computing vol. 4. IEEE Transactions on Emerging Topics in Computing, 5(1):1–9, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:RLa

[Ano17b] Anonymous. 2016 reviewers list. IEEE Transactions on Emerging Topics in Computing, 5(1):141–145, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:RLb

[Ano17c] Anonymous. 2017 reviewers list. IEEE Transactions on Emerging Topics in Computing, 5(4):603–606, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:ITEa

[Ano17d] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 5(1):C1, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:ITEb

[Ano17e] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 5(1):C2, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:ITEc

[Ano17f] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 5(2):C1, April/ June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:ITEd

[Ano17g] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 5(2):C2, April/ June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:ITEe

[Ano17h] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 5(3):C1, July/September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:ITEf

[Ano17i] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 5(3):C2, July/September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:ITEh

[Ano17j] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 5(4):C2, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:ITEg

[Ano17k] Anonymous. IEEE Transactions on Emerging Topics in Computing [front cover]. IEEE Transactions on Emerging Topics in Computing, 5(4):C1, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:TCa

[Ano17l] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 5(1):1–2, January/ March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:TCb

[Ano17m] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 5(2):146–147, April/ June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:TCc

[Ano17n] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 5(3):300–301, July/ September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2017:TCd

[Ano17o] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 5(4):449–i, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2018:IIT

[Ano18a]  Anonymous.  2017 in-

dex IEEE Transactions on Emerging Topics in Computing vol. 5. IEEE Transactions on Emerging Topics in Computing, 6(1):1–9, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2018:ITEa

[Ano18b] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 6(1):C1, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2018:ITEb

[Ano18c] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 6(1):C2, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2018:ITEd

[Ano18d] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 6(3):C1, July/September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2018:ITEe

[Ano18e] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 6(3):C2, July/September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2018:ITEf

[Ano18f] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 6(4):C1, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2018:ITEg

[Ano18g] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 6(4):C2, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2018:ITEc

[Ano18h] Anonymous. IEEE Transactions on Emerging Topics in Computing [front inside cover]. IEEE Transactions on Emerging Topics in Computing, 6(2):C2, April/ June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2018:FC

[Ano18i] Anonymous. Front cover. IEEE Transactions on Emerging Topics in Computing, 6(2):C1, April/June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2018:TCa

[Ano18j] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 6(1):1–2, January/ March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2018:TCb

[Ano18k] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 6(2):168–169, April/ June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2018:TCc

[Ano18l] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 6(3):301–302, July/ September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2018:TCd

[Ano18m] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 6(4):445–446, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:IIT

[Ano19a]  Anonymous.  2018 in-

dex IEEE Transactions on Emerging Topics in Computing vol. 6. IEEE Transactions on Emerging Topics in Computing, 7(1):1–11, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:RLa

[Ano19b] Anonymous. 2018 reviewers list. IEEE Transactions on Emerging Topics in Computing, 7(2):361–363, April/ June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:RL

[Ano19c] Anonymous. 2019 reviewers list. IEEE Transactions on Emerging Topics in Computing, 7(4):596–600, October/December 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:ITEa

[Ano19d] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 7(1):C1, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:ITEb

[Ano19e] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 7(1):C2, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:ITEc

[Ano19f] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 7(2):C1, April/ June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:ITEd

[Ano19g] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 7(2):C2, April/ June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:ITEe

[Ano19h] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 7(3):C1, July/September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:ITEf

[Ano19i] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 7(3):C2, July/September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:ITEg

[Ano19j] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 7(4):C1, October/December 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:ITEh

[Ano19k] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 7(4):C2, October/December 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:TCa

[Ano19l] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 7(1):1–2, January/ March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:TCb

[Ano19m] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 7(2):202–203, April/ June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:TCc

[Ano19n] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 7(3):364–365, July/ September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2019:TCd

[Ano19o] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 7(4):516–517, October/December 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2020:IIT

[Ano20a]  Anonymous.  2019 in-

dex IEEE Transactions on Emerging Topics in Computing vol. 7. IEEE Transactions on Emerging Topics in Computing, 8(1):1–11, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2020:RL

[Ano20b] Anonymous. 2019 reviewers list. IEEE Transactions on Emerging Topics in Computing, 8(4):998–1000, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2020:ITEa

[Ano20c] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 8(1):C1, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2020:ITEb

[Ano20d] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 8(1):C2, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2020:ITEc

[Ano20e] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 8(2):C1, April/ June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2020:ITEd

[Ano20f] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 8(2):C2, April/ June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2020:ITEe

[Ano20g] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 8(3):C1, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2020:ITEf

[Ano20h] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 8(3):C2, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2020:ITEg

[Ano20i] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 8(4):C1, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2020:ITEh

[Ano20j] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 8(4):C2, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2020:TCa

[Ano20k] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 8(1):1–2, January/ March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2020:TCb

[Ano20l] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 8(2):262–263, April/ June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2020:TCc

[Ano20m] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 8(3):559–560, July/ September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2020:TCd

[Ano20n] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 8(4):883–884, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2021:IIT

[Ano21a]  Anonymous.  2020 in-

dex IEEE Transactions on Emerging Topics in Computing vol. 8. IEEE Transactions on Emerging Topics in Computing, 9(1):1–16, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2021:RL

[Ano21b] Anonymous. 2021 reviewers list. IEEE Transactions on Emerging Topics in Computing, 9(4):2241–2243, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2021:ITEa

[Ano21c] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 9(1):C1, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2021:ITEb

[Ano21d] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 9(1):C1, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2021:ITEc

[Ano21e] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 9(2):C2, April/ June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2021:ITEd

[Ano21f] Anonymous. IEEE Transactions on Emerging Topics in Computing [front cover]. IEEE Transactions on Emerging Topics in Computing, 9(3):C1, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2021:ITEe

[Ano21g] Anonymous. IEEE Transactions on Emerging Topics in Computing [front cover]. IEEE Transactions on Emerging Topics in Computing, 9(3):C2, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2021:FCa

[Ano21h] Anonymous. [Front cover].

IEEE Transactions on Emerging Topics in Computing, 9(2):C1, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2021:FCb

[Ano21i] Anonymous. [Front cover].

IEEE Transactions on Emerging Topics in Computing, 9 (4):C1, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2021:FCc

[Ano21j] Anonymous. [Front cover]. IEEE Transactions on Emerging Topics in Computing, 9 (4):C2, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2021:TCa

[Ano21k] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 9(1):1–3, January/ March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2021:TCb

[Ano21l] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 9(2):i–iii, April/ June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2021:TCc

[Ano21m]  Anonymous.  Table of

contents. IEEE Transactions on Emerging Topics in Computing, 9(3):i–iii, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2021:TCd

[Ano21n] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 9(4):1637–1639, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2022:IIT

[Ano22a]  Anonymous.  2021 in-

dex IEEE Transactions on Emerging Topics in Computing vol. 9. IEEE Transactions on Emerging Topics in Computing, 10(1):1–34, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2022:ITEa

[Ano22b] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 10(1):C1, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2022:ITEb

[Ano22c] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 10(1):C2, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2022:ITEd

[Ano22d] Anonymous. IEEE Transactions on Emerging Topics in Computing. IEEE Transactions on Emerging Topics in Computing, 10(2):

C2, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2022:ITEc

[Ano22e] Anonymous. IEEE Transactions on Emerging Topics in Computing [front cover]. IEEE Transactions on Emerging Topics in Computing, 10(2): C1, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2022:ITEf

[Ano22f] Anonymous. IEEE Transactions on Emerging Topics in Computing information for authors. IEEE Transactions on Emerging Topics in Computing, 10(4):C2, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2022:ITEe

[Ano22g] Anonymous. IEEE Transactions on Emerging Topics in Computing publication information. IEEE Transactions on Emerging Topics in Computing, 10(3):C2, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2022:FCa

[Ano22h] Anonymous. Front cover. IEEE Transactions on Emerging Topics in Computing, 10(3):C1, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2022:FCb

[Ano22i] Anonymous. Front cover.

IEEE Transactions on Emerging Topics in Computing, 10 (4):C1, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2022:TCa

[Ano22j] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 10(1):1–3, January/ March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2022:TCb

[Ano22k] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 10(2):1–3, April/ June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2022:TCc

[Ano22l] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 10(3):1236– 1238, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2022:TCd

[Ano22m] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 10(4):1665– 1667, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2023:RL

[Ano23a] Anonymous. 2022 reviewers list*. IEEE Transactions on Emerging Topics in Computing, 11(1):272–275, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2023:ITEa

[Ano23b] Anonymous. IEEE Transactions on Emerging Topics in Computing information for authors. IEEE Transactions on Emerging Topics in Computing, 11(1):C2, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2023:ITEb

[Ano23c] Anonymous. IEEE Transactions on Emerging Topics in Computing information for authors. IEEE Transactions on Emerging Topics in Computing, 11(2): C2, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2023:ITEc

[Ano23d] Anonymous. IEEE Transactions on Emerging Topics in Computing information for authors. IEEE Transactions on Emerging Topics in Computing, 11(3):C2, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2023:ITEd

[Ano23e] Anonymous. IEEE Transactions on Emerging Topics in Computing information for authors. IEEE Transactions on Emerging Topics in Computing, 11(4):C2, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2023:FCa

[Ano23f] Anonymous. Front cover.

IEEE Transactions on Emerging Topics in Computing, 11(1):C1, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2023:FCb

[Ano23g] Anonymous. Front cover.

IEEE Transactions on Emerging Topics in Computing, 11(2):C1, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2023:FCc

[Ano23h] Anonymous. Front cover.

IEEE Transactions on Emerging Topics in Computing, 11(3):C1, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2023:FCd

[Ano23i] Anonymous. Front cover. IEEE Transactions on Emerging Topics in Computing, 11 (4):C1, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2023:TCa

[Ano23j] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 11(1):1–2, January/ March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2023:TCb

[Ano23k] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 11(2):276– 277, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2023:TCc

[Ano23l] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 11(3):548–549, July/ September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2023:TCd

[Ano23m] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 11(4):822–823, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2024:ITE

[Ano24a] Anonymous. IEEE Transactions on Emerging Topics in Computing information for authors. IEEE Transactions on Emerging Topics in Computing, 12(1):C2, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2024:FC

[Ano24b] Anonymous. Front cover.

IEEE Transactions on Emerging Topics in Computing, 12(1):C1, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Anonymous:2024:TC

[Ano24c] Anonymous. Table of contents. IEEE Transactions on Emerging Topics in Computing, 12(1):1–3, January/ March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

AlBadawi:2021:IPE

[APA^{+}21] Ahmad Al Badawi, Yuriy Polyakov, Khin Mi Mi Aung, Bharadwaj Veeravalli, and Kurt Rohloff. Implementation and performance evaluation of RNS variants of the BFV homomorphic encryption scheme. IEEE Transactions on Emerging Topics in Computing, 9(2):941–956, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Apruzzese:2020:DTP

[APCM20] Giovanni Apruzzese, Fabio Pierazzi, Michele Colajanni, and Mirco Marchetti. Detection and threat prioritization of pivoting attacks in large networks. IEEE Transactions on Emerging Topics in Computing, 8(2):404– 415, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Alrahis:2022:GSM

[APH^{+}22] Lilas Alrahis, Satwik Patnaik, Muhammad Abdullah Hanif, Hani Saleh, Muhammad Shafique, and Ozgur Sinanoglu. GNNUnlock+: a systematic methodology for designing graph neural networks-based oracleless unlocking schemes for provably secure logic locking. IEEE Transactions on Emerging Topics in Computing, 10(3):1575–1592, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Aksanli:2020:HBA

[AR20] Baris Aksanli and Tajana Simunic Rosing. Human behavior aware energy management in residential cyber-physical systems. IEEE Transactions on Emerging Topics in Computing, 8(1):45–57, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ahmed:2020:PMP

[AS20] Tanveer Ahmed and Abhishek Srivastava. A prototype model to predict human interest: Data based design to combine humans and machines. IEEE Transactions on Emerging Topics in Computing, 8(1):31–44, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Arevalo:2024:CMB

[AS24] Irina Arévalo and Jose L.

Salmeron. A chaotic mapsbased privacy-preserving distributed deep learning for incomplete and Non-IID datasets. IEEE Transactions on Emerging Topics in Computing, 12(1):357–367, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Akkan:2022:TDM

[ASA^{+}22] Nihat Akkan, Erzat Safaltin, Levent Aksoy, Ismail Çevik, Herman Sedef, Csaba A. Moritz, and Mustafa Altun. Technology development and modeling of switching lattices using square and H shaped four-terminal switches. IEEE Transactions on Emerging Topics in Computing, 10(1):351–360, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ash-Saki:2021:RDA

[ASKG21] Abdullah Ash-Saki, Mohammad Nasim Imtiaz Khan, and Swaroop Ghosh. Reconfigurable and dense analog circuit design using two terminal resistive memory. IEEE Transactions on Emerging Topics in Computing, 9(3):1596–1608, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ansari:2022:TAS

[ASYK^{+}22] Mohsen Ansari, Sepideh Safari, Sina Yari-Karin, Pourya Gohari-Nazari, Heba Khdr, Muhammad Shafique, Jörg Henkel, and Alireza Ejlali. Thermal-aware standby-sparing technique on heterogeneous real-time embedded systems. IEEE Transactions on Emerging Topics in Computing, 10(4):1883–1897, October/ December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Acheampong:2022:ATA

[AT22] Kingsley Nketia Acheampong and Wenhong Tian. Advancement of textual answer triggering: Cognitive boosting. IEEE Transactions on Emerging Topics in Computing, 10(1):361–372, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ahmadian:2021:ERS

[ATA21] Saba Ahmadian, Farhad Taheri, and Hossein Asadi. Evaluating reliability of SSD-based I/O caches in enterprise storage systems. IEEE Transactions on Emerging Topics in Computing, 9(4):1914–1929, October/ December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Angizi:2023:PNV

[ATPR23] Shaahin Angizi, Sepehr Tabrizchi, David Z. Pan, and Arman Roohi. PISA: a non-volatile processing-in-sensor accelerator for imaging systems. IEEE Transactions on Emerging Topics in Computing, 11(4):962–972, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Asabere:2021:SES

[AXAD21] Nana Yaw Asabere, Bo Xu, Amevi Acakpovi, and Nakema Deonauth. SARVE-2: Exploiting social venue recommendation in the context of smart conferences. IEEE Transactions on Emerging Topics in Computing, 9(1): 342–353, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Alawad:2021:PPD

[AYG^{+}21] Mohammed Alawad, Hong-Jun Yoon, Shang Gao, Brent Mumphrey, Xiao-Cheng Wu, Eric B. Durbin, Jong Cheol Jeong, Isaac Hands, David Rust, Linda Coyle, Lynne Penberthy, and Georgia Tourassi. Privacypreserving deep learning NLP models for cancer registries. IEEE Transactions on Emerging Topics in Computing, 9(3):1219–1230, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Alrahis:2021:SUS

[AYL^{+}21] Lilas Alrahis, Muhammad Yasin, Nimisha Limaye, Hani Saleh, Baker Mohammad, Mahmoud Al-Qutayri, and Ozgur Sinanoglu. ScanSAT: Unlocking static and dynamic scan obfuscation. IEEE Transactions on Emerging Topics in Computing, 9(4):1867–1882, October/ December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Abdulqadder:2021:DRS

[AZA^{+}21] Ihsan H. Abdulqadder, Deqing Zou, Israa T. Aziz, Bin Yuan, and Weiqi Dai. Deployment of robust security scheme in SDN based 5G network over NFV enabled cloud environment. IEEE Transactions on Emerging Topics in Computing, 9(2): 866–877, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bahoo:2023:EEG

[BAS23] Ali Akbar Bahoo, Omid Akbari, and Muhammad Shafique. An energyefficient generic accuracy configurable multiplier based on block-level voltage overscaling. IEEE Transactions on Emerging Topics in Computing, 11(4):851–867, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bouarfa:2016:ABM

[BBC16] Soufiane Bouarfa, Henk A. P. Blom, and Ricky Curran. Agent-based modeling and simulation of coordination by airline operations control. IEEE Transactions on Emerging Topics in Computing, 4(1):9–20, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Biasielli:2022:ABF

[BBC^{+}22] Matteo Biasielli, Cristiana Bolchini, Luca Cassano, Andrea Mazzeo, and Antonio Miele. Approximation-based fault tolerance in image processing applications. IEEE Transactions on Emerging Topics in Computing, 10(2):

648–661, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Brossard:2024:PDC

[BBE^{+}24] Mathias Brossard, Guilhem Bryant, Basma El Gaabouri, Xinxin Fan, Alexandre Ferreira, Edmund Grimley Evans, Christopher Haster, Evan Johnson, Derek Miller, Fan Mo, Dominic P. Mulligan, Nick Spinale, Eric van Hensbergen, Hugo J. M. Vincent, and Shale Xiong. Private delegated computations using strong isolation. IEEE Transactions on Emerging Topics in Computing, 12(1):386–398, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bombieri:2018:PPF

[BBF18] Nicola Bombieri, Federico Busato, and Franco Fummi. Pro++: a profiling framework for primitive-based GPU programming. IEEE Transactions on Emerging Topics in Computing, 6(3): 382–394, July/September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bellizia:2017:UPA

[BBM^{+}17] Davide Bellizia, Simone Bongiovanni, Pietro Monsurrò, Giuseppe Scotti, and Alessandro Trifiletti. Univariate power analysis attacks exploiting static dissipation of nanometer CMOS VLSI circuits for cryptographic applications. IEEE Transactions on Emerging Topics in Computing, 5(3): 329–339, July/September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Barbareschi:2023:CBA

[BBMM23] Mario Barbareschi, Salvatore Barone, Nicola Mazzocca, and Alberto Moriconi. A catalog-based AIG-Rewriting approach to the design of approximate components. IEEE Transactions on Emerging Topics in Computing, 11(1):70–81, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Blewitt:2015:TCS

[BBS^{+}15] William Blewitt, Matthew Brook, Craig Sharp, Gary Ushaw, and Graham Morgan. Toward consistency of state in MMOGs through semantically aware contention management. IEEE Transactions on Emerging Topics in Computing, 3(2):275–288, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bruguier:2016:CED

[BBT^{+}16] Florent Bruguier, Pascal Benoit, Lionel Torres, Lyonel Barthe, Morgan Bourree, and Victor Lomne. Cost-effective design strategies for securing embedded processors. IEEE Transactions on Emerging Topics in Computing, 4(1):60–72, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

BenSalem-Knapp:2022:BRE

[BBW22] Louise Ben Salem-Knapp, Sylvie Boldo, and William Weens. Bounding the roundoff error of the upwind scheme for advection. IEEE Transactions on Emerging Topics in Computing, 10(3): 1253–1262, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Baralis:2016:LSS

[BC16] Elena Baralis and Luca Cagliero. Learning from summaries: Supporting elearning activities by means of document summarization. IEEE Transactions on Emerging Topics in Computing, 4(3):416–428, July/ September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Baralis:2018:HAH

[BC18] Elena Baralis and Luca Cagliero. Highlighter: Automatic highlighting of electronic learning documents. IEEE Transactions on Emerging Topics in Computing, 6(1):7–19, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bernardi:2020:SBS

[BCD^{+}20] Paolo Bernardi, Riccardo Cantoro, Sergio De Luca, Ernesto Sanchez, Alessandro Sansonetti, and Giovanni Squillero. Softwarebased self-test techniques for dual-issue embedded processors. IEEE Transactions on Emerging Topics in Computing, 8(2):464–477, April/ June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Biswas:2021:ATW

[BCM^{+}21] Arnab Kumar Biswas, Navonil Chatterjee, Hemanta Kumar Mondal, Guy Gogniat, and Jean-Philippe Diguet. Attacks toward wireless network-on-chip and countermeasures. IEEE Transactions on Emerging Topics in Computing, 9(2):692– 706, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bhunia:2017:GEI

[BCSF17] Swarup Bhunia, An Chen, Ozgur Sinanoglu, and Jason M. Fung. Guest editors introduction: Security of beyond CMOS devices: Issues and opportunities. IEEE Transactions on Emerging Topics in Computing, 5(3):302–303, July/September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bai:2023:NAN

[BCX23] Yang Bai, Lixing Chen, and Jie Xu. NeuE: Automated neural network ensembles for edge intelligence. IEEE Transactions on Emerging Topics in Computing, 11(2): 485–496, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Brun:2021:SEBa

[BDd^{+}21a] Emeric Brun, David Defour, Pablo de Oliveira Castro, Matei Iştoan, Davide Mancusi, Eric Petit, and Alan Vaquet. A study of the effects and benefits of custom-precision mathematical libraries for HPC codes. IEEE Transactions on Emerging Topics in Computing, 9(3):1467– 1478, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic). See [BDD^{+}21b].

Brun:2021:SEBb

[BDD^{+}21b] Emeric Brun, David Defour, Pablo De Oliveira Castro, Matei Istoan, Davide Mancusi, Eric Petit, and Alan Vaquet. A study of the effects and benefits of custom-precision mathematical libraries for HPC codes. In IEEE [IEE21], page 62. ISBN 1-66542-293-9 (print), 1-66544-648-X (ebook). LCCN ???? See [BDd^{+}21a].

Bosio:2024:GEI

[BDFT24] Alberto Bosio, Ronald F.

DeMara, Deliang Fan, and Nima TaheriNejad. Guest editorial: IEEE Transactions on Emerging Topics in special section on emerging in-memory computing architectures and applications. IEEE Transactions on Emerging Topics in Computing, 12(1):4–6, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bisio:2013:GDE

[BDL^{+}13] Igor Bisio, Alessandro Delfino, Fabio Lavagetto, Mario Marchese, and Andrea Sciarrone. Gender-driven emotion recognition through speech signals for ambient intelligence applications. IEEE Transactions on Emerging Topics in Computing, 1(2):244–257, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bacciu:2024:DGN

[BEG^{+}24] Davide Bacciu, Federico Errica, Alessio Gravina, Lorenzo Madeddu, Marco Podda, and Giovanni Stilo. Deep graph networks for drug repurposing with multiprotein targets. IEEE Transactions on Emerging Topics in Computing, 12(1): 177–189, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bajard:2022:GVL

[BFPS22] Jean Claude Bajard, Kazuhide Fukushima, Thomas Plantard, and Arnaud Sipasseuth. Generating very large RNS bases. IEEE Transactions on Emerging Topics in Computing, 10(3):1289– 1301, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Belmonte:2021:VRS

[BGS^{+}21] Lidia M. Belmonte, Arturo S. García, Eva Segura, Paulo Novais, Rafael Morales, and Antonio Fernández-Caballero. Virtual reality simulation of a quadrotor to monitor dependent people at home. IEEE Transactions on Emerging Topics in Computing, 9(3):1301–1315, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Barenghi:2014:CDT

[BHB^{+}14] Alessandro Barenghi, Cédric Hocquet, David Bol, François-Xavier Standaert, Francesco Regazzoni, and Israel Koren. A combined designtime… test-time study of the vulnerability of subthreshold devices to low voltage fault attacks. IEEE Transactions on Emerging Topics in Computing, 2(2): 107–118, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Balaji:2023:NSI

[BHC^{+}23] Adarsha Balaji, Phu Khanh Huynh, Francky Catthoor, Nikil D. Dutt, Jeffrey L. Krichmar, and Anup Das. NeuSB: a scalable interconnect architecture for spiking neuromorphic hardware. IEEE Transactions on Emerging Topics in Computing, 11(2):373– 387, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bazzaz:2021:HPP

[BHPE21] Mostafa Bazzaz, Ali Hoseinghorban, Farimah Poursafaei, and Alireza Ejlali. High-performance predictable NVM-based instruction memory for real-time embedded systems. IEEE Transactions on Emerging Topics in Computing, 9(1):441–455, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Basilakis:2021:EPB

[BJ21] Jim Basilakis and Bahman Javadi. Efficient parallel binary operations on homomorphic encrypted real numbers. IEEE Transactions on Emerging Topics in Computing, 9(1):507–519, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bhamidipati:2022:HHR

[BK22] Padmaja Bhamidipati and Avinash Karanth. HREN: a hybrid reliable and energyefficient network-on-chip architecture. IEEE Transactions on Emerging Topics in Computing, 10(2):537– 548, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Branitskiy:2021:AML

[BKS21] Alexander Branitskiy, Igor Kotenko, and Igor Saenko. Applying machine learning and parallel data processing for attack detection in IoT. IEEE Transactions on Emerging Topics in Computing, 9(4):1642–1653, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Baharvand:2020:LLE

[BM20] Farshad Baharvand and Seyed Ghassem Miremadi. LEXACT: Low energy Nmodular redundancy using approximate computing for real-time multicore processors. IEEE Transactions on Emerging Topics in Computing, 8(2):431–441, April/ June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bortolotti:2018:EAB

[BMB^{+}18] Daniele Bortolotti, Mauro Mangia, Andrea Bartolini, Riccardo Rovatti, Gianluca Setti, and Luca Benini. Energy-aware bio-signal compressed sensing reconstruction on the WBSN-Gateway. IEEE Transactions on Emerging Topics in Computing, 6(3):370–381, July/ September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bagci:2021:RTD

[BMB^{+}21] Ibrahim Ethem Bagci, Thomas McGrath, Christine Barthelmes, Scott Dean, Ramón Bernardo Gavito, Robert James Young, and Utz Roedig. Resonanttunnelling diodes as PUF building blocks. IEEE Transactions on Emerging Topics in Computing, 9(2): 878–885, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Beckmann:2017:PET

[BMC17] Karsten Beckmann, Harika Manem, and Nathaniel C. Cady. Performance enhancement of a time-delay PUF design by utilizing integrated nanoscale ReRAM devices. IEEE Transactions on Emerging Topics in Computing, 5(3):304–316, July/September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Barrio:2016:CGI

[BMOS16] César Morillas Barrio, Mario Muñoz-Organero, and Joaquín Sánchez Soriano. Can gamification improve the benefits of student response systems in learning? An experimental study. IEEE Transactions on Emerging Topics in Computing, 4(3):429–438, July/September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Brisebarre:2023:EUM

[BMP23] Nicolas Brisebarre, Jean-Michel Muller, and Joris Picot. Error in ulps of the multiplication or division by a correctly-rounded function or constant in binary floating-point arithmetic. IEEE Transactions on Emerging Topics in Computing, ??(??):1– 11, 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bossuet:2014:PBT

[BNCF14] Lilian Bossuet, Xuan Thuy Ngo, Zouha Cherif, and Viktor Fischer. A PUF based on a transient effect ring oscillator and insensitive to locking phenomenon. IEEE Transactions on Emerging Topics in Computing, 2(1): 30–36, March 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Basu:2020:TUP

[BPB^{+}20] Prabal Basu, Pramesh Pandey, Aatreyi Bal, Chidhambaranathan Rajamanikkam, Koushik Chakraborty, and Sanghamitra Roy. TITAN: Uncovering the paradigm shift in security vulnerability at near-threshold computing. IEEE Transactions on Emerging Topics in Computing, 8(4):986–997, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bharti:2021:VIB

[BPB21] Sourabh Bharti, K. K. Pattanaik, and Paolo Bellavista. Value of information based sensor ranking for efficient sensor service allocation in service oriented wireless sensor networks. IEEE Transactions on Emerging Topics in Computing, 9(2): 823–838, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Barve:2018:PDI

[BPBG18] Yogesh D. Barve, Prithviraj Patil, Anirban Bhattacharjee, and Aniruddha Gokhale. PADS: Design and implementation of a cloudbased, immersive learning environment for distributed systems algorithms. IEEE Transactions on Emerging Topics in Computing, 6 (1):20–31, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Berral:2017:AFB

[BPC^{+}17] Josep Lluís Berral, Nicolas Poggi, David Carrera, Aaron Call, Rob Reinauer, and Daron Green. ALOJA: a framework for benchmarking and predictive analytics in Hadoop deployments. IEEE Transactions on Emerging Topics in Computing, 5(4):480–493, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bernardini:2016:TLH

[BR16] Riccardo Bernardini and Roberto Rinaldo. Theoretical limits of helperless stabilizers for physically unclonable constants. IEEE Transactions on Emerging Topics in Computing, 4(1):73–87, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bahrami:2022:PTM

[BRRE22] Fahimeh Bahrami, Behnaz Ranjbar, Nezam Rohbani, and Alireza Ejlali. PVMC: Task mapping and scheduling under process variation heterogeneity in mixedcriticality systems. IEEE Transactions on Emerging Topics in Computing, 10 (2):1166–1177, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Banadaki:2015:SES

[BS15] Yaser Mohammadi Banadaki and Ashok Srivastava. Scaling effects on static metrics and switching attributes of graphene nanoribbon FET for emerging technology. IEEE Transactions on Emerging Topics in Computing, 3(4):458– 469, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Brusilovsky:2016:OSS

[BSG^{+}16] Peter Brusilovsky, Sibel Somyürek, Julio Guerra, Roya Hosseini, Vladimir Zadorozhny, and Paula J. Durlach. Open social student modeling for personalized learning. IEEE Transactions on Emerging Topics in Computing, 4(3):450–461, July/September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Brown:2022:GGR

[BSJ22] Jacob Brown, Tanujay Saha, and Niraj K. Jha. GRAVI-TAS: Graphical reticulated attack vectors for Internetof-Things aggregate security. IEEE Transactions on Emerging Topics in Computing, 10(3):1331–1348, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bi:2017:TFC

[BSY^{+}17] Yu Bi, Kaveh Shamsi, Jiann-Shiun Yuan, Yier Jin, Michael Niemier, and Xiaobo Sharon Hu. Tunnel FET current mode logic for DPA-resilient circuit designs. IEEE Transactions on Emerging Topics in Computing, 5(3):340–352, July/ September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Blanzieri:2021:CSD

[BTC^{+}21] Enrico Blanzieri, Toma Tebaldi, Valter Cavecchia, Francesco Asnicar, Luca Masera, Gabriele Tomè, Eleonora Nigro, Enrica Colasurdo, Matteo Ciciani, Chiara Mazzoni, and Stefania Pilati. A computing system for discovering causal relationships among human genes to improve drug repositioning. IEEE Transactions on Emerging Topics in Computing, 9(4):1667– 1682, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Becattini:2024:TBG

[BTD24] Federico Becattini, Federico Maria Teotini, and Alberto Del Bimbo. Transformerbased graph neural networks for outfit generation. IEEE Transactions on Emerging Topics in Computing, 12(1):213–223, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bart:2020:DEB

[BTK^{+}20] Austin Cory Bart, Javier Tibau, Dennis Kafura, Clifford A. Shaffer, and Eli Tilevich. Design and evaluation of a block-based environment with a data science context. IEEE Transactions on Emerging Topics in Computing, 8(1):182–192, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Barchi:2021:FLR

[BUS^{+}21] Francesco Barchi, Gianvito Urgese, Alessandro Siino, Santa Di Cataldo, Enrico Macii, and Andrea Acquaviva. Flexible on-line reconfiguration of multi-core neuromorphic platforms. IEEE Transactions on Emerging Topics in Computing, 9(2): 915–927, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Burek:2022:AAB

[BWMM22] Elz.bieta Burek, Micha Wroński, Krzysztof Mańk, and Michał Misztal. Algebraic attacks on block ciphers using quantum annealing. IEEE Transactions on Emerging Topics in Computing, 10(2):678– 689, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bu:2013:GTS

[BY13] Shengrong Bu and F. Richard Yu. A game-theoretical scheme in the Smart Grid with demand-side management: Towards a smart cyber-physical power infrastructure. IEEE Transactions on Emerging Topics in Computing, 1(1):22–32, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

BenAhmed:2020:SPN

[BYB20] Achraf Ben Ahmed, Tsutomu Yoshinaga, and Abderazek Ben Abdallah. Scalable photonic networks-on-chip architecture based on a novel wavelength-shifting mechanism. IEEE Transactions on Emerging Topics in Computing, 8(2):533– 544, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Bi:2022:EMP

[BYZZ22] Jing Bi, Haitao Yuan, Kaiyi Zhang, and MengChu Zhou. Energy-minimized partial computation offloading for delay-sensitive applications in heterogeneous edge networks. IEEE Transactions on Emerging Topics in Computing, 10(4):1941– 1954, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cazzaniga:2021:FTN

[CBG^{+}21] Carlo Cazzaniga, Marta Bagatin, Simone Gerardin, Alessandra Costantino, and Christopher D. Frost. First tests of a new facility for device-level, board-level and system-level neutron irradiation of microelectronics. IEEE Transactions on Emerging Topics in Computing, 9(1):104–108, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Crepaldi:2021:LWL

[CBZ^{+}21] Marco Crepaldi, Alessandro Barcellona, Giorgio Zini, Alberto Ansaldo, Paolo Motto Ros, Alessandro Sanginario, Claudia Cuccu, Danilo Demarchi, and Luca Brayda. Live wire a low-complexity body channel communication system for landmark identification. IEEE Transactions on Emerging Topics in Computing, 9(3):1248– 1264, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cardone:2016:PLS

[CCFI16] Giuseppe Cardone, Antonio Corradi, Luca Foschini, and Raffaele Ianniello. ParticipAct: a large-scale crowdsensing platform. IEEE Transactions on Emerging Topics in Computing, 4 (1):21–32, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chang:2023:FES

[CCH^{+}23] Yu-Wei Chang, Hong-Yen Chen, Chansu Han, Tomohiro Morikawa, Takeshi Takahashi, and Tsung-Nan Lin. FINISH: Efficient and scalable NMF-Based federated learning for detecting malware activities. IEEE Transactions on Emerging Topics in Computing, 11(4):934–949, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Colonnelli:2021:SCB

[CCMA21] Iacopo Colonnelli, Barbara Cantalupo, Ivan Merelli, and Marco Aldinucci. Stream-Flow: Cross-breeding cloud with HPC. IEEE Transactions on Emerging Topics in Computing, 9(4):1723– 1737, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2014:VAL

[CCW^{+}14] Xiaodao Chen, Dan Chen, Lizhe Wang, Ze Deng, Rajiv Ranjan, Albert Y. Zomaya, and Shiyan Hu. Variationaware layer assignment with hierarchical stochastic optimization on a multicore platform. IEEE Transactions on Emerging Topics in Computing, 2(4):488– 500, December 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Charif:2020:DSC

[CCZN20] Amir Charif, Alexandre Coelho, Nacer-Eddine Zergainoh, and Michael Nicolaidis. A dynamic sufficient condition of deadlockfreedom for high-performance fault-tolerant routing in networks-on-chips. IEEE Transactions on Emerging Topics in Computing, 8(3): 642–654, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cui:2021:NPB

[CCZZ21] Aijiao Cui, Chip-Hong Chang, Wei Zhou, and Yue Zheng. A new PUF based lock and key solution for secure in-field testing of cryptographic chips. IEEE Transactions on Emerging Topics in Computing, 9 (2):1095–1105, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Conficconi:2023:EED

[CDC^{+}23] Davide Conficconi, Emanuele Del Sozzo, Filippo Carloni, Alessandro Comodi, Alberto Scolari, and Marco Domenico Santambrogio. An energyefficient domain-specific architecture for regular expressions. IEEE Transactions on Emerging Topics in Computing, 11(1):3–17, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cardarilli:2022:DSE

[CDF^{+}22a] Gian Carlo Cardarilli, Luca Di Nunzio, Rocco Fazzolari, Alberto Nannarelli, Massimo Petricca, and Marco Re. Design space exploration based methodology for residue number system digital filters implementation. IEEE Transactions on Emerging Topics in Computing, 10(1):186–198, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Corradi:2022:SSI

[CDF^{+}22b] Antonio Corradi, Giuseppe Di Modica, Luca Foschini, Lorenzo Patera, and Michele Solimando. SIRDAM4.0: a support infrastructure for reliable data acquisition and management in Industry 4.0. IEEE Transactions on Emerging Topics in Computing, 10(3):1605– 1620, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2013:PGR

[CDLS13] Xiao Chen, Zanxun Dai, Wenzhong Li, and Hongchi Shi. Performance guaranteed routing protocols for asymmetric sensor networks. IEEE Transactions on Emerging Topics in Computing, 1(1):111–120, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Conti:2020:NSA

[CDM20] Mauro Conti, Fabio De Gaspari, and Luigi V. Mancini. A novel stealthy attack to gather SDN configurationinformation. IEEE Transactions on Emerging Topics in Computing, 8(2):328– 340, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2019:PEC

[CF19] Kuan-Jung Chen and Shao-Yun Fang. Printability enhancement with color balancing for multiple patterning lithography. IEEE Transactions on Emerging Topics in Computing, 7(2): 244–252, April/June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2021:GSA

[CFL^{+}21] Ting Chen, Youzheng Feng, Zihao Li, Hao Zhou, Xiaopu Luo, Xiaoqi Li, Xiuzhuo Xiao, Jiachi Chen, and Xiaosong Zhang. GasChecker:

Scalable analysis for discovering gas-inefficient smart contracts. IEEE Transactions on Emerging Topics in Computing, 9(3):1433–1448, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chalmers:2022:DAD

[CFM^{+}22] C. Chalmers, P. Fergus, C. Aday Curbelo Montanez, S. Sikdar, F. Ball, and B. Kendall. Detecting activities of daily living and routine behaviours in dementia patients living alone using smart meter load disaggregation. IEEE Transactions on Emerging Topics in Computing, 10(1):157–169, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Copeland:2017:TEH

[CG17] Leana D. Copeland and Tom D. Gedeon. Tutorials in eLearning how presentation affects outcomes. IEEE Transactions on Emerging Topics in Computing, 5 (1):20–31, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cui:2024:SOM

[CGC^{+}24] Jia-Le Cui, Yanan Guo, Juntong Chen, Bo Liu, and Hao Cai. Sparsity-oriented MRAM-centric computing for efficient neural network inference. IEEE Transactions on Emerging Topics in Computing, 12(1):97–108, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cagliero:2021:ARA

[CGPB21] Luca Cagliero, Paolo Garza, Andrea Pasini, and Elena Baralis. Additional reviewer assignment by means of weighted association rules. IEEE Transactions on Emerging Topics in Computing, 9(1):329–341, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chang:2016:REP

[CH16] Yeim-Kuan Chang and Chun-Sheng Hsueh. Rangeenhanced packet classification design on FPGA. IEEE Transactions on Emerging Topics in Computing, 4(2): 214–224, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2019:EAS

[CH19] Te-Hsuan Chen and John P.

Hayes. Equivalence among stochastic logic circuits and its application to synthesis. IEEE Transactions on Emerging Topics in Computing, 7(1):67–79, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chu:2021:OSC

[CH21] Wei Chu and Shi-Yu Huang. Online safety checking for delay locked loops via embedded phase error monitor. IEEE Transactions on Emerging Topics in Computing, 9(2):735–744, April/ June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2023:LJF

[CH23] Shuo-Han Chen and Kuo-Hao Huang. Leveraging journaling file system for prompt secure deletion on interlaced recording drives. IEEE Transactions on Emerging Topics in Computing, 11(3):619–634, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chetto:2014:OSR

[Che14] Maryline Chetto. Optimal scheduling for real-time jobs in energy harvesting computing systems. IEEE Transactions on Emerging Topics in Computing, 2(2): 122–133, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cozzi:2018:ORL

[CKC^{+}18] Dario Cozzi, Sebastian Korf, Luca Cassano, Jens Hagemeyer, Andrea Domenici, Cinzia Bernardeschi, Luca Sterpone, and Mario Porrmann. OLT(RE)2: an online on-demand testing approach for permanent radiation effects in reconfigurable systems. IEEE Transactions on Emerging Topics in Computing, 6(4):511–523, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chang:2020:FRR

[CKKO20] Doohwang Chang, Jennifer N. Kitchen, Sayfe Kiaei, and Sule Ozev. In-field recovery of RF circuits from wearout based performance degradation. IEEE Transactions on Emerging Topics in Computing, 8(2):442– 452, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chowdhury:2022:CSA

[CKR^{+}22] Zamshed I. Chowdhury, S. Karen Khatamifard, Salonik Resch, Hüsrev Cılasun, Zhengyang Zhao, Masoud Zabihi, Meisam Razaviyayn, Jian-Ping Wang, Sachin S. Sapatnekar, and Ulya R. Karpuzcu. CRAM-Seq: Accelerating RNA-Seq abundance quantification using computational RAM. IEEE Transactions on Emerging Topics in Computing, 10(4):2055–2071, October/ December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2022:SPE

[CLA^{+}22] Junchao Chen, Thomas Lange, Marko Andjelkovic, Aleksandar Simevski, Li Lu, and Milos Krstic. Solar particle event and single event upset prediction from SRAM-based monitor and supervised machine learning. IEEE Transactions on Emerging Topics in Computing, 10(2):564– 580, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2024:GGP

[CLD^{+}24] Yiming Chen, Mingyen Lee, Guohao Dai, Mufeng Zhou, Nagadastagiri Challapalle, Tianyi Wang, Yao Yu, Yongpan Liu, Yu Wang, Huazhong Yang, Vijaykrishnan Narayanan, and Xueqing Li. GRAPHIC: Gather and process harmoniously in the cache with high parallelism and flexibility. IEEE Transactions on Emerging Topics in Computing, 12 (1):84–96, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2021:LPP

[CLL21] Jingxue Chen, Gao Liu, and Yining Liu. Lightweight privacy-preserving raw data publishing scheme. IEEE Transactions on Emerging Topics in Computing, 9(4):2170–2174, October/ December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2023:TAC

[CLLL23] Yuechen Chen, Shanshan Liu, Fabrizio Lombardi, and Ahmed Louri. A technique for approximate communication in network-on-chips for image classification. IEEE Transactions on Emerging Topics in Computing, 11(1):30–42, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chien:2018:NVC

[CLS18] Chen-Han Chien, Shih-Chii Liu, and Andreas Steimer. A neuromorphic VLSI circuit for spike-based random sampling. IEEE Transactions on Emerging Topics in Computing, 6(1):135–144, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chu:2016:IAL

[CLW16] Hung-Mao Chu, Tsung-Hsien Li, and Pi-Chung Wang. IP address lookup by using GPU. IEEE Transactions on Emerging Topics in Computing, 4(2):187– 198, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cui:2018:EAG

[CLW^{+}18] Tiansong Cui, Ji Li, Yanzhi Wang, Shahin Nazarian, and Massoud Pedram. An exploration of applying gate-length-biasing techniques to deeply-scaled Fin-FETs operating in multiple voltage regimes. IEEE Transactions on Emerging Topics in Computing, 6(2): 172–183, April/June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cheng:2015:JTC

[CLWG15] Zixue Cheng, Peng Li, Junbo Wang, and Song Guo. Just-in-time code offloading for wearable computing. IEEE Transactions on Emerging Topics in Computing, 3(1):74–83, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2018:AAE

[CLWX18] Honglong Chen, Wei Lou, Zhibo Wang, and Feng Xia. On achieving asynchronous energy-efficient neighbor discovery for mobile sensor networks. IEEE Transactions on Emerging Topics in Computing, 6(4):553–565, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cai:2023:EVF

[CLY^{+}23] Jianping Cai, Ximeng Liu, Zhiyong Yu, Kun Guo, and Jiayin Li. Efficient vertical federated learning method for ridge regression of largescale samples. IEEE Transactions on Emerging Topics in Computing, 11(2):511– 526, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Capotondi:2018:RSM

[CMB18] Alessandro Capotondi, Andrea Marongiu, and Luca Benini. Runtime support for multiple offload-based programming models on clustered manycore accelerators. IEEE Transactions on Emerging Topics in Computing, 6(3):330–342, July/ September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chandrakala:2021:BMB

[CMJ21] S. Chandrakala, S. Malini, and S. L. Jayalakshmi. Bag of models based embeddings for assessment of neurological disorders using speech intelligibility. IEEE Transactions on Emerging Topics in Computing, 9(3):1265– 1275, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2016:SDN

[CMK^{+}16] Jiann-Liang Chen, Yi-Wei Ma, Hung-Yi Kuo, Chu-Sing Yang, and Wen-Chien Hung. Software-defined network virtualization platform for enterprise network resource management. IEEE Transactions on Emerging Topics in Computing, 4(2): 179–186, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cheng:2021:ASP

[CMM^{+}21] Wu-Tung Cheng, Sylwester Milewski, Grzegorz Mrugalski, Janusz Rajski, Maciej Trawka, and Jerzy Tyszer. Autonomous scan patterns for laser voltage imaging. IEEE Transactions on Emerging Topics in Computing, 9(2):680–691, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2020:HPD

[CMMF20] Yong Chen, Emil Matus, Sadia Moriam, and Gerhard P. Fettweis. High performance dynamic resource allocation for guaranteed service in networkon-chips. IEEE Transactions on Emerging Topics in Computing, 8(2):503–516, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Carrabina:2017:ICG

[CMPT17] Jordi Carrabina, Mohammad Mashayekhi, Jofre Pallarès, and Lluis Terés. Inkjet-configurable gate arrays (IGA). IEEE Transactions on Emerging Topics in Computing, 5(2):238– 246, April/June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Conti:2021:NBI

[CMRV21] Vincenzo Conti, Carmelo Militello, Leonardo Rundo, and Salvatore Vitabile. A novel bio-inspired approach for high-performance management in service-oriented networks. IEEE Transactions on Emerging Topics in Computing, 9(4):1709– 1722, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cazzola:2016:GLP

[CO16] Walter Cazzola and Diego Mathias Olivares. Gradually learning programming supported by a growable programming language. IEEE Transactions on Emerging Topics in Computing, 4(3): 404–415, July/September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chang:2023:GEI

[CP23] Yuan-Hao Chang and Vincenzo Piuri. Guest editorial: IEEE Transactions on Emerging Topics in Computing thematic section on memory- centric designs: Processing-in-memory, in-memory computing, and near-memory computing for real-world applications. IEEE Transactions on Emerging Topics in Computing, 11(2):278– 280, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2015:PDS

[CPH^{+}15] Shaoming Chen, Lu Peng, Yue Hu, Zhou Zhao, Ashok Srivastava, Ying Zhang, Jin-Woo Choi, Bin Li, and Edward Song. Powering up dark silicon: Mitigating the limitation of power delivery via dynamic pin switching. IEEE Transactions on Emerging Topics in Computing, 3(4):489– 501, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Camara:2021:ACI

[CPLdFM21] Carmen Camara, Pedro Peris-Lopez, Jose Maria de Fuentes, and Samuel Marchal. Access control for implantable medical devices. IEEE Transactions on Emerging Topics in Computing, 9(3):1126–1138, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chakraborty:2017:FOC

[CPM^{+}17] Rajat Subhra Chakraborty, Samuel Pagliarini, Jimson Mathew, Sree Ranjani Rajendran, and M. Nirmala Devi. A flexible online checking technique to enhance hardware Trojan horse detectability by reliability analysis. IEEE Transactions on Emerging Topics in Computing, 5(2):260– 270, April/June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2017:IXB

[CQH17] Pengfei Chen, Yong Qi, and Di Hou. InvarNet-X: a black-box invariant-based approach to diagnosing Big Data systems. IEEE Transactions on Emerging Topics in Computing, 5(4):450–465, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2019:CSB

[CRi^{+}19] Bo-Wei Chen, Seungmin Rho, Muhammad imran, Mohsen Guizani, and Wei-Kang Fan. Cognitive sensors based on ridge phasesmoothing localization and multiregional histograms of oriented gradients. IEEE Transactions on Emerging Topics in Computing, 7(1): 123–134, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chatterjee:2023:FFT

[CRP^{+}23] Swetaki Chatterjee, Nikhil Rangarajan, Satwik Patnaik, Dinesh Rajasekharan, Ozgur Sinanoglu, and Yogesh Singh Chauhan. FerroCoin: Ferroelectric tunnel junction-based true random number generator. IEEE Transactions on Emerging Topics in Computing, 11(2): 541–547, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cococcioni:2022:LPP

[CRRS22] Marco Cococcioni, Federico Rossi, Emanuele Ruffaldi, and Sergio Saponara. A lightweight posit processing unit for RISC-V processors in deep neural network applications. IEEE Transactions on Emerging Topics in Computing, 10(4):1898– 1908, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cousins:2017:DFA

[CRS17] David Bruce Cousins, Kurt Rohloff, and Daniel Sumorok. Designing an FPGA-accelerated homomorphic encryption co-processor. IEEE Transactions on Emerging Topics in Computing, 5(2):193– 206, April/June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Callaghan:2017:MLG

[CSBME17] Michael Callaghan, Maggi Savin-Baden, Niall Mc-Shane, and Augusto Gómez Eguíluz. Mapping learning and game mechanics for serious games analysis in engineering education. IEEE Transactions on Emerging Topics in Computing, 5(1):77–83, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Canto:2023:EDS

[CSK^{+}23] Alvaro Cintas Canto, Ausmita Sarker, Jasmin Kaur, Mehran Mozaffari Kermani, and Reza Azarderakhsh. Error detection schemes assessed on FPGA for multipliers in lattice-based key encapsulation mechanisms in post-quantum cryptography. IEEE Transactions on Emerging Topics in Computing, 11(3):791–797, July/ September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chai:2022:DDW

[CSLG22] Xiaolin Chai, Yan Sun, Hong Luo, and Mohsen Guizani. DWES: a dynamic weighted evaluation system for scratch based on computational thinking. IEEE Transactions on Emerging Topics in Computing, 10(2): 917–932, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chamazcoti:2019:CEC

[CSM19] Saeideh Alinezhad Chamazcoti, Bardia Safaei, and Seyed Ghassem Miremadi.

Can erasure codes damage reliability in SSD-based storage systems? IEEE Transactions on Emerging Topics in Computing, 7(3): 435–446, July/September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cui:2020:DIT

[CSZ^{+}20] Xiaotong Cui, Samah Mohamed Saeed, Alwin Zulehner, Robert Wille, Kaijie Wu, Rolf Drechsler, and Ramesh Karri. On the difficulty of inserting Trojans in reversible computing architectures. IEEE Transactions on Emerging Topics in Computing, 8(4):960–972, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chou:2017:OSM

[CTC^{+}17] Chih-Yueh Chou, Shu-Fen Tseng, Wen-Chieh Chih, Zhi-Hong Chen, Po-Yao Chao, K. Robert Lai, Chien-Lung Chan, Liang-Chih Yu, and Yi-Lung Lin. Open student models of core competencies at the curriculum level: Using learning analytics for student reflection. IEEE Transactions on Emerging Topics in Computing, 5(1):32–44, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2023:RAK

[CTL^{+}23] Zheyi Chen, Pu Tian, Weixian Liao, Xuhui Chen, Guobin Xu, and Wei Yu. Resource-aware knowledge distillation for federated learning. IEEE Transactions on Emerging Topics in Computing, 11(3):706–719, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cannizzaro:2022:SDD

[CVP^{+}22] Davide Cannizzaro, Antonio Giuseppe Varrella, Stefano Paradiso, Roberta Sampieri, Yukai Chen, Alberto Macii, Edoardo Patti, and Santa Di Cataldo. Insitu defect detection of metal additive manufacturing: an integrated framework. IEEE Transactions on Emerging Topics in Computing, 10(1):74–86, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2021:LSP

[CWK^{+}21] Biwen Chen, Libing Wu, Neeraj Kumar, Kim-Kwang Raymond Choo, and Debiao He. Lightweight searchable public-key encryption with forward privacy over IIoT outsourced data. IEEE Transactions on Emerging Topics in Computing, 9(4):1753–1764, October/ December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2015:CCV

[CWZ^{+}15] Xiaodao Chen, Lizhe Wang, Albert Y. Zomaya, Lin Liu, and Shiyan Hu. Cloud computing for VLSI floorplanning considering peak temperature reduction. IEEE Transactions on Emerging Topics in Computing, 3(4): 534–543, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Cui:2022:SSA

[CWZ^{+}22] Aijiao Cui, Zhen Weng, Hui Zhang, Gang Qu, and Huawei Li. SATAM: a SAT Attack Resistant Active Metering against IC overbuilding. IEEE Transactions on Emerging Topics in Computing, 10(4):2025– 2041, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chang:2015:PVS

[CYBD15] Che-Yu Chang, Hsu-Chun Yen, Abderrahim Benslimane, and Der-Jiunn Deng. A pragmatic VBR stream scheduling policy for IEEE 802.11e HCCA access method. IEEE Transactions on Emerging Topics in Computing, 3(4):514–523, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2015:UMI

[CYL^{+}15] Cailian Chen, Jing Yan, Ning Lu, Yiyin Wang, Xian Yang, and Xinping Guan. Ubiquitous monitoring for industrial cyberphysical systems over relayassisted wireless sensor networks. IEEE Transactions on Emerging Topics in Computing, 3(3):352– 362, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chen:2021:TBH

[CYLJ21] Pai Chen, Jianhui Yue, Xiaofei Liao, and Hai Jin. Trade-off between hit rate and hit latency for optimizing DRAM cache. IEEE Transactions on Emerging Topics in Computing, 9 (1):55–64, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Chuang:2016:EMD

[CYP^{+}16] Ching-Chih Chuang, Ya-Ju Yu, Ai-Chun Pang, Hsueh-Wen Tseng, and Hsin-Peng Lin. Efficient multicast delivery for data redundancy minimization over wireless data centers. IEEE Transactions on Emerging Topics in Computing, 4(2):225– 241, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Dong:2020:GES

[DA20] Mianxiong Dong and Nirwan Ansari. Guest editorial: Special section on cyber-physical social systems integrating human into computing. IEEE Transactions on Emerging Topics in Computing, 8(1):4–5, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

DallOra:2022:DTP

[DAF^{+}22] Nicola Dall’Ora, Khaled Alamin, Enrico Fraccaroli, Massimo Poncino, Davide Quaglia, and Sara Vinco. Digital transformation of a production line: Network design, online data collection and energy monitoring. IEEE Transactions on Emerging Topics in Computing, 10(1):46–59, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

DeSio:2022:FNN

[DAS22] Corrado De Sio, Sarah Azimi, and Luca Sterpone. FireNN: Neural networks reliability evaluation on hybrid platforms. IEEE Transactions on Emerging Topics in Computing, 10(2):549– 563, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Darbani:2023:RRI

[DBLK23] Paria Darbani, Hakem Beitollahi, and Pejman Lotfi-Kamran. Rei: a reconfigurable interconnection unit for array-based CNN accelerators. IEEE Transactions on Emerging Topics in Computing, 11(4):895–906, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Duong-Ba:2014:DCS

[DBNBT14] Thuan Duong-Ba, Thinh Nguyen, Bella Bose, and Duc A. Tran. Distributed client-server assignment for online social network applications. IEEE Transactions on Emerging Topics in Computing, 2(4):422– 435, December 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Dang:2020:SDM

[DBOB20] Khanh N. Dang, Akram Ben Ahmed, Yuichi Okuyama, and Abderazek Ben Abdallah. Scalable design methodology and online algorithm for TSV-cluster defects recovery in highly reliable 3D-NoC systems. IEEE Transactions on Emerging Topics in Computing, 8(3):577–590, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Dang:2022:MMB

[DDB22] Khanh N. Dang, Nguyen Anh Vu Doan, and Abderazek Ben Abdallah. MigSpike: a migration based algorithms and architecture for scalable robust neuromorphic systems. IEEE Transactions on Emerging Topics in Computing, 10(2):602– 617, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

DaRolt:2014:TVS

[DDD^{+}14] Jean Da Rolt, Amitabh Das, Giorgio Di Natale, Marie-Lise Flottes, Bruno Rouzeyre, and Ingrid Verbauwhede. Test versus security: Past and present. IEEE Transactions on Emerging Topics in Computing, 2(1):50– 62, March 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Das:2020:NCT

[DDD20] Pranesh Das, Dushmanta Kumar Das, and Shouvik Dey. A new class topper optimization algorithm with an application to data clustering. IEEE Transactions on Emerging Topics in Computing, 8(4):948–959, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Delgado:2022:OEA

[DF22] Carmen Delgado and Jeroen Famaey. Optimal energyaware task scheduling for batteryless IoT devices. IEEE Transactions on Emerging Topics in Computing, 10(3):1374–1387, July/ September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Dehghani:2023:DAE

[DFRR23] Abbas Dehghani, Sadegh Fadaei, Bahman Ravaei, and Keyvan RahimiZadeh. Deadline-aware and energyefficient dynamic task mapping and scheduling for multicore systems based on wireless network-on-chip. IEEE Transactions on Emerging Topics in Computing, 11(4):1031–1044, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Drozdowicz:2022:SAC

[DGP22] Micha Drozdowicz, Maria Ganzha, and Marcin Paprzycki. Semantic access control for privacy management of personal sensing in smart cities. IEEE Transactions on Emerging Topics in Computing, 10(1):199–210, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Dou:2023:SAV

[DGW^{+}23] Yuqin Dou, Chongyan Gu, Chenghua Wang, Weiqiang Liu, and Fabrizio Lombardi. Security and approximation: Vulnerabilities in approximation-aware testing. IEEE Transactions on Emerging Topics in Computing, 11(1):265–271, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Dabbagh:2018:ETE

[DHGR18] Mehiar Dabbagh, Bechir Hamdaoui, Mohsen Guizani, and Ammar Rayes. Exploiting task elasticity and price heterogeneity for maximizing cloud computing profits. IEEE Transactions on Emerging Topics in Computing, 6(1):85–96, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Deng:2019:GEI

[DHPL19] Ruilong Deng, Zhu Han, Sangheon Pack, and Hao Liang. Guest editorial: Introduction to special section on Big Data computing for the Smart Grid. IEEE Transactions on Emerging Topics in Computing, 7(3): 366–368, July/September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

DAmelio:2024:GEE

[DLRL24] Alessandro D’Amelio, Jianyi Lin, Jean-Yves Ramel, and Raffaella Lanzarotti. Guest editorial: Emerging trends and advances in graphbased methods and applications. IEEE Transactions on Emerging Topics in Computing, 12(1):122–125, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Datta:2021:UID

[DLS21] Subhajit Datta, Rumana Lakdawala, and Santonu Sarkar. Understanding the inter-domain presence of research topics in the computing discipline. IEEE Transactions on Emerging Topics in Computing, 9(1):366–378, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Delgado-Lozano:2021:DAS

[DLTSNA21] Ignacio M. Delgado-Lozano, Erica Tena-Sánchez, Juan Núñez, and Antonio J. Acosta. Design and analysis of secure emerging cryptohardware using HyperFET devices. IEEE Transactions on Emerging Topics in Computing, 9(2):787–796, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

DAgostino:2021:GES

[DLTX21] Daniele D’Agostino, Francesco Leporati, Massimo Torquati, and Jingling Xue. Guest editorial: Special section on new trends in parallel and distributed computing for human sensible applications. IEEE Transactions on Emerging Topics in Computing, 9(4):1640–1641, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

DalBem:2017:SBF

[DMRR17] Vinícius Dal Bem, Felipe S. Marranghello, André I. Reis, and Renato P. Ribas. SATbased formulation for logical capacity evaluation of VIA-configurable structured ASIC. IEEE Transactions on Emerging Topics in Computing, 5(2):247–259, April/June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

DiNatale:2018:GEI

[DO18] Giorgio Di Natale and Marco Ottavi. Guest editorial: IEEE Transactions on Emerging Topics in Computing special issue on design & technology of integrated systems in deep submicron era. IEEE Transactions on Emerging Topics in Computing, 6(2):170– 171, April/June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

DeMara:2017:GEI

[DPO17] Ronald F. DeMara, Marco Platzner, and Marco Ottavi. Guest editorial: IEEE Transactions on Computers and IEEE Transactions on Emerging Topics in Computing joint special section on innovation in reconfigurable computing fabrics from devices to architectures. IEEE Transactions on Emerging Topics in Computing, 5(2):207– 209, April/June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Dimitrakopoulos:2021:SPAa

[DPP21a] Giorgos Dimitrakopoulos, Kleanthis Papachatzopoulos, and Vassilis Paliouras. Sum propagate adders. IEEE Transactions on Emerging Topics in Computing, 9(3):1479–1488, July/ September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic). See [DPP21b].

Dimitrakopoulos:2021:SPAb

[DPP21b] Giorgos Dimitrakopoulos, Kleanthis Papachatzopoulos, and Vassilis Paliouras. Sum propagate adders. In IEEE [IEE21], page 110. ISBN 1-66542-293-9 (print), 1-66544-648-X (ebook). LCCN ???? See [DPP21a].

Ding:2023:BIA

[DQB23] Andrew Ding, Ye Qiao, and Nader Bagherzadeh. BNN an ideal architecture for acceleration with resistive in memory computation. IEEE Transactions on Emerging Topics in Computing, 11(2): 281–291, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Dosso:2022:PEA

[DRV22] Fangan Yssouf Dosso, Jean-Marc Robert, and Pascal Véron. PMNS for efficient arithmetic and small memory cost. IEEE Transactions on Emerging Topics in Computing, 10(3):1263– 1277, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

DiCarlo:2021:GES

[DSS21] Stefano Di Carlo, Peilin Song, and Alessandro Savino. Guest editorial: Special section on emerging trends and computing paradigms for testing, reliability and security in future VLSI systems. IEEE Transactions on Emerging Topics in Computing, 9(2):649–650, April/ June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Demrozi:2023:LCW

[DTK^{+}23] Florenc Demrozi, Cristian Turetta, Philipp H.

Kindt, Fabio Chiarani, Ruggero Angelo Bacchin, Nicola Valè, Francesco Pascucci, Paola Cesari, Nicola Smania, Stefano Tamburin, and Graziano Pravadelli. A low-cost wireless body area network for human activity recognition in healthy life and medical applications. IEEE Transactions on Emerging Topics in Computing, 11(4):839–850, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Dullien:2020:WME

[Dul20] Thomas Dullien. Weird machines, exploitability, and provable unexploitability. IEEE Transactions on Emerging Topics in Computing, 8(2):391–403, April/ June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Denecke:2021:MHC

[DVA21] Kerstin Denecke, Sayan Vaaheesan, and Aaganya Arulnathan. A mental health chatbot for regulating emotions (SERMO)

— concept and usability test. IEEE Transactions on Emerging Topics in Computing, 9(3):1170–1182, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

DeGiovanni:2021:RTP

[DVPQ^{+}21] Elisabetta De Giovanni, Adriana Arza Valdés, Miguel Peón-Quirós, Amir Aminifar, and David Atienza. Real-time personalized atrial fibrillation prediction on multi-core wearable sensors. IEEE Transactions on Emerging Topics in Computing, 9(4):1654–1666, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Deng:2013:CSC

[DY13] Xi Deng and Yuanyuan Yang. Communication synchronization in clusterbased sensor networks for cyber-physical systems. IEEE Transactions on Emerging Topics in Computing, 1(1): 98–110, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Dai:2022:ILU

[DYJ22] Xiaoliang Dai, Hongxu Yin, and Niraj K. Jha. Incremental learning using a grow-and-prune paradigm with efficient neural networks. IEEE Transactions on Emerging Topics in Computing, 10(2):752– 762, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

DiNatale:2016:E

[DZ16] Giorgio Di Natale and Stefano Zanero. Editorial. IEEE Transactions on Emerging Topics in Computing, 4(1):33–34, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Drozdenko:2018:HSC

[DZD^{+}18] Benjamin Drozdenko, Matthew Zimmermann, Tuan Dao, Kaushik Chowdhury, and Miriam Leeser. Hardwaresoftware codesign of wireless transceivers on Zynq heterogeneous systems. IEEE Transactions on Emerging Topics in Computing, 6(4):566–578, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Deng:2023:CEM

[DZK^{+}23] Xiaoheng Deng, Jiahao Zhao, Zhufang Kuang, Xuechen Chen, Qi Guo, and Fengxiao Tang. Computation efficiency maximization in multi-UAV-enabled mobile edge computing systems based on 3D deployment optimization. IEEE Transactions on Emerging Topics in Computing, 11(3):778–790, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

DaiSenior:2021:SSS

[DZZ^{+}21] Hong-Ning Dai Senior, Zibin Zheng, Yan Zhang, Michael Rung Tsong Lyu, and Alberto Nannarelli. Special section on scalable computing for blockchain systems. IEEE Transactions on Emerging Topics in Computing, 9(3):1372, July/ September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ezeme:2021:PAD

[EAM21] Okwudili M. Ezeme, Akramul Azim, and Qusay H. Mahmoud. PESKEA: Anomaly detection framework for profiling kernel event attributes in embedded systems. IEEE Transactions on Emerging Topics in Computing, 9(2):957– 971, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

El-Derhalli:2021:DSE

[EDLT21] Hassnaa El-Derhalli, Sébastien Le Beux, and Sofiène Tahar. Design space exploration of stochastic computing architectures implemented using integrated optics. IEEE Transactions on Emerging Topics in Computing, 9(4):2158–2169, October/ December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Eguiluz:2020:EPE

[EGGOR20] Andoni Eguíluz, Mariluz Guenaga, Pablo Garaizar, and Cristian Olivares-Rodríguez. Exploring the progression of early programmers in a set of computational thinking challenges via clickstream analysis. IEEE Transactions on Emerging Topics in Computing, 8(1):256–261, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ebrahimipour:2021:SGS

[EGR21] S. M. Ebrahimipour, Behnam Ghavami, and Mohsen Raji. A statistical gate sizing method for timing yield and lifetime reliability optimization of integrated circuits. IEEE Transactions on Emerging Topics in Computing, 9(2):759–773, April/ June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Emura:2016:SAC

[EKO^{+}16] Keita Emura, Akira Kanaoka, Satoshi Ohta, Kazumasa Omote, and Takeshi Takahashi. Secure and anonymous communication technique: Formal model and its prototype implementation. IEEE Transactions on Emerging Topics in Computing, 4(1):88–101, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Enachescu:2018:LLS

[ELVC18] Marius Enachescu, Mihai Lefter, George Razvan Voicu, and Sorin Dan Cotofana. Low-leakage 3D stacked hybrid NEMFET-CMOS Dual Port memory. IEEE Transactions on Emerging Topics in Computing, 6(2):184–199, April/ June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Eslami:2023:FRR

[EM23] Nima Eslami and Mohammad Hossein Moaiyeri. A flexible and reliable RRAMbased in-memory computing architecture for dataintensive applications. IEEE Transactions on Emerging Topics in Computing, 11(3): 736–748, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

El-Moursy:2022:CCD

[EMDE^{+}22] Ali A. El-Moursy, Abdollah M. Darya, Ahmed S. Elwakil, Abhinand Jha, and Sohaib Majzoub. Chaotic clock driven cryptographic chip: Towards a DPA resistant AES processor. IEEE Transactions on Emerging Topics in Computing, 10(2): 792–805, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kundi:2022:UHS

[eSKZW^{+}22] Dur e Shahwar Kundi, Yuqing Zhang, Chenghua Wang, Ayesha Khalid, Máire O’Neill, and Weiqiang Liu. Ultra high-speed polynomial multiplications for lattice-based cryptography on FPGAs. IEEE Transactions on Emerging Topics in Computing, 10(4):1993– 2005, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ebrahimi:2022:RRR

[ESO^{+}22] Shahriar Ebrahimi, Reza Salkhordeh, Seyed Ali Osia, Ali Taheri, Hamid R. Rabiee, and Hossen Asadi. RC-RNN: Reconfigurable cache architecture for storage systems using recurrent neural networks. IEEE Transactions on Emerging Topics in Computing, 10(3):1492– 1506, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Edamatsu:2023:FMP

[ET23] Takuya Edamatsu and Daisuke Takahashi. Fast multiple-precision integer division using Intel AVX-512. IEEE Transactions on Emerging Topics in Computing, 11(1):224–236, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Fernandez:2021:SAE

[FABC21] Gabriel Fernandez, Jaume Abella, Guillem Bernat, and Francisco J. Cazorla. Surrogate applications for early design stage multicore contention modeling. IEEE Transactions on Emerging Topics in Computing, 9(1): 109–116, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ferretti:2021:CBH

[FAP21] Lorenzo Ferretti, Giovanni Ansaloni, and Laura Pozzi. Cluster-based heuristic for high level synthesis design space exploration. IEEE Transactions on Emerging Topics in Computing, 9 (1):35–43, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Fahad:2014:SCA

[FAT^{+}14] Adil Fahad, Najlaa Alshatri, Zahir Tari, Abdullah Alamri, Ibrahim Khalil, Albert Y. Zomaya, Sebti Foufou, and Abdelaziz Bouras. A survey of clustering algorithms for big data: Taxonomy and empirical analysis. IEEE Transactions on Emerging Topics in Computing, 2(3):267– 279, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Fricke:2022:AIM

[FBL^{+}22] Florian Fricke, Marcelo Brandalero, Sascha Liehr, Simon Kern, Klas Meyer, Stefan Kowarik, Robin Hierzegger, Stephan Westerdick, Michael Maiwald, and Michael Hübner. Artificial intelligence for mass spectrometry and nuclear magnetic resonance spectroscopy using a novel data augmentation method. IEEE Transactions on Emerging Topics in Computing, 10 (1):87–98, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Fadlullah:2022:HHC

[FK22] Zubair Md. Fadlullah and Nei Kato. HCP: Heterogeneous computing platform for federated learning based collaborative content caching towards 6G networks. IEEE Transactions on Emerging Topics in Computing, 10(1):112–123, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Furutani:2021:PPE

[FKNK21] Tohn Furutani, Yuichi Kawamoto, Hiroki Nishiyama, and Nei Kato. Proposal and performance evaluation of information diffusion technique with novel virtual-cell-based Wi-Fi direct. IEEE Transactions on Emerging Topics in Computing, 9(3): 1519–1528, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Faix:2019:DSM

[FLB^{+}19] Marvin Faix, Raphaël Laurent, Pierre Bessière, Emmanuel Mazer, and Jacques Droulez. Design of stochastic machines dedicated to approximate Bayesian inferences. IEEE Transactions on Emerging Topics in Computing, 7(1):60–66, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Fadaeinia:2021:PAL

[FM21a] Bijan Fadaeinia and Amir Moradi. 3-phase adiabatic logic and its sound SCA evaluation. IEEE Transactions on Emerging Topics in Computing, 9(4):2175– 2188, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Fasi:2021:ASRa

[FM21b] Massimiliano Fasi and Mantas Mikaitis. Algorithms for stochastically rounded elementary arithmetic operations in IEEE 754 floatingpoint arithmetic. IEEE Transactions on Emerging Topics in Computing, 9(3): 1451–1466, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic). See [FM21c].

Fasi:2021:ASRb

[FM21c] Massimiliano Fasi and Mantas Mikaitis. Algorithms for stochastically rounded elementary arithmetic operations in IEEE 754 floatingpoint arithmetic. In IEEE [IEE21], page 69. ISBN 1-66542-293-9 (print), 1-66544-648-X (e-book). LCCN ???? See [FM21b].

Fathalizadeh:2024:IGI

[FMA24] Amir Fathalizadeh, Vahideh Moghtadaiee, and Mina Alishahi. Indoor geoindistinguishability: Adopting differential privacy for indoor location data protection. IEEE Transactions on Emerging Topics in Computing, 12(1):293–306, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Fadlullah:2013:CQC

[FNK^{+}13] Zubair Md. Fadlullah, Hiroki Nishiyama, Yuichi Kawamoto, Hirotaka Ujikawa, Ken-Ichi Suzuki, and Naoto Yoshimoto. Cooperative QoS control scheme based on scheduling information in FiWi access network. IEEE Transactions on Emerging Topics in Computing, 1(2): 375–383, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Furano:2021:GEI

[FO21] Gianluca Furano and Marco Ottavi. Guest editorial: IEEE Transactions on Emerging Topics in Computing special issue on advanced command, control and on-board data processing for space avionic systems. IEEE Transactions on Emerging Topics in Computing, 9(1):65–66, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Feng:2017:RPI

[FSCX17] Ying Feng, Radu Stoleru, Chien-An Chen, and Geoffrey G. Xie. A routingprotocol-independent caching framework for mobile clouds. IEEE Transactions on Emerging Topics in Computing, 5(3):353–366, July/ September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Fehmel:2020:GAD

[FSK20] Thomas Fehmel, Dominik Stoffel, and Wolfgang Kunz. Generation of abstract driver models for IP integration verification. IEEE Transactions on Emerging Topics in Computing, 8(4):938–947, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Filtjens:2024:SBA

[FVS24] Benjamin Filtjens, Bart Vanrumste, and Peter Slaets. Skeleton-based action segmentation with multi-stage spatial-temporal graph convolutional neural networks. IEEE Transactions on Emerging Topics in Computing, 12(1):202–212, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Fleites:2015:EPD

[FWC15] Fausto C. Fleites, Haohong Wang, and Shu-Ching Chen. Enhancing product detection with multicue optimization for TV shopping applications. IEEE Transactions on Emerging Topics in Computing, 3(2):161–171, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ghanavati:2022:ABD

[GAI22] Sara Ghanavati, Jemal Abawajy, and Davood Izadi. Automata-based dynamic fault tolerant task scheduling approach in fog computing. IEEE Transactions on Emerging Topics in Computing, 10(1):488–499, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gorbil:2016:MAR

[GAPG16] Gokce Gorbil, Omer H. Abdelrahman, Mihajlo Pavloski, and Erol Gelenbe. Modeling and analysis of RRCbased signalling storms in 3G networks. IEEE Transactions on Emerging Topics in Computing, 4(1):113–127, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Grottke:2023:GES

[GAWT23] Michael Grottke, Alberto Avritzer, Hironori Washizaki, and Kishor Trivedi. Guest editorial special section on applied software aging and rejuvenation. IEEE Transactions on Emerging Topics in Computing, 11(3):550– 552, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gligorijevic:2021:ODC

[GBVS21] Milena Frtunić Gligorijević, Milo Bogdanović, Nataša Veljković, and Leonid Stoimenov. Open data categorization based on formal concept analysis. IEEE Transactions on Emerging Topics in Computing, 9(2):571–581, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Guo:2017:PEC

[GCBK17] Jinhong Guo, Zhi Chen, Yong-Ling Ban, and Yuejun Kang. Precise enumeration of circulating tumor cells using support vector machine algorithm on a microfluidic sensor. IEEE Transactions on Emerging Topics in Computing, 5(4):518–525, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gul:2020:JCA

[GCW20] Madiha Gul, Mohamed Chouikha, and Mamadou Wade. Joint crosstalk aware burst error fault tolerance mechanism for reliable on-chip communication. IEEE Transactions on Emerging Topics in Computing, 8(4):889–896, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Germiniani:2022:AGA

[GDP22] Samuele Germiniani, Alessandro Danese, and Graziano Pravadelli. Automatic generation of assertions for detection of firmware vulnerabilities through alignment of symbolic sequences. IEEE Transactions on Emerging Topics in Computing, 10(2): 728–739, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gelenbe:2015:EPW

[Gel15] Erol Gelenbe. Errors and power when communicating with spins. IEEE Transactions on Emerging Topics in Computing, 3(4):483– 488, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Geurkov:2020:TDP

[Geu20] Vadim Geurkov. On the theory and design of polynomial division circuits. IEEE Transactions on Emerging Topics in Computing, 8(3): 668–687, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Guo:2013:SICa

[GFKL13a] Song Guo, Hannes Frey, Nei Kato, and Yunhao Liu. Special issue on cyberphysical systems (CPS). Part I. IEEE Transactions on Emerging Topics in Computing, 1(1):6–9, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Guo:2013:SICb

[GFKL13b] Song Guo, Hannes Frey, Nei Kato, and Yunhao Liu. Special issue on cyberphysical systems (CPS). Part II. IEEE Transactions on Emerging Topics in Computing, 1(2):203– 206, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ghaffari:2015:DNS

[GHSA15] Mohsen Ghaffari, Behnoosh Hariri, Shervin Shirmohammadi, and Dewan Tanvir Ahmed. A dynamic networking substrate for distributed MMOGs. IEEE Transactions on Emerging Topics in Computing, 3(2): 289–302, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Garcia-Herrero:2021:CSE

[GHSMM21] Francisco Garcia-Herrero, Alfonso Sánchez-Macián, and Juan Antonio Maestro. Combined symbol error correction and spare throughsilicon vias for 3D memories. IEEE Transactions on Emerging Topics in Computing, 9(4):2139–2145, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Guo:2016:CAD

[GJZ^{+}16] Xiaofei Guo, Chenglu Jin, Chi Zhang, Athanasios Papadimitriou, David Hély, and Ramesh Karri. Can algorithm diversity in stream cipher implementation thwart (natural and) malicious faults? IEEE Transactions on Emerging Topics in Computing, 4(3):363–373, July/September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gong:2021:QIM

[GKC21] Young-Ho Gong, Joonho Kong, and Sung Woo Chung. Quantifying the impact of monolithic 3D (M3D) integration on L1 caches. IEEE Transactions on Emerging Topics in Computing, 9(2):854–865, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gangavarapu:2021:FLT

[GKSJ21] Tushaar Gangavarapu, Gokul S Krishnan, Sowmya Kamath S, and Jayakumar Jeganathan. FarSight: Longterm disease prediction using unstructured clinical nursing notes. IEEE Transactions on Emerging Topics in Computing, 9(3):1151– 1169, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gao:2013:CDR

[GLC^{+}13] Sheng Gao, Hao Luo, Da Chen, Shantao Li, Patrick Gallinari, Zhanyu Ma, and Jun Guo. A crossdomain recommendation model for cyber-physical systems. IEEE Transactions on Emerging Topics in Computing, 1(2):384– 393, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gu:2021:FFB

[GLC^{+}21] Chongyan Gu, Weiqiang Liu, Yijun Cui, Neil Hanley, Máire O’Neill, and Fabrizio Lombardi. A flip-flop based arbiter physical unclonable function (APUF) design with high entropy and uniqueness for FPGA implementation. IEEE Transactions on Emerging Topics in Computing, 9(4):1853– 1866, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Guo:2018:MEC

[GLFK18] Hongzhi Guo, Jiajia Liu, Zubair Md. Fadlullah, and Nei Kato. On minimizing energy consumption in FiWi enhanced LTE-A Het-Nets. IEEE Transactions on Emerging Topics in Computing, 6(4):579–591, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Goksoy:2023:CCA

[GLM^{+}23] A. Alper Goksoy, Guihong Li, Sumit K. Mandal, Umit Y. Ogras, and Radu Marculescu. CANNON: Communication-aware sparse neural network optimization. IEEE Transactions on Emerging Topics in Computing, 11(4):882–894, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Guo:2019:BDA

[GLZ19] Hongzhi Guo, Jiajia Liu, and Lei Zhao. Big data acquisition under failures in FiWi enhanced Smart Grid. IEEE Transactions on Emerging Topics in Computing, 7(3):420–432, July/ September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Garraghan:2014:AFR

[GMTX14] Peter Garraghan, Ismael Solis Moreno, Paul Townend, and Jie Xu. An analysis of failure-related energy waste in a large-scale cloud environment. IEEE Transactions on Emerging Topics in Computing, 2(2):166–180, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gochhayat:2015:EQS

[GP15a] Sarada Prasad Gochhayat and Venkatrarm Pallapa. An efficient QoS support for ubiquitous networks. IEEE Transactions on Emerging Topics in Computing, 3(4): 524–533, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gomes:2015:GEI

[GP15b] Abel J. P. Gomes and Edmond Prakash. Guest Editor’s introduction: Massively multiplayer online games technologies and applications. IEEE Transactions on Emerging Topics in Computing, 3(2):258–259, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gerontas:2021:PSM

[GPT^{+}21] Alexandros Gerontas, Vassilios Peristeras, Efthimios Tambouris, Eleni Kaliva, Ioannis Magnisalis, and Konstaninos Tarabanis. Public service models: a systematic literature review and synthesis. IEEE Transactions on Emerging Topics in Computing, 9(2):637– 648, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gutierrez-Santos:2017:SBG

[GSMGP17]  Sergio  Gutierrez-Santos,

Manolis Mavrikis, Eirini Geraniou, and Alexandra Poulovassilis. Similaritybased grouping to support teachers on collaborative activities in an exploratory mathematical microworld. IEEE Transactions on Emerging Topics in Computing, 5(1):56–68, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gupta:2023:EAB

[GSVA23] Eeshan Gupta, Shamik Sural, Jaideep Vaidya, and Vijayalakshmi Atluri. Enabling attribute-based access control in NoSQL databases. IEEE Transactions on Emerging Topics in Computing, 11(1):208–223, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gnawali:2022:HSM

[GT22] Krishna P. Gnawali and Spyros Tragoudas. Highspeed memristive ternary content addressable memory. IEEE Transactions on Emerging Topics in Computing, 10(3):1349–1360, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Garofalo:2021:XEEa

[GTC^{+}21a] Angelo Garofalo, Giuseppe Tagliavini, Francesco Conti, Luca Benini, and Davide Rossi. XpulpNN: Enabling energy efficient and flexible inference of quantized neural networks on RISC-V based IoT end nodes. IEEE Transactions on Emerging Topics in Computing, 9(3):1489– 1505, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic). See [GTC^{+}21b].

Garofalo:2021:XEEb

[GTC^{+}21b] Angelo Garofalo, Giuseppe Tagliavini, Francesco Conti, Luca Benini, and Davide Rossi. XpulpNN: Enabling energy efficient and flexible inference of quantized neural networks on RISC-V based IoT end nodes. In IEEE [IEE21], page 53. ISBN 1-66542-293-9 (print), 1-66544-648-X (ebook). LCCN ???? See [GTC^{+}21a].

Gadioli:2023:EES

[GVF^{+}23] Davide Gadioli, Emanuele Vitali, Federico Ficarelli, Chiara Latini, Candida Manelfi, Carmine Talarico, Cristina Silvano, Carlo Cavazzoni, Gianluca Palermo, and Andrea Rosario Beccari. EXSCALATE: an extreme-scale virtual screening platform for drug discovery targeting polypharmacology to fight SARS-CoV-2. IEEE Transactions on Emerging Topics in Computing, 11(1):170–181, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ghiasi:2023:AAC

[GVO^{+}23] Nika Mansouri Ghiasi, Nandita Vijaykumar, Geraldo F. Oliveira, Lois Orosa, Ivan Fernandez, Mohammad Sadrosadati, Konstantinos Kanellopoulos, Nastaran Hajinazar, Juan Gómez Luna, and Onur Mutlu. ALP: Alleviating CPU-Memory data movement overheads in memory-centric systems. IEEE Transactions on Emerging Topics in Computing, 11(2):388–403, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Garlando:2023:NMB

[GWD^{+}23] Umberto Garlando, Qi Wang, Oleksandr V. Dobrovolskiy, Andrii V. Chumak, and Fabrizio Riente. Numerical model for 32-bit magnonic ripple carry adder. IEEE Transactions on Emerging Topics in Computing, 11(3): 679–688, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ghosh:2017:MSO

[GWHG17] Mainak Ghosh, Wenting Wang, Gopalakrishna Holla, and Indranil Gupta. Morphus: Supporting online reconfigurations in sharded NoSQL systems. IEEE Transactions on Emerging Topics in Computing, 5(4):466–479, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gu:2021:SDS

[GWYJ21] Ke Gu, Na Wu, Bo Yin, and Weijia Jia. Secure data sequence query framework based on multiple fogs. IEEE Transactions on Emerging Topics in Computing, 9(4):1883–1900, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gharib:2022:PIB

[GZB22] Mohamad Gharib, Tommaso Zoppi, and Andrea Bondavalli. On the properness of incorporating binary classification machine learning algorithms into safetycritical systems. IEEE Transactions on Emerging Topics in Computing, 10(4):1671–1686, October/ December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gong:2018:PLP

[GZFS18] Yanmin Gong, Chi Zhang, Yuguang Fang, and Jinyuan Sun. Protecting location privacy for task allocation in ad hoc mobile cloud computing. IEEE Transactions on Emerging Topics in Computing, 6(1):110–121, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gu:2017:CER

[GZG^{+}17] Lin Gu, Deze Zeng, Song Guo, Ahmed Barnawi, and Yong Xiang. Cost efficient resource management in fog computing supported medical cyber-physical system. IEEE Transactions on Emerging Topics in Computing, 5(1):108–119, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gu:2014:CMB

[GZLG14] Lin Gu, Deze Zeng, Peng Li, and Song Guo. Cost minimization for Big Data processing in geo-distributed data centers. IEEE Transactions on Emerging Topics in Computing, 2(3):314– 323, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Grossi:2018:ATE

[GZO^{+}18] Alessandro Grossi, Cristian Zambelli, Piero Olivo, Paolo Pellati, Michele Ramponi, Christian Wenger, Jérémy Alvarez-Hérault, and Ken Mackay. An automated test equipment for characterization of emerging MRAM and RRAM arrays. IEEE Transactions on Emerging Topics in Computing, 6(2):269– 277, April/June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Gao:2022:FTP

[GZT^{+}22] Zhen Gao, Jinhua Zhu, Tong Yan Tyan, Anees Ullah, and Pedro Reviriego. Fault tolerant polyphase filters-based decimators for SRAM-based FPGA implementations. IEEE Transactions on Emerging Topics in Computing, 10(2):591– 601, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hoseinghorban:2022:PCR

[HAE22] Ali Hoseinghorban, Mohammad Abbasinia, and Alireza Ejlali. PROWL: a cache replacement policy for consistency aware renewable powered devices. IEEE Transactions on Emerging Topics in Computing, 10(1):476–487, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Huang:2022:INM

[HAKL22] Junqi Huang, Haider A. F.

Almurib, T. Nandha Kumar, and Fabrizio Lombardi. An inexact Newton method for unconstrained total variation-based image denoising by approximate addition. IEEE Transactions on Emerging Topics in Computing, 10 (2):1192–1207, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Herrero-Alvarez:2023:EPS

[HÁMLS23] Rafael Herrero-Álvarez, Gara Miranda, Coromoto León, and Eduardo Segredo. Engaging primary and secondary school students in computer science through computational thinking training. IEEE Transactions on Emerging Topics in Computing, 11(1):56–69, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Huang:2015:UOF

[HB15] Yuxia Huang and Ling Bian.

Using ontologies and formal concept analysis to integrate heterogeneous tourism information. IEEE Transactions on Emerging Topics in Computing, 3(2):172–184, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Han:2023:MPS

[HBY^{+}23] Honggui Han, Xing Bai, Hongyan Yang, Ying Hou, and Junfei Qiao. Multitask particle swarm optimization with dynamic transformation. IEEE Transactions on Emerging Topics in Computing, 11(3):749–763, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hu:2013:VCO

[HCCL13] Xiping Hu, Terry H. S. Chu, Henry C. B. Chan, and Victor C. M. Leung. Vita: a crowdsensing-oriented mobile cyber-physical system. IEEE Transactions on Emerging Topics in Computing, 1(1):148–165, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

He:2024:HDC

[HCGS24] Meixia He, Jianrui Chen, Maoguo Gong, and Zhongshi Shao. HDGCN: Dualchannel graph convolutional network with higher-order information for robust feature learning. IEEE Transactions on Emerging Topics in Computing, 12(1):126– 138, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Huang:2016:HFT

[HCWL16] Jen-Feng Huang, Guey-Yun Chang, Chun-Feng Wang, and Chih-Hao Lin. Heterogeneous flow table distribution in software-defined networks. IEEE Transactions on Emerging Topics in Computing, 4(2):252– 261, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Homayoun:2020:KAF

[HDA^{+}20] Sajad Homayoun, Ali Dehghantanha, Marzieh Ahmadzadeh, Sattar Hashemi, and Raouf Khayami. Know abnormal, find evil: Frequent pattern mining for ransomware threat hunting and intelligence. IEEE Transactions on Emerging Topics in Computing, 8(2): 341–351, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hu:2014:CCB

[HDL14] Rong Hu, Wanchun Dou, and Jianxun Liu. ClubCF: a clustering-based collaborative filtering approach for Big Data application. IEEE Transactions on Emerging Topics in Computing, 2(3): 302–313, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hosseini:2022:NVN

[HEYB22] Maryam S. Hosseini, Masoumeh Ebrahimi, Pooria Yaghini, and Nader Bagherzadeh. Near volatile and non-volatile memory processing in 3D systems. IEEE Transactions on Emerging Topics in Computing, 10(3):1657–1664, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

He:2017:EER

[HF17] Zhezhi He and Deliang Fan. Energy efficient reconfigurable threshold logic circuit with spintronic devices. IEEE Transactions on Emerging Topics in Computing, 5(2):223–237, April/ June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Harris:2016:UGT

[HH16] Rachel S. Harris and Charles B. Hodges. Using Google Tools for online coursework: Student perceptions. IEEE Transactions on Emerging Topics in Computing, 4(3): 385–391, July/September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hutton:2018:TRO

[HH18] Luke Hutton and Tristan Henderson. Toward reproducibility in online social network research. IEEE Transactions on Emerging Topics in Computing, 6(1): 156–167, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hsieh:2023:GDM

[HHC^{+}23] Yun-Shan Hsieh, Po-Chun Huang, Yuan-Hao Chang, Bo-Jun Chen, Wang Kang, and Wei-Kuan Shih. Granularitydriven management for reliable and efficient skyrmion racetrack memories. IEEE Transactions on Emerging Topics in Computing, 11 (1):95–111, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hsiao:2020:IPL

[HHM20] I-Han Hsiao, Po-Kai Huang, and Hannah Murphy. Integrating programming learning analytics across physical and digital space. IEEE Transactions on Emerging Topics in Computing, 8(1): 206–217, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hafid:2023:TPA

[HHS23] Abdelatif Hafid, Abdelhakim Senhaji Hafid, and Mustapha Samih. A tractable probabilistic approach to analyze Sybil attacks in sharding-based blockchain protocols. IEEE Transactions on Emerging Topics in Computing, 11(1): 126–136, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hussain:2023:LSE

[HHT23] Muhammad Awais Hussain, Shih-An Huang, and Tsung-Han Tsai. Learning with sharing: an edge-optimized incremental learning method for deep neural networks. IEEE Transactions on Emerging Topics in Computing, 11(2): 461–473, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hong:2021:GEI

[HINS21] Yuan Hong, Valerie Issarny, Surya Nepal, and Mudhakar Srivatsa. Guest editors introduction to the joint special section on secure and emerging collaborative computing and intelligent systems. IEEE Transactions on Emerging Topics in Computing, 9(3):1328–1329, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Huang:2021:ESI

[HJCK21] Shih-Chia Huang, Da-Wei Jaw, Bo-Hao Chen, and Sy-Yen Kuo. An efficient single image enhancement approach using luminance perception transformation. IEEE Transactions on Emerging Topics in Computing, 9(2):1083–1094, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Huang:2017:QTB

[HK17] Chin-Yu Huang and Tzu-Yu Kuo. Queueing-theory-based models for software reliability analysis and management. IEEE Transactions on Emerging Topics in Computing, 5(4):540–550, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hameed:2022:ANA

[HKC22] Fazal Hameed, Asif Ali Khan, and Jeronimo Castrillon. ALPHA: a novel algorithm-hardware co-design for accelerating DNA seed location filtering. IEEE Transactions on Emerging Topics in Computing, 10(3): 1464–1475, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hasan:2016:WWO

[HKZH16] Ragib Hasan, Rasib Khan, Shams Zawoad, and Md Munirul Haque. WORAL: a witness oriented secure location provenance framework for mobile devices. IEEE Transactions on Emerging Topics in Computing, 4(1): 128–141, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Han:2024:CHP

[HLK^{+}24] Kyung Hyun Han, Wai-Kong Lee, Angshuman Karmakar, Jose Maria Bermudo Mera, and Seong Oun Hwang. cuFE: High performance privacy preserving support vector machine with inner-product functional encryption. IEEE Transactions on Emerging Topics in Computing, 12(1):328–343, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Huang:2016:EAG

[HLL16] Chung-Ming Huang, Chao-Hsien Lee, and Hsin-Yi Lai. Energy-aware group LBS using D2D offloading and M2M-based mobile proxy handoff mechanisms over the mobile converged networks. IEEE Transactions on Emerging Topics in Computing, 4(4):528–540, October/December 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Halevi:2013:CAD

[HLM^{+}13] Tzipora Halevi, Haoyu Li, Di Ma, Nitesh Saxena, Jonathan Voris, and Tuo Xiang. Context-aware defenses to RFID unauthorized reading and relay attacks. IEEE Transactions on Emerging Topics in Computing, 1(2): 307–318, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hsu:2014:RLB

[HLW14] Roy Chaoming Hsu, Cheng-Ting Liu, and Hao-Li Wang. A reinforcement learningbased ToD provisioning dynamic power management for sustainable operation of energy harvesting wireless sensor node. IEEE Transactions on Emerging Topics in Computing, 2(2):181–191, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Huang:2017:DND

[HLX^{+}17] Jun Huang, Xiang Li, Cong-Cong Xing, Wei Wang, Kun Hua, and Song Guo. DTD: a novel double-track approach to clone detection for RFID-Enabled supply chains. IEEE Transactions on Emerging Topics in Computing, 5(1):134–140, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hoseinghorban:2021:CCA

[HMB^{+}21] Ali Hoseinghorban, Amir Mahdi Hosseini Monazzah, Mostafa Bazzaz, Bardia Safaei, and Alireza Ejlali. COACH: Consistency aware check-pointing for nonvolatile processor in energy harvesting systems. IEEE Transactions on Emerging Topics in Computing, 9(4):2076–2088, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Harutyunyan:2020:MPA

[HMSZ20] Gurgen Harutyunyan, Suren Martirosyan, Samvel Shoukourian, and Yervant Zorian. Memory physical aware multi-level fault diagnosis flow. IEEE Transactions on Emerging Topics in Computing, 8(3):700–711, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hou:2019:TFS

[HNGZ19] Weigang Hou, Zhaolong Ning, Lei Guo, and Xu Zhang. Temporal, functional and spatial Big Data computing framework for large-scale Smart Grid. IEEE Transactions on Emerging Topics in Computing, 7(3):369–379, July/September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Huang:2021:SPN

[HQP^{+}21] Feihu Huang, Shaojie Qiao, Jian Peng, Bing Guo, and Nan Han. STPR: a personalized next pointof-interest recommendation model with spatio-temporal effects based on purpose ranking. IEEE Transactions on Emerging Topics in Computing, 9(2):994–1005, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hui:2019:UBD

[HSG19] Yilong Hui, Zhou Su, and Song Guo. Utility based data computing scheme to provide sensing service in Internet of Things. IEEE Transactions on Emerging Topics in Computing, 7(2): 337–348, April/June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Han:2024:GEI

[HSS24] Jinguang Han, Patrick Schaumont, and Willy Susilo. Guest editorial: IEEE Transactions on Emerging Topics in Computing special section on advances in emerging privacypreserving computing. IEEE Transactions on Emerging Topics in Computing, 12(1): 266–268, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Huang:2017:SIT

[HSzXZ17] Chung-Ming Huang, Chih-Hsiang Shao, Shou zhi Xu, and Huan Zhou. The Social Internet of Thing (S-IOT)-based mobile group handoff architecture and schemes for proximity service. IEEE Transactions on Emerging Topics in Computing, 5(3):425–437, July/ September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hao:2022:SDS

[HTH^{+}22] Jialu Hao, Wenjuan Tang, Cheng Huang, Jian Liu, Huimei Wang, and Ming Xian. Secure data sharing with flexible user access privilege update in cloudassisted IoMT. IEEE Transactions on Emerging Topics in Computing, 10(2):933– 947, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hossain:2023:SRS

[HTS^{+}23] Mousam Hossain, Adrian Tatulian, Shadi Sheikhfaal, Harshavardhana R. Thummala, and Ronald F. De-Mara. Scalable reasoning and sensing using processing-in-memory with hybrid spin/CMOS-based analog/digital blocks. IEEE Transactions on Emerging Topics in Computing, 11(2): 343–357, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Harrison:2015:SAT

[HWFR15] Brent Harrison, Stephen G. Ware, Matthew W. Fendt, and David L. Roberts. A survey and analysis of techniques for player behavior prediction in massively multiplayer online role-playing games. IEEE Transactions on Emerging Topics in Computing, 3(2):260–274, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

He:2020:GRA

[HWH^{+}20] Xiaoming He, Kun Wang, Huawei Huang, Toshiaki Miyazaki, Yixuan Wang, and Song Guo. Green resource allocation based on deep reinforcement learning in content-centric IoT. IEEE Transactions on Emerging Topics in Computing, 8(3):781–796, July/ September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

He:2015:LCM

[HWL15] Zhiyang He, Ji Wu, and Tao Li. Label correlation mixture model: a supervised generative approach to multilabel spoken document categorization. IEEE Transactions on Emerging Topics in Computing, 3(2):235–245, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Huang:2013:BSS

[HWSN13] Zhen Huang, Cheng Wang, Milos Stojmenovic, and Amiya Nayak. Balancing system survivability and cost of Smart Grid via modeling cascading failures. IEEE Transactions on Emerging Topics in Computing, 1(1):45–56, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hung:2017:IRS

[HWW^{+}17] Jui-Long Hung, Morgan C. Wang, Shuyan Wang, Maha Abdelrasoul, Yaohang Li, and Wu He. Identifying at-risk students for early interventions a time-series clustering approach. IEEE Transactions on Emerging Topics in Computing, 5 (1):45–55, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hu:2014:SLN

[HXL^{+}14] Chuanping Hu, Zheng Xu, Yunhuai Liu, Lin Mei, Lan Chen, and Xiangfeng Luo. Semantic link networkbased model for organizing multimedia Big Data. IEEE Transactions on Emerging Topics in Computing, 2(3): 376–387, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Han:2020:DMO

[HYL^{+}20] Guangjie Han, Xuan Yang, Li Liu, Wenbo Zhang, and Mohsen Guizani. A disaster management-oriented path planning for mobile anchor node-based localization in wireless sensor networks. IEEE Transactions on Emerging Topics in Computing, 8(1):115–125, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hu:2019:SIL

[HYW19] Nianhang Hu, Mengmei Ye, and Sheng Wei. Surviving information leakage hardware Trojan attacks using hardware isolation. IEEE Transactions on Emerging Topics in Computing, 7(2): 253–261, April/June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Huang:2016:GTR

[HYZ^{+}16] Jun Huang, Ying Yin, Yanxiao Zhao, Qiang Duan, Wei Wang, and Shui Yu. A game-theoretic resource allocation approach for intercell device-to-device communications in cellular networks. IEEE Transactions on Emerging Topics in Computing, 4(4):475–486, October/December 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hou:2018:PRD

[HZQ^{+}18] Weigang Hou, Rui Zhang, Wen Qi, Kejie Lu, Jianping Wang, and Lei Guo. A provident resource defragmentation framework for mobile cloud computing. IEEE Transactions on Emerging Topics in Computing, 6 (1):32–44, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Hu:2015:AAB

[HZS^{+}15] Xiping Hu, Jidi Zhao, Boon-Chong Seet, Victor C. M. Leung, Terry H. S. Chu, and Henry Chan. S-Aframe: Agentbased multilayer framework with context-aware semantic service for vehicular social networks. IEEE Transactions on Emerging Topics in Computing, 3(1):44– 63, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Huang:2021:SAE

[HZY21] Shijia Huang, Jinghui Zhong, and Wei-Jie Yu. Surrogateassisted evolutionary framework with adaptive knowledge transfer for multi-task optimization. IEEE Transactions on Emerging Topics in Computing, 9(4):1930– 1944, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

IEEE:2021:ISC

[IEE21] IEEE, editor. 2021 IEEE 28th Symposium on Computer Arithmetic: ARITH 2021: virtual conference, 14–16 June 2021: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2021. ISBN 1-66542-293-9 (print), 1-66544-648-X (ebook). LCCN ????

IEEE:2022:ISC

[IEE22] IEEE, editor. 2022 IEEE 29th Symposium on Computer Arithmetic: ARITH 2022: virtual conference, 12–14 September 2022: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2022. ISBN 1-66547-827-6, 1-66547-828-4. LCCN ????

Iijima:2021:AHA

[IMZ^{+}21] Ryo Iijima, Shota Minami, Yunao Zhou, Tatsuya Takehisa, Takeshi Takahashi, Yasuhiro Oikawa, and Tatsuya Mori. Audio hotspot attack: an attack on voice assistance systems using directional sound beams and its feasibility. IEEE Transactions on Emerging Topics in Computing, 9(4):2004–2018, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Impedovo:2021:ASV

[IP21] Donato Impedovo and Giuseppe Pirlo. Automatic signature verification in the mobile cloud scenario: Survey and way ahead. IEEE Transactions on Emerging Topics in Computing, 9(1):554–568, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Imani:2018:ACU

[IPiR18] Mohsen Imani, Shruti Patil, and Tajana imuni Rosing. Approximate computing using multiple-access singlecharge associative memory. IEEE Transactions on Emerging Topics in Computing, 6(3):305–316, July/ September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Imani:2019:RCA

[IPRR19] Mohsen Imani, Daniel Peroni, Abbas Rahimi, and Tajana Simunic Rosing. Resistive CAM acceleration for tunable approximate computing. IEEE Transactions on Emerging Topics in Computing, 7(2):271–280, April/June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ichihara:2019:CAD

[ISI^{+}19] Hideyuki Ichihara, Tatsuyoshi Sugino, Shota Ishii, Tsuyoshi Iwagaki, and Tomoo Inoue. Compact and accurate digital filters based on stochastic computing. IEEE Transactions on Emerging Topics in Computing, 7(1):31–43, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Islam:2021:NCC

[Isl21]  Riadul Islam.  Negative

capacitance clock distribution. IEEE Transactions on Emerging Topics in Computing, 9(1):547–553, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Jayasankaran:2022:TPS

[JBSS^{+}22] N. G. Jayasankaran, A. Sanabria Bórbon, E. Sánchez-Sinencio, J. Hu, and J. Rajendran. Towards provablysecure analog and mixedsignal locking against overproduction. IEEE Transactions on Emerging Topics in Computing, 10(1):386–403, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Jiang:2021:OFC

[JCM^{+}21] Qi Jiang, Zhiren Chen, Jianfeng Ma, Xindi Ma, Jian Shen, and Dapeng Wu. Optimized fuzzy commitment based key agreement protocol for wireless body area network. IEEE Transactions on Emerging Topics in Computing, 9(2):839–853, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Joardar:2023:RRC

[JDL^{+}23] Biresh Kumar Joardar, Janardhan Rao Doppa, Hai Li, Krishnendu Chakrabarty, and Partha Pratim Pande. ReaLPrune: ReRAM crossbaraware lottery ticket pruning for CNNs. IEEE Transactions on Emerging Topics in Computing, 11(2):303– 317, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

JahierPagliari:2021:LOA

[JDP^{+}21] Daniele Jahier Pagliari, Santa Di Cataldo, Edoardo Patti, Alberto Macii, Enrico Macii, and Massimo Poncino. Low-overhead adaptive brightness scaling for energy reduction in OLED displays. IEEE Transactions on Emerging Topics in Computing, 9(3):1625–1636, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Jasemi:2021:ERE

[JHB21] Masoomeh Jasemi, Shaahin Hessabi, and Nader Bagherzadeh. Enhancing reliability of emerging memory technology for machine learning accelerators. IEEE Transactions on Emerging Topics in Computing, 9(4):2234– 2240, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Jiang:2021:ERM

[JHL21] Wei-Cheng Jiang, Kao-Shing Hwang, and Jin-Ling Lin. An experience replay method based on tree structure for reinforcement learning. IEEE Transactions on Emerging Topics in Computing, 9(2):972–982, April/ June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Josyula:2021:PCM

[JKL21] Sai Prashanth Josyula, Johanna Törnquist Krasemann, and Lars Lundberg. Parallel computing for multi-objective train rescheduling. IEEE Transactions on Emerging Topics in Computing, 9(4):1683– 1696, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Junejo:2021:TTM

[JKSC21] Aisha Kanwal Junejo, Nikos Komninos, Mithileysh Sathiyanarayanan, and Bhawani Shankar Chowdhry. Trustee: a trust management system for fogenabled cyber physical systems. IEEE Transactions on Emerging Topics in Computing, 9(4):2030–2041, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Joldes:2021:SSE

[JLN21] Mioara Joldes, Fabrizio Lamberti, and Alberto Nannarelli. Special section on emerging and impacting trends on computer arithmetic. IEEE Transactions on Emerging Topics in Computing, 9(3):1449–1450, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Jedda:2016:DRC

[JM16] Ahmed Jedda and Hussein T. Mouftah. Decentralized RFID coverage algorithms with applications for the reader collisions avoidance problem. IEEE Transactions on Emerging Topics in Computing, 4(4):502–515, October/December 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Jourjon:2017:FTL

[JMBR^{+}17] Guillaume Jourjon, Johann M. Marquez-Barja, Thierry Rakotoarivelo, Alexander Mikroyannidis, Kostas Lampropoulos, Spyros Denazis, Christos Tranoris, Daan Pareit, John Domingue, Luiz A. Dasilva, and Max Ott. FORGE Toolkit: Leveraging distributed systems in eLearning platforms. IEEE Transactions on Emerging Topics in Computing, 5(1):7–19, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Jia:2022:REA

[JMLH22] Yan Jia, John McDermid, Tom Lawton, and Ibrahim Habli. The role of explainability in assuring safety of machine learning in healthcare. IEEE Transactions on Emerging Topics in Computing, 10(4):1746– 1760, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Joseph:2014:TES

[JND14] Siny Joseph, Vinod Namboodiri, and Vishnu Cherusola Dev. Toward environmentally sustainable mobile computing through an economic framework. IEEE Transactions on Emerging Topics in Computing, 2(2): 212–224, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

JahierPagliari:2022:GET

[JSBM22] Daniele Jahier Pagliari, Frank Schirrmeister, Nader Bagherzadeh, and Enrico Macii. Guest editorial: Thematic section on applications of emerging computing technologies in smart manufacturing and Industry 4.0. IEEE Transactions on Emerging Topics in Computing, 10(1):6–8, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Jin:2018:ABR

[JSZ18] A-Long Jin, Wei Song, and Weihua Zhuang. Auctionbased resource allocation for sharing cloudlets in mobile cloud computing. IEEE Transactions on Emerging Topics in Computing, 6 (1):45–57, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Jiang:2021:DMC

[JSZ^{+}21] Weiwen Jiang, Edwin Hsing-Mean Sha, Qingfeng Zhuge, Lei Yang, Hailiang Dong, and Xianzhang Chen. On the design of minimal-cost pipeline systems satisfying hard/soft real-time constraints. IEEE Transactions on Emerging Topics in Computing, 9(1):24–34, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Jia:2023:SAP

[JYZ^{+}23] Kai Jia, Xiao Yu, Chen Zhang, Wenhua Hu, Dongdong Zhao, and Jianwen Xiang. Software aging prediction for cloud services using a gate recurrent unit neural network model based on time series decomposition. IEEE Transactions on Emerging Topics in Computing, 11(3):580–593, July/ September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Khobreh:2016:OBA

[KAF^{+}16] Marjan Khobreh, Fazel Ansari, Madjid Fathi, Réka Vas, Stefan T. Mol, Hannah A. Berkers, and Krisztián Varga. An ontology-based approach for the semantic representation of job knowledge. IEEE Transactions on Emerging Topics in Computing, 4(3):462–473, July/September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kiaei:2023:ASB

[KCS23] Pantea Kiaei, Thomas Conroy, and Patrick Schaumont. Architecture support for bitslicing. IEEE Transactions on Emerging Topics in Computing, 11(2):497– 510, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kim:2022:EAM

[KDKB22] Min Soo Kim, Alberto A.

Del Barrio, HyunJin Kim, and Nader Bagherzadeh. The effects of approximate multiplication on convolutional neural networks. IEEE Transactions on Emerging Topics in Computing, 10(2):904– 916, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kassab:2021:ERB

[KDM^{+}21] Mohamad Kassab, Joanna DeFranco, Tarek Malas, Phillip Laplante, Giuseppe Destefanis, and Valdemar Vicente Graciano Neto. Exploring research in blockchain for healthcare and a roadmap for the future. IEEE Transactions on Emerging Topics in Computing, 9(4):1835– 1852, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kato:2020:GEI

[KG20] Nei Kato and Song Guo. Guest Editor’s introduction: Special section on emerging technologies for disaster management. IEEE Transactions on Emerging Topics in Computing, 8(1):104–105, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kato:2015:GES

[KGM15a] Nei Kato, Song Guo, and Vojislav B. Mišić. Guest editorial: Special issue of IEEE Transactions on Emerging Topics in Computing on emerging mobile and ubiquitous systems. Part II. IEEE Transactions on Emerging Topics in Computing, 3(3): 305–306, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kato:2015:SII

[KGM15b] Nei Kato, Song Guo, and Vojislav B. Mišić. Special issue of the IEEE Transactions on Emerging Topics in Computing on emerging mobile and ubiquitous systems. Part I. IEEE Transactions on Emerging Topics in Computing, 3(1): 5–7, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kuang:2014:TBA

[KHY^{+}14] Liwei Kuang, Fei Hao, Laurence T. Yang, Man Lin, Changqing Luo, and Geyong Min. A tensor-based approach for Big Data representation and dimensionality reduction. IEEE Transactions on Emerging Topics in Computing, 2(3):280– 291, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kumar:2018:BCC

[KIM^{+}18] Neeraj Kumar, Rahat Iqbal, Sudip Misra, Joel J. P. C. Rodrigues, and Mohammad S. Obaidat. Bayesian cooperative coalition game as a service for RFID-based secure QoS management in mobile cloud. IEEE Transactions on Emerging Topics in Computing, 6(1):58–71, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Karmakar:2022:CAG

[KJC22] Rajit Karmakar, Suman Sekhar Jana, and Santanu Chattopadhyay. A cellular automata guided finitestate-machine watermarking strategy for IP protection of sequential circuits. IEEE Transactions on Emerging Topics in Computing, 10(2): 806–823, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kong:2014:PBS

[KK14] Joonho Kong and Farinaz Koushanfar. Processorbased strong physical unclonable functions with aging-based response tuning. IEEE Transactions on Emerging Topics in Computing, 2(1):16–29, March 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Khodamoradi:2021:SSF

[KK21] Alireza Khodamoradi and Ryan Kastner. O(N)-space spatiotemporal filter for reducing noise in neuromorphic vision sensors. IEEE Transactions on Emerging Topics in Computing, 9 (1):15–23, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kang:2022:DER

[KK22] Geon Kang and Seung-Chan Kim. DeepEcho: Echoacoustic recognition of materials using returning echoes with deep neural networks. IEEE Transactions on Emerging Topics in Computing, 10(1):450–462, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kim:2023:RRE

[KK23] Jiho Kim and Tae-Hwan Kim. ROSETTA: a resource and energy-efficient inference processor for recurrent neural networks based on programmable data formats and fine activation pruning. IEEE Transactions on Emerging Topics in Computing, 11(3):650–663, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kaur:2022:HCL

[KKA22] Jasmin Kaur, Mehran Mozaffari Kermani, and Reza Azarderakhsh. Hardware constructions for lightweight cryptographic block cipher QARMA with error detection mechanisms. IEEE Transactions on Emerging Topics in Computing, 10(1): 514–519, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Karmakar:2021:EKG

[KKC21] Rajit Karmakar, Harshit Kumar, and Santanu Chattopadhyay. Efficient keygate placement and dynamic scan obfuscation towards robust logic encryption. IEEE Transactions on Emerging Topics in Computing, 9(4):2109–2124, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kim:2017:PAD

[KKM17] Duckhwan Kim, Jaeha Kung, and Saibal Mukhopadhyay. A power-aware digital multilayer perceptron accelerator with on-chip training based on approximate computing. IEEE Transactions on Emerging Topics in Computing, 5(2):164– 178, April/June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kalapaaking:2024:BBF

[KKY24] Aditya Pribadi Kalapaaking, Ibrahim Khalil, and Xun Yi. Blockchainbased federated learning with SMPC model verification against poisoning attack for healthcare systems. IEEE Transactions on Emerging Topics in Computing, 12(1):269–280, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kansakar:2020:SMC

[KM20] Prasanna Kansakar and Arslan Munir. Selecting microarchitecture configuration of processors for Internet of Things. IEEE Transactions on Emerging Topics in Computing, 8(4):973–985, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Karkar:2018:NCM

[KMD^{+}18] Ammar Karkar, Terrence Mak, Nizar Dahir, Ra ed Al-Dujaily, Kin-Fai Tong, and Alex Yakovlev. Networkon-chip multicast architectures using hybrid wire and surface-wave interconnects. IEEE Transactions on Emerging Topics in Computing, 6(3):357–369, July/September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kawamoto:2022:UAI

[KMK22] Yuichi Kawamoto, Takuto Mitsuhashi, and Nei Kato. UAV-Aided information diffusion for vehicle-to-vehicle (V2V) in disaster scenarios. IEEE Transactions on Emerging Topics in Computing, 10(4):1909–1917, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Khan:2015:INL

[KMM15] Mohammad Shahnoor Islam Khan, Jelena Mišić, and Vojislav B. Mišić. Impact of network load on the performance of a polling MAC with wireless recharging of nodes. IEEE Transactions on Emerging Topics in Computing, 3(3):307– 316, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kong:2021:VVR

[KMW^{+}21] Xiangjie Kong, Mengyi Mao, Wei Wang, Jiaying Liu, and Bo Xu. VO-PRec: Vector representation learning of papers with text information and structural identity for recommendation. IEEE Transactions on Emerging Topics in Computing, 9(1):226–237, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kawamoto:2013:MLN

[KNK13] Yuichi Kawamoto, Hiroki Nishiyama, and Nei Kato. MA-LTRT: a novel method to improve network connectivity and power consumption in mobile ad-hoc based cyber-physical systems. IEEE Transactions on Emerging Topics in Computing, 1(2):366–374, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Karri:2014:SIE

[KP14] Ramesh Karri and Miodrag Potkonjak. Special issue on emerging nanoscale architectures for hardware security, trust, and reliability: Part 1. IEEE Transactions on Emerging Topics in Computing, 2(1): 2–3, March 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kashi:2021:MLM

[KPL^{+}21] Shir Kashi, Ronit Feingold Polak, Boaz Lerner, Lior Rokach, and Shelly Levy-Tzedek. A machine-learning model for automatic detection of movement compensations in stroke patients. IEEE Transactions on Emerging Topics in Computing, 9(3):1234–1247, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Khodabandehloo:2021:CTM

[KR21] Elham Khodabandehloo and Daniele Riboni. Collaborative trajectory mining in smart-homes to support early diagnosis of cognitive decline. IEEE Transactions on Emerging Topics in Computing, 9(3):1194–1205, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Khoshkbarforoushha:2017:DBW

[KRG^{+}17] Alireza Khoshkbarforoushha, Rajiv Ranjan, Raj Gaire, Ehsan Abbasnejad, Lizhe Wang, and Albert Y. Zomaya. Distribution based workload modelling of continuous queries in clouds. IEEE Transactions on Emerging Topics in Computing, 5(1):120–133, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kumar:2018:AMR

[KS18] Somesh Kumar and Rohit Sharma. Analytical model for resistivity and mean free path in on-chip interconnects with rough surfaces. IEEE Transactions on Emerging Topics in Computing, 6(2):233–243, April/ June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kaur:2022:HCE

[KSKA22] Jasmin Kaur, Ausmita Sarker, Mehran Mozaffari Kermani, and Reza Azarderakhsh. Hardware constructions for error detection in lightweight welch-gong (WG)-Oriented streamcipher WAGE benchmarked on FPGA. IEEE Transactions on Emerging Topics in Computing, 10 (2):1208–1215, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kermani:2016:GEI

[KSU16] Mehran Mozaffari Kermani, Erkay Savas, and Shambhu J. Upadhyaya. Guest editorial: Introduction to the special issue on emerging security trends for deeply-embedded computing systems. IEEE Transactions on Emerging Topics in Computing, 4(3):318–320, July/September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kumar:2019:ESN

[KTM19] S. Dinesh Kumar, Himanshu Thapliyal, and Azhar Mohammad. EE-SPFAL: a novel energy-efficient secure positive feedback adiabatic logic for DPA resistant RFID and smart card. IEEE Transactions on Emerging Topics in Computing, 7(2):281–293, April/ June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kinseher:2019:ITR

[KVP19] Josef Kinseher, Moritz Voelker, and Ilia Polian. Improving testability and reliability of advanced SRAM architectures. IEEE Transactions on Emerging Topics in Computing, 7(3):456–467, July/September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kato:2020:FSP

[KWS^{+}20] Takaaki Kato, Senling Wang, Yasuo Sato, Seiji Kajihara, and Xiaoqing Wen. A flexible scan-in power control method in logic BIST and its evaluation with TEG chips. IEEE Transactions on Emerging Topics in Computing, 8(3):591–601, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kesan:2021:EIR

[KZ21] Jay P. Kesan and Linfeng Zhang. An empirical investigation of the relationship between local government budgets, IT expenditures, and cyber losses. IEEE Transactions on Emerging Topics in Computing, 9(2):582– 596, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Kasuya:2020:CEW

[KZT^{+}20] Seiji Kasuya, Xiaokang Zhou, Kiichi Tago, Shoji Nishimura, and Qun Jin. Cyber-enabled well-being oriented daily living support based on personal data analysis. IEEE Transactions on Emerging Topics in Computing, 8(2):493–502, April/ June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lo:2013:CNH

[LA13] Chun-Hao Lo and Nirwan Ansari. CONSUMER: a novel hybrid intrusion detection system for distribution networks in Smart Grid. IEEE Transactions on Emerging Topics in Computing, 1(1):33–44, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2023:SMC

[LBX^{+}23] Yueting Li, Tianshuo Bai, Xinyi Xu, Yundong Zhang, Bi Wu, Hao Cai, Biao Pan, and Weisheng Zhao. A survey of MRAM-Centric computing: From near memory to in memory. IEEE Transactions on Emerging Topics in Computing, 11(2):318– 330, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Laurenciu:2020:RAD

[LC20] Nicoleta Cucu Laurenciu and Sorin Dan Cotofana. Reliability aware design and lifetime management of computing platforms. IEEE Transactions on Emerging Topics in Computing, 8(3): 602–615, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2020:GEI

[LCLK20] Zhe Liu, Kim-Kwang Raymond Choo, Weiqiang Liu, and Muhammad Khurram Khan. Guest editorial: Introduction to the special section on cyber security threats and defense advance. IEEE Transactions on Emerging Topics in Computing, 8(2):264–266, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2018:CRA

[LCM18] Taizhi Liu, Chang-Chih Chen, and Linda Milor. Comprehensive reliabilityaware statistical timing analysis using a unified gatedelay model for microprocessors. IEEE Transactions on Emerging Topics in Computing, 6(2):219–232, April/June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2023:AHC

[LCM^{+}23] Guangyan Li, Donglong Chen, Gaoyu Mao, Wangchen Dai, Abdurrashid Ibrahim Sanka, and Ray C. C. Cheung. Algorithm-hardware co-design of split-radix discrete Galois transformation for KyberKEM. IEEE Transactions on Emerging Topics in Computing, 11(4):824–838, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2021:RPR

[LCR^{+}21] Shanshan Liu, Ke Chen, Pedro Reviriego, Weiqiang Liu, Ahmed Louri, and Fabrizio Lombardi. Reduced precision redundancy for reliable processing of data. IEEE Transactions on Emerging Topics in Computing, 9(4):1960–1971, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lyu:2019:SEO

[LCY^{+}19] Ling Lyu, Cailian Chen, Jing Yan, Feilong Lin, Cunqing Hua, and Xinping Guan. State estimation oriented wireless transmission for ubiquitous monitoring in industrial cyber-physical systems. IEEE Transactions on Emerging Topics in Computing, 7(1):187–201, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2020:SCP

[LCY20] Chen Liu, Patrick Cronin, and Chengmo Yang. Securing cyber-physical systems from hardware Trojan collusion. IEEE Transactions on Emerging Topics in Computing, 8(3):655–667, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2021:EFH

[LCZ21] Jiguo Li, Ningyu Chen, and Yichen Zhang. Extended file hierarchy access control scheme with attribute-based encryption in cloud computing. IEEE Transactions on Emerging Topics in Computing, 9(2):983–993, April/ June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

L:2021:WMD

[LD21] Shiva Darshan S. L. and Jaidhar C. D. Windows malware detector using convolutional neural network based on visualization images. IEEE Transactions on Emerging Topics in Computing, 9(2):1057–1069, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2021:PPR

[LDCY21] Ximeng Liu, Robert H. Deng, Kim-Kwang Raymond Choo, and Yang Yang. Privacy-preserving reinforcement learning design for patient-centric dynamic treatment regimes. IEEE Transactions on Emerging Topics in Computing, 9(1): 456–470, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lin:2020:TNV

[LDJ20] Chun-Cheng Lin, Der-Jiunn Deng, and Shun-Yu Jhong. A triangular NodeTrix visualization interface for overlapping social community structures of cyber-physical-social systems in smart factories. IEEE Transactions on Emerging Topics in Computing, 8(1):58–68, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Laplante:2016:GES

[LDLN16] Phillip A. Laplante, Claudio Demartini, Fabrizio Lamberti, and Colin J. Neill. Guest editorial: Special issue on emerging trends in education part I. IEEE Transactions on Emerging Topics in Computing, 4(3): 382–384, July/September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Laplante:2017:GES

[LDLN17] Phillip A. Laplante, Claudio Demartini, Fabrizio Lamberti, and Colin J. Neill. Guest editorial: Special issue on emerging trends in education part II. IEEE Transactions on Emerging Topics in Computing, 5(1): 5–6, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2016:PRB

[LDOG16] He Li, Mianxiong Dong, Kaoru Ota, and Minyi Guo. Pricing and repurchasing for Big Data processing in multi-clouds. IEEE Transactions on Emerging Topics in Computing, 4(2):266– 277, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2014:TQE

[LFBL14] Chi Harold Liu, Jun Fan, Joel W. Branch, and Kin K. Leung. Toward QoI and energy-efficiency in Internetof-Things sensory environments. IEEE Transactions on Emerging Topics in Computing, 2(4):473– 487, December 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2018:TME

[LFL^{+}18] Hui Li, Alain Fourmigue, Sébastien Le Beux, Ian O’Connor, and Gabriela Nicolescu. Towards maximum energy efficiency in nanophotonic interconnects with thermal-aware on-chip laser tuning. IEEE Transactions on Emerging Topics in Computing, 6(3):343–356, July/September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2013:JOE

[LGC13] Peng Li, Song Guo, and Zixue Cheng. Joint optimization of electricity and communication cost for meter data collection in Smart Grid. IEEE Transactions on Emerging Topics in Computing, 1(2):297–306, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2019:DDE

[LGT^{+}19] Wen-Tai Li, Sai Ram Gubba, Wayes Tushar, Chau Yuen, Naveed Ul Hassan, H. Vincent Poor, Kristin L. Wood, and Chao-Kai Wen. Data driven electricity management for residential air conditioning systems: an experimental approach. IEEE Transactions on Emerging Topics in Computing, 7(3):380–391, July/September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ledy:2015:IRS

[LHB^{+}15] Jonathan Ledy, Benoît Hilt, Hervé Boeglen, Anne-Marie Poussard, Frédéric Drouhin, Rodolphe Vauzelle, and Pascal Lorenz. Impact of realistic simulation on the evaluation of mobile ad hoc routing protocols. IEEE Transactions on Emerging Topics in Computing, 3(3):317– 334, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lorenz:2018:GEN

[LHJ18] Pascal Lorenz, Sofiane Hamrioui, and Abbas Jamalipour. Guest editorial: Next generation wireless computing systems. IEEE Transactions on Emerging Topics in Computing, 6(4):551–552, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lindquist:2021:IAR

[LHKH21] Wyatt Lindquist, Sumi Helal, Ahmed Khaled, and Wesley Hutchinson. IoTility: Architectural requirements for enabling health IoT ecosystems. IEEE Transactions on Emerging Topics in Computing, 9(3): 1206–1218, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lamberti:2020:GEJ

[LHMW20] Fabrizio Lamberti, Gwo-Jen Hwang, Baltasar Fernández Manjón, and Wenping Wang. Guest editorial: Joint special issue on innovation in technologies for educational computing. IEEE Transactions on Emerging Topics in Computing, 8(1): 179–181, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2018:FOD

[LHP^{+}18] Jingyu Liu, Yongzhen Huang, Junran Peng, Jun Yao, and Liang Wang. Fast object detection at constrained energy. IEEE Transactions on Emerging Topics in Computing, 6(3):409–416, July/ September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2022:MIS

[LHS^{+}22] Fangqing Liu, Han Huang, Junpeng Su, Stuart D. Semujju, Zhongming Yang, and Zhifeng Hao. Manifoldinspired search-based algorithm for automated test case generation. IEEE Transactions on Emerging Topics in Computing, 10 (2):1075–1090, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2020:LPM

[LHZS20] Hao Liu, Linpeng Huang, Yanmin Zhu, and Yanyan Shen. LibreKV: a persistent in-memory key-value store. IEEE Transactions on Emerging Topics in Computing, 8(4):916–927, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Luo:2015:DGS

[LJD^{+}15] Feng Luo, Chunxiao Jiang, Jun Du, Jian Yuan, Yong Ren, Shui Yu, and Mohsen Guizani. A distributed gateway selection algorithm for UAV networks. IEEE Transactions on Emerging Topics in Computing, 3(1): 22–33, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liufu:2022:RNI

[LJX^{+}22] Ying Liufu, Long Jin, Jinqiang Xu, Xiuchun Xiao, and Dongyang Fu. Reformative noise-immune neural network for equalityconstrained optimization applied to image target detection. IEEE Transactions on Emerging Topics in Computing, 10(2):973– 984, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lin:2019:WLC

[LK19] Sheng-En David Lin and Dae Hyun Kim. Wire length characteristics of multi-tier gate-level monolithic 3D ICs. IEEE Transactions on Emerging Topics in Computing, 7(2):301–310, April/ June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lu:2022:GED

[LK22] Ching-Hu Lu and Hao-Chung Ku. Goal and elitedata-driven anthropomorphic learning for streamingimage analytics. IEEE Transactions on Emerging Topics in Computing, 10(3): 1532–1545, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lu:2023:NCB

[LK23] Xiaozhou Lu and Sunghwan Kim. New construction of balanced codes based on weights of data for DNA storage. IEEE Transactions on Emerging Topics in Computing, 11(4):973–984, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2013:CCP

[LL13] Ming Li and Pan Li. Crowdsourcing in cyber-physical systems: Stochastic optimization with strong stability. IEEE Transactions on Emerging Topics in Computing, 1(2):218–231, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2015:EEM

[LLD^{+}15] Hongwei Li, Dongxiao Liu, Yuanshun Dai, Tom H. Luan, and Xuemin Sherman Shen. Enabling efficient multi-keyword ranked search over encrypted mobile cloud data through blind storage. IEEE Transactions on Emerging Topics in Computing, 3(1):127– 138, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2018:PSE

[LLD^{+}18] Hongwei Li, Dongxiao Liu, Yuanshun Dai, Tom H. Luan, and Shui Yu. Personalized search over encrypted data with efficient and secure updates in mobile clouds. IEEE Transactions on Emerging Topics in Computing, 6(1):97–109, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Leitersdorf:2023:AHT

[LLG^{+}23] Orian Leitersdorf, Dean Leitersdorf, Jonathan Gal, Mor Dahan, Ronny Ronen, and Shahar Kvatinsky. AritPIM: High-throughput inmemory arithmetic. IEEE Transactions on Emerging Topics in Computing, 11(3): 720–735, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2019:MED

[LLL19] Jun Li, Baochun Li, and Bo Li. Mist: Efficient dissemination of erasurecoded data in data centers. IEEE Transactions on Emerging Topics in Computing, 7(3):468–480, July/ September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2020:MCT

[LLL^{+}20] Zhetao Li, Yuxin Liu, Anfeng Liu, Shiguo Wang, and Haolin Liu. Minimizing convergecast time and energy consumption in green Internet of Things. IEEE Transactions on Emerging Topics in Computing, 8(3):797–813, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2023:WBE

[LLLG23] Yanhua Liu, Zhihuang Liu, Ximeng Liu, and Wenzhong Guo. A web back-end database leakage incident reconstruction framework over unlabeled logs. IEEE Transactions on Emerging Topics in Computing, 11(1):237– 252, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liao:2016:ESA

[LLS^{+}16] Weixian Liao, Ming Li, Sergio Salinas, Pan Li, and Miao Pan. Energysource-aware cost optimization for green cellular networks with strong stability. IEEE Transactions on Emerging Topics in Computing, 4(4):541–555, October/December 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2015:EBE

[LLW^{+}15] Lin Liu, Yang Liu, Lizhe Wang, Albert Zomaya, and Shiyan Hu. Economical and balanced energy usage in the smart home infrastructure: a tutorial and new results. IEEE Transactions on Emerging Topics in Computing, 3(4):556– 570, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lee:2022:DKB

[LLW22] Kuen-Jong Lee, Ching-An Liu, and Chia-Chi Wu. A dynamic-key based secure scan architecture for manufacturing and in-field IC testing. IEEE Transactions on Emerging Topics in Computing, 10(1):373–385, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lamberti:2018:GES

[LMC18] Fabrizio Lamberti, Tiziana Margaria, and Henry C. B. Chan. Guest editorial: Special section on computing education & learning technologies. IEEE Transactions on Emerging Topics in Computing, 6(1):5–6, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2013:FMC

[LMK^{+}13] Shan Liu, Salman Mashayekh, Deepa Kundur, Takis Zourntos, and Karen Butler-Purry. A framework for modeling cyber-physical switching attacks in Smart Grid. IEEE Transactions on Emerging Topics in Computing, 1(2):273–285, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lamberti:2014:COF

[LMS^{+}14] Fabrizio Lamberti, Federico Manuri, Andrea Sanna, Gianluca Paravati, Pietro Pezzolla, and Paolo Montuschi. Challenges, opportunities, and future trends of emerging techniques for augmented reality-based maintenance. IEEE Transactions on Emerging Topics in Computing, 2(4):411– 421, December 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2017:VAR

[LMW^{+}17] Peng Li, Toshiaki Miyazaki, Kun Wang, Song Guo, and Weihua Zhuang. Vehicleassist resilient information and network system for disaster management. IEEE Transactions on Emerging Topics in Computing, 5(3): 438–448, July/September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2015:EET

[LNK^{+}15] Meng Li, Hiroki Nishiyama, Nei Kato, Yasunori Owada, and Kiyoshi Hamaguchi. On the energy-efficient of throughput-based scheme using renewable energy for wireless mesh networks in disaster area. IEEE Transactions on Emerging Topics in Computing, 3(3):420– 431, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lombardi:2013:EM

[Lom13] Fabrizio Lombardi. EIC message. IEEE Transactions on Emerging Topics in Computing, 1(1):3–5, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lombardi:2014:EM

[Lom14] Fabrizio Lombardi. EiC message. IEEE Transactions on Emerging Topics in Computing, 2(3):252– 253, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lombardi:2015:EM

[Lom15] Fabrizio Lombardi. EIC message. IEEE Transactions on Emerging Topics in Computing, 3(1):3–4, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lombardi:2016:MEC

[Lom16] Fabrizio Lombardi. Message from the Editor-in-Chief. IEEE Transactions on Emerging Topics in Computing, 4(1):3–4, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lombardi:2017:MEC

[Lom17] Fabrizio Lombardi. Message from the Editor-in-Chief. IEEE Transactions on Emerging Topics in Computing, 5(1):3–4, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lin:2013:SCD

[LPY^{+}13] Man Lin, Yongwen Pan, Laurence T. Yang, Minyi Guo, and Nenggan Zheng. Scheduling co-design for reliability and energy in cyberphysical systems. IEEE Transactions on Emerging Topics in Computing, 1(2):

353–365, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Luo:2014:COU

[LPZ^{+}14] Yu Luo, Lina Pu, Michael Zuba, Zheng Peng, and Jun-Hong Cui. Challenges and opportunities of underwater cognitive acoustic networks. IEEE Transactions on Emerging Topics in Computing, 2(2):198–211, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2023:HBN

[LQT^{+}23] Zhibo Li, Yurong Qian, Fengxiao Tang, Ming Zhao, and Yusen Zhu. H-BILSTM: a novel bidirectional long short term memory network based intelligent early warning scheme in mobile edge computing (MEC). IEEE Transactions on Emerging Topics in Computing, 11(1): 253–264, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2021:EAE

[LRG^{+}21] Shanshan Liu, Pedro Reviriego, Jing Guo, Jie Han, and Fabrizio Lombardi. Exploiting asymmetry in eDRAM errors for redundancy-free errortolerant design. IEEE Transactions on Emerging Topics in Computing, 9(4):2064–2075, October/ December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2021:VMS

[LRHL21] Shanshan Liu, Pedro Reviriego, José Alberto Hernández, and Fabrizio Lombardi. Voting margin: a scheme for error-tolerant k nearest neighbors classifiers for machine learning. IEEE Transactions on Emerging Topics in Computing, 9(4):2089– 2098, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2021:DLM

[LRL21] Shanshan Liu, Pedro Reviriego, and Fabrizio Lombardi. Detection of limited magnitude errors in emerging multilevel cell memories by one-bit parity (OBP) or two-bit parity (TBP). IEEE Transactions on Emerging Topics in Computing, 9(4):1792–1802, October/ December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2022:GES

[LRLG22] Shanshan Liu, Pedro Reviriego, Fabrizio Lombardi, and Patrick Girard. Guest editorial: Special section on to be safe and dependable in the era of artificial intelligence: Emerging techniques for trusted and reliable machine learning. IEEE Transactions on Emerging Topics in Computing, 10(4):1668– 1670, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Limaye:2022:PLP

[LRP^{+}22] Nimisha Limaye, Nikhil Rangarajan, Satwik Patnaik, Ozgur Sinanoglu, and Kanad Basu. PolyWorm: Leveraging polymorphic behavior to implant hardware Trojans. IEEE Transactions on Emerging Topics in Computing, 10(3):1443– 1455, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Luza:2022:EER

[LRS^{+}22] Lucas Matana Luza, Annachiara Ruospo, Daniel Söderström, Carlo Cazzaniga, Maria Kastriotou, Ernesto Sanchez, Alberto Bosio, and Luigi Dilillo. Emulating the effects of radiation-induced soft-errors for the reliability assessment of neural networks. IEEE Transactions on Emerging Topics in Computing, 10(4):1867–1882, October/ December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2021:PMA

[LRXW21] Jiaqiang Li, Pedro Reviriego, Liyi Xiao, and Haotian Wu. Protecting memories against soft errors: The case for customizable error correction codes. IEEE Transactions on Emerging Topics in Computing, 9(2): 651–663, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2014:SHM

[LRYK14] Chen Liu, Jeyavijayan Rajendran, Chengmo Yang, and Ramesh Karri. Shielding heterogeneous MPSoCs from untrustworthy 3PIPs through security-driven task scheduling. IEEE Transactions on Emerging Topics in Computing, 2(4):461– 472, December 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lin:2021:SLM

[LS21] Jianyi Lin and Emiliano Sironi. Sparse logistic maximum likelihood estimation for optimal well-being determinants. IEEE Transactions on Emerging Topics in Computing, 9(3):1316–1327, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2017:GEI

[LSO17] Weiqiang Liu, Earl E. Swartzlander, and Máire O’Neill. Guest editorial: Introduction to the special issue on emerging technologies and designs for application-specific computing. IEEE Transactions on Emerging Topics in Computing, 5(2):148–150, April/ June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2020:QPV

[LSP^{+}20] Xiang Li, J. David Smith, Tianyi Pan, Thang N. Dinh, and My T. Thai. Quantifying privacy vulnerability to socialbot attacks: an adaptive non-submodular model. IEEE Transactions on Emerging Topics in Computing, 8(3):855–868, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Luo:2021:WYP

[LSR^{+}21] Feiheng Luo, Aixin Sun, Aravind Sesagiri Raamkumar, Mojisola Erdt, and Yin-Leng Theng. Will your paper get promoted by a citation? A case study of citation promoter in computer science discipline. IEEE Transactions on Emerging Topics in Computing, 9(1):238–245, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lee:2021:FRC

[LT21] Ivan Lee and Yun Tie.

Fitness and research complexity among researchactive universities in the world. IEEE Transactions on Emerging Topics in Computing, 9(1):293–301, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2022:GLB

[LTL^{+}22] Xiao Liu, Zaiyang Tang, Peng Li, Song Guo, Xuepeng Fan, and Jinbo Zhang. A graph learning based approach for identity inference in DApp platform blockchain. IEEE Transactions on Emerging Topics in Computing, 10(1):438–449, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lu:2022:CAB

[LVJ22] Jie Lu, Naveen Verma, and Niraj K. Jha. Convolutional autoencoder-based transfer learning for multitask image inferences. IEEE Transactions on Emerging Topics in Computing, 10 (2):1045–1057, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Luo:2024:FBA

[LVR^{+}24] Yandong Luo, Johan Vanderhaegen, Oleg Rybakov, Martin Kraemer, Niel Warren, and Shimeng Yu. A FeFET-based ADC offset robust compute-in-memory architecture for streaming keyword spotting (KWS). IEEE Transactions on Emerging Topics in Computing, 12(1):23–34, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lin:2022:ICF

[LWC22] Chia-Yu Lin, Li-Chun Wang, and Shu-Ping Chang. Incremental checkpointing for fault-tolerant stream processing systems: a data structure approach. IEEE Transactions on Emerging Topics in Computing, 10(1): 124–136, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lin:2021:BBD

[LWM^{+}21] Xi Lin, Jun Wu, Shahid Mumtaz, Sahil Garg, Jianhua Li, and Mohsen Guizani. Blockchain-based on-demand computing resource trading in IoV-Assisted smart city. IEEE Transactions on Emerging Topics in Computing, 9(3):1373–1385, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Limaye:2022:APD

[LWN^{+}22] Nimisha Limaye, Christian Wachsmann, Mohammed Nabeel, Mohammed Ashraf, Arun Kanuparthi, and Ozgur Sinanoglu. Anti-DOTE: Protecting debug against outsourced test entities. IEEE Transactions on Emerging Topics in Computing, 10(3):1507– 1518, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2022:APA

[LXL^{+}22a] Anran Li, Shuangshuang Xue, Xiang-Yang Li, Lan Zhang, and Jianwei Qian. AppDNA: Profiling app behavior via deep-learning function call graphs. IEEE Transactions on Emerging Topics in Computing, 10(1): 414–427, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2022:DUA

[LXL^{+}22b] Weiqiang Liu, Tao Xu, Jing Li, Chenghua Wang, Paolo Montuschi, and Fabrizio Lombardi. Design of unsigned approximate hybrid dividers based on restoring array and logarithmic dividers. IEEE Transactions on Emerging Topics in Computing, 10(1):339–350, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Lu:2021:MMG

[LXX^{+}21] Ye Lu, Kunpeng Xie, Guanbin Xu, Han Dong, Cheng Li, and Tao Li. MTFC: a multi-GPU training framework for cube-CNN-based hyperspectral image classification. IEEE Transactions on Emerging Topics in Computing, 9(4):1738– 1752, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2024:GRP

[LXX^{+}24] Xiaolin Li, Qikui Xu, Zhenyu Xu, Hongyan Zhang, and Li Xu. Graph reconfigurable pooling for graph representation learning. IEEE Transactions on Emerging Topics in Computing, 12(1): 139–149, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2018:EEN

[LYA18] Cheng Li, Mo Yang, and Paul Ampadu. An energyefficient NoC router with adaptive fault-tolerance using channel slicing and ondemand TMR. IEEE Transactions on Emerging Topics in Computing, 6(4):538–550, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Luo:2020:CDF

[LYBZ20] Wenjian Luo, Zhenglong Yan, Chenyang Bu, and Daofu Zhang. Community detection by fuzzy relations. IEEE Transactions on Emerging Topics in Computing, 8(2):478–492, April/ June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2022:FRU

[LYK^{+}22] Jing Li, Dong Hye Ye, Mathias Kolsch, Juan P. Wachs, and Charles A. Bouman. Fast and robust UAV to UAV detection and tracking from video. IEEE Transactions on Emerging Topics in Computing, 10(3):1519– 1531, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2013:RCT

[LYS13] Junliang Liu, Zheng Yang, and Ivan Stojmenovic. Receiver consensus: On-time warning delivery for vehicular ad-hoc networks. IEEE Transactions on Emerging Topics in Computing, 1(1): 57–68, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2021:DAP

[LYW^{+}21] Peng Liu, Zhiqiang You, Jigang Wu, Michael Elimu, Weizheng Wang, Shuo Cai, and Yinhe Han. Defect analysis and parallel testing for 3D hybrid CMOS-Memristor memory. IEEE Transactions on Emerging Topics in Computing, 9(2): 745–758, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2022:PAE

[LZD^{+}22] Dingding Li, Niyang Zhang, Mianxiong Dong, Hao Chen, Kaoru Ota, and Yong Tang. PM-AIO: an effective asynchronous I/O system for persistent memory. IEEE Transactions on Emerging Topics in Computing, 10(3): 1558–1574, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2019:SBB

[LZH19] Lin Liu, Yuchen Zhou, and Shiyan Hu. Stochastic buffering for bundled SWCNT interconnects considering unidimensional fabrication variation. IEEE Transactions on Emerging Topics in Computing, 7(4):585–595, October/December 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liang:2021:CCB

[LZL^{+}21] Wei Liang, Dafang Zhang, Xia Lei, Mingdong Tang, Kuan-Ching Li, and Albert Y. Zomaya. Circuit copyright blockchain: Blockchain-based homomorphic encryption for IP circuit protection. IEEE Transactions on Emerging Topics in Computing, 9(3): 1410–1420, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2022:MLB

[LZLC22] Junxiu Liu, Shunsheng Zhang, Yuling Luo, and Lvchen Cao. Machine learning-based similarity attacks for chaos-based cryptosystems. IEEE Transactions on Emerging Topics in Computing, 10(2):824– 837, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Liu:2021:DAM

[LZM^{+}21] Weiqiang Liu, Tingting Zhang, Emma McLarnon, Maire O’Neill, Paolo Montuschi, and Fabrizio Lombardi. Design and analysis of majority logic-based approximate adders and multipliers. IEEE Transactions on Emerging Topics in Computing, 9(3):1609–1624, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2021:HEP

[LZW21] Xingxin Li, Youwen Zhu, and Jian Wang. Highly efficient privacy preserving location-based services with enhanced one-round blind filter. IEEE Transactions on Emerging Topics in Computing, 9(4):1803–1814, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Li:2022:QMB

[LZX^{+}22] YaoChong Li, Ri-Gui Zhou, RuiQing Xu, Jia Luo, and She-Xiang Jiang. A quantum mechanics-based framework for EEG signal feature extraction and classification. IEEE Transactions on Emerging Topics in Computing, 10(1):211–222, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Macii:2018:GES

[MA18] Alberto Macii and Andrea Acquaviva. Guest editorial for the special section on emerging computational paradigms. IEEE Transactions on Emerging Topics in Computing, 6(3):303–304, July/September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ma:2023:BAM

[MBW23] Siming Ma, David Brooks, and Gu-Yeon Wei. A binary-activation, multilevel weight RNN and training algorithm for ADC-/ DAC-Free and noise-resilient processing-in-memory inference with eNVM. IEEE Transactions on Emerging Topics in Computing, 11(2): 292–302, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Migliorini:2021:DTA

[MCB21] Sara Migliorini, Damiano Carra, and Alberto Belussi. Distributing tourists among POIs with an adaptive trip recommendation system. IEEE Transactions on Emerging Topics in Computing, 9(4):1765–1779, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Magan-Carrion:2022:IRN

[MCUDCD22] Roberto Magán-Carrión, Daniel Urda, Ignacio Diaz-Cano, and Bernabé Dorronsoro. Improving the reliability of network intrusion detection systems through dataset integration. IEEE Transactions on Emerging Topics in Computing, 10(4):1717–1732, October/ December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Musci:2021:OFD

[MDB^{+}21] Mirto Musci, Daniele De Martini, Nicola Blago, Tullio Facchinetti, and Marco Piastra. Online fall detection using recurrent neural networks on smart wearable devices. IEEE Transactions on Emerging Topics in Computing, 9(3):1276–1289, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Murillo:2022:PPL

[MDB^{+}22] Raul Murillo, Alberto A.

Del Barrio, Guillermo Botella, Min Soo Kim, HyunJin Kim, and Nader Bagherzadeh. PLAM: a posit logarithmapproximate multiplier. IEEE Transactions on Emerging Topics in Computing, 10(4):2079–2085, October/ December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Manna:2023:PPL

[MDB^{+}23] Siladittya Manna, Dipayan Das, Saumik Bhattacharya, Umapada Pal, and Sukalpa Chanda. PLSM: a parallelized liquid state machine for unintentional action detection. IEEE Transactions on Emerging Topics in Computing, 11(2):474– 484, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Meloni:2022:AAP

[Mél22] Nicolas Méloni. An alternative approach to polynomial modular number system internal reduction. IEEE Transactions on Emerging Topics in Computing, 10(3): 1278–1288, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Metra:2018:MEC

[Met18] Cecilia Metra. Message from the Editor-in-Chief. IEEE Transactions on Emerging Topics in Computing, 6(1): 3–4, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Metra:2019:MEC

[Met19] Cecilia Metra. Message from the Editor-in-Chief. IEEE Transactions on Emerging Topics in Computing, 7(1): 3–4, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Metra:2020:MEC

[Met20] Cecilia Metra. Message from the Editor-in-Chief. IEEE Transactions on Emerging Topics in Computing, 8(4):885–886, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Morra:2021:GES

[MGMC21] Lia Morra, Valentina Gatteschi, Saraju P. Mohanty, and Yuan-Hao Chang. Guest editorial: Special section on new frontiers in computing for next-generation healthcare systems. IEEE Transactions on Emerging Topics in Computing, 9(3):1106– 1108, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Motaman:2019:CBC

[MGR19] Seyedhamidreza Motaman, Swaroop Ghosh, and Nitin Rathi. Cache bypassing and checkpointing to circumvent data security attacks on STTRAM. IEEE Transactions on Emerging Topics in Computing, 7(2):262– 270, April/June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Melero:2017:DIL

[MHL17] Javier Melero and Davinia Hernández-Leo. Design and implementation of locationbased learning games: Four case studies with QuesTIn-Situ: The game. IEEE Transactions on Emerging Topics in Computing, 5 (1):84–94, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Maestro-Izquierdo:2019:EVM

[MIMMY^{+}19] Marcos Maestro-Izquierdo, Javier Martin-Martinez, Albert Crespo Yepes, Manel Escudero, Rosana Rodriguez, Montserrat Nafria, Xavier Aymerich, and Antonio Rubio. Experimental verification of memristorbased material implication NAND operation. IEEE Transactions on Emerging Topics in Computing, 7(4):545–552, October/December 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mosenia:2017:CSS

[MJ17] Arsalan Mosenia and Niraj K. Jha. A comprehensive study of security of Internetof-Things. IEEE Transactions on Emerging Topics in Computing, 5(4):586–602, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mohanty:2015:GES

[MK15] Saraju P. Mohanty and Sandip Kundu. Guest editorial: Special section on circuit and system design methodologies for emerging technologies. IEEE Transactions on Emerging Topics in Computing, 3(4):456– 457, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Martinez:2022:ORU

[MKAR22] Antonio Leonel Hernández Martínez, Saqib Khursheed, Turki Alnuayri, and Daniele Rossi. Online remaining useful lifetime prediction using support vector regression. IEEE Transactions on Emerging Topics in Computing, 10(3):1546– 1557, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Majeed:2024:TOP

[MKH24] Abdul Majeed, Safiullah Khan, and Seong Oun Hwang. Towards optimization of privacy-utility trade-off using similarity and diversity based clustering. IEEE Transactions on Emerging Topics in Computing, 12(1):368–385, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ma:2015:EEE

[MLG^{+}15] Ruhui Ma, Jian Li, Haibing Guan, Mingyuan Xia, and Xue Liu. EnDAS: Efficient encrypted data search as a mobile cloud service. IEEE Transactions on Emerging Topics in Computing, 3(3): 372–383, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ma:2020:ADR

[MLH^{+}20] Jinhua Ma, Jianghua Liu, Xinyi Huang, Yang Xiang, and Wei Wu. Authenticated data redaction with fine-grained control. IEEE Transactions on Emerging Topics in Computing, 8(2): 291–302, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mehmood:2020:EAN

[MLLU20] Amjad Mehmood, Zhihan Lv, Jaime Lloret, and Muhammad Muneer Umar. ELDC: an artificial neural network based energyefficient and robust routing scheme for pollution monitoring in WSNs. IEEE Transactions on Emerging Topics in Computing, 8(1): 106–114, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Moreno-Leon:2020:TDD

[MLRRG20] Jesús Moreno-León, Gregorio Robles, and Marcos Román-González. Towards data-driven learning paths to develop computational thinking with scratch. IEEE Transactions on Emerging Topics in Computing, 8(1):193–205, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mayhew:2014:CNC

[MM14] Matthew Mayhew and Radu Muresan. On-chip nanoscale capacitor decoupling architectures for hardware security. IEEE Transactions on Emerging Topics in Computing, 2(1):4–15, March 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mozafari:2018:EPE

[MM18] Seyyed Hasan Mozafari and Brett H. Meyer. Efficient performance evaluation of multi-core SIMT processors with hot redundancy. IEEE Transactions on Emerging Topics in Computing, 6(4):498–510, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mallasen:2022:POSa

[MMD^{+}22a] David Mallasén, Raul Murillo, Alberto A. Del Barrio, Guillermo Botella, Luis Piñuel, and Manuel Prieto-Matias. PERCIVAL: Opensource posit RISC-V core with quire capability. IEEE Transactions on Emerging Topics in Computing, 10(3): 1241–1252, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic). See [MMD^{+}22b].

Mallasen:2022:POSb

[MMD^{+}22b] David Mallasén, Raul Murillo, Alberto A. Del Barrio, Guillermo Botella, Luis Piñuel, and Manuel Prieto-Matias. PERCIVAL: Opensource posit RISC-V core with quire capability. In IEEE [IEE22], page 66. ISBN 1-66547-827-6, 1-66547-828-4. LCCN ???? Authors and title only. See [MMD^{+}22a].

Mercorio:2021:DGD

[MMM^{+}21] Fabio Mercorio, Mario Mezzanzanica, Vincenzo Moscato, Antonio Picariello, and Giancarlo Sperlí. DICO: a Graph-DB framework for community detection on big scholarly data. IEEE Transactions on Emerging Topics in Computing, 9(4):1987– 2003, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Markatopoulou:2015:LFT

[MMPP15] Foteini Markatopoulou, Vasileios Mezaris, Nikiforos Pittaras, and Ioannis Patras. Local features and a two-layer stacking architecture for semantic concept detection in video. IEEE Transactions on Emerging Topics in Computing, 3(2):193–204, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mortlock:2022:GLC

[MMY^{+}22] Trier Mortlock, Deepan Muthirayan, Shih-Yuan Yu, Pramod P. Khargonekar, and Mohammad Abdullah Al Faruque. Graph learning for cognitive digital twins in manufacturing systems. IEEE Transactions on Emerging Topics in Computing, 10(1):34–45, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Montuschi:2020:TYS

[Mon20] Paolo Montuschi. Thankyou state of the journal editorial by the 2019 outgoing (acting) Editor-in-Chief. IEEE Transactions on Emerging Topics in Computing, 8(1):3, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Montuschi:2021:ENE

[Mon21] Paolo Montuschi. Editorial from the new Editor in Chief. IEEE Transactions on Emerging Topics in Computing, 9(1):4, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Montuschi:2022:TJE

[Mon22] Paolo Montuschi. TETC’s January 2022 Editor-in-Chief’s report on the state of the journal. IEEE Transactions on Emerging Topics in Computing, 10(1):4–5, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Michael:2018:GES

[MPK18] Maria K. Michael, Salvatore Pontarelli, and Omer Khan. Guest editorial: Special section on defect and fault tolerance in VLSI and nanotechnology. IEEE Transactions on Emerging Topics in Computing, 6(4):447–449, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mazumdar:2021:DAE

[MR21] Subhra Mazumdar and Sushmita Ruj. Design of anonymous endorsement system in hyperledger fabric. IEEE Transactions on Emerging Topics in Computing, 9(4):1780–1791, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mayor:2022:MTR

[MRBS22] Jesus Mayor, Laura Raya, Sofia Bayona, and Alberto Sanchez. Multitechnique redirected walking method. IEEE Transactions on Emerging Topics in Computing, 10(2):997–1008, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mayor:2021:CSV

[MRS21] Jesus Mayor, Laura Raya, and Alberto Sanchez. A comparative study of virtual reality methods of interaction and locomotion based on presence, cybersickness, and usability. IEEE Transactions on Emerging Topics in Computing, 9(3):1542– 1553, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Matthews:2019:LMS

[MS19] Suzanne J. Matthews and Aaron St. Leger. Leveraging MapReduce and synchrophasors for real-time anomaly detection in the Smart Grid. IEEE Transactions on Emerging Topics in Computing, 7(3):392–403, July/September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Metra:2020:GEI

[MS20] Cecilia Metra and Matteo Sonza Reorda. Guest Editor’s introduction: Special section on high dependability systems. IEEE Transactions on Emerging Topics in Computing, 8(2):416– 417, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mazumdar:2017:TAC

[MSAS17] Bodhisatwa Mazumdar, Samah Mohamed Saeed, Sk Subidh Ali, and Ozgur Sinanoglu. Timing attack and countermeasure on NEMS relay based design of block ciphers. IEEE Transactions on Emerging Topics in Computing, 5(3):317–328, July/September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Munir:2014:REW

[MSLL14] Sirajum Munir, John A.

Stankovic, Chieh-Jan Mike Liang, and Shan Lin. Reducing energy waste for computers by human-in-the-loop control. IEEE Transactions on Emerging Topics in Computing, 2(4):448– 460, December 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mahalakshmi:2021:EPS

[MSS21] G. S. Mahalakshmi, R. Siva, and S. Sendhilkumar. On the expressive power of scientific manuscripts. IEEE Transactions on Emerging Topics in Computing, 9(1): 269–279, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Meier:2019:GES

[MT19] Andreas Meier and Luis Terán. Guest editorial: Special section on eDemocracy and eGovernment (SIEDEG). IEEE Transactions on Emerging Topics in Computing, 7(2):204– 205, April/June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Meier:2021:GESa

[MT21a] Andreas Meier and Luis Terán. Guest editorial: Special Issue on eGovernment Development and Applications (SIEGDA). IEEE Transactions on Emerging Topics in Computing, 9(1): 117–118, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Meier:2021:GESb

[MT21b] Andreas Meier and Luis Terán. Guest editorial: Special Section on eGovernment Research, Management and Innovation (SIEGRMI). IEEE Transactions on Emerging Topics in Computing, 9(2):569– 570, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mao:2021:IRC

[MTFK21] Bomin Mao, FengXiao Tang, Zubair Md. Fadlullah, and Nei Kato. An intelligent route computation approach based on real-time deep learning strategy for software defined communication systems. IEEE Transactions on Emerging Topics in Computing, 9(3):1554–1565, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mahmoud:2022:SWB

[MVC^{+}22] Abdulqader Mahmoud, Frederic Vanderveken, Florin Ciubotaru, Christoph Adelmann, Said Hamdioui, and Sorin Cotofana. Spin wave based approximate computing. IEEE Transactions on Emerging Topics in Computing, 10(4):1932– 1940, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Milic:2021:CAM

[MVS21] Petar Milić, Nataša Veljković, and Leonid Stoimenov. Comparative analysis of metadata models on egovernment open data platforms. IEEE Transactions on Emerging Topics in Computing, 9(1):119–130, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mendis:2021:BPD

[MWW^{+}21] Gihan J. Mendis, Yifu Wu, Jin Wei, Moein Sabounchi, and Rigoberto Roche. A blockchain-powered decentralized and secure computing paradigm. IEEE Transactions on Emerging Topics in Computing, 9(4):2201– 2222, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Miele:2020:GEI

[MYM20] Antonio Miele, Qiaoyan Yu, and Maria K. Michael. Guest Editor’s introduction: Special section on reliability-aware design and analysis methods for digital systems: From gate to system level. IEEE Transactions on Emerging Topics in Computing, 8(3):561–563, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Maekawa:2017:HWC

[MYS17] Takuya Maekawa, Naomi Yamashita, and Yasushi Sakurai. How well can a user’s location privacy preferences be determined without using GPS location data? IEEE Transactions on Emerging Topics in Computing, 5(4):526–539, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Mao:2018:TRT

[MYT^{+}18] Huizi Mao, Song Yao, Tianqi Tang, Boxun Li, Jun Yao, and Yu Wang. Towards real-time object detection on embedded systems. IEEE Transactions on Emerging Topics in Computing, 6(3):417–431, July/ September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Maharjan:2017:UCD

[MZGT17] Sabita Maharjan, Yan Zhang, Stein Gjessing, and Danny H. K. Tsang. User-centric demand response management in the Smart Grid with multiple providers. IEEE Transactions on Emerging Topics in Computing, 5(4):494–505, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ma:2016:LRP

[MZY^{+}16] Xiaojing Ma, Wenjun Kevin Zeng, Laurence T. Yang, Deqing Zou, and Hai Jin. Lossless ROI privacy protection of H.264/ AVC compressed surveillance videos. IEEE Transactions on Emerging Topics in Computing, 4(3):349–362, July/September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Nittala:2017:STV

[NAM17] Ramakrishna Nittala, Andrea Acquaviva, and Enrico Macii. A software toolchain for variability awareness on heterogeneous multicore platforms. IEEE Transactions on Emerging Topics in Computing, 5(1):95–107, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Nasab:2023:HPR

[NAMJ23] Milad Tanavardi Nasab, Abdolah Amirany, Mohammad Hossein Moaiyeri, and Kian Jafari. Highperformance and robust spintronic/CNTFET-based binarized neural network hardware accelerator. IEEE Transactions on Emerging Topics in Computing, 11(2): 527–533, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Najaflou:2021:FDT

[NB21] Yashar Najaflou and Kris Bubendorfer. Forming dream teams: a chemistryoriented approach in social networks. IEEE Transactions on Emerging Topics in Computing, 9(1):204–215, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Nemati:2018:CLD

[NBRF18] Nastaran Nemati, Paul Beckett, Mark C. Reed, and Karl Fant. Clockless DFT-Less test strategy for null convention logic. IEEE Transactions on Emerging Topics in Computing, 6(4):460–473, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Nugraha:2016:AWD

[NBS16] Yudhistira Nugraha, Ian Brown, and Ashwin Sasongko Sastrosubroto. An adaptive wideband Delphi method to study state cyber-defence requirements. IEEE Transactions on Emerging Topics in Computing, 4(1):47–59, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Nogueira:2023:LSC

[NdCFB^{+}23] Mateus Nogueira, Erica da Cunha Ferreira, Pedro Tubenchlak Boechat, Felipe Assis, Estevão Rabello, Rafael Nascimento, Daniel Sadoc Menasché, Geraldo Xexéo, Abhishek Ramchandran, and Katinka Wolter. A large scale characterization of device uptimes. IEEE Transactions on Emerging Topics in Computing, 11(3):553–565, July/ September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Nepal:2019:AHL

[NHT^{+}19] Kumud Nepal, Soheil Hashemi, Hokchhay Tann, R. Iris Bahar, and Sherief Reda. Automated high-level generation of low-power approximate computing circuits. IEEE Transactions on Emerging Topics in Computing, 7(1):18–30, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Najafi:2021:HQS

[NL21] M. Hassan Najafi and David J. Lilja. High quality down-sampling for deterministic approaches to stochastic computing. IEEE Transactions on Emerging Topics in Computing, 9(1): 7–14, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Nan:2022:DPM

[NLF^{+}22] Yucen Nan, Wei Li, Lu Feng, Chengwen Luo, Jianqiang Li, and Albert Y. Zomaya. Developing practical multiview learning for clinical analytics in P4 medicine. IEEE Transactions on Emerging Topics in Computing, 10(2): 948–961, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Nogues:2019:ALA

[NMP19] Erwan Nogues, Daniel Menard, and Maxime Pelcat. Algorithmic-level approximate computing applied to energy efficient HEVC decoding. IEEE Transactions on Emerging Topics in Computing, 7(1): 5–17, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Nishad:2015:PIS

[NS15] Atul Kumar Nishad and Rohit Sharma. Performance improvement in SC-MLGNRs interconnects using interlayer dielectric insertion. IEEE Transactions on Emerging Topics in Computing, 3(4):470– 482, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Naghavi:2022:TPF

[NSH22] Amin Naghavi, Sepideh Safari, and Shaahin Hessabi. Tolerating permanent faults with low-energy overhead in multicore mixed-criticality systems. IEEE Transactions on Emerging Topics in Computing, 10(2):985– 996, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Nia:2016:PIL

[NSKRJ16] Arsalan Mohsen Nia, Susmita Sur-Kolay, Anand Raghunathan, and Niraj K. Jha. Physiological information leakage: a new frontier in health information security. IEEE Transactions on Emerging Topics in Computing, 4(3):321–334, July/September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ntalianis:2016:RAB

[NT16] Klimis Ntalianis and Nicolas Tsapatsoulis. Remote authentication via biometrics: a robust videoobject steganographic mechanism over wireless networks. IEEE Transactions on Emerging Topics in Computing, 4(1):156–174, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Nayak:2021:SSE

[NT21] Sanjeet Kumar Nayak and Somanath Tripathy. SEMKC: Secure and efficient computation over outsourced data encrypted under multiple keys. IEEE Transactions on Emerging Topics in Computing, 9(1):414–428, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Nicanfar:2013:ENC

[NTAL13] Hasen Nicanfar, Peyman Talebifard, Amr Alasaad, and Victor C. M. Leung. Enhanced network coding to maintain privacy in Smart Grid communication. IEEE Transactions on Emerging Topics in Computing, 1(2):

286–296, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Narayanan:2014:LKI

[NVS^{+}14] Iyswarya Narayanan, Arunchandar Vasan, Venkatesh Sarangan, Jamsheeda Kadengal, and Anand Sivasubramaniam. Little knowledge isn’t always dangerous understanding water distribution networks using centrality metrics. IEEE Transactions on Emerging Topics in Computing, 2(2):225–238, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Naveed:2020:AMB

[NW20] Kashif Naveed and Hui Wu.

Aster: Multi-bit soft error recovery using idempotent processing. IEEE Transactions on Emerging Topics in Computing, 8(4):928–937, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Niu:2017:TFA

[NWSG17] Jianwei Niu, Shihao Wang, Yiming Su, and Song Guo. Temporal factor-aware video affective analysis and recommendation for cyber-based social media. IEEE Transactions on Emerging Topics in Computing, 5(3):412–424, July/September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ni:2021:NTB

[NYC^{+}21] Tianming Ni, Zhao Yang, Hao Chang, Xiaoqiang Zhang, Lin Lu, Aibin Yan, Zhengfeng Huang, and Xiaoqing Wen. A novel TDMAbased fault tolerance technique for the TSVs in 3D-ICs using honeycomb topology. IEEE Transactions on Emerging Topics in Computing, 9(2):724–734, April/ June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Osorio:2022:BFA

[OAP^{+}22] John Osorio, Adrià Armejach, Eric Petit, Greg Henry, and Marc Casas. A BF16 FMA is all you need for DNN training. IEEE Transactions on Emerging Topics in Computing, 10(3):1302– 1314, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Omana:2022:RRD

[OBM22] Martin Omaña, Sejuti Bardhan, and Cecilia Metra. Reliability risks due to faults affecting selectors of ReRAMs and possible solutions. IEEE Transactions on Emerging Topics in Computing, 10(4):2086– 2091, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ota:2015:MCD

[ODCZ15] Kaoru Ota, Mianxiong Dong, Shan Chang, and Hongzi Zhu. MMCD: Cooperative downloading for highway VANETs. IEEE Transactions on Emerging Topics in Computing, 3(1): 34–43, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Oden:2021:CDS

[Ode21]  Lena Oden.  Compar-

ing data staging techniques for large scale brain images. IEEE Transactions on Emerging Topics in Computing, 9(4):1697–1708, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Omana:2018:LCS

[OEM18] M. Omaña, T. Edara, and C. Metra. Low-cost strategy to mitigate the impact of aging on latches robustness. IEEE Transactions on Emerging Topics in Computing, 6(4):488–497, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Odone:2021:GEA

[OGL^{+}21] Francesca Odone, Giuliano Grossi, Raffaella Lanzarotti, Henry Medeiros, and Nicoletta Noceti. Guest editorial assistive computing technologies for human wellbeing. IEEE Transactions on Emerging Topics in Computing, 9(3):1231–1233, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ohyama:2022:IRS

[OKK22] Takahiro Ohyama, Yuichi Kawamoto, and Nei Kato. Intelligent reflecting surface (IRS) allocation scheduling method using combinatorial optimization by quantum computing. IEEE Transactions on Emerging Topics in Computing, 10(3):1633– 1644, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ohyama:2023:QCB

[OKK23a] Takahiro Ohyama, Yuichi Kawamoto, and Nei Kato. Quantum computing based optimization for intelligent reflecting surface (IRS)- aided cell-free network. IEEE Transactions on Emerging Topics in Computing, 11 (1):18–29, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ohyama:2023:RAO

[OKK23b] Takahiro Ohyama, Yuichi Kawamoto, and Nei Kato. Resource allocation optimization by quantum computing for shared use of standalone IRS. IEEE Transactions on Emerging Topics in Computing, 11(4):950–961, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Oubbati:2021:LCU

[OLL^{+}21] Omar Sami Oubbati, Abderrahmane Lakas, Pascal Lorenz, Mohammed Atiquzzaman, and Abbas Jamalipour. Leveraging communicating UAVs for emergency vehicle guidance in urban areas. IEEE Transactions on Emerging Topics in Computing, 9(2):1070–1082, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ogasawara:2020:OBV

[OMSO20] Kazuya Ogasawara, Tomo Miyazaki, Yoshihiro Sugaya, and Shinichiro Omachi. Object-based video coding by visual saliency and temporal correlation. IEEE Transactions on Emerging Topics in Computing, 8(1): 168–178, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Onizawa:2017:SPO

[OMTH17] Naoya Onizawa, Akira Mochizuki, Akira Tamakoshi, and Takahiro Hanyu. Sudden power-outage resilient in-processor checkpointing for energy-harvesting nonvolatile processors. IEEE Transactions on Emerging Topics in Computing, 5(2): 151–163, April/June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Olague:2023:AAA

[OPIV^{+}23] Gustavo Olague, Roberto Pineda, Gerardo Ibarra-Vazquez, Matthieu Olague, Axel Martinez, Sambit Bakshi, Jonathan Vargas, and Isnardo Reducindo. Adversarial attacks assessment of salient object detection via symbolic learning. IEEE Transactions on Emerging Topics in Computing, 11(4):1018–1030, October/ December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Oberman:2022:GES

[OSPN22] Stuart Oberman, Leonel Sousa, Bogdan Pasca, and Alberto Nannarelli. Guest editorial: Special section on emerging and impacting trends on computer arithmetic. IEEE Transactions on Emerging Topics in Computing, 10(3):1239– 1240, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Omar:2013:PEV

[OZAL13] Hassan Aboubakr Omar, Weihua Zhuang, Atef Abdrabou, and Li Li. Performance evaluation of VeMAC supporting safety applications in vehicular networks. IEEE Transactions on Emerging Topics in Computing, 1(1):69–83, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Omar:2015:GPP

[OZL15] Hassan Aboubakr Omar, Weihua Zhuang, and Li Li. Gateway placement and packet routing for multihop in-vehicle Internet access. IEEE Transactions on Emerging Topics in Computing, 3(3):335–351, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Pouyan:2018:MCM

[PAR18] Peyman Pouyan, Esteve Amat, and Antonio Rubio. Memristive crossbar memory lifetime evaluation and reconfiguration strategies. IEEE Transactions on Emerging Topics in Computing, 6(2):207–218, April/ June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Patnaik:2021:MAI

[PASK21] Satwik Patnaik, Mohammed Ashraf, Ozgur Sinanoglu, and Johann Knechtel. A modern approach to IP protection and Trojan prevention: Split manufacturing for 3D ICs and obfuscation of vertical interconnects. IEEE Transactions on Emerging Topics in Computing, 9(4):1815– 1834, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Perera:2019:GEI

[PBKL19] Charith Perera, Athman Bouguettaya, Salil Kanhere, and Chi Harold Liu. Guest editorial: Introduction to the special section on sensor data computing as a service in Internet of Things. IEEE Transactions on Emerging Topics in Computing, 7(2): 311–313, April/June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Patel:2021:NQI

[PBTP21] Om Prakash Patel, Neha Bharill, Aruna Tiwari, and Mukesh Prasad. A novel quantum-inspired fuzzy based neural network for data classification. IEEE Transactions on Emerging Topics in Computing, 9(2):1031–1044, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Politou:2021:BMC

[PCAP21] Eugenia Politou, Fran Casino, Efthimios Alepis, and Constantinos Patsakis. Blockchain mutability: Challenges and proposed solutions. IEEE Transactions on Emerging Topics in Computing, 9(4):1972–1986, October/ December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Pedroli:2021:NAM

[PCG^{+}21] Elisa Pedroli, Pietro Cipresso, Luca Greci, Sara Arlati, Atieh Mahroo, Valentina Mancuso, Lorenzo Boilini, Monica Rossi, Laura Stefanelli, Karine Goulene, Marco Sacco, Marco Stramba-Badiale, Giuseppe Riva, and Andrea Gaggioli. A new application for the motor rehabilitation at home: Structure and usability of bal-app. IEEE Transactions on Emerging Topics in Computing, 9(3):1290–1300, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Pu:2019:CRE

[PCXF19] Lingjun Pu, Xu Chen, Jingdong Xu, and Xiaoming Fu. Content retrieval at the edge: a social-aware and named data cooperative framework. IEEE Transactions on Emerging Topics in Computing, 7(1):135–148, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Papadimitriou:2023:ACM

[PG23] George Papadimitriou and Dimitris Gizopoulos. Anatomy of on-chip memory hardware fault effects across the layers. IEEE Transactions on Emerging Topics in Computing, 11(2):420– 431, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

PanizoAlonso:2021:VSE

[PGM^{+}21] Luis Panizo Alonso, Mila Gascó, David Y. Marcos del Blanco, José Á. Hermida Alonso, Jordi Barrat, and Héctor Aláiz Moreton. E-voting system evaluation based on The Council of Europe recommendations: Helios voting. IEEE Transactions on Emerging Topics in Computing, 9(1):161–173, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Pendyala:2020:GLN

[PIK20] Shilpa Pendyala, Sheikh Ariful Islam, and Srinivas Katkoori. Gate level NBTI and leakage cooptimization in combinational circuits with input vector cycling. IEEE Transactions on Emerging Topics in Computing, 8(3):738–749, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Pajouh:2019:TLD

[PJK^{+}19] Hamed Haddad Pajouh, Reza Javidan, Raouf Khayami, Ali Dehghantanha, and Kim-Kwang Raymond Choo. A two-layer dimension reduction and two-tier classification model for anomalybased intrusion detection in IoT backbone networks. IEEE Transactions on Emerging Topics in Computing, 7(2):314–323, April/June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Park:2022:EER

[PKK22a] Jong Kang Park, Duckyong Kim, and Jong Tae Kim. Efficient error-resilient bus coding method using bit-basis orthogonal integrative multiplexing. IEEE Transactions on Emerging Topics in Computing, 10 (2):1178–1191, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Park:2022:AHS

[PKK22b] Sehun Park, Jae-Joon Kim, and Jaeha Kung. AutoRelax: HW-SW cooptimization for efficient SpGEMM operations with automated relaxation in deep learning. IEEE Transactions on Emerging Topics in Computing, 10(3):1428– 1442, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Piliposyan:2022:HTD

[PKR22] Gor Piliposyan, Saqib Khursheed, and Daniele Rossi. Hardware Trojan detection on a PCB through differential power monitoring. IEEE Transactions on Emerging Topics in Computing, 10(2): 740–751, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Parhi:2019:CAF

[PL19] Keshab K. Parhi and Yin Liu. Computing arithmetic functions using stochastic logic by series expansion. IEEE Transactions on Emerging Topics in Computing, 7(1):44–59, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Panneerselvam:2020:AOR

[PLA20] John Panneerselvam, Lu Liu, and Nick Antonopoulos. An approach to optimise resource provision with energy-awareness in datacentres by combating task heterogeneity. IEEE Transactions on Emerging Topics in Computing, 8(3):762–780, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Plantard:2021:EWSa

[Pla21a] Thomas Plantard. Efficient word size modular arithmetic. IEEE Transactions on Emerging Topics in Computing, 9(3):1506– 1518, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic). See [Pla21b].

Plantard:2021:EWSb

[Pla21b] Thomas Plantard. Efficient word size modular arithmetic. In IEEE [IEE21], page 139. ISBN 1-66542-293-9 (print), 1-66544-648-X (e-book). LCCN ???? See [Pla21a].

Perera:2015:EIT

[PLJ15] Charith Perera, Chi Harold Liu, and Srimal Jayawardena. The emerging Internet of Things marketplace from an industrial perspective: a survey. IEEE Transactions on Emerging Topics in Computing, 3(4):585– 598, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Piumatti:2020:RRT

[PLSM20] Giovanni Piumatti, Fabrizio Lamberti, Andrea Sanna, and Paolo Montuschi. Robust robot tracking for next-generation collaborative robotics-based gaming environments. IEEE Transactions on Emerging Topics in Computing, 8(3):869–882, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Poslad:2015:SIE

[PMC^{+}15] Stefan Poslad, Stuart E.

Middleton, Fernando Chaves, Ran Tao, Ocal Necmioglu, and Ulrich Bügel. A semantic IoT early warning system for natural environment crisis management. IEEE Transactions on Emerging Topics in Computing, 3(2): 246–257, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Prieto:2021:FPO

[PMLT21] Alvaro E. Prieto, Jose-Norberto Mazón, and Adolfo Lozano-Tello. Framework for prioritization of open data publication: an application to smart cities. IEEE Transactions on Emerging Topics in Computing, 9(1): 131–143, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Papachatzopoulos:2023:NSB

[PP23a] Kleanthis Papachatzopoulos and Vassilis Paliouras. Noise-shaping binary-to-stochastic converters for reduced-length bit-streams. IEEE Transactions on Emerging Topics in Computing, 11(4):1002–1017, October/ December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Papachatzopoulos:2023:PBD

[PP23b] Kleanthis Papachatzopoulos and Vassilis Paliouras. Path-based delay variation models for parallel-prefix adders. IEEE Transactions on Emerging Topics in Computing, 11(3):689–705, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Parra:2021:PAM

[PPF^{+}21] Pablo Parra, Óscar R. Polo, Javier Fernández, Antonio Da Silva, Sebastián Sánchez, and Agustín Martínez. A platform-aware modeldriven embedded software engineering process based on annotated analysis models. IEEE Transactions on Emerging Topics in Computing, 9(1):78–89, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Park:2023:BCB

[PPKN23] Seong-Joon Park, Hosung Park, Hee-Youl Kwak, and Jong-Seon No. BIC codes: Bit insertion-based constrained codes with error correction for DNA storage. IEEE Transactions on Emerging Topics in Computing, 11(3):764–777, July/ September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Perach:2024:UBB

[PRKK24] Ben Perach, Ronny Ronen, Benny Kimelfeld, and Shahar Kvatinsky. Understanding bulk-bitwise processing in-memory through database analytics. IEEE Transactions on Emerging Topics in Computing, 12(1): 7–22, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Pandey:2021:PAT

[PS21] Rakesh Pandey and Aryabartta Sahu. Performance and area trade-off of 3d-stacked DRAM based chip multiprocessor with hybrid interconnect. IEEE Transactions on Emerging Topics in Computing, 9(4):1945–1959, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Plantard:2022:TBN

[PSSZ22] Thomas Plantard, Arnaud Sipasseuth, Willy Susilo, and Vincent Zucca. Tight bound on NewHope failure probability. IEEE Transactions on Emerging Topics in Computing, 10(4):1955– 1965, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Pourmeidani:2021:PIR

[PSZD21] Hossein Pourmeidani, Shadi Sheikhfaal, Ramtin Zand, and Ronald F. DeMara. Probabilistic interpolation recoder for energy-error-product efficient DBNs with p-bit devices. IEEE Transactions on Emerging Topics in Computing, 9(4):2146– 2157, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Pappel:2021:LMI

[PTD21] Ingrid Pappel, Valentyna Tsap, and Dirk Draheim. The e-LocGov model for introducing e-governance into local governments: an Estonian case study. IEEE Transactions on Emerging Topics in Computing, 9(2): 597–611, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Plusquellic:2023:PPA

[PTM23] Jim Plusquellic, Eirini Eleni Tsiropoulou, and Cyrus Minwalla. Privacy-preserving authentication protocols for IoT devices using the SiRF PUF. IEEE Transactions on Emerging Topics in Computing, 11(4):918–933, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Panopoulou:2021:EAM

[PTT21] Eleni Panopoulou, Efthimios Tambouris, and Konstantinos Tarabanis. An eParticipation acceptance model. IEEE Transactions on Emerging Topics in Computing, 9(1):188–199, January/ March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Pilato:2015:TSE

[PV15] Giovanni Pilato and Giorgio Vassallo. TSVD as a statistical estimator in the latent semantic analysis paradigm. IEEE Transactions on Emerging Topics in Computing, 3(2):185–192, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Qiu:2015:DAH

[QCN^{+}15] Meikang Qiu, Zhi Chen, Jianwei Niu, Ziliang Zong, Gang Quan, Xiao Qin, and Laurence T. Yang. Data allocation for hybrid memory with genetic algorithm. IEEE Transactions on Emerging Topics in Computing, 3(4):544– 555, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Qiao:2023:THP

[QGF^{+}23] Weikang Qiao, Licheng Guo, Zhenman Fang, Mau-Chung Frank Chang, and Jason Cong. TopSort: a high-performance twophase sorting accelerator optimized on HBM-based FPGAs. IEEE Transactions on Emerging Topics in Computing, 11(2):404– 419, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Qi:2022:LLR

[QLL^{+}22] Qi Qi, Yan Lu, Jiashi Li, Jingyu Wang, Haifeng Sun, and Jianxin Liao. Learning low resource consumption CNN through pruning and quantization. IEEE Transactions on Emerging Topics in Computing, 10(2):886– 903, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Qiang:2017:UCM

[QLT17] Li Qiang, Jie Li, and Corinne Touati. A user centered multi-objective handoff scheme for hybrid 5G environments. IEEE Transactions on Emerging Topics in Computing, 5(3):380–390, July/September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Qi:2022:SDF

[QSLG22] Peng Qi, Yan Sun, Hong Luo, and Mohsen Guizani. Scratch-DKG: a framework for constructing scratch domain knowledge graph. IEEE Transactions on Emerging Topics in Computing, 10(1):170–185, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Qin:2018:DIP

[QWC^{+}18] Zhen Qin, Yilei Wang, Hongrong Cheng, Yingjie Zhou, Zhengguo Sheng, and Victor C. M. Leung. Demographic information prediction: a portrait of smartphone application users. IEEE Transactions on Emerging Topics in Computing, 6(3):432–444, July/September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Qin:2023:PAR

[QZW^{+}23] Fangyun Qin, Zheng Zheng, Xiaohui Wan, Zhihao Liu, and Zhiping Shi. Predicting aging-related bugs using network analysis on aging-related dependency networks. IEEE Transactions on Emerging Topics in Computing, 11(3):566–579, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Roy:2022:AAC

[RAD22] Avishek Sinha Roy, Hardik Agrawal, and Anindya Sundar Dhar. ACBAMaccuracy-configurable sign inclusive broken array Booth multiplier design. IEEE Transactions on Emerging Topics in Computing, 10(4):2072–2078, October/ December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rzecki:2022:AES

[RB22] Krzysztof Rzecki and Mateusz Baran. Application of elastic shape analysis to user authentication and identification. IEEE Transactions on Emerging Topics in Computing, 10 (2):1157–1165, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rojo:2023:APF

[RdPF^{+}23] Javier Rojo, Lara Guedes de Pinho, César Fonseca, Manuel José Lopes, Sumi Helal, Juan Hernández, Jose Garcia-Alonso, and Juan Manuel Murillo. Analyzing the performance of feature selection on regression problems: a case study on older adults functional profile. IEEE Transactions on Emerging Topics in Computing, 11(1):137–152, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rambo:2022:SAI

[RDS^{+}22] Eberle A. Rambo, Bryan Donyanavard, Minjun Seo, Florian Maurer, Thawra Kadeed, Caio B. de Melo, Biswadip Maity, Anmol Surhonne, Andreas Herkersdorf, Fadi Kurdahi, Nikil Dutt, and Rolf Ernst. The self-aware information processing factory paradigm for mixed-critical multiprocessing. IEEE Transactions on Emerging Topics in Computing, 10(1):250–266, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rakka:2023:DDT

[RFKK23] Mariam Rakka, Mohammed E. Fouda, Rouwaida Kanj, and Fadi Kurdahi. DT2CAM: a decision tree to content addressable memory framework. IEEE Transactions on Emerging Topics in Computing, 11(3):805–810, July/ September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rathore:2021:MGB

[RGP^{+}21] Muhammad Mazhar Ullah Rathore, Malik Junaid Jami Gul, Anand Paul, Ashraf Ali Khan, Raja Wasim Ahmad, Joel J. P. C. Rodrigues, and Spiridon Bakiras. Multilevel graph-based decision making in big scholarly data: an approach to identify expert reviewer, finding quality impact factor, ranking journals and researchers. IEEE Transactions on Emerging Topics in Computing, 9(1):280–292, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rathor:2020:NLC

[RGS20] Vijaypal Singh Rathor, Bharat Garg, and G. K. Sharma. A novel low complexity logic encryption technique for designfor-trust. IEEE Transactions on Emerging Topics in Computing, 8(3):688–699, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Riente:2016:UCT

[RHRV16] Fabrizio Riente, Izhar Hussain, Massimo Ruo Roch, and Marco Vacca. Understanding CMOS technology through TAMTAMS web. IEEE Transactions on Emerging Topics in Computing, 4(3):392–403, July/ September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Romagnoli:2023:RSS

[RKdN^{+}23] Raffaele Romagnoli, Bruce H.

Krogh, Dionisio de Niz, Anton D. Hristozov, and Bruno Sinopoli. Runtime system support for CPS software rejuvenation. IEEE Transactions on Emerging Topics in Computing, 11(3):594–604, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Reviriego:2022:CSE

[RLE^{+}22] Pedro Reviriego, Shanshan Liu, Otmar Ertl, Farzad Niknia, and Fabrizio Lombardi. Computing the similarity estimate using approximate memory. IEEE Transactions on Emerging Topics in Computing, 10(3): 1593–1604, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rodrigues:2022:ODM

[RLK22] Tiago Koketsu Rodrigues, Jiajia Liu, and Nei Kato. Offloading decision for mobile multi-access edge computing in a multi-tiered 6G network. IEEE Transactions on Emerging Topics in Computing, 10(3):1414– 1427, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ren:2015:QAT

[RLX15] Xiaojiang Ren, Weifa Liang, and Wenzheng Xu. Qualityaware target coverage in energy harvesting sensor networks. IEEE Transactions on Emerging Topics in Computing, 3(1):8– 21, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rahmani:2020:FBS

[RM20] Kamran Rahmani and Prabhat Mishra. Feature-based signal selection for postsilicon debug using machine learning. IEEE Transactions on Emerging Topics in Computing, 8(4):907–915, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Recalde:2019:WYS

[RMB^{+}19] Lorena Recalde, Jonathan Mendieta, Ludovico Boratto, Luis Terán, Carmen Vaca, and Gabriela Baquerizo. Who you should not follow: Extracting word embeddings from tweets to identify groups of interest and hijackers in demonstrations. IEEE Transactions on Emerging Topics in Computing, 7(2):206–217, April/ June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rostami:2014:RRE

[RMK^{+}14] Masoud Rostami, Mehrdad Majzoobi, Farinaz Koushanfar, Dan S. Wallach, and Srinivas Devadas. Robust and reverse-engineering resilient PUF authentication and key-exchange by substring matching. IEEE Transactions on Emerging Topics in Computing, 2(1): 37–49, March 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rout:2023:RRD

[RMSD23] Sidhartha Sankar Rout, Badri M, Mitali Sinha, and Sujay Deb. ReDeSIGN: Reuse of debug structures for improvement in performance gain of NoC based MPSoCs. IEEE Transactions on Emerging Topics in Computing, 11(2):432– 447, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rasheed:2022:EZK

[RMVN22] Amar Rasheed, Rabi N. Mahapatra, Cihan Varol, and Karpoor Narashimha. Exploiting zero knowledge proof and blockchains towards the enforcement of anonymity, data integrity and privacy (ADIP) in the IoT. IEEE Transactions on Emerging Topics in Computing, 10(3):1476– 1491, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rangarajan:2022:ODD

[RPK^{+}22] Nikhil Rangarajan, Satwik Patnaik, Johann Knechtel, Ramesh Karri, Ozgur Sinanoglu, and Shaloo Rakheja. Opening the doors to dynamic camouflaging: Harnessing the power of polymorphic devices. IEEE Transactions on Emerging Topics in Computing, 10(1):137–156, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rios:2023:BLC

[RPL^{+}23] Marco Rios, Flavio Ponzina, Alexandre Levisse, Giovanni Ansaloni, and David Atienza. Bit-line computing for CNN accelerators co-design in edge AI inference. IEEE Transactions on Emerging Topics in Computing, 11(2):358– 372, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Reviriego:2023:BSC

[RPM23] Pedro Reviriego, Salvatore Pontarelli, and Jorge Martínez. Bitwise signature comparison: Enabling more efficient similarity estimation. IEEE Transactions on Emerging Topics in Computing, 11(3):798–804, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rai:2022:SPV

[RPR^{+}22] Shubham Rai, Satwik Patnaik, Ansh Rupani, Johann Knechtel, Ozgur Sinanoglu, and Akash Kumar. Security promises and vulnerabilities in emerging reconfigurable nanotechnologybased circuits. IEEE Transactions on Emerging Topics in Computing, 10(2):763– 778, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rangel-Patino:2020:HFS

[RPVWRS^{+}20] Francisco Elias Rangel-Patiño, Andres Viveros-Wacher, José Ernesto Rayas-Sánchez, Ismael Duron-Rosales, Edgar Andrei Vega-Ochoa, Nagib Hakim, and Enrique Lopez-Miralrio. A holistic formulation for system margining and jitter tolerance optimization in industrial post-silicon validation. IEEE Transactions on Emerging Topics in Computing, 8(2):453–463, April/ June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rahman:2016:ARR

[RRFT16] MD. Tauhidur Rahman, Fahim Rahman, Domenic Forte, and Mark Tehranipoor. An aging-resistant RO-PUF for reliable key generation. IEEE Transactions on Emerging Topics in Computing, 4(3):335–348, July/September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Roussou:2020:CEI

[RS20] Maria Roussou and Mel Slater. Comparison of the effect of interactive versus passive virtual reality learning activities in evoking and sustaining conceptual change. IEEE Transactions on Emerging Topics in Computing, 8(1):233–244, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rathor:2021:LRL

[RS21] Vijaypal Singh Rathor and G. K. Sharma. A lightweight robust logic locking technique to thwart sensitization and cone-based attacks. IEEE Transactions on Emerging Topics in Computing, 9(2):811–822, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rodrigues:2013:ISS

[RSB13] André Rodrigues, Jorge Sá Silva, and Fernando Boavida. iSenior a support system for elderly citizens. IEEE Transactions on Emerging Topics in Computing, 1(2):207–217, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rohani:2021:RAM

[RSH^{+}21] Darius A. Rohani, Aaron Springer, Victoria Hollis, Jakob E. Bardram, and Steve Whittaker. Recommending activities for mental health and wellbeing: Insights from two user studies. IEEE Transactions on Emerging Topics in Computing, 9(3):1183–1193, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rodrigues:2021:ECS

[RSK21] Tiago Koketsu Rodrigues, Katsuya Suto, and Nei Kato. Edge cloud server deployment with transmission power control through machine learning for 6G Internet of Things. IEEE Transactions on Emerging Topics in Computing, 9(4):2099– 2108, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Roy:2021:EWO

[RSM^{+}21] Arindam Roy, Shamik Sural, Arun Kumar Majumdar, Jaideep Vaidya, and Vijayalakshmi Atluri. Enabling workforce optimization in constrained attribute-based access control systems. IEEE Transactions on Emerging Topics in Computing, 9(4):1901– 1913, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Rambo:2020:BGB

[RSSE20] Eberle A. Rambo, Christoph Seitz, Selma Saidi, and Rolf Ernst. Bridging the gap between resilient networkson-chip and real-time systems. IEEE Transactions on Emerging Topics in Computing, 8(2):418–430, April/ June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Riente:2022:PCR

[RTVG22] Fabrizio Riente, Giovanna Turvani, Marco Vacca, and Mariagrazia Graziano. Parallel computation in the racetrack memory. IEEE Transactions on Emerging Topics in Computing, 10 (2):1216–1221, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ren:2022:MLD

[RWD22] Zijie Ren, Jiafu Wan, and Pan Deng. Machinelearning-driven digital twin for lifecycle management of complex equipment. IEEE Transactions on Emerging Topics in Computing, 10(1): 9–22, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ruan:2023:SCO

[RWJZ23] Chang Ruan, Jianxin Wang, Wanchun Jiang, and Tao Zhang. Scheduling coflows by online identification in data center network. IEEE Transactions on Emerging Topics in Computing, 11(4):1057–1069, October/ December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ranjan:2016:AMT

[RWZ^{+}16] Rajiv Ranjan, Lizhe Wang, Albert Y. Zomaya, Jie Tao, Prem Prakash Jayaraman, and Dimitrios Georgakopoulos. Advances in methods and techniques for processing streaming Big Data in datacentre clouds. IEEE Transactions on Emerging Topics in Computing, 4(2): 262–265, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Roohi:2018:PPR

[RZAD18] Arman Roohi, Ramtin Zand, Shaahin Angizi, and Ronald F. DeMara. A parity-preserving reversible QCA gate with self-checking cascadable resiliency. IEEE Transactions on Emerging Topics in Computing, 6(4):450–459, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ren:2019:JCA

[RZD^{+}19] Ju Ren, Yaoxue Zhang, Ruilong Deng, Ning Zhang, Deyu Zhang, and Xuemin Shen. Joint channel access and sampling rate control in energy harvesting cognitive radio sensor networks. IEEE Transactions on Emerging Topics in Computing, 7(1):149–161, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Saha:2022:SSH

[SAAJ22] Tanujay Saha, Najwa Aaraj, Neel Ajjarapu, and Niraj K. Jha. SHARKS: Smart hacking approaches for RisK scanning in Internet-of-Things and cyber-physical systems based on machine learning. IEEE Transactions on Emerging Topics in Computing, 10(2):870– 885, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sheikhfaal:2023:EER

[SAD23] Shadi Sheikhfaal, Shaahin Angizi, and Ronald F. De-Mara. Energy-efficient recurrent neural network with MRAM-based probabilistic activation functions. IEEE Transactions on Emerging Topics in Computing, 11(2): 534–540, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sagar:2019:VSC

[Sag19] B. S. Daya Sagar. Variablespecific classification of zones, pairs of zones, and clusters of a spatial system via modified gravity model. IEEE Transactions on Emerging Topics in Computing, 7(2):230–241, April/June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Syarif:2019:NPC

[SAI^{+}19] Abdusy Syarif, Abdelhafid Abouaissa, Lhassane Idoumghar, Pascal Lorenz, René Schott, and G. Stacey Staples. New path centrality based on operator calculus approach for wireless sensor network deployment. IEEE Transactions on Emerging Topics in Computing, 7(1):162–173, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Saha:2022:MLA

[SAJ22] Tanujay Saha, Najwa Aaraj, and Niraj K. Jha. Machine learning assisted security analysis of 5g-network-connected systems. IEEE Transactions on Emerging Topics in Computing, 10(4):2006–2024, October/ December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sajjad:2020:CMA

[SAM^{+}20] Maryam Sajjad, Aakash Ahmad, Asad Waqar Malik, Ahmed B. Altamimi, and Ibrahim Alseadoon. Classification and mapping of adaptive security for mobile computing. IEEE Transactions on Emerging Topics in Computing, 8(3):814–832, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sedjelmaci:2021:CSF

[SBAR21] Hichem Sedjelmaci, Imane Horiya Brahmi, Nirwan Ansari, and Mubashir Husain Rehmani. Cyber security framework for vehicular network based on a hierarchical game. IEEE Transactions on Emerging Topics in Computing, 9(1):429–440, January/ March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Samiee:2021:LEA

[SBD^{+}21] Ashkan Samiee, Payal Borulkar, Ronald F. DeMara, Peiyi Zhao, and Yu Bai. Lowenergy acceleration of binarized convolutional neural networks using a spin hall effect based logic-in-memory architecture. IEEE Transactions on Emerging Topics in Computing, 9(2):928– 940, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sutradhar:2024:PLT

[SBD^{+}24] Purab Ranjan Sutradhar, Sathwika Bavikadi, Sai Manoj Pudukotai Dinakarrao, Mark A. Indovina, and Amlan Ganguly. 3DL-PIM: a look-up table oriented programmable processing in memory architecture based on the 3-D stacked memory for data-intensive applications. IEEE Transactions on Emerging Topics in Computing, 12(1):60–72, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sazeides:2022:RTE

[SBGC22] Yiannakis Sazeides, Arkady Bramnik, Ron Gabor, and Ramon Canal. A real-time error detection (RTD) architecture and its use for reliability and post-silicon validation for F/F based memory arrays. IEEE Transactions on Emerging Topics in Computing, 10(2):524– 536, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Song:2021:FKV

[SBHL21] Jie Song, Yuanguo Bi, Guangjie Han, and Tiantian Li. FacetsBase: a key-value store optimized for querying on scholarly data. IEEE Transactions on Emerging Topics in Computing, 9(1): 302–315, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Strada:2021:HBH

[SBL^{+}21] Francesco Strada, Andrea Bottino, Fabrizio Lamberti, Giulia Mormando, and Pier Luigi Ingrassia. Holo-BLSD — a holographic tool for self-training and selfevaluation of emergency response skills. IEEE Transactions on Emerging Topics in Computing, 9(3):1581–1595, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Saha:2021:STA

[SBM21] Niloy Saha, Samaresh Bera, and Sudip Misra. Sway: Traffic-aware QoS routing in software-defined IoT. IEEE Transactions on Emerging Topics in Computing, 9(1): 390–401, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sinha:2022:SAR

[SBR^{+}22] Mitali Sinha, Pramit Bhattacharyya, Sidhartha Sankar Rout, Neha Bhairavi Prakriya, and Sujay Deb. Securing an accelerator-rich system from flooding-based denialof-service attacks. IEEE Transactions on Emerging Topics in Computing, 10(2): 855–869, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sadhya:2022:QEA

[SC22] Debanjan Sadhya and Bodhi Chakraborty. Quantifying the effects of anonymization techniques over microdatabases. IEEE Transactions on Emerging Topics in Computing, 10(4):1979– 1992, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Soor:2021:IWC

[SCD^{+}21] Sampriti Soor, Aditya Challa, Sravan Danda, B. S. Daya Sagar, and Laurent Najman. Iterated watersheds, a connected variation of Kmeans for clustering GIS data. IEEE Transactions on Emerging Topics in Computing, 9(2):626–636, April/ June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

She:2013:SSD

[SCFH13] James She, Jon Crowcroft, Hao Fu, and Pin-Han Ho. Smart signage: a draggable cyber-physical broadcast/multicast media system. IEEE Transactions on Emerging Topics in Computing, 1(2):232–243, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Shu:2020:DDA

[SCS^{+}20] Lei Shu, Yuanfang Chen, Zhihong Sun, Fei Tong, and Mithun Mukherjee. Detecting the dangerous area of toxic gases with wireless sensor networks. IEEE Transactions on Emerging Topics in Computing, 8(1):137–147, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Suma:2021:SEH

[SDW^{+}21] Rossella Suma, Kurt Debattista, Derrick Watson, Elisabeth Blagrove, and Alan Chalmers. Subjective evaluation of high dynamic range imaging for face matching. IEEE Transactions on Emerging Topics in Computing, 9(4):2042–2052, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Seoane:2021:CIP

[SDZ^{+}21] Laura Seoane, Carlos Díaz, Jesús Zafra, Sergio Ibarmia, César Quintana, Carlos Pérez Canora, Andoni G. Moral, and A. Araujo. CAN implementation and performance for Raman Laser Spectrometer (RLS) instrument on Exomars 2020 Mission. IEEE Transactions on Emerging Topics in Computing, 9(1):67–77, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Shi:2017:ODV

[SFL17] Yang Shi, Hongfei Fan, and Qin Liu. An obfuscatable designated verifier signature scheme. IEEE Transactions on Emerging Topics in Computing, 5(2):271– 285, April/June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sun:2016:NIA

[SFZ^{+}16] Daniel Sun, Min Fu, Liming Zhu, Guoqiang Li, and Qinghua Lu. Nonintrusive anomaly detection with streaming performance metrics and logs for DevOps in public clouds: a case study in AWS. IEEE Transactions on Emerging Topics in Computing, 4(2):278– 289, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sheu:2015:GEI

[SG15] Phillip C.-Y. Sheu and Arif Ghafoor. Guest editorial: Introduction to the special issue on advances in semantic computing. IEEE Transactions on Emerging Topics in Computing, 3(2):148–151, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Shoukry:2020:RRM

[SG20] Laila Shoukry and Stefan Göbel. Reasons and responses: a multimodal serious games evaluation framework. IEEE Transactions on Emerging Topics in Computing, 8(1):245–255, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Siaterlis:2013:ETS

[SGH13] Christos Siaterlis, Béla Genge, and Marc Hohenadel. EPIC: a testbed for scientifically rigorous cyber-physical security experimentation. IEEE Transactions on Emerging Topics in Computing, 1(2):319– 330, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sinanoglu:2019:GEI

[SK19] Ozgur Sinanoglu and Omer Khan. Guest Editors introduction: Special section on emerging technologies in computer design. IEEE Transactions on Emerging Topics in Computing, 7(2):

242–243, April/June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Seo:2023:LLA

[SK23a] Hyoju Seo and Yongtae Kim. A low latency approximate adder design based on dual sub-adders with error recovery. IEEE Transactions on Emerging Topics in Computing, 11(3):811–816, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Shahbazi:2023:OHI

[SK23b] Karim Shahbazi and Seok-Bum Ko. An optimized hardware implementation of modular multiplication of binary ring LWE. IEEE Transactions on Emerging Topics in Computing, 11(3): 817–821, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Salehi:2020:MPV

[SKD20] Soheil Salehi, Navid Khoshavi, and Ronald F. DeMara. Mitigating process variability for non-volatile cache resilience and yield. IEEE Transactions on Emerging Topics in Computing, 8(3): 724–737, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sadeghi-Kohan:2020:SAM

[SKKN20] Somayeh Sadeghi-Kohan, Mehdi Kamal, and Zainalabedin Navabi. Selfadjusting monitor for measuring aging rate and advancement. IEEE Transactions on Emerging Topics in Computing, 8(3):627–641, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Salahvarzi:2023:WWL

[SKM^{+}23] Arash Salahvarzi, Mohsen Khosroanjam, Amir Mahdi Hosseini Monazzah, Hakem Beitollahi, Umit Y. Ogras, and Mahdi Fazeli. WiSE: When learning assists resolving STT-MRAM efficiency challenges. IEEE Transactions on Emerging Topics in Computing, 11 (1):43–55, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Shen:2020:SVD

[SLB^{+}20] Jian Shen, Dengzhi Liu, Md Zakirul Alam Bhuiyan, Jun Shen, Xingming Sun, and Aniello Castiglione. Secure verifiable database supporting efficient dynamic operations in cloud computing. IEEE Transactions on Emerging Topics in Computing, 8(2):280–290, April/ June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Shen:2013:CMA

[SLC^{+}13] Wenlong Shen, Lu Liu, Xianghui Cao, Yong Hao, and Yu Cheng. Cooperative message authentication in vehicular cyber-physical systems. IEEE Transactions on Emerging Topics in Computing, 1(1):84–97, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Shi:2017:NDS

[SLR^{+}17] Jiajun Shi, Mingyu Li, Mostafizur Rahman, Santosh Khasanvis, and Csaba Andras Moritz. NP-Dynamic skybridge: a fine-grained 3D IC technology with NP-Dynamic logic. IEEE Transactions on Emerging Topics in Computing, 5(2):286– 299, April/June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sanchez-Macian:2022:RID

[SMMTBM^{+}22] Alfonso Sánchez-Macián, Alonso Martín-Toledano, Jefferson Andres Bravo-Montes, Francisco García-Herrero, and Juan Antonio Maestro. Reducing the impact of defects in quantum-dot cellular automata (QCA) approximate adders at nano scale. IEEE Transactions on Emerging Topics in Computing, 10(2): 635–647, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Shirakawa:2015:WBS

[SNHN15] Masumi Shirakawa, Kotaro Nakayama, Takahiro Hara, and Shojiro Nishio. Wikipedia-based semantic similarity measurements for noisy short texts using Extended Naive Bayes. IEEE Transactions on Emerging Topics in Computing, 3(2): 205–219, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Suto:2014:OBD

[SNK^{+}14] Katsuya Suto, Hiroki Nishiyama, Nei Kato, Kimihiro Mizutani, Osamu Akashi, and Atsushi Takahara. An overlaybased data mining architecture tolerant to physical network disruptions. IEEE Transactions on Emerging Topics in Computing, 2(3): 292–301, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sinanoglu:2020:GEI

[SO20] Ozgur Sinanoglu and Umit Ogras. Guest Editors’ introduction: Special issue on emerging technologies in computer design. IEEE Transactions on Emerging Topics in Computing, 8(4):887–888, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sinanoglu:2021:GES

[SO21] Ozgur Sinanoglu and Umit Ogras. Guest editorial: Special issue on emerging technologies in computer design. IEEE Transactions on Emerging Topics in Computing, 9(1):5–6, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sari:2020:SAP

[SP20] Aitzan Sari and Mihalis Psarakis. Scrubbing-aware placement for reliable FPGA systems. IEEE Transactions on Emerging Topics in Computing, 8(3):564–576, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sylla:2016:DMS

[SPCB16] C. Sylla, I. S. P. Pereira, C. P. Coutinho, and P. Branco. Digital manipulatives as scaffolds for preschoolers language development. IEEE Transactions on Emerging Topics in Computing, 4(3):439–449, July/ September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Su:2020:ISC

[SQX^{+}20] Zhou Su, Qifan Qi, Qichao Xu, Song Guo, and Xiaowei Wang. Incentive scheme for cyber physical social systems based on user behaviors. IEEE Transactions on Emerging Topics in Computing, 8(1):92–103, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Singh:2014:ELL

[SR14] Nidhi Singh and Shrisha Rao. Ensemble learning for large-scale workload prediction. IEEE Transactions on Emerging Topics in Computing, 2(2):149–165, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Saleem:2022:IIR

[SSF^{+}22] Yasir Saleem, Pablo Sotres, Samuel Fricker, Carmen López de la Torre, Noel Crespi, Gyu Myoung Lee, Roberto Minerva, and Luis Sánchez. IoTRec: The IoT recommender for smart parking system. IEEE Transactions on Emerging Topics in Computing, 10(1):280–296, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Shi:2013:WAR

[SSL^{+}13] Zhiguo Shi, Ruixue Sun, Rongxing Lu, Jian Qiao, Jiming Chen, and Xuemin Shen. A wormhole attack resistant neighbor discovery scheme with RDMA protocol for 60 GHz directional network. IEEE Transactions on Emerging Topics in Computing, 1(2):341– 352, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Steinmetz:2022:KCD

[SSR^{+}22] Charles Steinmetz, Greyce N. Schroeder, Ricardo N. Rodrigues, Achim Rettberg, and Carlos E. Pereira. Key-components for digital twin modeling with granularity: Use case car-as-a-service. IEEE Transactions on Emerging Topics in Computing, 10 (1):23–33, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Scekic:2020:PMH

[SSV^{+}20] Ognjen Scekic, Tommaso Schiavinotto, Svetoslav Videnov, Michael Rovatsos, Hong-Linh Truong, Daniele Miorandi, and Schahram Dustdar. A programming model for hybrid collaborative adaptive systems. IEEE Transactions on Emerging Topics in Computing, 8(1): 6–19, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Shivashankar:2014:NVI

[SSVJ14] Shivashankar, Hosahalli Narayanagowda Suresh, Golla Varaprasad, and Guruswamy Jayanthi. Notice of violation of IEEE publication principles: Designing energy routing protocol with power consumption optimization in MANET. IEEE Transactions on Emerging Topics in Computing, 2(2):192–197, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Subramanyan:2014:RED

[STL^{+}14] Pramod Subramanyan, Nestan Tsiskaridze, Wenchao Li, Adrià Gascón, Wei Yang Tan, Ashish Tiwari, Natarajan Shankar, Sanjit A. Seshia, and Sharad Malik. Reverse engineering digital circuits using structural and functional analyses. IEEE Transactions on Emerging Topics in Computing, 2(1): 63–80, March 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Shin:2024:EGU

[STSC24] Yong-Min Shin, Cong Tran, Won-Yong Shin, and Xin Cao. Edgeless-GNN: Unsupervised representation learning for edgeless nodes. IEEE Transactions on Emerging Topics in Computing, 12(1):150–162, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Shewaga:2020:CSR

[SUQKA20] Robert Shewaga, Alvaro Uribe-Quevedo, Bill Kapralos, and Fahad Alam. A comparison of seated and room-scale virtual reality in a serious game for epidural preparation. IEEE Transactions on Emerging Topics in Computing, 8(1):218–232, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sunaga:2017:RBC

[SWK^{+}17] Yusuke Sunaga, Hironori Washizaki, Katsuhiko Kakehi, Yoshiaki Fukazawa, Shoso Yamato, and Masashi Okubo. Relation between combinations of personal characteristic types and educational effectiveness for a controlled project-based learning course. IEEE Transactions on Emerging Topics in Computing, 5(1):69–76, January/March 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Seto:2015:UHO

[SWL15] Jamie Seto, Ye Wang, and Xiaodong Lin. Userhabit-oriented authentication model: Toward secure, user-friendly authentication for mobile devices. IEEE Transactions on Emerging Topics in Computing, 3(1): 107–118, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sun:2017:RVS

[SWLG17] Yan Sun, Tin-Yu Wu, Xinming Li, and Mohsen Guizani. A rule verification system for smart buildings. IEEE Transactions on Emerging Topics in Computing, 5(3):367–379, July/ September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Shen:2020:CCG

[SWW^{+}20] Jian Shen, Anxi Wang, Chen Wang, Jiguo Li, and Yan Zhang. Contentcentric group user authentication for secure social networks. IEEE Transactions on Emerging Topics in Computing, 8(3):833–844, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Shen:2021:SVD

[SWW^{+}21] Jian Shen, Chen Wang, Anxi Wang, Sai Ji, and Yan Zhang. A searchable and verifiable data protection scheme for scholarly Big Data. IEEE Transactions on Emerging Topics in Computing, 9(1):216–225, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sun:2022:TSD

[SWY^{+}22] Jianyuan Sun, Zidong Wang, Hui Yu, Shu Zhang, Junyu Dong, and Pengxiang Gao. Two-stage deep regression enhanced depth estimation from a single RGB image. IEEE Transactions on Emerging Topics in Computing, 10(2):719– 727, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Singh:2019:MEC

[SY19] Shailendra Singh and Abdulsalam Yassine. Mining energy consumption behavior patterns for households in Smart Grid. IEEE Transactions on Emerging Topics in Computing, 7(3):404–419, July/September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Spantidi:2023:TDI

[SZA^{+}23] Ourania Spantidi, Georgios Zervakis, Sami Alsalamin, Isai Roman-Ballesteros, Jörg Henkel, Hussam Amrouch, and Iraklis Anagnostopoulos. Targeting DNN inference via efficient utilization of heterogeneous precision DNN accelerators. IEEE Transactions on Emerging Topics in Computing, 11(1): 112–125, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Sun:2020:PML

[SZL^{+}20] Xiaoqiang Sun, Peng Zhang, Joseph K. Liu, Jianping Yu, and Weixin Xie. Private machine learning classification based on fully homomorphic encryption. IEEE Transactions on Emerging Topics in Computing, 8(2):352– 364, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Song:2022:GEC

[SZL^{+}22] Tengfei Song, Wenming Zheng, Suyuan Liu, Yuan Zong, Zhen Cui, and Yang Li. Graph-embedded convolutional neural network for image-based EEG emotion recognition. IEEE Transactions on Emerging Topics in Computing, 10(3):1399– 1413, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tunali:2019:FLM

[TA19] Onur Tunali and Mustafa Altun. A fast logic mapping algorithm for multiple-type-defect tolerance in reconfigurable nano-crossbar arrays. IEEE Transactions on Emerging Topics in Computing, 7(4):518–529, October/December 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Torrens:2019:NRC

[TAC^{+}19] Gabriel Torrens, Bartomeu Alorda, Cristian Carmona, Daniel Malagón-Periánez, Jaume Segura, and Sebastia Bota. A 65-nm reliable 6T CMOS SRAM cell with minimum size transistors. IEEE Transactions on Emerging Topics in Computing, 7(3):447–455, July/September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tarihi:2022:QGS

[TAT^{+}22] Mojtaba Tarihi, Soheil Azadvar, Arash Tavakkol, Hossein Asadi, and Hamid Sarbazi-Azad. Quick generation of SSD performance models using machine learning. IEEE Transactions on Emerging Topics in Computing, 10(4):1821–1836, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tekin:2015:DEC

[TAV15] Cem Tekin, Onur Atan, and Mihaela Van Der Schaar. Discover the expert: Context-adaptive expert selection for medical diagnosis. IEEE Transactions on Emerging Topics in Computing, 3(2):220–234, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Touati:2018:SCI

[TBG^{+}18] Aymen Touati, Alberto Bosio, Patrick Girard, Arnaud Virazel, Paolo Bernardi, Matteo Sonza Reorda, and Etienne Auvray. Scanchain intra-cell aware testing. IEEE Transactions on Emerging Topics in Computing, 6(2):278–287, April/ June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Thilakanathan:2016:SCD

[TCNC16] Danan Thilakanathan, Shiping Chen, Surya Nepal, and Rafael Calvo. SafeProtect: Controlled data sharing with user-defined policies in cloud-based collaborative environment. IEEE Transactions on Emerging Topics in Computing, 4(2): 301–315, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tzitzikas:2019:PEF

[TD19] Yannis Tzitzikas and Eleftherios Dimitrakis. Preferenceenriched faceted search for voting aid applications. IEEE Transactions on Emerging Topics in Computing, 7(2):218–229, April/ June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tang:2017:HPF

[TDG17] Xifan Tang, Giovanni De Micheli, and Pierre-Emmanuel Gaillardon. A highperformance FPGA architecture using one-level RRAM-based multiplexers. IEEE Transactions on Emerging Topics in Computing, 5(2):210–222, April/ June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tsimpos:2018:LPF

[TDVS18] Andreas Tsimpos, Andreas Christos Demartinos, Spyridon Vlassis, and George Souliotis. A lowpower frequency multiplier for multi-GHz applications. IEEE Transactions on Emerging Topics in Computing, 6(2):200–206, April/ June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tsimpos:2021:MRP

[TDVS21] Andreas Tsimpos, Andreas Christos Demartinos, Spyridon Vlassis, and George Souliotis. Multirate programmable equalizer for M-PHY serial interface. IEEE Transactions on Emerging Topics in Computing, 9(1):379–389, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tran:2021:SGP

[TDZ21] Duc A. Tran, Thuy T. Do, and Ting Zhang. A stochastic geo-partitioning problem for mobile edge computing. IEEE Transactions on Emerging Topics in Computing, 9(4):2189–2200, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tavana:2020:NMP

[TFK20] Mohammad Khavari Tavana, Yunsi Fei, and David Kaeli. Nacre**Nacre, or mother-of-pearl, is one of nature’s remarkable examples of a durable and break-resistant structure: Durable, secure and energyefficient non-volatile memory utilizing data versioning. IEEE Transactions on Emerging Topics in Computing, 8(4):897–906, October/December 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tang:2019:NAU

[TFM^{+}19] Fengxiao Tang, Zubair Md. Fadlullah, Bomin Mao, Nei Kato, Fumie Ono, and Ryu Miura. On a novel adaptive UAV-Mounted cloudletaided recommendation system for LBSNs. IEEE Transactions on Emerging Topics in Computing, 7(4):565–577, October/December 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tavella:2021:DMS

[TGDC^{+}21] Federico Tavella, Alberto Giaretta, Triona Marie Dooley-Cullinane, Mauro Conti, Lee Coffey, and Sasitharan Balasubramaniam. DNA molecular storage system: Transferring digitally encoded information through bacterial nanonetworks. IEEE Transactions on Emerging Topics in Computing, 9(3):1566–1580, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tseng:2016:GES

[TH16] Yu-Chee Tseng and Sun-Yuan Hsieh. Guest editorial: Special section on emerging topics in the design of high performance Internet routers. IEEE Transactions on Emerging Topics in Computing, 4(2):177– 178, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tseng:2016:LFI

[THTK16] Yuh-Min Tseng, Sen-Shan Huang, Tung-Tso Tsai, and Jia-Hua Ke. List-free IDbased mutual authentication and key agreement protocol for multiserver architectures. IEEE Transactions on Emerging Topics in Computing, 4(1):102–112, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tsuchida:2022:IBL

[TKK^{+}22] Hikaru Tsuchida, Yuichi Kawamoto, Nei Kato, Kazuma Kaneko, Shigenori Tani, Masatake Hangai, and Hiroshi Aruga. Improvement of battery lifetime based on communication resource control in lowearth-orbit satellite constellations. IEEE Transactions on Emerging Topics in Computing, 10(3):1388– 1398, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tsigkanos:2021:GCB

[TKTP21] Antonis Tsigkanos, Nektarios Kranitis, George Theodorou, and Antonis Paschalis. A 3.3 Gbps CCSDS 123.0-B-1 multispectral & hyperspectral image compression hardware accelerator on a space-grade SRAM FPGA. IEEE Transactions on Emerging Topics in Computing, 9(1):90–103, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tsoutsos:2014:FAZ

[TM14] Nektarios Georgios Tsoutsos and Michail Maniatakos. Fabrication attacks: Zerooverhead malicious modifications enabling modern microprocessor privilege escalation. IEEE Transactions on Emerging Topics in Computing, 2(1):81– 93, March 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Thapliyal:2021:QCD

[TMCVH21] Himanshu Thapliyal, Edgard Muñoz-Coreas, T. S. S. Varun, and Travis S. Humble. Quantum circuit designs of integer division optimizing T-count and Tdepth. IEEE Transactions on Emerging Topics in Computing, 9(2):1045–1056, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tani:2019:ABC

[TMS^{+}19] Shigenori Tani, Katsuyuki Motoyoshi, Hiroyasu Sano, Atsushi Okamura, Hiroki Nishiyama, and Nei Kato. An adaptive beam control technique for Q band satellite to maximize diversity gain and mitigate interference to terrestrial networks. IEEE Transactions on Emerging Topics in Computing, 7(1):115–122, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tani:2020:FEH

[TMS^{+}20] Shigenori Tani, Katsuyuki Motoyoshi, Hiroyasu Sano, Atsushi Okamura, Hiroki Nishiyama, and Nei Kato. Flexibility-enhanced HTS system for disaster management: Responding to communication demand explosion in a disaster. IEEE Transactions on Emerging Topics in Computing, 8(1): 159–167, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Takaishi:2014:TEE

[TNKM14] Daisuke Takaishi, Hiroki Nishiyama, Nei Kato, and Ryu Miura. Toward energy efficient Big Data gathering in densely distributed sensor networks. IEEE Transactions on Emerging Topics in Computing, 2(3):388– 397, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tsikoudis:2016:LLL

[TPM16] Nikos Tsikoudis, Antonis Papadogiannakis, and Evangelos P. Markatos. LEoNIDS: a low-latency and energy-efficient networklevel intrusion detection system. IEEE Transactions on Emerging Topics in Computing, 4(1):142–155, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Teji:2024:GET

[TRD^{+}24] Binon Teji, Swarup Roy, Devendra Singh Dhami, Dinabandhu Bhandari, and Pietro Hiram Guzzi. Graph embedding techniques for predicting missing links in biological networks: an empirical evaluation. IEEE Transactions on Emerging Topics in Computing, 12(1): 190–201, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tolosana:2022:CCI

[TRGVR^{+}22] Ruben Tolosana, Juan Carlos Ruiz-Garcia, Ruben Vera-Rodriguez, Jaime Herreros-Rodriguez, Sergio Romero-Tapiador, Aythami Morales, and Julian Fierrez. Child– computer interaction with mobile devices: Recent works, new dataset, and age detection. IEEE Transactions on Emerging Topics in Computing, 10(4):2042– 2054, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tahir:2019:NSL

[TRR^{+}19] Shahzaib Tahir, Sushmita Ruj, Yogachandran Rahulamathavan, Muttukrishnan Rajarajan, and Cornelius Glackin. A new secure and lightweight searchable encryption scheme over encrypted cloud data. IEEE Transactions on Emerging Topics in Computing, 7(4):530–544, October/December 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tchendjou:2021:DLC

[TS21] Ghislain Takam Tchendjou and Emmanuel Simeu. Detection, location and concealment of defective pixels in image sensors. IEEE Transactions on Emerging Topics in Computing, 9(2): 664–679, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tan:2017:DSE

[TSH^{+}17] Song Tan, Wenzhan Song, Dan Huang, Qifen Dong, and Lang Tong. Distributed software emulator for cyberphysical analysis in Smart Grid. IEEE Transactions on Emerging Topics in Computing, 5(4):506–517, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Torra:2016:DPO

[TSS16] Vicenç Torra, Termeh Shafie, and Julián Salas. Data protection for online social networks and P -stability for graphs. IEEE Transactions on Emerging Topics in Computing, 4(3):374–381, July/September 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Tong:2022:AIM

[TW22] Guangmo Tong and Ruiqi Wang. On adaptive influence maximization under general feedback models. IEEE Transactions on Emerging Topics in Computing, 10(1):463–475, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Talafy:2021:HPM

[TZZB21] Javad Talafy, Farzaneh Zokaee, Hamid R. Zarandi, and Nader Bagherzadeh. A high performance, multibit output logic-in-memory adder. IEEE Transactions on Emerging Topics in Computing, 9(4):2223– 2233, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Urgese:2018:ONT

[UBMA18] Gianvito Urgese, Francesco Barchi, Enrico Macii, and Andrea Acquaviva. Optimizing network traffic for spiking neural network simulations on densely interconnected many-core neuromorphic platforms. IEEE Transactions on Emerging Topics in Computing, 6(3):317–329, July/September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Voicu:2017:HPC

[VC17] George Razvan Voicu and Sorin Dan Cotofana. Highperformance, cost-effective 3D stacked wide-operand adders. IEEE Transactions on Emerging Topics in Computing, 5(2):179–192, April/June 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Vitali:2021:EMC

[VGP^{+}21] Emanuele Vitali, Davide Gadioli, Gianluca Palermo, Martin Golasowski, João Bispo, Pedro Pinto, Jan Martinovič, Kateřina Slaninová, João M. P. Cardoso, and Cristina Silvano. An efficient Monte Carlobased probabilistic timedependent routing calculation targeting a server-side car navigation system. IEEE Transactions on Emerging Topics in Computing, 9 (2):1006–1019, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Vogel-Heuser:2022:BEF

[VHFH^{+}22] Birgit Vogel-Heuser, Juliane Fischer, Dieter Hess, Eva-Maria Neumann, and Marcus Würr. Boosting extrafunctional code reusability in cyber-physical production systems: The error handling case study. IEEE Transactions on Emerging Topics in Computing, 10(1):60–73, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Vimalajeewa:2022:SBJ

[VKBB22] Dixon Vimalajeewa, Chamil Kulatunga, Donagh P. Berry, and Sasitharan Balasubramaniam. A servicebased joint model used for distributed learning: Application for smart agriculture. IEEE Transactions on Emerging Topics in Computing, 10(2):838– 854, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Vatajelu:2019:CSE

[VPTH19] Elena Ioana Vatajelu, Paolo Prinetto, Mottaqiallah Taouil, and Said Hamdioui. Challenges and solutions in emerging memory testing. IEEE Transactions on Emerging Topics in Computing, 7(3):493–506, July/ September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Veeranna:2017:HTD

[VS17] Nandeesha Veeranna and Benjamin Carrion Schafer. Hardware Trojan detection in behavioral intellectual properties (IP’s) using property checking techniques. IEEE Transactions on Emerging Topics in Computing, 5(4):576–585, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Venugopalan:2023:AVF

[VSH23] Sarad Venugopalan, Ivana Stančoková, and Ivan Homoliak. Always on voting: a framework for repetitive voting on the blockchain. IEEE Transactions on Emerging Topics in Computing, 11(4):1082– 1092, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Vourkas:2018:MPA

[VSS18] Ioannis Vourkas, Dimitrios Stathis, and Georgios Ch. Sirakoulis. Massively parallel analog computing: Ariadne’s Thread was made of memristors. IEEE Transactions on Emerging Topics in Computing, 6(1):145–155, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2022:BBE

[Wan22] Zhiwei Wang. Blockchainbased edge computing data storage protocol under simplified group signature. IEEE Transactions on Emerging Topics in Computing, 10(2):1009–1019, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wu:2022:EEA

[WASW22] Yun Wu, Andreas Aßmann, Brian D. Stewart, and Andrew M. Wallace. Energy efficient approximate 3D image reconstruction. IEEE Transactions on Emerging Topics in Computing, 10(4):1854–1866, October/ December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2022:MPA

[WC22] Tinghui Wang and Diane J. Cook. Multi-person activity recognition in continuously monitored smart homes. IEEE Transactions on Emerging Topics in Computing, 10(2):1130– 1141, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2020:HPC

[WCC^{+}20] Gai-Ge Wang, Xingjuan Cai, Zhihua Cui, Geyong Min, and Jinjun Chen. High performance computing for cyber physical social systems by using evolutionary multi-objective optimization algorithm. IEEE Transactions on Emerging Topics in Computing, 8(1):20–30, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2023:IPH

[WCF23] Tinghui Wang, Diane J.

Cook, and Thomas R. Fischer. The indoor predictability of human mobility: Estimating mobility with smart home sensors. IEEE Transactions on Emerging Topics in Computing, 11(1):182–193, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2021:OSP

[WCG^{+}21] Jian Wang, Zhe Chen, Shize Guo, Yubai Li, and Zhonghai Lu. Optimal sprinting pattern in thermal constrained CMPs. IEEE Transactions on Emerging Topics in Computing, 9(1): 484–495, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wu:2019:FFC

[WDO^{+}19] Jun Wu, Mianxiong Dong, Kaoru Ota, Jianhua Li, and Zhitao Guan. FCSS: Fogcomputing-based contentaware filtering for security services in informationcentric social networks. IEEE Transactions on Emerging Topics in Computing, 7(4):553–564, October/December 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wachter:2018:HDF

[WFB^{+}18] Eduardo W. Wächter, Vinicius Fochi, Francisco Barreto, Alexandre M. Amory, and Fernando G. Moraes. A hierarchical and distributed fault tolerant proposal for NoC-based MPSoCs. IEEE Transactions on Emerging Topics in Computing, 6(4):524–537, October/December 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2020:EPB

[WGYL20] Qianlong Wang, Yifan Guo, Lixing Yu, and Pan Li. Earthquake prediction based on spatio-temporal data mining: an LSTM network approach. IEEE Transactions on Emerging Topics in Computing, 8(1): 148–158, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Waidyasooriya:2021:HPF

[WH21] Hasitha Muthumala Waidyasooriya and Masanori Hariyama. Highly-parallel FPGA accelerator for simulated quantum annealing. IEEE Transactions on Emerging Topics in Computing, 9(4):2019– 2029, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2016:DAA

[WHC16] Hung-Hsiang Wang, Hsin-An Hou, and Jyh-Cheng Chen. Design and analysis of an antenna control mechanism for time division duplexing distributed antenna systems over high-speed rail communications. IEEE Transactions on Emerging Topics in Computing, 4(4):516–527, October/December 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2023:QBA

[WHL23a] Huiwei Wang, Yaqian Huang, and Huaqing Li. Quadtreebased adaptive spatial decomposition for range queries under local differential privacy. IEEE Transactions on Emerging Topics in Computing, 11(4):1045–1056, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wu:2023:GIL

[WHL23b] Di Wu, Yi He, and Xin Luo. A graph-incorporated latent factor analysis model for high-dimensional and sparse data. IEEE Transactions on Emerging Topics in Computing, 11(4):907–917, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2015:MBA

[WJ15] Wanyuan Wang and Yichuan Jiang. Multiagent-based allocation of complex tasks in social networks. IEEE Transactions on Emerging Topics in Computing, 3(4): 571–584, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2019:WPA

[WJ19] Xiaofeng Wang and Jiulei Jiang. Warning propagation algorithm for the MAX-3-SAT problem. IEEE Transactions on Emerging Topics in Computing, 7(4):578–584, October/December 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Won:2023:PGP

[WJL23] Dong-Ok Won, Yong-Nam Jang, and Seong-Whan Lee. PlausMal-GAN: Plausible malware training based on generative adversarial networks for analogous zero-day malware detection. IEEE Transactions on Emerging Topics in Computing, 11(1):82–94, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2014:SIC

[WK14] Lei Wang and Krishna Kant.

Special issue on computational sustainability. IEEE Transactions on Emerging Topics in Computing, 2(2): 119–121, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2014:EAD

[WLC^{+}14] Yan Wang, Kenli Li, Hao Chen, Ligang He, and Keqin Li. Energy-aware data allocation and task scheduling on heterogeneous multiprocessor systems with time constraints. IEEE Transactions on Emerging Topics in Computing, 2(2):134–148, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2021:AMW

[WLG^{+}21] Shaodi Wang, Hochul Lee, Cecile Grezes, Pedram Khalili Amiri, Kang L. Wang, and Puneet Gupta. Adaptive MRAM write and read with MTJ variation monitor. IEEE Transactions on Emerging Topics in Computing, 9(1):402–413, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wei:2021:SSD

[WLL^{+}21] Xiaohui Wei, Zijian Li, Yuanyuan Liu, Shang Gao, and Hengshan Yue. SDLSC-TA: Subarea division learning based task allocation in sparse mobile crowdsensing. IEEE Transactions on Emerging Topics in Computing, 9(3):1344–1358, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2022:HID

[WLL^{+}22] Jilong Wang, Rui Li, Renfa Li, Bin Fu, and Danny Z. Chen. HMCKRAutoEncoder: an interpretable deep learning framework for time series analysis. IEEE Transactions on Emerging Topics in Computing, 10(1):99–111, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wan:2021:TBA

[WLO^{+}21] Jie Wan, MingSong Li, Michael J. O’Grady, Xiang Gu, Munassar A. A. H. Alawlaqi, and Gregory M. P. O’Hare. Time-bounded activity recognition for ambient assisted living. IEEE Transactions on Emerging Topics in Computing, 9(1): 471–483, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2022:SDF

[WLWQ22] Tianshun Wang, Yang Li, Yuan Wu, and Tony Q. S. Quek. Secrecy driven federated learning via cooperative jamming: an approach of latency minimization. IEEE Transactions on Emerging Topics in Computing, 10(4):1687– 1703, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2016:DSP

[WLX^{+}16] Dan Wang, Jiangchuan Liu, Jianliang Xu, Hongbo Jiang, and Chonggang Wang. Data sweeper: a proactive filtering framework for errorbounded sensor data collection. IEEE Transactions on Emerging Topics in Computing, 4(4):487–501, October/December 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wei:2021:DPB

[WLY^{+}21] Jianhao Wei, Yaping Lin, Xin Yao, Jin Zhang, and Xinbo Liu. Differential privacy-based genetic matching in personalized medicine. IEEE Transactions on Emerging Topics in Computing, 9(3):1109–1125, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wen:2013:PPP

[WLZ^{+}13] Mi Wen, Rongxing Lu, Kuan Zhang, Jingsheng Lei, Xiaohui Liang, and Xuemin Shen. PaRQ: a privacy-preserving range query scheme over encrypted metering data for Smart Grid. IEEE Transactions on Emerging Topics in Computing, 1(1):178–191, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wei:2022:MCH

[WLZG22] Tingyang Wei, Wei-Li Liu, Jinghui Zhong, and Yue-Jiao Gong. Multiclass classification on high dimension and low sample size data using genetic programming. IEEE Transactions on Emerging Topics in Computing, 10(2):704– 718, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Watson:2017:NHA

[WMAB17] David Watson, Gordon Morison, Ali Ahmadinia, and Tom Buggy. A novel hardware accelerator for embedded object detection applications. IEEE Transactions on Emerging Topics in Computing, 5(4):551–562, October/December 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2013:OAO

[WMN13] Yu Wang, Shiwen Mao, and R. M. Nelms. Online algorithm for optimal real-time energy distribution in the Smart Grid. IEEE Transactions on Emerging Topics in Computing, 1(1):10–21, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2024:SCA

[WMP^{+}24] Ziyu Wang, Fan-Hsuan Meng, Yongmo Park, Jason K. Eshraghian, and Wei D. Lu. Side-channel attack analysis on inmemory computing architectures. IEEE Transactions on Emerging Topics in Computing, 12(1):109–121, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2022:IPL

[WQG^{+}22] Haozhao Wang, Zhihao Qu, Song Guo, Xin Gao, Ruixuan Li, and Baoliu Ye. Intermittent pulling with local compensation for communication-efficient distributed learning. IEEE Transactions on Emerging Topics in Computing, 10(2): 779–791, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wu:2021:DFM

[WRT^{+}21] Lizhou Wu, Siddharth Rao, Mottaqiallah Taouil, Guilherme Cardoso Medeiros, Moritz Fieback, Erik Jan Marinissen, Gouri Sankar Kar, and Said Hamdioui. Defect and fault modeling framework for STT-MRAM testing. IEEE Transactions on Emerging Topics in Computing, 9(2):707–723, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2021:BTB

[WSLL21] Chen Wang, Jian Shen, Jin-Feng Lai, and Jianwei Liu. B-TSCA: Blockchain assisted trustworthiness scalable computation for V2I authentication in VANETs. IEEE Transactions on Emerging Topics in Computing, 9(3):1386–1396, July/ September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Watkins:2020:RHL

[WT20] Adam Watkins and Spyros Tragoudas. Radiation hardened latch designs for double and triple node upsets. IEEE Transactions on Emerging Topics in Computing, 8(3):616–626, July/ September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wen:2015:CLS

[WTW^{+}15] Hong Wen, Jie Tang, Jinsong Wu, Huanhuan Song, Tingyong Wu, Bin Wu, Pin-Han Ho, Shi-Chao Lv, and Li-Min Sun. A crosslayer secure communication model based on Discrete Fractional Fourier Fransform (DFRFT). IEEE Transactions on Emerging Topics in Computing, 3(1): 119–126, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wu:2021:PMA

[WVC21] Yan Wu, Srinivasan Venkatramanan, and Dah Ming Chiu. A population model for academia: Case study of the computer science community using DBLP bibliography 1960–2016. IEEE Transactions on Emerging Topics in Computing, 9(1): 258–268, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2022:DCP

[WWG^{+}22] Yale Wang, Chenghua Wang, Chongyan Gu, Yijun Cui, Máire O’Neill, and Weiqiang Liu. A dynamically configurable PUF and dynamic matching authentication protocol. IEEE Transactions on Emerging Topics in Computing, 10 (2):1091–1104, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Waris:2022:HPP

[WWL^{+}22] Haroon Waris, Chenghua Wang, Weiqiang Liu, Jie Han, and Fabrizio Lombardi. Hybrid partial product-based high-performance approximate recursive multipliers. IEEE Transactions on Emerging Topics in Computing, 10(1):507–513, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Waris:2022:AAR

[WWXL22] Haroon Waris, Chenghua Wang, Chenyu Xu, and Weiqiang Liu. AxRMs: Approximate recursive multipliers using high-performance building blocks. IEEE Transactions on Emerging Topics in Computing, 10 (2):1229–1235, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wei:2023:VDC

[WX23] Xiangye Wei and Liming Xiu. A VLSI digital circuit platform for performing deterministic stochastic computing in the time dimension using fraction operations on rational numbers. IEEE Transactions on Emerging Topics in Computing, 11(1):194–207, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2024:PDP

[WXT^{+}24] Fangwei Wang, Meiyun Xie, Zhiyuan Tan, Qingru Li, and Changguang Wang. Preserving differential privacy in deep learning based on feature relevance region segmentation. IEEE Transactions on Emerging Topics in Computing, 12(1):307– 315, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2016:PDI

[WZH^{+}16] Shangguang Wang, Ao Zhou, Ching-Hsien Hsu, Xuanyu Xiao, and Fangchun Yang. Provision of data-intensive services through energy- and QoS-aware virtual machine placement in national cloud data centers. IEEE Transactions on Emerging Topics in Computing, 4(2):290– 300, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wei:2021:BTF

[WZL21] Wenqi Wei, Qi Zhang, and Ling Liu. Bitcoin transaction forecasting with deep network representation learning. IEEE Transactions on Emerging Topics in Computing, 9(3):1359–1371, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2021:SRA

[WZLK21] Jiadai Wang, Lei Zhao, Jiajia Liu, and Nei Kato. Smart resource allocation for mobile edge computing: a deep reinforcement learning approach. IEEE Transactions on Emerging Topics in Computing, 9(3):1529–1541, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2014:ESC

[WZR^{+}14] Lizhe Wang, Hui Zhong, Rajiv Ranjan, Albert Zomaya, and Peng Liu. Estimating the statistical characteristics of remote sensing Big Data in the wavelet transform domain. IEEE Transactions on Emerging Topics in Computing, 2(3):324– 337, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wu:2023:EEC

[WZR^{+}23] Bi Wu, Haonan Zhu, Dayane Reis, Zhaohao Wang, Ying Wang, Ke Chen, Weiqiang Liu, Fabrizio Lombardi, and Xiaobo Sharon Hu. An energy-efficient computingin-memory (CiM) scheme using field-free spin-orbit torque (SOT) magnetic RAMs. IEEE Transactions on Emerging Topics in Computing, 11(2):331– 342, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2013:PAO

[WZRR13] Cong Wang, Bingsheng Zhang, Kui Ren, and Janet M. Roveda. Privacyassured outsourcing of image reconstruction service in cloud. IEEE Transactions on Emerging Topics in Computing, 1(1):166–177, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Wang:2020:SCC

[WZY^{+}20] Shangguang Wang, Ao Zhou, Mingzhe Yang, Lei Sun, Ching-Hsien Hsu, and Fangchun Yang. Service composition in cyber-physical-social systems. IEEE Transactions on Emerging Topics in Computing, 8(1):82–91, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xu:2022:AMB

[XAQ22] Qian Xu, Md Tanvir Arafin, and Gang Qu. An approximate memory based defense against model inversion attacks to neural networks. IEEE Transactions on Emerging Topics in Computing, 10(4):1733– 1745, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xia:2014:MRW

[XCW^{+}14] Feng Xia, Zhen Chen, Wei Wang, Jing Li, and Laurence T. Yang. MVCWalker: Random walk-based most valuable collaborators recommendation exploiting academic factors. IEEE Transactions on Emerging Topics in Computing, 2(3): 364–375, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xiao:2020:SSE

[XCX^{+}20] Fu Xiao, Jing Chen, Xiaohui Xie, Linqing Gui, Lijuan Sun, and Ruchuan Wang. SEARE: a system for exercise activity recognition and quality evaluation based on green sensing. IEEE Transactions on Emerging Topics in Computing, 8(3):752–761, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xia:2024:CAN

[XCY^{+}24] Feng Xia, Xin Chen, Shuo Yu, Mingliang Hou, Mujie Liu, and Linlin You. Coupled attention networks for multivariate time series anomaly detection. IEEE Transactions on Emerging Topics in Computing, 12(1): 240–253, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xia:2021:GES

[XGLW21] Feng Xia, C. Lee Giles, Huan Liu, and Kuansan Wang. Guest editorial:

Scholarly Big Data. IEEE Transactions on Emerging Topics in Computing, 9(1): 200–203, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xie:2022:EHI

[XHWI22] Jiafeng Xie, Pengzhou He, Xiaofang Wang, and José L. Imaña. Efficient hardware implementation of finite field arithmetic AB + CAB + C for binary ring-LWE based post-quantum cryptography. IEEE Transactions on Emerging Topics in Computing, 10(2):1222– 1228, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xu:2019:SRR

[XL19] Ming Xu and Ling Liu.

Sender-receiver role-based energy-aware scheduling for Internet of Underwater Things. IEEE Transactions on Emerging Topics in Computing, 7(2):324–336, April/June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xiao:2022:RAR

[XLQ^{+}22] Yunpeng Xiao, Wen Li, Shuai Qiang, Qian Li, Hanchun Xiao, and Yanbing Liu. A rumor & anti-rumor propagation model based on data enhancement and evolutionary game. IEEE Transactions on Emerging Topics in Computing, 10(2): 690–703, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xu:2020:FND

[XOD20] Jianwen Xu, Kaoru Ota, and Mianxiong Dong. Fast networking for disaster recovery. IEEE Transactions on Emerging Topics in Computing, 8(3):845–854, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xu:2021:EAE

[XSY^{+}21] Lei Xu, Shifeng Sun, Xingliang Yuan, Joseph K. Liu, Cong Zuo, and Chungen Xu. Enabling authorized encrypted search for multi-authority medical databases. IEEE Transactions on Emerging Topics in Computing, 9(1):534–546, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xiao:2020:VDU

[XSYW20] Fu Xiao, Le-Tian Sha, Zai-Ping Yuan, and Ru-Chuan Wang. VulHunter: a discovery for unknown bugs based on analysis for known patches in Industry Internet of Things. IEEE Transactions on Emerging Topics in Computing, 8(2):267– 279, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xu:2015:EID

[XSZ^{+}15] Qichao Xu, Zhou Su, Kuan Zhang, Pinyi Ren, and Xuemin Sherman Shen. Epidemic information dissemination in mobile social networks with opportunistic links. IEEE Transactions on Emerging Topics in Computing, 3(3):399– 409, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xiao:2016:QRU

[XTXY16] Ying Xiao, Krishnaiyan Thulasiraman, Guoliang Xue, and Mamta Yadav. QoS routing under multiple additive constraints: a generalization of the LARAC algorithm. IEEE Transactions on Emerging Topics in Computing, 4(2):242– 251, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xue:2023:AAD

[XWZ^{+}23] Mingfu Xue, Zhiyu Wu, Yushu Zhang, Jian Wang, and Weiqiang Liu. Adv-Params: an active DNN intellectual property protection technique via adversarial perturbation based parameter encryption. IEEE Transactions on Emerging Topics in Computing, 11(3): 664–678, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xia:2022:FDI

[XYDJ22] Wenhan Xia, Hongxu Yin, Xiaoliang Dai, and Niraj K. Jha. Fully dynamic inference with deep neural networks. IEEE Transactions on Emerging Topics in Computing, 10(2):962– 972, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xue:2023:USS

[XYH^{+}23] Mingfu Xue, Chengxiang Yuan, Can He, Yinghao Wu, Zhiyu Wu, Yushu Zhang, Zhe Liu, and Weiqiang Liu. Use the spear as a shield: an adversarial example based privacy-preserving technique against membership inference attacks. IEEE Transactions on Emerging Topics in Computing, 11(1): 153–169, January/March 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xu:2017:SSB

[XYZ^{+}17] Zheng Xu, Neil Y. Yen, Hui Zhang, Xiao Wei, Zhihan Lv, Kim-Kwang Raymond Choo, Lin Mei, and Xiangfeng Luo. Social sensors based online attention computing of public safety events. IEEE Transactions on Emerging Topics in Computing, 5(3):403–411, July/September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xu:2021:BED

[XZW^{+}21a] Yang Xu, Cheng Zhang, Guojun Wang, Zheng Qin, and Quanrun Zeng. A blockchain-enabled deduplicatable data auditing mechanism for network storage services. IEEE Transactions on Emerging Topics in Computing, 9(3):1421–1432, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Xue:2021:JDE

[XZW21b] Qiao Xue, Youwen Zhu, and Jian Wang. Joint distribution estimation and Naïve Bayes classification under local differential privacy. IEEE Transactions on Emerging Topics in Computing, 9(4):2053–2063, October/December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yang:2018:DTM

[YATR18] Jinseok Yang, Alper Sinan Akyurek, Sameer Tilak, and Tajana Simunic Rosing. Design of transmission manager in heterogeneous WSNs. IEEE Transactions on Emerging Topics in Computing, 6(3):395–408, July/September 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yan:2023:TDN

[YCH^{+}23] Aibin Yan, Aoran Cao, Zhengfeng Huang, Jie Cui, Tianming Ni, Patrick Girard, Xiaoqing Wen, and Jiliang Zhang. Two doublenode-upset-hardened flipflop designs for high-performance applications. IEEE Transactions on Emerging Topics in Computing, 11(4):1070– 1081, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yin:2022:TEE

[YCL^{+}22] Hongxu Yin, Guoyang Chen, Yingmin Li, Shuai Che, Weifeng Zhang, and Niraj K. Jha. Towards execution-efficient LSTMs via hardware-guided growand-prune paradigm. IEEE Transactions on Emerging Topics in Computing, 10(4):1799–1809, October/ December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yu:2014:DQA

[YCLW14] Yang Yu, Jian Chen, Shangquan Lin, and Ying Wang. A dynamic QoSaware logistics service composition algorithm based on social network. IEEE Transactions on Emerging Topics in Computing, 2(4):399– 410, December 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yadav:2022:BDA

[YCS22] Akash Yadav, Joydeep Chandra, and Ashok Singh Sairam. A budget and deadline aware task assignment scheme for crowdsourcing environment. IEEE Transactions on Emerging Topics in Computing, 10 (2):1020–1034, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yilmaz:2021:AAC

[YDH21] Yildiran Yilmaz, Viet-Hoa Do, and Basel Halak. AR-MOR: an anti-counterfeit security mechanism for low cost radio frequency identification systems. IEEE Transactions on Emerging Topics in Computing, 9(4):2125–2138, October/ December 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yu:2018:DOR

[YDM^{+}18] Rong Yu, Jiefei Ding, Sabita Maharjan, Stein Gjessing, Yan Zhang, and Danny H. K. Tsang. Decentralized and optimal resource cooperation in geodistributed mobile cloud computing. IEEE Transactions on Emerging Topics in Computing, 6(1):72–84, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yu:2022:DLB

[YGLO22] Shichao Yu, Chongyan Gu, Weiqiang Liu, and Máire O’Neill. Deep learningbased hardware Trojan detection with block-based netlist information extraction. IEEE Transactions on Emerging Topics in Computing, 10(4):1837–1853, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yu:2022:SSA

[YJ22] Ye Yu and Niraj K. Jha. SPRING: a sparsity-aware reduced-precision monolithic 3D CNN accelerator architecture for training and inference. IEEE Transactions on Emerging Topics in Computing, 10(1):237–249, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yu:2018:EVR

[YK18] Weize Yu and Selçuk Köse. Exploiting voltage regulators to enhance various power attack countermeasures. IEEE Transactions on Emerging Topics in Computing, 6(2):244–257, April/June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yeganeh-Khaksar:2022:RRM

[YKAE22]  Amir  Yeganeh-Khaksar,

Mohsen Ansari, and Alireza Ejlali. ReMap: Reliability management of peak-power-aware real-time embedded systems through task replication. IEEE Transactions on Emerging Topics in Computing, 10(1):312–323, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yoo:2022:FCF

[YKP22] Eungeun Yoo, Haneul Ko, and Sangheon Pack. Fuzzy clustered federated learning algorithm for solar power generation forecasting. IEEE Transactions on Emerging Topics in Computing, 10(4):2092–2098, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Ye:2020:ADM

[YKW^{+}20] Jiaxing Ye, Takumi Kobayashi, Xiaoyan Wang, Hiroshi Tsuda, and Masahiro Murakawa. Audio data mining for anthropogenic disaster identification: an automatic taxonomy approach. IEEE Transactions on Emerging Topics in Computing, 8(1): 126–136, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yan:2023:SEE

[YLW23] Zihui Yan, Cong Liang, and Huaming Wu. A segmentededit error-correcting code with re-synchronization function for DNA-based storage systems. IEEE Transactions on Emerging Topics in Computing, 11(3):605–618, July/September 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yu:2020:ABC

[YLY^{+}20] Yong Yu, Yannan Li, Bo Yang, Willy Susilo, Guomin Yang, and Jian Bai. Attribute-based cloud data integrity auditing for secure outsourced storage. IEEE Transactions on Emerging Topics in Computing, 8(2): 377–390, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yan:2021:NLC

[YLZ^{+}21] Aibin Yan, Chaoping Lai, Yinlei Zhang, Jie Cui, Zhengfeng Huang, Jie Song, Jing Guo, and Xiaoqing Wen. Novel low cost, double-and-triple-node-upset-tolerant latch designs for nano-scale CMOS. IEEE Transactions on Emerging Topics in Computing, 9(1): 520–533, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yin:2021:DPD

[YMDJ21] Hongxu Yin, Bilal Mukadam, Xiaoliang Dai, and Niraj K. Jha. DiabDeep: Pervasive diabetes diagnosis based on wearable medical sensors and efficient neural networks. IEEE Transactions on Emerging Topics in Computing, 9(3):1139–1150, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yasin:2020:RAL

[YMSR20] Muhammad Yasin, Bodhisatwa Mazumdar, Ozgur Sinanoglu, and Jeyavijayan Rajendran. Removal attacks on logic locking and camouflaging techniques. IEEE Transactions on Emerging Topics in Computing, 8(2): 517–532, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yang:2022:WSC

[YMT22] Songxiang Yang, Lin Ma, and Xuezhi Tan. Weakly supervised class-agnostic image similarity search based on convolutional neural network. IEEE Transactions on Emerging Topics in Computing, 10(4):1789– 1798, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yang:2022:IIL

[YMX22] Bo Yang, Chaofan Ma, and Xiaofang Xia. An interrelated imitation learning method for heterogeneous drone swarm coordination. IEEE Transactions on Emerging Topics in Computing, 10(4):1704– 1716, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yu:2020:SBS

[YNA^{+}20] Jintao Yu, Răzvan Nane, Imran Ashraf, Mottaqiallah Taouil, Said Hamdioui, Henk Corporaal, and Koen Bertels. Skeletonbased synthesis flow for computation-in-memory architectures. IEEE Transactions on Emerging Topics in Computing, 8(2):545– 558, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yao:2022:DER

[YPS^{+}22] Jianrong Yao, Lihui Pang, Yang Su, Zhi Zhang, Wei Yang, Anmin Fu, and Yansong Gao. Design and evaluate recomposited OR-AND-XOR-PUF. IEEE Transactions on Emerging Topics in Computing, 10(2):662– 677, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yin:2014:OSR

[YQ14] Chi-En Yin and Gang Qu.

Obtaining statistically random information from silicon physical unclonable functions. IEEE Transactions on Emerging Topics in Computing, 2(2):96–106, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yang:2015:CUB

[YQZ^{+}15] Jie Yang, Yuanyuan Qiao, Xinyu Zhang, Haiyang He, Fang Liu, and Gang Cheng. Characterizing user behavior in mobile Internet. IEEE Transactions on Emerging Topics in Computing, 3(1): 95–106, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yan:2022:RLB

[YSS^{+}22] Kunpeng Yan, Hangguan Shan, Tengxu Sun, Haoji Hu, Yingxiao Wu, Lu Yu, Zhaoyang Zhang, and Tony Q. S. Quek. Reinforcement learning-based mobile edge computing and transmission scheduling for video surveillance. IEEE Transactions on Emerging Topics in Computing, 10 (2):1142–1156, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yin:2022:ECP

[YW22] Bo Yin and Xuetao Wei. Efficient crowdsourced Paretooptimal queries over partial orders with quality guarantee. IEEE Transactions on Emerging Topics in Computing, 10(1):297–311, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yan:2022:NQN

[YXF^{+}22] Aibin Yan, Zhelong Xu, Xiangfeng Feng, Jie Cui, Zhili Chen, Tianming Ni, Zhengfeng Huang, Patrick Girard, and Xiaoqing Wen. Novel quadruple-node-upset-tolerant latch designs with optimized overhead for reliable computing in harsh radiation environments. IEEE Transactions on Emerging Topics in Computing, 10(1): 404–413, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yang:2015:SCS

[YYX15] Zuyuan Yang, Wei Yan, and Yong Xiang. On the security of compressed sensing-based signal cryptosystem. IEEE Transactions on Emerging Topics in Computing, 3(3): 363–371, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yang:2024:AIO

[YZG^{+}24] Yaoqi Yang, Bangning Zhang, Daoxing Guo, Renhui Xu, Chunhua Su, and Weizheng Wang. Age of information optimization for privacy-preserving mobile crowdsensing. IEEE Transactions on Emerging Topics in Computing, 12(1):281– 292, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Yao:2021:RRL

[YZZ^{+}21] Haipeng Yao, Bo Zhang, Peiying Zhang, Sheng Wu, Chunxiao Jiang, and Song Guo. RDAM: a reinforcement learning based dynamic attribute matrix representation for virtual network embedding. IEEE Transactions on Emerging Topics in Computing, 9(2): 901–914, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zheng:2015:AAF

[ZC15] Lei Zheng and Lin Cai.

AFDA: Asynchronous flipped diversity ALOHA for emerging wireless networks with long and heterogeneous delay. IEEE Transactions on Emerging Topics in Computing, 3(1):64–73, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2014:ESD

[ZCT^{+}14] Fan Zhang, Junwei Cao, Wei Tan, Samee U. Khan, Keqin Li, and Albert Y. Zomaya. Evolutionary scheduling of dynamic multitasking workloads for big-data analytics in elastic cloud. IEEE Transactions on Emerging Topics in Computing, 2(3): 338–351, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhou:2015:DPA

[ZCZ^{+}15] Yuchen Zhou, Xiaodao Chen, Albert Y. Zomaya, Lizhe Wang, and Shiyan Hu. A dynamic programming algorithm for leveraging probabilistic detection of energy theft in smart home. IEEE Transactions on Emerging Topics in Computing, 3(4): 502–513, December 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2016:GTB

[ZCZY16] Aiqing Zhang, Jianxin Chen, Liang Zhou, and Shui Yu. Graph theorybased QoE-driven cooperation stimulation for content dissemination in device-to-device communication. IEEE Transactions on Emerging Topics in Computing, 4(4):556–567, October/December 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhu:2013:FAP

[ZDLG13] Haojin Zhu, Suguo Du, Muyuan Li, and Zhaoyu Gao. Fairness-aware and privacy-preserving friend matching protocol in mobile social networks. IEEE Transactions on Emerging Topics in Computing, 1(1): 192–200, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2020:GEI

[ZEL20] Haijun Zhang, Didier El Baz, and Victor C. M. Leung. Guest Editor’s introduction: Special section on green computing in Internet of Things. IEEE Transactions on Emerging Topics in Computing, 8(3):750–751, July/September 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2014:CSR

[ZHC^{+}14] Daqiang Zhang, Ching-Hsien Hsu, Min Chen, Quan Chen, Naixue Xiong, and Jaime Lloret. Cold-start recommendation using biclustering and fusion for large-scale social recommender systems. IEEE Transactions on Emerging Topics in Computing, 2(2): 239–250, June 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2014:CRS

[ZHL^{+}14] Desheng Zhang, Tian He, Yunhuai Liu, Shan Lin, and John A. Stankovic. A carpooling recommendation system for taxicab services. IEEE Transactions on Emerging Topics in Computing, 2(3):254– 266, September 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2021:DRM

[ZK21] Da Zhang and Mansur R. Kabuka. Distributed relationship mining over big scholar data. IEEE Transactions on Emerging Topics in Computing, 9(1):354–365, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhou:2023:ECC

[ZK23] Jun Zhou and Masaaki Kondo. An edge–cloud collaboration framework for graph processing in smart society. IEEE Transactions on Emerging Topics in Computing, 11(4):985– 1001, October/December 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zheng:2022:ALE

[ZL22] Hao Zheng and Ahmed Louri. Agile: a learningenabled power and performanceefficient network-on-chip design. IEEE Transactions on Emerging Topics in Computing, 10(1):223–236, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2016:DMC

[ZLC16] Hongke Zhang, Hongbin Luo, and Han-Chieh Chao. Dealing with mobilitycaused outdated mappings in networks with identifier/ locator separation. IEEE Transactions on Emerging Topics in Computing, 4(2): 199–213, April/June 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhou:2022:SSA

[ZLG^{+}22] Xiaoping Zhou, Zhenlong Liu, Maozu Guo, Jichao Zhao, and Jialin Wang. SACC: a size adaptive content caching algorithm in fog/edge computing using deep reinforcement learning. IEEE Transactions on Emerging Topics in Computing, 10(4):1810–1820, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhou:2021:AIA

[ZLW^{+}21] Xiaokang Zhou, Wei Liang, Kevin I-Kai Wang, Runhe Huang, and Qun Jin. Academic influence aware and multidimensional network analysis for research collaboration navigation based on scholarly Big Data. IEEE Transactions on Emerging Topics in Computing, 9(1): 246–257, January/March 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2022:DRA

[ZLW^{+}22] Chenyan Zhang, Jing Li, Jia Wu, Donghua Liu, Jun Chang, and Rong Gao. Deep recommendation with adversarial training. IEEE Transactions on Emerging Topics in Computing, 10(4):1966–1978, October/ December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhu:2014:MRB

[ZLX^{+}14] Yuqing Zhu, Deying Li, Wen Xu, Weili Wu, Lidan Fan, and James Willson. Mutualrelationship-based community partitioning for social networks. IEEE Transactions on Emerging Topics in Computing, 2(4):436– 447, December 2014. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zheng:2018:DRT

[ZLX18] Yu-Jun Zheng, Hai-Feng Ling, and Jin-Yun Xue. Disaster rescue task scheduling: an evolutionary multiobjective optimization approach. IEEE Transactions on Emerging Topics in Computing, 6(2):288–300, April/June 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zeng:2022:DAN

[ZLXL22] Yuejie Zeng, Kenli Li, Lin Xiao, and Qing Liao. Design and analysis of a novel integral design scheme for finding finite-time solution of time-varying matrix inequalities. IEEE Transactions on Emerging Topics in Computing, 10(1):267–279, January/March 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhong:2021:SFA

[ZLY^{+}21] Xiaogang Zhong, Mingzhen Li, Hailong Yang, Yi Liu, and Depei Qian. swMR: a framework for accelerating MapReduce applications on Sunway Taihulight. IEEE Transactions on Emerging Topics in Computing, 9 (2):1020–1030, April/June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2017:ICS

[ZLZ^{+}17] Xing Zhang, Yongjing Li, Yan Zhang, Jiaxin Zhang, Hailing Li, Shuo Wang, and Danvang Wang. Information caching strategy for cyber social computing based wireless networks. IEEE Transactions on Emerging Topics in Computing, 5(3): 391–402, July/September 2017. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2021:BBT

[ZLZ^{+}21] Haibin Zhang, Jiajia Liu, Huanlei Zhao, Peng Wang, and Nei Kato. Blockchainbased trust management for Internet of Vehicles. IEEE Transactions on Emerging Topics in Computing, 9(3): 1397–1409, July/September 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhou:2015:AAC

[ZMA15] Hui Zhou, Shiwen Mao, and Prathima Agrawal. Approximation algorithms for cell association and scheduling in femtocell networks. IEEE Transactions on Emerging Topics in Computing, 3(3): 432–443, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhu:2022:DTS

[ZMK22] Yishi Zhu, Bomin Mao, and Nei Kato. A dynamic task scheduling strategy for multi-access edge computing in IRS-aided vehicular networks. IEEE Transactions on Emerging Topics in Computing, 10(4):1761– 1771, October/December 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zeinali:2019:NNB

[ZMRM19] Behzad Zeinali, Jens Kargaard Madsen, Praveen Raghavan, and Farshad Moradi. A novel nondestructive bit-line discharging scheme for deep submicrometer STT-RAMs. IEEE Transactions on Emerging Topics in Computing, 7(2): 294–300, April/June 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhao:2022:SFK

[ZMX^{+}22] Haoyu Zhao, Weidong Min, Jianqiang Xu, Qing Han, Wei Li, Qi Wang, Ziyuan Yang, and Linghua Zhou. SPACE: Finding keyspeaker in complex multiperson scenes. IEEE Transactions on Emerging Topics in Computing, 10(3):1645– 1656, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zacharelos:2022:ARM

[ZNS^{+}22] Efstratios Zacharelos, Italo Nunziata, Gerardo Saggese, Antonio G. M. Strollo, and Ettore Napoli. Approximate recursive multipliers using low power building blocks. IEEE Transactions on Emerging Topics in Computing, 10(3):1315– 1330, July/September 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zambonelli:2016:CLS

[ZOS16] Franco Zambonelli, Andrea Omicini, and Paul Scerri. Coordination in large-scale socio-technical systems: Introduction to the special section. IEEE Transactions on Emerging Topics in Computing, 4(1):5–8, January/March 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhu:2015:SLI

[ZS15] Qiusha Zhu and Mei-Ling Shyu. Sparse linear integration of content and context modalities for semantic concept retrieval. IEEE Transactions on Emerging Topics in Computing, 3(2):152–160, June 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhou:2020:YTT

[ZSCY20] Lu Zhou, Chunhua Su, Wayne Chiu, and Kuo-Hui Yeh. You think, therefore you are: Transparent authentication system with brainwave-oriented bio-features for IoT networks. IEEE Transactions on Emerging Topics in Computing, 8(2):303–312, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhu:2015:TOM

[ZSL^{+}15] Chunsheng Zhu, Zhengguo Sheng, Victor C. M. Leung, Lei Shu, and Laurence T. Yang. Toward offering more useful data reliably to mobile cloud from wireless sensor network. IEEE Transactions on Emerging Topics in Computing, 3(1): 84–94, March 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhao:2024:FHL

[ZWH^{+}24] Bo-Wei Zhao, Lei Wang, Peng-Wei Hu, Leon Wong, Xiao-Rui Su, Bao-Quan Wang, Zhu-Hong You, and Lun Hu. Fusing higher and lower-order biological information for drug repositioning via graph representation learning. IEEE Transactions on Emerging Topics in Computing, 12(1):163–176, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhu:2023:TLR

[ZWPL23] Xiubin Zhu, Dan Wang, Witold Pedrycz, and Zhiwu Li. Transfer learning realized with nonlinearly transformed input space. IEEE Transactions on Emerging Topics in Computing, 11(2): 448–460, April/June 2023. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2013:IMC

[ZWWF13] Zhaoyang Zhang, Honggang Wang, Chonggang Wang, and Hua Fang. Interference mitigation for cyberphysical wireless body area network system using social networks. IEEE Transactions on Emerging Topics in Computing, 1(1):121–132, June 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2015:MES

[ZWWF15] Zhaoyang Zhang, Honggang Wang, Chonggang Wang, and Hua Fang. Modeling epidemics spreading on social contact networks. IEEE Transactions on Emerging Topics in Computing, 3(3): 410–419, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2018:TVS

[ZWZ18] Weiwen Zhang, Yonggang Wen, and Xinwen Zhang. Towards virus scanning as a service in mobile cloud computing: Energy-efficient dispatching policy under Nversion protection. IEEE Transactions on Emerging Topics in Computing, 6(1): 122–134, January/March 2018. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2021:FPP

[ZXLS21] Pengcheng Zhang, Fang Xiong, Hareton Leung, and Wei Song. FunkR-pDAE: Personalized project recommendation using deep learning. IEEE Transactions on Emerging Topics in Computing, 9(2):886–900, April/ June 2021. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2013:MBP

[ZXXH13] Yuexin Zhang, Li Xu, Yang Xiang, and Xinyi Huang. A matrix-based pairwise key establishment scheme for wireless mesh networks using pre deployment knowledge. IEEE Transactions on Emerging Topics in Computing, 1(2):331–340, December 2013. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhong:2022:MIE

[ZYC^{+}22] Jinghui Zhong, Jiaquan Yang, Yongliang Chen, Wei-Li Liu, and Liang Feng. Mining implicit equations from data using gene expression programming. IEEE Transactions on Emerging Topics in Computing, 10 (2):1058–1074, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhuang:2020:EED

[ZYH^{+}20] Yunhui Zhuang, Anjia Yang, Gerhard P. Hancke, Duncan S. Wong, and Guomin Yang. Energyefficient distance-bounding with residual charge computation. IEEE Transactions on Emerging Topics in Computing, 8(2):365–376, April/June 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2019:RTA

[ZYLT19] Xinglin Zhang, Zheng Yang, Yunhao Liu, and Shaohua Tang. On reliable task assignment for spatial crowdsourcing. IEEE Transactions on Emerging Topics in Computing, 7(1):174–186, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zeng:2016:HSL

[ZYMG16] Jing Zeng, Laurence T. Yang, Jianhua Ma, and Minyi Guo. Hyperspace-Flow: a system-level design methodology for smart space. IEEE Transactions on Emerging Topics in Computing, 4(4):568–583, October/December 2016. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zheng:2022:CEF

[ZYS^{+}22] Yu-Jun Zheng, Si-Lan Yu, Qin Song, Yu-Jiao Huang, Wei-Guo Sheng, and Sheng-Yong Chen. Co-evolutionary fuzzy deep transfer learning for disaster relief demand forecasting. IEEE Transactions on Emerging Topics in Computing, 10(3):1361– 1373, July/September 2022.

ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhou:2024:MSP

[ZYY^{+}24] Dongbo Zhou, Hongwei Yu, Jie Yu, Shuai Zhao, Wenhui Xu, Qianqian Li, and Fengyin Cai. MFDS-STGCN: Predicting the behaviors of college students with fine-grained spatial-temporal activities data. IEEE Transactions on Emerging Topics in Computing, 12(1):254–265, January/March 2024. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhao:2020:TBM

[ZYZ20] Yaliang Zhao, Laurence T.

Yang, and Ronghao Zhang. Tensor-based multiple clustering approaches for cyberphysical-social applications. IEEE Transactions on Emerging Topics in Computing, 8(1):69–81, January/March 2020. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2015:CBS

[ZYZZ15] Sihai Zhang, Dandan Yin, Yanqin Zhang, and Wuyang Zhou. Computing on base station behavior using Erlang measurement and call detail record. IEEE Transactions on Emerging Topics in Computing, 3(3):444– 453, September 2015. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhan:2022:SIM

[ZZH^{+}22] Yufeng Zhan, Jie Zhang, Zicong Hong, Leijie Wu, Peng Li, and Song Guo. A survey of incentive mechanism design for federated learning. IEEE Transactions on Emerging Topics in Computing, 10(2):1035–1044, April/June 2022. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zuolo:2019:LSD

[ZZM^{+}19] Lorenzo Zuolo, Cristian Zambelli, Alessia Marelli, Rino Micheloni, and Piero Olivo. LDPC soft decoding with improved performance in 1X-2X MLC and TLC NAND flash-based solid state drives. IEEE Transactions on Emerging Topics in Computing, 7(3):507–515, July/September 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Zhang:2024:FEM

[ZZZY24] Yue Zhang, Youwen Zhu, Yuqian Zhou, and Jiabin Yuan. Frequency estimation mechanisms under ϵδutility-optimized local differential privacy. IEEE Transactions on Emerging Topics in Computing, 12(1): 316–327, January/March 2024. ISSN 2168-6750 REFERENCES 203

(print), 2376-4562 (electronic).